### v1

# SLRQA: A Sparse Low-Rank Quaternion Model for Color Image Processing with Convergence Analysis

Abstract
In this paper, we propose a Sparse Low-rank Quaternion Approximation (SLRQA) model for color image processing problems with noisy observations. The proposed SLRQA is a quaternion model that combines low-rankness and sparsity priors. Furthermore, it does not need an initial rank estimate. To solve the SLRQA model, a nonconvex proximal linearized ADMM (PL-ADMM) algorithm is proposed. Furthermore, the global convergence analysis of the PL-ADMM under mild assumptions is presented. When the observation is noise-free, an SLRQA-NF model of the limiting case of the SLRQA is proposed. Subsequently, a nonconvex proximal linearized ADMM (PL-ADMM-NF) algorithm for the SLRQA-NF is given. In numerical experiments, we verify the effectiveness of quaternion representation. Furthermore, for color image denoising and color image inpainting problems, SLRQA and SLRQA-NF models demonstrate superior performance both quantitatively and visually when compared with some state-of-the-art methods.
Key words: Color image denoising, Quaternion matrix completion, Nonconvex linearized ADMM.
1 Introduction
Low-rankness plays an important role in image processing and has been used in image denoising [10], image inpainting [15], image deblurring [26], and image filtering [7]. A commonly encountered approach for exploiting the low-rankness is to find a matrix with the lowest rank such that certain matching errors are minimized. Due to the existence of the rank function, such problems are known as NP-hard [25]. To overcome this difficulty, Candès et al. [6] have proved that the nuclear norm is the tightest convex relaxation of the NP-hard rank minimization function. In addition to the nuclear norm, other surrogates have also been proposed, such as weighted nuclear norm (WNNM) [14], Schatten $p$-norm [24], weighted Schatten $p$-norm [32], and log-determinant penalty [19], which have shown competitive effectiveness in various applications. In addition to low-rankness, sparsity is also a significant property of natural images that has been considered in image recovery [35, 5], video recovery [34], texture repairing [21], and face recognition [40]. The intuition behind sparsity is that images under certain transforms are usually sparse, such as wavelet or discrete cosine transform (DCT). Moreover, quaternion representation has shown great potential for color image processing problems. A color image can be encoded as a pure quaternion matrix. As claimed in [10], the main advantage of quaternion representation is that a color image is treated as a pure quaternion matrix naturally and holistically, meaning that the correlation information among the three RGB channels is used. It is empirically shown in [10] that methods using quaternion representation outperform methods with monochromatic techniques. Moreover, quaternion has been increasingly used for color image processing in the past two decades [10].
These three priors, i.e. low-rankness, sparsity, and quaternion representation, all play important roles in image processing. However, models based on all these three priors are still limited. In the following subsection, we propose a quaternion-based model that combines low-rankness and sparsity properties to address problems arising in image processing.
1.1 A novel quaternion based model
Let $\mathbf{\dot{Y}}\in\mathbb{H}^{m\times n}$ denote an observed color image satisfying $\mathbf{\dot{Y}}=\mathcal{A}({\mathbf{\dot{X}}})+\bm{\varepsilon}$, where ${\mathbf{\dot{X}}}\in\mathbb{H}^{m\times n}$ is the desired noiseless color image, $\mathbb{H}^{m\times n}$ denotes the set of all $m\times n$ quaternion matrices, $\mathcal{A}$ denotes a linear operator, and $\bm{\varepsilon}$ is a random noise [^1] The definition of quaternion matrices is clarified in Section 2.. According to the principle of maximizing the posterior probability, an approach to recover the image ${\mathbf{\dot{X}}}$ is to maximize likelihood estimation function, i.e., ${\mathbf{\dot{X}}}_{*}=\operatorname*{arg\,max}_{\mathbf{\dot{X}}}p({\mathbf{\dot{X}}}|\mathbf{\dot{Y}}).$ It follows from the Bayes rule that $p({\mathbf{\dot{X}}}|\mathbf{\dot{Y}})=\frac{p(\mathbf{\dot{Y}}|{\mathbf{\dot{X}}})p({\mathbf{\dot{X}}})}{p(\mathbf{\dot{Y}})}\propto p(\mathbf{\dot{Y}}|{\mathbf{\dot{X}}})p({\mathbf{\dot{X}}}).$ By taking the logarithm, an equivalent optimization problem is given as follows:
${\mathbf{\dot{X}}}_{*}$ $=\mathop{\mathrm{arg\,min}}_{{\mathbf{\dot{X}}}}-\log p(\mathbf{\dot{Y}}|{\mathbf{\dot{X}}})-\log p({\mathbf{\dot{X}}})=\mathop{\mathrm{arg\,min}}_{{\mathbf{\dot{X}}}}E(\mathbf{\dot{Y}};{\mathbf{\dot{X}}})+R({\mathbf{\dot{X}}}),$ (1)
where $E(\mathbf{\dot{Y}};{\mathbf{\dot{X}}})$ is the negative log-likelihood function and is also called the data fitting term, and $R({\mathbf{\dot{X}}})$ is the prior term. It is assumed throughout this paper that the entries of the noise $\bm{\varepsilon}$ are drawn from the Gaussian distribution with mean zero and variance $\tau^{2}$. It follows that $E(\mathbf{\dot{Y}};{\mathbf{\dot{X}}})=\frac{1}{2\tau^{2}}\|\mathcal{A}({\mathbf{\dot{X}}})-\mathbf{\dot{Y}}\|_{\mathrm{F}}^{2}$ if all the entries of $\mathbf{\dot{Y}}$ are observed, and $E(\mathbf{\dot{Y}};{\mathbf{\dot{X}}})=\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2}$ if only partial entries of $\mathbf{\dot{Y}}$ are observed, where $\Omega$ denotes the indices of the observed entries, that is, $(\mathcal{P}_{\Omega}(\mathbf{\dot{Y}}))_{ij}=\mathbf{\dot{Y}}_{ij}$ if $(i,j)\in\Omega$ and $(\mathcal{P}_{\Omega}(\mathbf{\dot{Y}}))_{ij}=0$ if $(i,j)\notin\Omega$.
By integrating the low-rankness and sparsity into the prior term, i.e., $R({\mathbf{\dot{X}}})=\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)+\lambda p(\mathcal{W}({\mathbf{\dot{X}}}))$, and letting $\mathbf{\dot{W}}=\mathcal{W}({\mathbf{\dot{X}}})$, we propose the following Sparse Low-rank Quaternion Approximation (SLRQA) model for color image processing:
$\min_{{\mathbf{\dot{X}}},\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}}$ $\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)+\lambda p(\mathbf{\dot{W}})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2},$ (2)
$\mathrm{s.t.}$ $\mathcal{W}({\mathbf{\dot{X}}})=\mathbf{\dot{W}},$
where ${\mathbf{\dot{X}}}\in\mathbb{H}^{m\times n}$ denotes a color image, $\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}$ is the coefficient matrix under an orthogonal transform $\mathcal{W}$ such as quaternion wavelet transform or quaternion discrete cosine transform (QDCT), $\mathcal{A}$ is a known operator such as a total variation operator, a blurring operator or the Radon transform, $p(\mathbf{\dot{W}})$ is a function that has Lipschitz continuously differentiable and promotes the sparsity of $\mathbf{\dot{W}}$, $\phi$ is a nonconvex surrogate function that substitutes the rank function, $\sigma_{i}({\mathbf{\dot{X}}})$ is the $i$-th singular value of ${\mathbf{\dot{X}}}$ defined in Section 2, $\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)$ promotes the low-rankness of ${\mathbf{\dot{X}}}$. Some nonconvex surrogate functions for $\phi$ have been given in [10, 23] and listed in Table 1.
Table 1: nonconvex surrogate functions $\phi(x,\gamma)$
Nonconvex function Log-determinant Schatten-$\gamma$ Logarithm Laplace [9] Weighted Schatten-$\gamma$ [32] ETP
$\phi(x,\gamma)$ log$(1+x^{2})$ $x^{\gamma}$ log($\gamma$+$x$) $1-e^{-{\frac{x}{\gamma}}}$ $wx^{\gamma}$ $\frac{\lambda(1-\exp(-\gamma x))}{1-\exp(-\gamma)}$
The $\ell_{1}$ norm has been widely used to promote sparsity. However, the Lipschitz continuous differentiability of $p(\mathbf{\dot{W}})$ is assumed in our convergence analysis [^2] We use a variant of the ADMM method to solve SLRQA. Note that such an assumption has been made for all the ADMM methods with global convergence guaranteed for nonconvex problems as far as we know, see, e.g., a review in [22]. Therefore, a smoothing function such as the Huber function can be used as a substitution for $p(\mathbf{\dot{W}})$. It is empirically shown in Section 6 that such a substitution yields satisfactory performance. This strategy has been used in [38] to guarantee convergence theoretically.
When the observed color image is noise-free, or the variance $\tau^{2}$ in the noise $\bm{\varepsilon}$ is zero, Model (2) is undefined due to the existence of a zero in the denominator of the data fitting term. In this case, one can reformulate the data fitting term as a constraint. It follows that a variant of SLRQA for the noise-free problem is given by
$\min_{{\mathbf{\dot{X}}},\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}}$ $\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)+\lambda p(\mathbf{\dot{W}}),$ (3)
$\mathrm{s.t.}$ $\mathcal{W}({\mathbf{\dot{X}}})=\mathbf{\dot{W}},\quad\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}))=\mathcal{P}_{\Omega}(\mathbf{\dot{Y}}).$
Since $\mathcal{W}$ is an orthogonal transform such as quaternion wavelet transform or quaternion discrete cosine transform (QDCT) [12], we have $\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)=\sum_{i}\phi(\sigma_{i}(\mathbf{\dot{Q}}_{1}\mathbf{\dot{W}}\mathbf{\dot{Q}}_{2}^{*}),\gamma)=\sum_{i}\phi(\sigma_{i}(\mathbf{\dot{W}}),\gamma)$, the superscript * denotes the conjugate transpose operator. Let $\mathbf{\dot{Z}}=\mathcal{W}({\mathbf{\dot{X}}})$. We propose a model named SLRQA-NF which is equivalent to (3) for noise-free problems:
$\min_{\mathbf{\dot{Z}},\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}}$ $\sum_{i}\phi(\sigma_{i}(\mathbf{\dot{Z}}),\gamma)+\lambda p(\mathbf{\dot{W}}),$ (4)
$\mathrm{s.t.}$ $\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}))=\mathcal{P}_{\Omega}(\mathbf{\dot{Y}}),\quad\mathbf{\dot{Z}}=\mathbf{\dot{W}},$
where the superscript ##\## denotes the adjoint operator which is defined in Section 2.
1.2 Related work
The derivations of SLRQA and SLRQA-NF models are motivated by LADM [21] and LRQA [10] models. To be specific, LADM is a convex model that utilizes the sparsity and low-rankness properties while quaternion representation and nonconvex surrogate rank function have not been considered. LRQA is a quaternion-based model that makes use of the low-rankness property by replacing the nuclear norm with nonconvex surrogate functions, but it has not taken advantage of sparsity. SLRQA and SLRQA-NF models take advantage of LADM and LRQA by using quaternion representation, low-rankness, and sparsity. Next, we discuss some image processing models related to SLRQA and SLRQA-NF.
Image denoising (noisy observation and identity linear operator $\mathcal{A}$):
If the low-rankness prior is omitted, i.e. $\phi\equiv 0$, the Lipschitz continuous gradients assumption is dropped, and $p(\mathbf{\dot{W}})=\|\mathbf{\dot{W}}\|_{1}$, then the classical analysis based approach [5] with an extension to quaternion representation is presented:
$\min_{{\mathbf{\dot{X}}}\in\mathbb{H}^{m\times n}}$ $\lambda\|\mathcal{W}({\mathbf{\dot{X}}})\|_{1}+\frac{1}{2\tau^{2}}\|{\mathbf{\dot{X}}}-\mathbf{\dot{Y}}\|_{\mathrm{F}}^{2},$ (5)
where the variable $\mathbf{\dot{W}}$ is eliminated by $\mathbf{\dot{W}}=\mathcal{W}({\mathbf{\dot{X}}})$. If the sparse term $p(\mathbf{\dot{W}})$ is not used, then SLRQA becomes the LRQA denoising model in [10, Equation 20]
$\min_{{\mathbf{\dot{X}}}\in\mathbb{H}^{m\times n}}\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)+\frac{1}{2\tau^{2}}\|{\mathbf{\dot{X}}}-\mathbf{\dot{Y}}\|_{\mathrm{F}}^{2}.$
As noted in [10], LRQA is a generalization of other models such as the WNNM model.
Image deblurring (noisy observation and blurring kernel operator $\mathcal{A}$):
If the low-rankness prior is not used, i.e., $\phi\equiv 0$ and let $p(\mathbf{\dot{W}})=\|\mathbf{\dot{W}}\|_{1}$, then the standard analysis based model for image deblurring in [5] with extensions to quaternion representation is presented
$\min_{{\mathbf{\dot{X}}}\in\mathbb{H}^{m\times n}}\lambda\|\mathcal{W}({\mathbf{\dot{X}}})\|_{1}+\frac{1}{2\tau^{2}}\|\mathcal{A}({\mathbf{\dot{X}}})-\mathbf{\dot{Y}}\|_{\mathrm{F}}^{2}.$
Image inpainting (partially observed images): If the observed image is noise-free, the linear operator $\mathcal{A}$ is the identity, $p(\mathbf{\dot{W}})=\|\mathbf{\dot{W}}\|_{1}$, and $\phi$ is the nuclear norm, then the SLRQA-NF model becomes the LADM model in [21] with an extension to quaternion representation:
$\min_{\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}}$ $\|\mathbf{\dot{W}}\|_{*}+\lambda\|\mathbf{\dot{W}}\|_{1},\quad\mathrm{s.t.}\,\mathcal{P}_{\Omega}(\mathcal{W}^{\#}(\mathcal{\mathbf{\dot{W}}}))=\mathcal{P}_{\Omega}(\mathbf{\dot{Y}}).$ (6)
Furthermore, if the sparsity term $\phi$ is dropped and a nonconvex surrogate is used, then SLRQA-NF model becomes the LRQA [10, Equation 22] given by
$\min_{{\mathbf{\dot{X}}}\in\mathbb{H}^{m\times n}}$ $\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma),\quad\mathrm{s.t.}\,\mathcal{P}_{\Omega}(\mathcal{\mathcal{{\mathbf{\dot{X}}}}})=\mathcal{P}_{\Omega}(\mathbf{\dot{Y}}).$ (7)
The above-reviewed models do not simultaneously use the low-rankness, sparsity, and quaternion representation. Recently, a quaternion based model combined with low-rankness and sparsity has been proposed in [17]. Specifically, the model is given by:
$\min_{\mathbf{\dot{L}}\in\mathbb{H}^{m\times r},\mathbf{\dot{D}}\in\mathbb{H}^{r\times r},\mathbf{\dot{R}}\in\mathbb{H}^{r\times n},{\mathbf{\dot{X}}}\in\mathbb{H}^{m\times n},\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}}\,\|\mathbf{\dot{D}}\|_{*}+\lambda\|\mathbf{\dot{W}}\|_{1},$ (8)
$\mathrm{s.t.}$ $\mathbf{\dot{L}}^{*}\mathbf{\dot{L}}=\mathbf{I}_{r},\,\mathbf{\dot{R}}\mathbf{\dot{R}}^{*}=\mathbf{I}_{r},~{}~{}\mathcal{W}({\mathbf{\dot{X}}})=\mathbf{\dot{W}},~{}~{}{\mathbf{\dot{X}}}=\mathbf{\dot{L}}\mathbf{\dot{D}}\mathbf{\dot{R}},~{}~{}\mathcal{P}_{\Omega}(\mathbf{\dot{Y}}-\mathbf{\dot{L}}\mathbf{\dot{D}}\mathbf{\dot{R}})=\mathbf{0}.$
where $r$ is a prescribed number and the superscript * denotes the conjugate transpose operator, see its definition in Section 2. Different from SLRQA models, Model (8) uses the nuclear norm to promote low-rank instead of nonconvex surrogate functions $\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)$. Furthermore, it needs to estimate the rank $r$ in the initialization phase while SLRQA does not. Another model that uses the three priors is proposed in [34, 35] for color image and video recovery.
The model is given by
$\min_{{\mathbf{\dot{X}}}}$ $\|{\mathbf{\dot{X}}}\|_{*}-\sum_{i=r}^{\min(m,n)}\sigma_{i}({\mathbf{\dot{X}}})+\lambda\|\mathbf{\dot{W}}\|_{1},$ (9)
$\mathrm{s.t.}$ $\mathcal{P}_{\Omega}({\mathbf{\dot{X}}}-\mathbf{\dot{Y}})=0,\,\mathcal{W}({\mathbf{\dot{X}}})=\mathbf{\dot{W}}.$
Note that Model (9) also needs to estimate the rank $r$ and the low-rankness is promoted by penalizing a truncated nuclear norm, which is convex. Moreover, convergence analyses of the proposed algorithms for solving the existing models in [17, 34, 35] with three priors are not given.
In this paper, proximal linearized ADMM algorithms are proposed to solve Models (2) and (4). Note that the convergence of nonconvex ADMM and its variants has been studied widely. However, as far as we know, the existing ADMM and its variants either consider different optimization problems or use different optimization techniques. A nonexhaustive review is given as follows. Wang et al. [29] comprehensively study the convergence of ADMM for nonconvex and nonsmooth objective functions subject to coupled linear equality constraints. Their assumptions require that the solution of the subproblem is Lipschitz continuous with respect to the input, which is often overly stringent and hard to verify for practical applications. Hong et al. [18] study the convergence of ADMM for solving nonconvex consensus problems, but it is under the case where every constraint matrix has full column rank. Compared with the above analyses, our convergence analysis only needs one constraint matrix that has full column rank. In [36], Yashtini proposes a proximal linearized ADMM for solving linearly constrained nonconvex and possibly nonsmooth optimization problems. The proposed method employs a linearization strategy that targets both the quadratic penalty term and the objective function. Above all, all the nonconvex ADMM and their variant algorithms for nonconvex situations are different from PL-ADMM where only one variable is linearized in the situation. For the reviewed algorithms, the first two algorithms are traditional ADMM algorithms for nonconvex problems, while the ADMM of the last algorithm uses the technique of linearization for both two variables. For a detailed review of other non-convex ADMM algorithms, we refer the readers to [22].
1.3 Contribution
The main contributions of this paper are listed as follows:
  • We propose a novel SLRQA model to address image processing problems with noise by maximizing the a posteriori probability estimate. The proposed SLRQA takes advantage of quaternion representation, low-rankness, and sparsity without requiring an estimation of the rank.
  • We give a proximal linearized alternating direction method of multipliers (PL-ADMM) algorithm. Under mild conditions, we prove that the sequence generated by the algorithm converges to a stationary point. To the best of our knowledge, this is the first convergence analysis result of an algorithm for a quaternion-based image processing model.
  • For noise-free observations, we propose a variant of SLRQA called SLRQA-NF. Inspired by PL-ADMM, an algorithm named PL-ADMM-NF is presented to solve the SLRQA-NF model.
  • In the numerical experiments, we verify that the quaternion representation outperforms the RGB monochromatic representation. Moreover, compared with some representative state-of-the-art image processing models, the promising performance of SLRQA and SLRQA-NF demonstrates their superiority and robustness numerically and visually.
1.4 Organization
The rest of this paper is organized as follows. In Section 2, the preliminaries and notations are introduced. The PL-ADMM algorithm framework for solving SLRQA model is given in Section 3. The global convergence is established in Section 4. In Section 5, the PL-ADMM-NF algorithm for solving SLRQA-NF model is presented. In Section 6, extensive numerical experiments are conducted to demonstrate the performance of the models and algorithms by comparing them with other state-of-the-art methods. Finally, the conclusion is given in Section 7.
2 Preliminaries
Throughout this paper, scalars, real vectors, real matrices, and real tensor are respectively denoted by lowercase letters $x$, boldface lowercase letters $\mathbf{x}$, boldface capital letters $\mathbf{X}$, and boldface capital letters with a tilde $\widetilde{\mathbf{X}}$ respectively. $\mathbb{H}$ denotes the quaternion algebra proposed in [16]. A dot above a variable $\dot{q}$ represents a quaternion number. A quaternion number consists of one real part and three imaginary parts:
$\dot{q}=x_{s}+x_{i}\mathbf{i}+x_{j}\mathbf{j}+x_{k}\mathbf{k}.$
where $x_{s},x_{i},x_{j},x_{k}\in\mathbb{R}$, and $\mathbf{i},\mathbf{j},\mathbf{k}$ are three imaginary units. The quaternion $\dot{q}$ is called a pure quaternion if and only if it has a zero real part, $\mathbf{\;i.e.\;}x_{s}=0$. The real part of $\dot{q}$ is $\text{Re}(\dot{q})=x_{s}$. The conjugate $\overline{\dot{q}}$ is defined as $\overline{\dot{q}}=x_{s}-x_{i}\mathbf{i}-x_{j}\mathbf{j}-x_{k}\mathbf{k}$. The modulus $|\dot{q}|$ is defined as $|\dot{q}|=\sqrt{x_{s}^{2}+x_{i}^{2}+x_{j}^{2}+x_{k}^{2}}$. The inverse of $\dot{q}$ is given by $\dot{q}^{-1}=\dfrac{\overline{\dot{q}}}{|\dot{q}|^{2}}$. For a quaternion matrix ${\mathbf{\dot{X}}},$ ${\mathbf{\dot{X}}}_{ij}$ denotes the $i$-th row, $j$-th coloum element of ${\mathbf{\dot{X}}}_{ij}$ and ${\mathbf{\dot{X}}}_{ijk}$ to denote the $i$-th row, $j$-th and $k$-th imaginary units. The Frobenius norm of ${\mathbf{\dot{X}}}$ is defined as $\|{\mathbf{\dot{X}}}\|_{\mathrm{F}}:=\sqrt{\sum_{i=1}^{n}\sum_{j=1}^{m}|{\mathbf{\dot{X}}}_{ij}|^{2}}$. Accordingly, the conjugate operator $\overline{\mathbf{\dot{A}}}$ , transpose operator $\mathbf{\dot{A}}^{\mathrm{T}}$ and the conjugate transpose operator $\mathbf{\dot{A}}^{*}$ are defined as $\overline{\mathbf{\dot{A}}}=(\overline{\dot{a}})_{ij},\mathbf{\dot{A}}^{\mathrm{T}}=(\dot{a})_{ji}$ and $\mathbf{\dot{A}}^{*}=(\overline{\dot{a}})_{ji}$, respectively. Furthermore, the multiplication rule of quaternion is given by: $\mathbf{i}\cdot 1=\mathbf{i},\quad\mathbf{j}\cdot 1=\mathbf{j},\quad\mathbf{k}\cdot 1=\mathbf{k}.\quad$ $\mathbf{i}^{2}=\mathbf{j}^{2}=\mathbf{k}^{2}=-1,\mathbf{i}\mathbf{j}=-\mathbf{j}\mathbf{i}=\mathbf{k},\,\mathbf{j}\mathbf{k}=-\mathbf{k}\mathbf{j}=\mathbf{i},\,\mathbf{k}\mathbf{i}=-\mathbf{i}\mathbf{k}=\mathbf{j}$. Note that the multiplication of two quaternions $\dot{q}_{1}$ and $\dot{q}_{2}$ is not commutative, i.e., $\dot{q}_{1}\dot{q}_{2}\neq\dot{q}_{2}\dot{q}_{1}$ in general. $\mathbf{\dot{Q}}$ is called a unitary matrix if and only if $\mathbf{\dot{Q}}\mathbf{\dot{Q}}^{*}=\mathbf{\dot{Q}}^{*}\mathbf{\dot{Q}}=\mathbf{I},$ where $\mathbf{I}$ denotes a real identity matrix. The inner product of two quaternion matrices $\mathbf{\dot{G}}_{1},\mathbf{\dot{G}}_{2}\in\mathbb{H}^{m\times n}$ is defined as $\left\langle\mathbf{\dot{G}}_{1},\mathbf{\dot{G}}_{2}\right\rangle:=\text{tr}(\mathbf{\dot{G}}_{1}^{*}\mathbf{\dot{G}}_{2})$. The image of a given matrix $\mathbf{\dot{G}}_{1}$ is denoted by ${\rm range}(\mathbf{\dot{G}}_{1})$. Given an operator $\mathcal{W}:\mathbb{H}^{m\times n}\rightarrow\mathbb{H}^{m\times n}$, the weighted norm is defined as $\|{\mathbf{\dot{X}}}\|_{\mathcal{W}}:=\left\langle\mathcal{W}({\mathbf{\dot{X}}}),{\mathbf{\dot{X}}}\right\rangle$. The operator norm is defined as $\|\mathcal{W}\|:=\sup_{{\mathbf{\dot{X}}}}\frac{\|\mathcal{W}({\mathbf{\dot{X}}})\|_{\mathrm{F}}}{\|{\mathbf{\dot{X}}}\|_{\mathrm{F}}}$. Let $\mathcal{W}^{\#}$ denote the adjoint operator of $\mathcal{W}$, i.e., $\left\langle\mathcal{W}(\mathbf{\dot{G}}_{1}),\mathbf{\dot{G}}_{2}\right\rangle=\left\langle\mathbf{\dot{G}}_{1},\mathcal{W}^{\#}(\mathbf{\dot{G}}_{2})\right\rangle$ holds for any $\mathbf{\dot{G}}_{1},\mathbf{\dot{G}}_{2}\in\mathbb{H}^{m\times n}$. If $\mathcal{W}$ is given by a quaternion matrix, then $\mathcal{W}^{\#}$ is the conjugate transpose of the matrix. For $\widetilde{\mathbf{\dot{G}}}:=(\mathbf{\dot{G}}_{1},\cdots,\mathbf{\dot{G}}_{p})\in\mathbb{H}^{n_{1}\times m_{1}}\times\cdots\mathbb{H}^{n_{p}\times m_{p}},$ $|||\widetilde{\mathbf{\dot{G}}}\||_{\mathrm{F}}:=\sqrt{\sum_{i=1}^{p}\|\mathbf{\dot{G}}_{i}\|_{\mathrm{F}}^{2}}$. Hence we have $\sum_{i=1}^{p}\frac{1}{\sqrt{p}}\|\mathbf{\dot{G}}_{i}\|_{\mathrm{F}}\leq\||\widetilde{\mathbf{\dot{G}}}\||_{\mathrm{F}}\leq\sum_{i=1}^{p}\|\mathbf{\dot{G}}_{i}\|_{\mathrm{F}},$ where the first equality follows from Cauchy-Schwarz inequality.
The quaternion singular value decomposition theorem will be used in the solving process of our algorithm. We state it as follows.
Theorem 2.1 (Quaternion singular value decomposition (QSVD)[37]).
For any given quaternion matrix $\mathbf{\dot{A}}\in\mathbb{H}^{m\times n},$ there exist two unitary quaternion matrices $\mathbf{\dot{U}}\in\mathbb{H}^{m\times m},\mathbf{\dot{V}}\in\mathbb{H}^{n\times n}$ such that
$\mathbf{\dot{A}}=\mathbf{\dot{U}}\left(\begin{matrix}\mathbf{\Sigma}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}\\ \end{matrix}\right)\mathbf{\dot{V}}^{*},$
where $\mathbf{\Sigma}\in\mathbb{R}^{r\times r}$ is a diagonal matrix with $\mathbf{\Sigma}_{1,1}\geq\mathbf{\Sigma}_{2,2}\geq\cdots\geq\mathbf{\Sigma}_{r,r}>0$ being the singular values of $\mathbf{\dot{A}}.$ We denote $\sigma_{i}(\mathbf{\dot{A}})$ to be the $i$-th singular value of $\mathbf{\dot{A}}$ and $\bm{\sigma}_{\mathbf{\dot{A}}}=(\mathbf{\Sigma}_{1,1},\cdots,\mathbf{\Sigma}_{r,r},0,\cdots,0)\in\mathbb{R}^{l}$ as the singular value vector of $\mathbf{\dot{A}}$, where $l=\min\{m,n\}$.
According to the QSVD, the rank of a quaternion matrix is defined as the number of nonzero singular values while the nuclear norm $\|\mathbf{\dot{A}}\|_{*}$ is defined as the sum of all nonzero singular values.
3 An Algorithm for SLRQA model
The proposed proximal linearized alternating direction method of multipliers (PL-ADMM) is stated in Algorithm 1.
Algorithm 1 Proixmal linearized-ADMM(PL-ADMM) for SLRQA in (2)
1:A noisy observed image $\mathbf{\dot{Y}}$; an initial iterate ${\mathbf{\dot{X}}}_{0}:=\mathbf{\dot{Y}}$; two parameters $\mu\in(0,2)$ and $\beta>0$; a tolerance $\eta>0$; two sequences of regularization parameters $\{L_{1,k}\}$ and $\{L_{2,k}\}$ satisfying $\sup_{k\geq 0}{L_{i,k}}<\infty$ and $\inf_{k>0}{L_{i,k}}>0$ for $i=1,2$;
2:A recovered image ${\mathbf{\dot{X}}}_{*}$;
3:Set $\mathbf{\dot{W}}_{0}=\bm{0}$ and $\mathbf{\dot{\Lambda}}_{0}=\bm{0}$;
4:for $k=0,1,2,\ldots$ do
5:      ${\mathbf{\dot{X}}}$-subproblem:
${\mathbf{\dot{X}}}_{k+1}\in$ $\mathop{\mathrm{arg\,min}}_{{\mathbf{\dot{X}}}}\;\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)$
$+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})-\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}))\|_{\mathrm{F}}^{2}+\frac{\beta}{2}\|\mathcal{W}({\mathbf{\dot{X}}})-\mathbf{\dot{W}}_{k}+\frac{\mathbf{\dot{\Lambda}}_{k}}{\beta}\|_{\mathrm{F}}^{2}+\frac{L_{1,k}}{2}\|{\mathbf{\dot{X}}}-{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}^{2};$
6:      $\mathbf{\dot{W}}$-subproblem:
$\mathbf{\dot{W}}_{k+1}=\mathop{\mathrm{arg\,min}}_{\mathbf{\dot{W}}}$ $\;\lambda\left\langle\nabla p(\mathbf{\dot{W}}_{k}),\mathbf{\dot{W}}\right\rangle+\frac{\beta}{2}\|\mathbf{\dot{W}}-(\mathcal{W}({\mathbf{\dot{X}}}_{k+1})+\frac{\mathbf{\dot{\Lambda}}_{k}}{\beta})\|_{\mathrm{F}}^{2}+\frac{L_{2,k}}{2}\|\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}^{2};$
7:      Update Lagrange multipliers: $\mathbf{\dot{\Lambda}}_{k+1}=\mathbf{\dot{\Lambda}}_{k}+\mu\beta(\mathcal{W}({\mathbf{\dot{X}}}_{k+1})-\mathbf{\dot{W}}_{k+1});$
8:     if $\|{\mathbf{\dot{X}}}_{k+1}-{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}+\|\mathbf{\dot{W}}_{k+1}-\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}+\|\mathcal{W}({\mathbf{\dot{X}}}_{k+1})-\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}<\eta$ then
9:         Return ${\mathbf{\dot{X}}}_{*}={\mathbf{\dot{X}}}_{k+1}$;
10:     end if
11:end for
Each iteration of PL-ADMM minimizes the augmented Lagrangian function by using the alternating direction method. Specifically, the augmented Lagrangian function of SLRQA in (2) is given by
$\mathcal{L}_{\beta}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})=\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)+$ $\lambda p(\mathbf{\dot{W}})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})-\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}))\|_{\mathrm{F}}^{2}$
$+\left\langle\mathbf{\dot{\Lambda}},\mathcal{W}({\mathbf{\dot{X}}})-\mathbf{\dot{W}}\right\rangle+\frac{\beta}{2}\|\mathcal{W}({\mathbf{\dot{X}}})-\mathbf{\dot{W}}\|^{2}_{\mathrm{F}},$
where $\mathbf{\dot{\Lambda}}\in\mathbb{H}^{m\times n}$ denotes the Lagrange multipliers and $\beta>0$.
In Step 5, ${\mathbf{\dot{X}}}_{k+1}$ is obtained by minimizing $\mathcal{L}_{\beta}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})+\frac{L_{1,k}}{2}\|{\mathbf{\dot{X}}}-{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}^{2}$ for a fixed $\mathbf{\dot{W}}_{k}$ and $\mathbf{\dot{\Lambda}}_{k}$. This subproblem does not yield a closed-form solution in general and therefore may be solved by nonsmooth algorithms such as a subgradient method or the linearized Bregman algorithm in [4]. If the linear operator $\mathcal{A}$ in (2) is the identity and all pixels of the noisy image $\mathbf{\dot{Y}}$ are observed, i.e., $\Omega$ indicates the whole image, then the ${\mathbf{\dot{X}}}$-subproblem can be solved efficiently. Specifically, the subproblem can be reformulated as
${\mathbf{\dot{X}}}_{k+1}=$ $\mathop{\mathrm{arg\,min}}_{{\mathbf{\dot{X}}}}\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)+$ (10)
$\qquad\qquad\frac{\tau^{2}(\beta+L_{1,k})+1}{2\tau^{2}}\left\|{\mathbf{\dot{X}}}-\frac{\tau^{2}}{\tau^{2}(\beta+L_{1,k})+1}\left(\frac{1}{\tau^{2}}\mathbf{\dot{Y}}+\beta\mathcal{W}^{\#}(\mathbf{\dot{W}}_{k}-\frac{\mathbf{\dot{\Lambda}}_{k}}{\beta})+L_{1,k}{\mathbf{\dot{X}}}_{k}\right)\right\|_{\mathrm{F}}^{2},$
If $\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)=\sigma_{i}({\mathbf{\dot{X}}}),$ the ${\mathbf{\dot{X}}}$-subproblem has a closed-form solution by using the soft thresholding method on the singular values [4]. When a nonconvex surrogate function $\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)$ is used, the subproblem of ${\mathbf{\dot{X}}}$ can be solved via a convex-concave procedure(CCP) in [20]. For completeness, CCP is stated in Algorithm 2. Lemma 3.1 is crucial in CCP and it transforms the matrix optimization problem of ${\mathbf{\dot{X}}}$ into an optimization problem of singular values.
Lemma 3.1 ([10]).
Let $\mathbf{\hat{{\mathbf{\dot{X}}}}}=\mathbf{\dot{U}}\bm{\Sigma}\mathbf{\dot{V}}^{*}\in\mathbb{H}^{m\times n}$ be a singular value decomposition of $\mathbf{\hat{{\mathbf{\dot{X}}}}}$. The solution of
$\mathop{\mathrm{arg\,min}}_{{\mathbf{\dot{X}}}\in\mathbb{H}^{m\times n}}\,\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)+\frac{\mu}{2}\|\mathbf{\hat{{\mathbf{\dot{X}}}}}-{\mathbf{\dot{X}}}\|_{\mathrm{F}}^{2}$ (11)
can be represented by ${\mathbf{\dot{X}}}_{*}=\mathbf{\dot{U}}\bm{\Sigma_{*}}\mathbf{\dot{V}}^{*},$ where $\bm{(\Sigma_{*})}_{i,i}=(\bm{\sigma_{*}})_{i},$ for $i=1,\cdots,\min\{m,n\}$ where $\bm{\sigma_{*}}$ is obtained by
$\bm{\sigma_{*}}=\mathop{\mathrm{arg\,min}}_{\bm{\sigma}\,\geq 0}\phi(\bm{\sigma},\gamma)+\frac{\mu}{2}\|\bm{\sigma}-\bm{\sigma_{\bm{\hat{{\mathbf{\dot{X}}}}}}}\|_{2}^{2}.$ (12)
In the $k$-th iteration of Algorithm 2, the function $\phi(\bm{\sigma},\gamma)$ is approximated by a linearized function $\phi(\bm{\sigma}_{k},\gamma)+\nabla_{\bm{\sigma}}\phi(\bm{\sigma}_{k},\gamma)^{\mathrm{T}}\bm{\sigma}$. It follows that the update of $\bm{\sigma}$ in Problem (12) at $\bm{\sigma}_{k}$ is given by
$\bm{\sigma}_{k+1}=\max\left\{\bm{\sigma}_{\bm{\hat{{\mathbf{\dot{X}}}}}}-\frac{\nabla_{\bm{\sigma}}\phi(\bm{\sigma}_{k},\gamma)}{\mu},\bm{0}\right\}.$ (13)
It has been shown in [20, Theorem 4] that any limit points of $\{\bm{\sigma}_{k}\}_{k=1}^{\infty}$ are stationary points of (12).
Algorithm 2 CCP to solve Problem (11)
1:A quaternion matrix $\bm{\hat{{\mathbf{\dot{X}}}}}\in\mathbb{H}^{m\times n}$; a nonconvex function $\phi$; a constant $\mu>0$; a tolerance $\eta_{C}$; and an initial iterate $\bm{\sigma}_{0}=\mathbf{0}$;
2:$\bm{{\mathbf{\dot{X}}}}_{*}$;
3:Compute QSVD $\mathbf{\hat{{\mathbf{\dot{X}}}}}=\mathbf{\dot{U}}\bm{\Sigma}\mathbf{\dot{V}}^{*}$ and let $\bm{\sigma_{\hat{{\mathbf{\dot{X}}}}}}$ satisfy $\bm{\Sigma}=\mathrm{diag}(\bm{\sigma_{\hat{{\mathbf{\dot{X}}}}}})$;
4:for $k=1,2,\ldots$ do
5:     Compute $\bm{\sigma}_{k}$ by (13);
6:     if $\|\bm{\sigma}_{k-1}-\bm{\sigma}_{k}\|_{2}<\eta_{C}$ then
7:         Set $\bm{\sigma}_{*}=\bm{\sigma}_{k}$ and return ${\mathbf{\dot{X}}}_{*}=\mathbf{\dot{U}}\bm{\Sigma_{*}}\mathbf{\dot{V}}^{*}$, where $\bm{\Sigma_{*}}=\mathrm{diag}(\bm{\sigma_{*}})$;
8:     end if
9:end for
In Step 6, the $\mathbf{\dot{W}}$-subproblem approximately minimizes $\mathcal{L}_{\beta}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})+\frac{L_{2,k}}{2}\|\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}^{2}$ for a fixed ${\mathbf{\dot{X}}}_{k+1}$ and $\mathbf{\dot{\Lambda}}_{k}$. Specifically, the objective function of the $\mathbf{\dot{W}}$-subproblem is obtained by using a linearized strategy for $p(\mathbf{\dot{W}})$ in $\mathcal{L}_{\beta}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})+\frac{L_{2,k}}{2}\|\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}^{2}$. It follows that $\mathbf{\dot{W}}_{k+1}$ has a closed form solution
$\mathbf{\dot{W}}_{k+1}=\frac{1}{\beta+L_{2,k}}\left(\beta\mathcal{W}({\mathbf{\dot{X}}}_{k+1})+\mathbf{\dot{\Lambda}}_{k}+L_{2,k}\mathbf{\dot{W}}_{k}\right)-\frac{\lambda}{\beta+L_{2,k}}\nabla p(\mathbf{\dot{W}}_{k}).$ (14)
Step 7 is used to update the Lagrange multipliers. The difference between the classic ADMM for convex function and PL-ADMM is that $\mu\in(0,2)$ in PL-ADMM rather than $\mu\in(0,\frac{\sqrt{5}+1}{2})$ [8].
4 Convergence analysis
The global convergence is established for a generalization of SLRQA in (2):
$\min_{{\mathbf{\dot{X}}},\mathbf{\dot{W}}}$ $f({\mathbf{\dot{X}}})+g(\mathbf{\dot{W}})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2},$ (15)
$\mathrm{s.t.}$ $\mathcal{C}({\mathbf{\dot{X}}})+\mathcal{B}(\mathbf{\dot{W}})=\mathbf{\dot{B}},$
where $f$ is continuous and $g$ is continuously differentiable, $\mathcal{C},\mathcal{B}:\mathbb{H}^{m\times n}\rightarrow\mathbb{H}^{m\times n}$ are given linear maps and $\mathbf{\dot{B}}\in\mathbb{H}^{m\times n}$. Specifically, if $f({\mathbf{\dot{X}}})=\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma),g(\mathbf{\dot{W}})=p(\mathbf{\dot{W}})$, $\mathcal{B}=-\mathbf{I}$, $\mathcal{C}=\mathcal{W}$, and $\mathbf{\dot{B}}=0$, then Problem (15) becomes the SLRQA in (2).
The augmented Lagrangian function of (15) is
$\mathcal{L}_{\beta}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})=f({\mathbf{\dot{X}}})+g(\mathbf{\dot{W}})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2}+\left\langle\mathbf{\dot{\Lambda}},\mathcal{C}({\mathbf{\dot{X}}})+\mathcal{B}(\mathbf{\dot{W}})-\mathbf{\dot{B}}\right\rangle$ (16)
$+\frac{\beta}{2}\|\mathcal{C}({\mathbf{\dot{X}}})+\mathcal{B}(\mathbf{\dot{W}})-\mathbf{\dot{B}}\|^{2}_{\mathrm{F}}$
$=f({\mathbf{\dot{X}}})+g(\mathbf{\dot{W}})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2}+\frac{\beta}{2}\left\|\mathcal{C}({\mathbf{\dot{X}}})+\mathcal{B}(\mathbf{\dot{W}})-\mathbf{\dot{B}}+\frac{\mathbf{\dot{\Lambda}}}{\beta}\right\|_{\mathrm{F}}^{2}-\frac{1}{2\beta}\left\|\mathbf{\dot{\Lambda}}\right\|_{\mathrm{F}}^{2}.$
The proposed PL-ADMM for (15) is
${\mathbf{\dot{X}}}_{k+1}$ $\in\mathop{\mathrm{arg\,min}}_{{\mathbf{\dot{X}}}}f({\mathbf{\dot{X}}})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2}$ (17)
$+\frac{\beta}{2}\left\|\mathcal{C}({\mathbf{\dot{X}}})+\mathcal{B}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}+\frac{\mathbf{\dot{\Lambda}}_{k}}{\beta}\right\|_{\mathrm{F}}^{2}+\frac{L_{1,k}}{2}\|{\mathbf{\dot{X}}}-{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}^{2},$
$\mathbf{\dot{W}}_{k+1}$ $\in\mathop{\mathrm{arg\,min}}_{\mathbf{\dot{W}}}\left\langle\nabla g(\mathbf{\dot{W}}_{k}),\mathbf{\dot{W}}\right\rangle+\frac{\beta}{2}\left\|\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}})-\mathbf{\dot{B}}+\frac{\mathbf{\dot{\Lambda}}_{k}}{\beta}\right\|_{\mathrm{F}}^{2}+\frac{L_{2,k}}{2}\left\|\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\right\|^{2}_{\mathrm{F}},$
$\mathbf{\dot{\Lambda}}_{k+1}$ $=\mathbf{\dot{\Lambda}}_{k}+\mu\beta\cdot(\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}),$
which is a generalization of Algorithm 1.
The choices of $\beta,L_{1,k}$ and $L_{2,k}$ are crucial for the convergence of (17). Generally speaking, $\beta$ should be chosen large enough, and $L_{1,k}$, $L_{2,k}$ should be chosen in a suitable range. In order to give the conditions of $\beta$ and $L_{1,k},L_{2,k}$, we introduce the following notations. Let $q_{i}:=\sup_{k\geq 0}{L_{i,k}}<\infty$ and $q_{i}^{-}:=\inf_{k>0}{L_{i,k}}>0$ for $i=1,2$. Define $\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}$ as the smallest positive eigenvalue of $\mathcal{B}^{\#}\mathcal{B}$, and $\rho(\mu):=1-|1-\mu|$. For any sequence $\{\mathbf{\dot{U}}_{k}\}_{k\geq 0}$, define $\Delta$ operator by $\Delta\mathbf{\dot{U}}_{k}:=\mathbf{\dot{U}}_{k}-\mathbf{\dot{U}}_{k-1}.$
The assumptions for the convergence analysis are given in Assumption 4.1, including the conditions of $\beta$, $L_{1,k}$, and $L_{2,k}$.
Assumption 4.1.
  1. A1
    The continuous function $f({\mathbf{\dot{X}}})$ is coercive, i.e. $f({\mathbf{\dot{X}}})\rightarrow\infty$ if $\|{\mathbf{\dot{X}}}\|_{\mathrm{F}}\rightarrow\infty$, and $g(\mathbf{\dot{W}})$ is bounded from below.
  2. A2
    $g(\mathbf{\dot{W}})$ has $L_{g}$ Lipschitz continuous gradient, i.e., for every $\mathbf{\dot{W}}_{1},\mathbf{\dot{W}}_{2}$ , it holds that
    $\|\nabla g(\mathbf{\dot{W}}_{1})-\nabla g(\mathbf{\dot{W}}_{2})\|_{\mathrm{F}}\leq L_{g}\|\mathbf{\dot{W}}_{1}-\mathbf{\dot{W}}_{2}\|_{\mathrm{F}}.$ (18)
    It has been shown in [2, Lemma 5.7] that (18) implies that for every $\mathbf{\dot{W}}_{1}$ and $\mathbf{\dot{W}}_{2}$, it holds that
    $g(\mathbf{\dot{W}}_{1})\leq g(\mathbf{\dot{W}}_{2})+\left\langle\nabla g(\mathbf{\dot{W}}_{2}),\mathbf{\dot{W}}_{1}-\mathbf{\dot{W}}_{2}\right\rangle+\frac{L_{g}}{2}\|\mathbf{\dot{W}}_{1}-\mathbf{\dot{W}}_{2}\|_{\mathrm{F}}^{2}.$ (19)
  3. A3
    The matrix $\mathcal{B}^{\#}\mathcal{B}$ is full rank, ${\rm range}(\mathcal{C})\subseteq{\rm range}(\mathcal{B})$, and $\mathbf{\dot{B}}\in{\rm range}(\mathcal{B})$.
  4. A4
    The parameters $\beta>0$, $\mu\in(0,2)$, and there exist three constants $a_{1}>0,a_{2}>0$, and $r>1$ such that
    $q_{1}^{-}\mathbf{I}\succeq a_{1}\mathbf{I}\,\quad\mbox{and}\,\quad q_{2}^{-}\mathbf{I}+\beta\mathcal{B}^{\#}\mathcal{B}-(r\theta_{0}+r\theta_{1}+L_{g})\mathbf{I}\succeq a_{2}\mathbf{I},$ (20)
    where $\mathbf{I}$ is the identity matrix with correct size,
    $\theta_{0}:=\frac{2\mu(q_{2}+L_{g})^{2}}{\beta\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}(\rho(\mu))^{2}},\,\hbox{ and }\theta_{1}:=\frac{2\mu q_{2}^{2}}{\beta\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}(\rho(\mu))^{2}}.$ (21)
  5. A5
    The parameter $\beta$ satisfies $\beta>\frac{2L_{g}}{\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}\rho(\mu)}.$
One can verify that A1, A2, and A3 in Assumption 4.1 are satisfied for SLRQA in (2). If $\beta$ is chosen sufficiently large and the sequences $\{L_{1,k}\}$ and $\{L_{2,k}\}$ are chosen such that $q_{i}^{-},q_{i}$, $i=1,2$ in suitable range, then A4 and A5 in Assumption 4.1 are also satisfied. For example, if set $L_{1,k}=L_{2,k}=1$ for all $k$ and let
$\beta>\max\left\{\frac{2L_{g}}{\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}\rho(\mu)},\frac{r(2\mu(1+L_{g})^{2}+2\mu)}{(\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}})^{2}(\rho(\mu))^{2}}+\frac{L_{g}}{\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}},1\right\},$
then A4 and A5 in Assumption 4.1 hold. Therefore, Assumption 4.1 is reasonable.
Lemmas 4.1 and 4.2 show that the augmented Lagrangian function $\mathcal{L}_{\beta}$ is sufficiently descent when updating $\mathbf{\dot{W}}$ and ${\mathbf{\dot{X}}}$. They are used in Lemma 4.3 for bounding $\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1})-\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})$.
Lemma 4.1 (Sufficient descent of $\mathcal{L}_{\beta}$ for ${\mathbf{\dot{X}}}$ update).
The sequence $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ generated by (17) satisfies:
$\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})-\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\geq\frac{L_{1,k}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2}_{\mathrm{F}}.$
Proof.
According to the ${\mathbf{\dot{X}}}_{k+1}$ update in (17), the following inequality holds:
$f({\mathbf{\dot{X}}}_{k+1})-f({\mathbf{\dot{X}}}_{k})+\frac{\beta}{2}\|\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta\|_{\mathrm{F}}^{2}+\frac{L_{1,k}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}$ (22)
$+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}((\mathcal{A}({\mathbf{\dot{X}}}_{k+1})-\mathbf{\dot{Y}}))\|_{\mathrm{F}}^{2}\leq\frac{\beta}{2}\left\|\mathcal{C}({\mathbf{\dot{X}}}_{k})+\mathcal{B}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta\right\|_{\mathrm{F}}^{2}+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}((\mathcal{A}({\mathbf{\dot{X}}}_{k})-\mathbf{\dot{Y}}))\|_{\mathrm{F}}^{2}.$
It follows that
$\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})-\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})$
$=f({\mathbf{\dot{X}}}_{k})-f({\mathbf{\dot{X}}}_{k+1})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}((\mathcal{A}({\mathbf{\dot{X}}}_{k})-\mathbf{\dot{Y}}))\|_{\mathrm{F}}^{2}-\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}((\mathcal{A}({\mathbf{\dot{X}}}_{k+1})-\mathbf{\dot{Y}}))\|_{\mathrm{F}}^{2}$
$-\frac{\beta}{2}\|\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta\|_{\mathrm{F}}^{2}+\frac{\beta}{2}\|\mathcal{C}({\mathbf{\dot{X}}}_{k})+\mathcal{B}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta\|_{\mathrm{F}}^{2}\geq\frac{L_{1,k}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2},$
which completes the proof. ∎
Lemma 4.2 (Sufficient descent of $\mathcal{L}_{\beta}$ for $\mathbf{\dot{W}}$ update).
Suppose A2 in Assumption 4.1 holds. The sequence $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ generated by (17) satisfies:
$\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})-\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k})\geq\|\Delta\mathbf{\dot{W}}_{k+1}\|^{2}_{B_{k}},$
where $B_{k}:=L_{2,k}\mathbf{I}-L_{g}\mathbf{I}+\frac{\beta}{2}\mathcal{B}\mathcal{B}^{\#}$.
Proof.
According to the optimality condition of $\mathbf{\dot{W}}_{k+1},$ the following equality holds:
$\nabla g(\mathbf{\dot{W}}_{k})+\beta\mathcal{B}^{\#}(\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta)+L_{2,k}\Delta\mathbf{\dot{W}}_{k+1}=0.$ (23)
It follows that
$\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})-\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k})$
$=g(\mathbf{\dot{W}}_{k})-g(\mathbf{\dot{W}}_{k+1})+\frac{\beta}{2}\|\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}+\frac{\mathbf{\dot{\Lambda}}_{k}}{\beta}\|_{\mathrm{F}}^{2}-\frac{\beta}{2}\|\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}+\frac{\mathbf{\dot{\Lambda}}_{k}}{\beta}\|_{\mathrm{F}}^{2}$
$=g(\mathbf{\dot{W}}_{k})-g(\mathbf{\dot{W}}_{k+1})+\frac{\beta}{2}\|\mathcal{B}(\Delta\mathbf{\dot{W}}_{k+1})\|_{\mathrm{F}}^{2}-\beta\left\langle\mathcal{B}^{\#}(\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}+\frac{\mathbf{\dot{\Lambda}}_{k}}{\beta}),\Delta\mathbf{\dot{W}}_{k+1}\right\rangle$
[…]
which completes the proof. ∎
Lemma 4.3 is used in Lemma 4.4, which is crucial for constructing a merit function.
Lemma 4.3.
Suppose A2 in Assumption 4.1 holds. The sequence $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ generated by (17) satisfies:
$\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1})\leq\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})-\|\Delta\mathbf{\dot{W}}_{k+1}\|^{2}_{B_{k}}-\frac{L_{1,k}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2}_{\mathrm{F}}+\frac{1}{\beta\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}.$
Proof.
According to the update of $\mathbf{\dot{\Lambda}}$ in (17), Lemma 4.1, and Lemma 4.2, we have
$\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1})$ $=\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k})+\frac{1}{\beta\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}$
$\leq\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})-\|\Delta\mathbf{\dot{W}}_{k+1}\|_{B_{k}}^{2}-\frac{L_{1,k}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}+\frac{1}{\beta\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}.$
Lemma 4.4.
Suppose A2,A3 in Assumption 4.1 hold, the sequence $\{({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1})\}_{k\geq 0}$ generated by (17) satisfies:
$\frac{1}{\beta\mu}\|\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}\leq\theta_{0}\|\Delta\mathbf{\dot{W}}_{k}\|^{2}_{\mathrm{F}}+\theta_{1}\|\Delta\mathbf{\dot{W}}_{k+1}\|^{2}_{\mathrm{F}}+\theta_{2}\|\mathcal{B}^{\#}(\Delta\mathbf{\dot{\Lambda}}_{k})\|^{2}_{\mathrm{F}}-\theta_{2}\|\mathcal{B}^{\#}(\Delta\mathbf{\dot{\Lambda}}_{k+1})\|^{2}_{\mathrm{F}},$ (24)
$\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1})+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{B_{k}-r\theta_{1}\mathbf{I}}^{2}+\frac{L_{1,k}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2}_{\mathrm{F}}+\frac{r-1}{\beta\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}$
$+r\theta_{2}\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}\leq\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})+r\theta_{0}\|\Delta\mathbf{\dot{W}}_{k}\|^{2}_{\mathrm{F}}+r\theta_{2}\|\mathcal{B}^{\#}(\Delta\mathbf{\dot{\Lambda}}_{k})\|^{2}_{\mathrm{F}},$ (25)
where $r>1,\theta_{0},\theta_{1}$ are defined in (21), and $\theta_{2}:=\frac{|1-\mu|}{\beta\mu\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}\rho(\mu)}.$
Proof.
Let $k\geq 1$ be fixed and define the matrix
$\mathbf{\dot{G}}_{k+1}:=-L_{2,k}\Delta\mathbf{\dot{W}}_{k+1}-\nabla g(\mathbf{\dot{W}}_{k}).$ (26)
It follows that $\Delta\mathbf{\dot{G}}_{k+1}=L_{2,k-1}\Delta\mathbf{\dot{W}}_{k}-L_{2,k}\Delta\mathbf{\dot{W}}_{k+1}+\nabla g(\mathbf{\dot{W}}_{k-1})-\nabla g(\mathbf{\dot{W}}_{k}).$ By the triangle inequality, the following inequality holds:
$\|\Delta\mathbf{\dot{G}}_{k+1}\|_{\mathrm{F}}\leq\|\nabla g(\mathbf{\dot{W}}_{k-1})-\nabla g(\mathbf{\dot{W}}_{k})\|_{\mathrm{F}}+L_{2,k}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+L_{2,k-1}\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}.$
By A2 in Assumption 4.1, $\nabla g(\mathbf{\dot{W}})$ is $L_{g}$ Lipschitz continuous and $q_{2}=\sup_{k\geq 0}L_{2,k}<\infty,$ we have $\|\Delta\mathbf{\dot{G}}_{k+1}\|_{\mathrm{F}}\leq(L_{g}+q_{2})\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}+q_{2}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}.$ Hence it follows that:
$\|\Delta\mathbf{\dot{G}}_{k+1}\|^{2}_{\mathrm{F}}\leq 2(L_{g}+q_{2})^{2}\|\Delta\mathbf{\dot{W}}_{k}\|^{2}_{\mathrm{F}}+2q_{2}^{2}\|\Delta\mathbf{\dot{W}}_{k+1}\|^{2}_{\mathrm{F}}.$ (27)
Expressing the optimality condition of $\mathbf{\dot{W}}$ subproblem using $\mathbf{\dot{G}}_{k+1}$, we have $\mathbf{\dot{G}}_{k+1}=\beta\mathcal{B}^{\#}(\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta).$ Combining this with the $\mathbf{\dot{\Lambda}}$ update, we obtain
#$\mathcal{B}^{\#}\mathbf{\dot{\Lambda}}_{k+1}=\mu\mathbf{\dot{G}}_{k+1}+(1-\mu)\mathcal{B}^{\#}\mathbf{\dot{\Lambda}}_{k}.$ (28)
It follows that #$\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}=\mu\Delta\mathbf{\dot{G}}_{k+1}+(1-\mu)\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k}.$ Since $\mu\in(0,2),$ we have
#$\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}=\rho(\mu)\cdot\frac{\mu\Delta\mathbf{\dot{G}}_{k+1}}{\rho(\mu)}+|1-\mu|\cdot(\text{sign}(1-\mu)\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k}),$
By the convexity of $\|\cdot\|^{2}_{\mathrm{F}}$, the update of $\mathbf{\dot{\Lambda}}$ in (17) and A3 in Assumption 4.1, it follows that $\Delta\mathbf{\dot{\Lambda}}_{k+1}\in\text{{\rm range}}(\mathcal{B})$, Hence, the following inequalities hold
$\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}\rho(\mu)\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}\leq$ $\rho(\mu)\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2}=\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2}-|1-\mu|\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}$
$\leq$ $\frac{\mu^{2}}{\rho(\mu)}\|\Delta\mathbf{\dot{G}}_{k+1}\|^{2}_{\mathrm{F}}+|1-\mu|\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k}\|^{2}_{\mathrm{F}}-|1-\mu|\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}},$ (29)
where the first inequality follows from A3 in Assumption 4.1 and the definition of $\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}$. Dividing both sides of (29) by $\beta\mu\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}\rho(\mu)$ and using (27) yield (24). Finally, Inequality (25) follows by multiplying the inequality (24) by $r>1$ and combining it with Lemma 4.3. ∎
Let $\mathcal{R}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}},\mathbf{\dot{W}}^{\prime},\mathbf{\dot{\Lambda}}^{\prime}):=\mathcal{L}_{\beta}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})+r\theta_{0}\|\mathbf{\dot{W}}-\mathbf{\dot{W}}^{\prime}\|^{2}_{\mathrm{F}}+r\theta_{2}\|\mathcal{B}^{\#}(\mathbf{\dot{\Lambda}}-\mathbf{\dot{\Lambda}}^{\prime})\|^{2}_{\mathrm{F}}$. The merit function $\mathcal{R}_{k}$ is defined by
$\mathcal{R}_{k}:=\mathcal{R}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},\mathbf{\dot{\Lambda}}_{k-1})=\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})+r\theta_{0}\|\Delta\mathbf{\dot{W}}_{k}\|^{2}_{\mathrm{F}}+r\theta_{2}\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k}\|^{2}_{\mathrm{F}}.$ (30)
If Assumption 4.1 holds, then it can be shown that the merit function is descent. Specifically, we have
$\mathcal{R}_{k+1}+a\left(\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}\right)$ (31)
$\leq\mathcal{R}_{k+1}+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{B_{k}-r(\theta_{1}+\theta_{0})\mathbf{I}}^{2}+\frac{L_{1,k}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2}_{\mathrm{F}}+\frac{r-1}{\beta\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}\leq\mathcal{R}_{k}\leq\mathcal{R}_{k_{0}},$
where $a=\min\{a_{1},a_{2},\frac{r-1}{\beta\mu}\}$ and $a_{1},a_{2}$ are defined in Assumption 4.1, the first inequality follows from A4, A5 in Assumption 4.1, the second inequality follows from Lemma 4.4, and the third inequality is due to induction of $\mathcal{R}_{k}\leq\mathcal{R}_{k+1}$ for any $k\geq k_{0}$.
Theorem 4.1 (Bounded sequence of $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$).
Assume Assumption 4.1 holds. The sequence $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ generated by (17) is bounded.
Proof.
According to (31), there exists $k_{0}\geq 1$ such that $\mathcal{R}_{k+1}\leq\mathcal{R}_{k_{0}}$ for all $k\geq k_{0}.$ Hence the following inequality holds:
$f({\mathbf{\dot{X}}}_{k+1})+g(\mathbf{\dot{W}}_{k+1})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}_{k}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2}+\frac{\beta}{2}\left\|\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}+\frac{\mathbf{\dot{\Lambda}}_{k+1}}{\beta}\right\|_{\mathrm{F}}^{2}$ (32)
$-\frac{1}{2\beta}\left\|\mathbf{\dot{\Lambda}}_{k+1}\right\|_{\mathrm{F}}^{2}+(r\theta_{0}+a)\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}+a(\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2})+r\theta_{2}\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2}\leq\mathcal{R}_{k_{0}}.$
According to (28), we have
$\mu\mathcal{B}^{\#}\mathbf{\dot{\Lambda}}_{k+1}=\mu\mathbf{\dot{G}}_{k+1}+(1-\mu)\mathcal{B}^{\#}(\mathbf{\dot{\Lambda}}_{k}-\mathbf{\dot{\Lambda}}_{k+1}).$ (33)
Since $\mu\in(0,2),$ Equation (33) can be rewritten as:
$\mu\mathcal{B}^{\#}\mathbf{\dot{\Lambda}}_{k+1}=\rho(\mu)\frac{\mu\mathbf{\dot{G}}_{k+1}}{\rho(\mu)}+|1-\mu|\left(\text{sign}(1-\mu)\mathcal{B}^{\#}(\mathbf{\dot{\Lambda}}_{k}-\mathbf{\dot{\Lambda}}_{k+1})\right).$
By the convexity of $\|\cdot\|^{2}$, we have $\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}\mu^{2}\|\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2}\leq\frac{\mu^{2}}{\rho(\mu)}\|\mathbf{\dot{G}}_{k+1}\|_{\mathrm{F}}^{2}+|1-\mu|\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2}.$ According to definition (26), we have $\|\mathbf{\dot{G}}_{k+1}\|_{\mathrm{F}}^{2}\leq 2q_{2}^{2}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}+2\|\nabla g(\mathbf{\dot{W}}_{k})\|_{\mathrm{F}}^{2}\leq 2(q_{2}+L_{g})^{2}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}+2\|\nabla g(\mathbf{\dot{W}}_{k+1})\|_{\mathrm{F}}^{2}.$ It follows that:
$-\frac{1}{2\beta}\|\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2}\geq-\theta_{3}\|\nabla g(\mathbf{\dot{W}}_{k+1})\|_{\mathrm{F}}^{2}-\theta_{4}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}-\theta_{5}\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2},$ (34)
where
$\theta_{3}:=\frac{1}{\beta\rho(\mu)\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}},\quad\theta_{4}:=\frac{(q_{2}+L_{g})^{2}}{\beta\rho(\mu)\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}},\quad\theta_{5}:=\frac{|1-\mu|}{2\beta\mu^{2}\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}}.$
Using Inequalities (34) and (32), we obtain
$f({\mathbf{\dot{X}}}_{k+1})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}_{k}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2}+\frac{\beta}{2}\left\|\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}+\frac{\mathbf{\dot{\Lambda}}_{k+1}}{\beta}\right\|_{\mathrm{F}}^{2}$ (35)
$+a(\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2})+(r\theta_{0}+a-\theta_{4})\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}+(r\theta_{2}-\theta_{5})\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2}$
$\leq\mathcal{R}_{k_{0}}-\inf_{\mathbf{\dot{W}}}\left\{g(\mathbf{\dot{W}})-\theta_{3}\|\nabla g(\mathbf{\dot{W}})\|_{\mathrm{F}}^{2}\right\}.$
According to (19) in Assumption A2, setting $\mathbf{\dot{W}}_{1}=\mathbf{\dot{W}}-\delta\nabla g(\mathbf{\dot{W}})$ and $\mathbf{\dot{W}}_{2}=\mathbf{\dot{W}}$, it follows that $g(\mathbf{\dot{W}}_{k}-\delta\nabla g(\mathbf{\dot{W}}_{k}))\leq g(\mathbf{\dot{W}}_{k})-(\delta-\frac{L_{g}\delta^{2}}{2})\|\nabla g(\mathbf{\dot{W}}_{k})\|_{\mathrm{F}}^{2}.$ Since $g$ is bounded from below, there exist $M$ such that
$-M<\inf\{g(\mathbf{\dot{W}})-(\delta-\frac{L_{g}\delta^{2}}{2})\|\nabla g(\mathbf{\dot{W}})\|_{\mathrm{F}}^{2}:\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}\}.$ (36)
We choose $\delta=\frac{1}{L_{g}}$. According to Assumption A5, we have $\theta_{3}<\frac{1}{2L_{g}}=\delta-\frac{L_{g}\delta^{2}}{2}.$ Since $r>1$ and $\mu\in(0,2),$ according to the denifition of $\theta_{0},\theta_{2},\theta_{4}$ and $\theta_{5}$, it holds that $r\theta_{2}-\theta_{5}>0$ and $r\theta_{0}+a-\theta_{4}>0$. It follows from (35) that
$f({\mathbf{\dot{X}}}_{k+1})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}_{k}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2}$ (37)
$+\frac{\beta}{2}\left\|\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}+\frac{\mathbf{\dot{\Lambda}}_{k+1}}{\beta}\right\|_{\mathrm{F}}^{2}+a(\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2})<\mathcal{R}_{k_{0}}+M.$
Since $f$ is coercive, the sequence $\{{\mathbf{\dot{X}}}_{k}\}_{k\geq k_{0}}$ is bounded and hence $\{\mathcal{C}{\mathbf{\dot{X}}}_{k}\}_{k\geq k_{0}}$ is bounded. According to (37), $\Delta\mathbf{\dot{\Lambda}}_{k+1}$ is bounded. By the $\mathbf{\dot{\Lambda}}$ update in (17), we have $\mathcal{B}\mathbf{\dot{W}}_{k+1}=\frac{1}{\beta\mu}\Delta\mathbf{\dot{\Lambda}}_{k+1}-\mathcal{C}{\mathbf{\dot{X}}}_{k+1}+\mathbf{\dot{B}}.$ Since $\{\mathcal{C}{\mathbf{\dot{X}}}_{k}\}_{k\geq k_{0}}$ is bounded and $\mathcal{B}^{\#}\mathcal{B}$ is full rank, it follows that $\{\mathbf{\dot{W}}_{k}\}_{k\geq k_{0}}$ is bounded. From the fact that $\frac{\beta}{2}\left\|\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}+\frac{\mathbf{\dot{\Lambda}}_{k+1}}{\beta}\right\|_{\mathrm{F}}^{2}$ is bounded, it follows that $\{\mathbf{\dot{\Lambda}}_{k}\}_{k\geq k_{0}}$ is bounded. As a consequence, $\{{\mathbf{\dot{X}}}_{k}\}_{k\geq 1},\{\mathbf{\dot{W}}_{k}\}_{k\geq 1},\{\mathbf{\dot{\Lambda}}_{k}\}_{k\geq 1}$ is bounded. ∎
Using Theorem 4.1, we have the following convergence result of $\mathcal{R}_{k}$.
Lemma 4.5.
Suppose Assumption 4.1 holds. The sequence $\{\mathcal{R}_{k}\}_{k\geq 1}$ is bounded from below and converges.
Proof.
It follows from Theorem 4.1 that $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 1}$ is bounded. Since $\mathcal{L}_{\beta}$ is continous respect to ${\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k}$, it follows that $\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})$ is bounded from below. Hence $\mathcal{R}_{k}$ is bounded from below. According to Assumption 4.1 and (31), $\{\mathcal{R}_{k}\}_{k\geq 1}$ is monotonically decreasing. As a consequence, $\{\mathcal{R}_{k}\}_{k\geq 1}$ is bounded from below and converges. ∎
The following convergence result of sequence $\{{\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k}\}$ is given by Lemma 4.5.
Theorem 4.2.
Suppose Assumption 4.1 holds. It follows that
$\lim_{k\rightarrow\infty}\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}=0,\lim_{k\rightarrow\infty}\|\Delta{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}=0,\lim_{k\rightarrow\infty}\|\Delta\mathbf{\dot{\Lambda}}_{k}\|_{\mathrm{F}}=0.$
Proof.
According to Theorem 4.1, summing up (31) from $k_{0}$ to $K\geq k_{0}$, the following inequality holds
$\sum_{k=k_{0}}^{K}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}+\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2}_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}<\frac{1}{a}(\mathcal{R}_{k_{0}}-\inf_{k\geq 1}\mathcal{R}_{k}).$
According to Lemma 4.5, it follows that $-\infty<\inf_{k\geq 1}\mathcal{R}_{k}$. Let $K\rightarrow+\infty$, since $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ is bounded, then $\mathcal{R}_{k_{0}}$ is bounded and hence $\sum_{k\geq 0}\|\Delta\mathbf{\dot{W}}_{k+1}\|^{2}_{\mathrm{F}}+\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2}_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}<\infty,$ which completes the proof. ∎
Theorem 4.2 has proven the limiting behavior of $\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}},\|\Delta{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}$ and $\|\Delta\mathbf{\dot{\Lambda}}_{k}\|_{\mathrm{F}}$, which will be used to prove the limiting behavior of $\mathcal{R}_{k}.$ We next prove that the subgradient of $\mathcal{L}_{\beta}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})$ can be bounded by $\Delta{\mathbf{\dot{X}}}_{k+1},\Delta\mathbf{\dot{W}}_{k+1}$ and $\Delta\mathbf{\dot{\Lambda}}_{k+1}.$
Lemma 4.6 (Subgradient bound).
Suppose that Assumption A2 holds. Let $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ be a sequence generated by (17). Then $\widetilde{\mathbf{\dot{D}}}_{k+1}:=(\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k+1}},\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k+1}},\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k+1}})\in\partial\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1})$, where
$\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k+1}}$ $:=\mathcal{C}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}-L_{2,k}\Delta{\mathbf{\dot{X}}}_{k+1}-\beta\mathcal{C}^{\#}\mathcal{B}\Delta\mathbf{\dot{W}}_{k+1},\quad\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k+1}}:=\frac{1}{\beta\mu}\Delta\mathbf{\dot{\Lambda}}_{k+1},\,$
$\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k+1}}$ $:=\nabla g(\mathbf{\dot{W}}_{k+1})-\nabla g(\mathbf{\dot{W}}_{k})+\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}-L_{1,k}\Delta\mathbf{\dot{W}}_{k+1}.$
Furthermore, there exists $\rho>0$ such that
$|\|\widetilde{\mathbf{\dot{D}}}_{k+1}\||\leq\pi\left(\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}\right),$
where
$\pi:=\max\left\{q_{1},\beta\|\mathcal{C}\|\|\mathcal{B}\|+q_{2}+L_{g},\|\mathcal{C}\|+\|\mathcal{B}\|+\frac{1}{\beta\mu}\right\}.$ (38)
Proof.
According to the optimality condition of the update of ${\mathbf{\dot{X}}}_{k+1},$ it follows that
$-\frac{1}{\tau^{2}}\mathcal{A}^{\#}\mathcal{P}^{\#}_{\Omega}(\mathcal{P}_{\Omega}(\mathcal{A}{\mathbf{\dot{X}}}_{k+1}-\mathbf{\dot{Y}}))-\beta\mathcal{C}^{\#}(\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta)-L_{1,k}\Delta{\mathbf{\dot{X}}}_{k+1}\in\partial f({\mathbf{\dot{X}}}_{k+1}),$
which yields
$\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k+1}}=-\mathcal{C}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}-L_{1,k}\Delta{\mathbf{\dot{X}}}_{k+1}+\beta\mathcal{C}^{\#}\mathcal{B}\Delta\mathbf{\dot{W}}_{k+1}\in\partial_{{\mathbf{\dot{X}}}}\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1}).$ (39)
According to the optimality condition of the update of $\mathbf{\dot{W}}_{k+1},$ the following equality holds:
$\nabla g(\mathbf{\dot{W}}_{k})+\beta\mathcal{B}^{\#}(\mathcal{C}{\mathbf{\dot{X}}}_{k+1}-\mathcal{B}\mathbf{\dot{W}}_{k+1}-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta)+L_{2,k}\Delta\mathbf{\dot{W}}_{k+1}=0.$ (40)
It follows from (40) and the update of $\mathbf{\dot{\Lambda}}_{k+1}$ that
$\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k+1}}$ $=\nabla g(\mathbf{\dot{W}}_{k+1})-\nabla g(\mathbf{\dot{W}}_{k})+\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}-L_{2,k}\Delta\mathbf{\dot{W}}_{k+1}\in\partial_{\mathbf{\dot{W}}}\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1}).$ (41)
$\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k+1}}$ $=\frac{1}{\beta\mu}\Delta\mathbf{\dot{\Lambda}}_{k+1}\in\partial_{\mathbf{\dot{\Lambda}}}\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1}).$
Combining (39), (41), we obtain
$\|\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k+1}}\|$ $\leq\|\mathcal{C}\|\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}+q_{1}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}+\beta\|\mathcal{C}\|\|\mathcal{B}\|\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}},$
$\|\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k+1}}\|$ $\leq\|\mathcal{B}\|\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}+(q_{2}+L_{g})\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}},\quad\|\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k+1}}\|=\frac{1}{\beta\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}.$
Therefore, the following inequality holds:
$|\|\widetilde{\mathbf{\dot{D}}}_{k+1}\||\leq\|\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k+1}}\|_{\mathrm{F}}+\|\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k+1}}\|_{\mathrm{F}}+\|\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k+1}}\|_{\mathrm{F}}\leq\rho(\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}),$ (42)
which completes the proof. ∎
Lemma 4.7.
Suppose that Assumption 4.1 holds. Any limit point $({\mathbf{\dot{X}}}_{*},\mathbf{\dot{W}}_{*},\mathbf{\dot{\Lambda}}_{*})$ of the sequence $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{j\geq 0}$ generated by (17) is a stationary point of $\mathcal{L}_{\beta}({\mathbf{\dot{X}}},\mathbf{\dot{W}};\mathbf{\dot{\Lambda}})$, i.e.
$0=\partial f({\mathbf{\dot{X}}}_{*})+\frac{1}{\tau^{2}}\mathcal{A}^{\#}\mathcal{P}^{\#}_{\Omega}(\mathcal{P}_{\Omega}(\mathcal{A}{\mathbf{\dot{X}}}_{k+1}-\mathbf{\dot{Y}}))+\mathcal{C}^{\#}\mathbf{\dot{\Lambda}}_{*},\quad 0\in\partial g(\mathbf{\dot{W}}_{*})+\mathcal{B}^{\#}\mathbf{\dot{\Lambda}}_{*},\quad\mathcal{C}{\mathbf{\dot{X}}}_{*}+\mathcal{B}\mathbf{\dot{W}}_{*}=\mathbf{\dot{B}}.$
Proof.
Let $\{({\mathbf{\dot{X}}}_{k_{j}},\mathbf{\dot{W}}_{k_{j}},\mathbf{\dot{\Lambda}}_{k_{j}})\}_{j\geq 0}$ be a subsequence of $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ such that $({\mathbf{\dot{X}}}_{*},\mathbf{\dot{W}}_{*},\mathbf{\dot{\Lambda}}_{*})=\lim_{j\rightarrow\infty}({\mathbf{\dot{X}}}_{k_{j}},\mathbf{\dot{W}}_{k_{j}},\mathbf{\dot{\Lambda}}_{k_{j}})$. By the continuity of $\mathcal{L}_{\beta}$, $\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k_{j}},\mathbf{\dot{W}}_{k_{j}},\mathbf{\dot{\Lambda}}_{k_{j}})\rightarrow\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{*},\mathbf{\dot{W}}_{*},\mathbf{\dot{\Lambda}}_{*})$ as $j\rightarrow\infty$. Let $\mathbf{\dot{D}}_{k_{j}}\in\partial\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k_{j}},\mathbf{\dot{W}}_{k_{j}},\mathbf{\dot{\Lambda}}_{k_{j}})$ and according to Lemma 4.6, it follows that $\|\mathbf{\dot{D}}_{k_{j}}\|_{\mathrm{F}}\leq\rho\left(\|\Delta\mathbf{\dot{W}}_{k_{j}}\|_{\mathrm{F}}+\|\Delta{\mathbf{\dot{X}}}_{k_{j}}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k_{j}}\|_{\mathrm{F}}\right).$ According to Theorem 4.2, it follows that $\mathbf{\dot{D}}_{k_{j}}\rightarrow 0.$ By the closeness criterion of the limiting sub-differential, $({\mathbf{\dot{X}}}_{*},\mathbf{\dot{W}}_{*},\mathbf{\dot{\Lambda}}_{*})\in\text{crit}\,\mathcal{L}_{\beta}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})$. The proof is completed. ∎
The following Lemma states the behavior of limit points of $\{{\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k}\}.$
Lemma 4.8.
Suppose Assumption 4.1 holds. If $({\mathbf{\dot{X}}}_{*},\mathbf{\dot{W}}_{*},\mathbf{\dot{\Lambda}}_{*})$ is a limit point of a converging subsequence $\{{\mathbf{\dot{X}}}_{k_{j}},\mathbf{\dot{W}}_{k_{j}},\mathbf{\dot{\Lambda}}_{k_{j}}\}$, then it follows that
$\lim_{j\rightarrow\infty}\mathcal{R}_{k_{j}}=\mathcal{R}({\mathbf{\dot{X}}}_{*},\mathbf{\dot{W}}_{*},\mathbf{\dot{\Lambda}}_{*},\mathbf{\dot{W}}_{*},\mathbf{\dot{\Lambda}}_{*})=\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{*},\mathbf{\dot{W}}_{*},\mathbf{\dot{\Lambda}}_{*})=g(\mathbf{\dot{W}}_{*})+f({\mathbf{\dot{X}}}_{*})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}_{*}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2}.$
Proof.
Let $\{({\mathbf{\dot{X}}}_{k_{j}},\mathbf{\dot{W}}_{k_{j}},\mathbf{\dot{\Lambda}}_{k_{j}})\}_{j\geq 0}$ be a subsequence generated by (17) such that $({\mathbf{\dot{X}}}_{k_{j}},\mathbf{\dot{W}}_{k_{j}},\mathbf{\dot{\Lambda}}_{k_{j}})\rightarrow({\mathbf{\dot{X}}}_{*},\mathbf{\dot{W}}_{*},\mathbf{\dot{\Lambda}}_{*})$ as $j\rightarrow\infty$. According to Lemma 4.2, $\|\Delta{\mathbf{\dot{X}}}_{k_{j}}\|_{\mathrm{F}}\rightarrow 0$ and $\|\mathcal{C}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k_{j}}\|_{\mathrm{F}}\leq\|\mathcal{C}\|\|\Delta\mathbf{\dot{\Lambda}}_{k_{j}}\|_{\mathrm{F}}\rightarrow 0$ when $j\rightarrow\infty$. According to the definition of $\mathcal{R}_{k}$, it follows that
$\lim_{j\rightarrow\infty}\mathcal{R}_{k_{j}}=\lim_{j\rightarrow\infty}\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k_{j}},\mathbf{\dot{W}}_{k_{j}},\mathbf{\dot{\Lambda}}_{k_{j}}).$
Since $\|\Delta\mathbf{\dot{\Lambda}}_{k_{j}+1}\|_{\mathrm{F}}\rightarrow 0$ as $j\rightarrow\infty$, $\|\mathcal{C}{\mathbf{\dot{X}}}_{k_{j}}+\mathcal{B}\mathbf{\dot{W}}_{k_{j}}-\mathbf{\dot{B}}\|_{\mathrm{F}}\rightarrow 0$. According to the definition of $\mathcal{L}_{\beta}$ and the fact that $\{\mathbf{\dot{\Lambda}}_{k_{j}}\}_{j\geq 0}$ is a bounded sequence we also have $\left\langle\mathbf{\dot{\Lambda}}_{k_{j}},\mathcal{C}{\mathbf{\dot{X}}}_{k_{j}}+\mathcal{B}\mathbf{\dot{W}}_{k_{j}}-b\right\rangle\rightarrow 0$ as $j\rightarrow\infty$. According to the fact that $f$ is continuous, it follows that:
$\lim_{j\rightarrow\infty}\mathcal{R}({\mathbf{\dot{X}}}_{k_{j}},\mathbf{\dot{W}}_{k_{j}},\mathbf{\dot{\Lambda}}_{k_{j}},\mathbf{\dot{W}}_{k_{j}},\mathbf{\dot{\Lambda}}_{k_{j}})=\lim_{j\rightarrow\infty}\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k_{j}},\mathbf{\dot{W}}_{k_{j}},\mathbf{\dot{\Lambda}}_{k_{j}})$
$=$ $f({\mathbf{\dot{X}}}_{*})+g(\mathbf{\dot{W}}_{*})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}_{*}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2}.$
Theorem 4.3 (Properties of limit point set).
Suppose Assumption 4.1 holds. Sequence $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ generated by (17) satisfies the following properties:
  1. 1.
    The limit point set of the sequence $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}$, denoted by $\omega\left(\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}\right)$ is nonempty, and compact.
  2. 2.
    $\lim_{k\rightarrow\infty}\text{dist}\left[({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k}),\omega\left(\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}\right)\right]=0.$
  3. 3.
    $\omega\left(\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}\right)\subseteq\text{crit}\,\mathcal{L}_{\beta}$.
Proof.
Since $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ is bounded by Theorem 4.1. There exists at least one limit point of $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ and $\omega\left(\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}\right)$ is bounded. According to [27, Chapter 2, Exercise 6], it follows that $\omega\left(\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}\right)$ is closed. Hence $\omega\left(\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}\right)$ is nonempty and compact. As a consequence, $\lim_{k\rightarrow\infty}\text{dist}\left[({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k}),\omega\left(\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}\right)\right]=0.$ It follows from Lemma 4.7 that $\omega\left(\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}\right)\subseteq\text{crit}\,\mathcal{L}_{\beta}$, which completes the proof.
Lemma 4.9.
Suppose that the Assumption 4.1 holds. Let $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ be a sequence generated by (17). Define
$\mathbf{\dot{S}}_{{\mathbf{\dot{X}}}_{k}}:=\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k}},\quad\mathbf{\dot{S}}_{\mathbf{\dot{W}}_{k}}:=\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k}}+2r\theta_{0}\Delta\mathbf{\dot{W}}_{k},\quad\mathbf{\dot{S}}_{\mathbf{\dot{\Lambda}}_{k}}:=\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k}}+2r\theta_{2}\mathcal{B}\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k},$
$\mathbf{\dot{S}}_{{\mathbf{\dot{W}}^{\prime}}_{k}}:=-2r\theta_{0}\Delta\mathbf{\dot{W}}_{k},\quad\mathbf{\dot{S}}_{{\mathbf{\dot{\Lambda}}^{\prime}}_{k}}:=-2\theta_{2}\mathcal{B}\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k}$
where $(\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k}},\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k}},\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k}})\in\partial\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})$. Then
$\widetilde{\mathbf{\dot{S}}}_{k}:=(\mathbf{\dot{S}}_{{\mathbf{\dot{X}}}_{k}},\mathbf{\dot{S}}_{\mathbf{\dot{W}}_{k}},\mathbf{\dot{S}}_{\mathbf{\dot{\Lambda}}_{k}},\mathbf{\dot{S}}_{{\mathbf{\dot{W}}^{\prime}}_{k}},\mathbf{\dot{S}}_{{\mathbf{\dot{\Lambda}}^{\prime}}_{k}})\in\partial\mathcal{R}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},\mathbf{\dot{\Lambda}}_{k-1})$
for $k\geq 1$, and it holds that
$|||\widetilde{\mathbf{\dot{S}}}_{k}|||\leq\tilde{\pi}\Big{(}\|\Delta{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k}\|_{\mathrm{F}}\Big{)},$ (43)
where
$\tilde{\pi}=\sqrt{3}\pi+4r\max\{\theta_{0},\theta_{2}\|\mathcal{B}\|^{2}\},$ (44)
$r>1$, $\pi$ is given in (38).
Proof.
Let $k\geq 1$ fixed, and $(\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k}},\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k}},\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k}})\in\partial\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})$. By taking partial derivatives of $\mathcal{R}_{k}$ with respect to ${\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}},\mathbf{\dot{W}}^{\prime},\mathbf{\dot{\Lambda}}^{\prime}$ we obtain
$\mathbf{\dot{S}}_{{\mathbf{\dot{X}}}_{k}}$ ::=:= $\partial_{\mathbf{\dot{X}}}\mathcal{R}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},\mathbf{\dot{\Lambda}}_{k-1})=\partial_{\mathbf{\dot{X}}}\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})=\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k}},$
$\mathbf{\dot{S}}_{\mathbf{\dot{W}}_{k}}$ ::=:= $\nabla_{\mathbf{\dot{W}}}\mathcal{R}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},\mathbf{\dot{\Lambda}}_{k-1})=\nabla_{\mathbf{\dot{W}}}\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})+2r\theta_{0}\Delta\mathbf{\dot{W}}_{k}=\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k}}+2r\theta_{0}\Delta\mathbf{\dot{W}}_{k},$
$\mathbf{\dot{S}}_{\mathbf{\dot{\Lambda}}_{k}}$ ::=:= $\nabla_{\mathbf{\dot{\Lambda}}}\mathcal{R}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},\mathbf{\dot{\Lambda}}_{k-1})=\nabla_{\mathbf{\dot{\Lambda}}}\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})+2r\theta_{2}\mathcal{B}\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k}=\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k}}+2r\theta_{2}\mathcal{B}\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k},$
$\mathbf{\dot{S}}_{{\mathbf{\dot{W}}^{\prime}}_{k}}$ ::=:= $\nabla_{\mathbf{\dot{W}}^{\prime}}\mathcal{R}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},\mathbf{\dot{\Lambda}}_{k-1})=-2r\theta_{0}\Delta\mathbf{\dot{W}}_{k},$
$\mathbf{\dot{S}}_{{\mathbf{\dot{\Lambda}}^{\prime}}_{k}}$ ::=:= $\nabla_{\mathbf{\dot{\Lambda}}^{\prime}}\mathcal{R}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},\mathbf{\dot{\Lambda}}_{k-1})=-2r\theta_{2}\mathcal{B}\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k}.$
By the triangle inequality, we obtain
$\|\mathbf{\dot{S}}_{{\mathbf{\dot{X}}}_{k}}\|_{\mathrm{F}}=\|\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k}}\|_{\mathrm{F}},\quad\|\mathbf{\dot{S}}_{\mathbf{\dot{W}}_{k}}\|_{\mathrm{F}}\leq\|\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k}}\|_{\mathrm{F}}+2r\theta_{0}\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}},$
$\|\mathbf{\dot{S}}_{\mathbf{\dot{\Lambda}}_{k}}\|_{\mathrm{F}}\leq\|\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k}}\|_{\mathrm{F}}+2r\theta_{2}\|\mathcal{B}\|^{2}\|\Delta\mathbf{\dot{\Lambda}}_{k}\|_{\mathrm{F}},\quad\|\mathbf{\dot{S}}_{{\mathbf{\dot{W}}^{\prime}}_{k}}\|_{\mathrm{F}}=2r\theta_{0}\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}},\quad\|\mathbf{\dot{S}}_{{\mathbf{\dot{\Lambda}}^{\prime}}_{k}}\|_{\mathrm{F}}=2r\theta_{2}\|\mathcal{B}\|^{2}\|\Delta\mathbf{\dot{\Lambda}}_{k}\|_{\mathrm{F}}.$
By Lemma 4.6, it follows that
$|||\widetilde{\mathbf{\dot{S}}}_{k}|||$$\leq$ $\|\mathbf{\dot{S}}_{{\mathbf{\dot{X}}}_{k}}\|_{\mathrm{F}}+\|\mathbf{\dot{S}}_{\mathbf{\dot{W}}_{k}}\|_{\mathrm{F}}+\|\mathbf{\dot{S}}_{\mathbf{\dot{\Lambda}}_{k}}\|_{\mathrm{F}}+\|\mathbf{\dot{S}}_{{\mathbf{\dot{W}}^{\prime}}_{k}}\|_{\mathrm{F}}+\|\mathbf{\dot{S}}_{{\mathbf{\dot{\Lambda}}^{\prime}}_{k}}\|_{\mathrm{F}}$
$\leq$ $\|\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k}}\|_{\mathrm{F}}+\|\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k}}\|_{\mathrm{F}}+\|\mathbf{\dot{D}}^{k}_{\mathbf{\dot{\Lambda}}}\|_{\mathrm{F}}+4r\theta_{0}\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}+4r\theta_{2}\|\mathcal{B}\|^{2}\|\Delta\mathbf{\dot{\Lambda}}_{k}\|_{\mathrm{F}}$
$\leq$ $\sqrt{3}|||\mathbf{\dot{D}}_{k}|||+4r\theta_{0}\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}+4r\theta_{2}\|\mathcal{B}\|^{2}\|\Delta\mathbf{\dot{\Lambda}}_{k}\|_{\mathrm{F}}$
$\leq$ $\sqrt{3}\rho\|\Delta{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}+(\sqrt{3}\rho+4r\theta_{0})\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}+(\sqrt{3}\rho+4r\theta_{2}\|\mathcal{B}\|^{2})\|\Delta\mathbf{\dot{\Lambda}}_{k}\|_{\mathrm{F}},$
which completes the proof. ∎
We next prove the limiting behavior of $\mathcal{R}_{k}$, the detailed proof is similar to Theorem 4.3 and we omit it.
Lemma 4.10.
Suppose the Assumption 4.1 holds. If $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ is a sequence generated by Algorithm 1, then the following statements hold.
  • (i)
    The set $\varGamma:=\omega\Big{(}\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},\mathbf{\dot{\Lambda}}_{k-1})\}_{k\geq 1}\Big{)}$ is nonempty, and compact.
  • (ii)
    $\varGamma\subseteq\{({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})\in\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{p}\times\mathbb{R}^{m}\times\mathbb{R}^{p}:\;({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})\in{\rm crit}(\mathcal{L}_{\beta})\Big{\}}.$
  • (iii)
    $\lim_{k\to\infty}{\rm dist}\Big{[}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},{\mathbf{\dot{X}}}_{k-1}),\varGamma\Big{]}=0$.
  • (iv)
    The sequences $\{\mathcal{R}_{k}\}_{k\geq 0}$, $\{\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ approach to the same limit and if $({\mathbf{\dot{X}}}_{*},\mathbf{\dot{W}}_{*},\mathbf{\dot{\Lambda}}_{*},\mathbf{\dot{W}}_{*},\mathbf{\dot{\Lambda}}_{*})\in\varGamma$, then
    $\mathcal{R}({\mathbf{\dot{X}}}_{*},\mathbf{\dot{W}}_{*},\mathbf{\dot{\Lambda}}_{*},\mathbf{\dot{W}}_{*},\mathbf{\dot{\Lambda}}_{*})=\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{*},\mathbf{\dot{W}}_{*},\mathbf{\dot{\Lambda}}_{*})=g(\mathbf{\dot{W}}_{*})+f({\mathbf{\dot{X}}}_{*})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}_{*}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2}.$
Proof.
The results follow immediately from Theorem 4.1, 4.2 and Lemma 4.9. ∎
The following convergence result of the whole sequence, established under the Kurdyka-Łojasiewicz property, follows from [36, Theorem 2]. The difference lies in the construction of $\mathcal{R}_{k}$. The proofs are given for completeness.
Theorem 4.4 (Global Convergence).
Suppose that Assumption 4.1 holds and $\mathcal{R}$ defined in (30) satisfies the KL property on the limit point set $\varGamma$, i.e. for every $\mathbf{\dot{V}}_{*}=({\mathbf{\dot{X}}}_{*},\mathbf{\dot{W}}_{*},\mathbf{\dot{\Lambda}}_{*},\mathbf{\dot{W}}_{*},\mathbf{\dot{\Lambda}}_{*})\in\varGamma$, there exists $\varepsilon>0$ and desingularizing function $\psi:[0,\eta]\rightarrow[0,\infty),$ for some $\eta\in[0,\infty)$ such that for all $\mathbf{\dot{V}}=({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}},\mathbf{\dot{W}}^{\prime},\mathbf{\dot{\Lambda}}^{\prime})$ in the following set:
$\mathcal{S}:=\{\mathbf{\dot{V}}:\text{dist}(\mathbf{\dot{V}},\varGamma)<\varepsilon\,\,\text{and}\,\,\mathcal{R}(\mathbf{\dot{V}}^{*})<\mathcal{R}(\mathbf{\dot{V}})<\mathcal{R}(\mathbf{\dot{V}}^{*})+\eta\},$ (45)
the inequality
$\psi^{\prime}(\mathcal{R}(\mathbf{\dot{V}})-\mathcal{R}(\mathbf{\dot{V}}^{*}))\text{dist}(0,\partial\mathcal{R}(\mathbf{\dot{V}}))\geq 1$ (46)
holds, then $\{\mathbf{\dot{T}}_{k}\}_{k\geq 0}:=\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ satisfies the finite length property: $\sum_{k=0}^{\infty}\|{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}+\|\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}+\|\mathbf{\dot{\Lambda}}_{k}\|_{\mathrm{F}}<\infty,$ and consequently converges to a stationary point of (15).
Proof.
According to Lemma 4.5, $\lim_{k\rightarrow\infty}\mathcal{R}_{k}:=\mathcal{R}_{\infty}$. It follows that the sequence $\{\mathcal{E}_{k}\}_{k\geq k_{0}}$ defined by
$\mathcal{E}_{k}:=\mathcal{R}_{k}-\mathcal{R}_{\infty},$
is non-negative, monotonically decreasing, and converges to 0. We consider the following two cases:
Case 1. There exists $k_{1}\geq k_{0}$ such that $\mathcal{E}_{k_{1}}=0$. Hence $\mathcal{E}_{k}=0$ for all $k\geq k_{1}$ and, it follows that $\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2}\leq\frac{1}{a}(\mathcal{E}_{k}-\mathcal{E}_{k+1})=0,\quad\forall k\geq k_{1}$ by (31). This gives rise to
$\sum_{k\geq 0}\Big{(}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}\Big{)}\leq\sum_{k=1}^{k_{1}}\Big{(}\|\Delta{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k}\|_{\mathrm{F}}\Big{)}<+\infty.$
The latter conclusion is due to the fact that the sequence is bounded from Theorem 4.1.
Case 2. The error sequence $\mathcal{E}_{k}=\mathcal{R}_{k}-\mathcal{R}_{\infty}>0$ for all $k\geq k_{0}$. Then by (31), it follows that
$|||\Delta\mathbf{\dot{T}}_{k+1}\||_{\mathrm{F}}^{2}=\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2}_{\mathrm{F}}+\|\Delta\mathbf{\dot{W}}_{k+1}\|^{2}_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}\leq\frac{1}{a}(\mathcal{E}_{k}-\mathcal{E}_{k+1}),\quad\forall k\geq k_{0}.$ (47)
By Lemma 4.10, $\varGamma$ is nonempty, compact. Furthermore, $\mathcal{R}_{k}$ takes on a constant value $\mathcal{R}_{\infty}$ on $\varGamma$.
Since the sequence $\{\mathcal{R}_{k}\}_{k\geq k_{0}}$ is monotonically decreasing to $\mathcal{R}_{\infty}$, there exists $k_{1}\geq k_{0}\geq 1$ such that
$\mathcal{R}_{\infty}<\mathcal{R}_{k}<\mathcal{R}_{\infty}+\eta,\quad\forall k\geq k_{1}.$
By Theorem 4.10, it follows that $\lim_{k\to\infty}{\rm dist}\Big{[}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},{\mathbf{\dot{X}}}_{k-1}),\varGamma\Big{]}=0.$ Thus there exists $k_{2}\geq 1$ such that
${\rm dist}\Big{[}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},{\mathbf{\dot{X}}}_{k-1}),\varGamma\Big{]}<\varepsilon,\quad\forall k\geq k_{2}.$
Let $\tilde{k}=\max\{k_{1},k_{2},3\}$. It follows that $({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},{\mathbf{\dot{X}}}_{k-1})\in\mathcal{S}$ for $k\geq\tilde{k}$, where $\mathcal{S}$ defined in (45). Thus, we have
$\psi^{\prime}(\mathcal{E}_{k})\cdot{\rm dist}\Big{(}0,\partial\mathcal{R}_{k}\Big{)}\geq 1.$ (48)
Since $\psi$ is concave, it holds that $\psi(\mathcal{E}_{k})-\psi(\mathcal{E}_{k+1})\geq\psi^{\prime}(\mathcal{E}_{k})(\mathcal{E}_{k}-\mathcal{E}_{k+1}).$ Together with (47) and (48), it follows that
$|||\Delta\mathbf{\dot{T}}_{k+1}|||^{2}_{\mathrm{F}}$$\leq$ $\psi^{\prime}(\mathcal{E}_{k})|||\Delta\mathbf{\dot{T}}_{k+1}|||^{2}_{\mathrm{F}}\cdot{\rm dist}(0,\partial\mathcal{R}_{k})$
$\leq$ $\frac{1}{a}\psi^{\prime}(\mathcal{E}_{k})(\mathcal{E}_{k}-\mathcal{E}_{k+1})\cdot{\rm dist}(0,\partial\mathcal{R}_{k})$
$\leq$ $\frac{1}{a}\Big{(}\psi(\mathcal{E}_{k})-\psi(\mathcal{E}_{k+1})\Big{)}\cdot{\rm dist}(0,\partial\mathcal{R}_{k}).$
By the arithmetic mean-geometric mean inequality, for any $\gamma>0$, we have
$|\|\Delta\mathbf{\dot{T}}_{k+1}\||_{\mathrm{F}}\leq\frac{\gamma}{2a}\Big{(}\psi(\mathcal{E}_{k})-\psi(\mathcal{E}_{k+1})\Big{)}+\frac{1}{2\gamma}{\rm dist}(0,\partial\mathcal{R}_{k}).$
It follows that
$\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}\leq\frac{\sqrt{3}\gamma}{2a}\Big{(}\psi(\mathcal{E}_{k})-\psi(\mathcal{E}_{k+1})\Big{)}+\frac{\sqrt{3}}{2\gamma}{\rm dist}(0,\partial\mathcal{R}_{k}).$ (49)
Then according to Lemma 4.9, we obtain
$\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}\leq\frac{\sqrt{3}\gamma}{2a}\Big{(}\psi(\mathcal{E}_{k})-\psi(\mathcal{E}_{k+1})\Big{)}+\frac{\sqrt{3}\tilde{\rho}}{2\gamma}\Big{(}\|\Delta{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k}\|_{\mathrm{F}}\Big{)}.$ (50)
According to the equality $\sum_{k=\underline{k}}^{K}\|\Delta{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}=\sum_{k=\underline{k}}^{K}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}+\|\Delta{\mathbf{\dot{X}}}_{\underline{k}}\|-\|\Delta{\mathbf{\dot{X}}}_{K}\|_{\mathrm{F}}$, we choose $\gamma>0$ large enough such that $1>\sqrt{3}\tilde{\rho}/2\gamma$, and let $\delta_{0}=1-\frac{\sqrt{3}\tilde{\rho}}{2\gamma}$. Summing up (50) from $k=\underline{k}\geq\tilde{k}$ to $K\geq\underline{k}$ gives
$\delta_{0}(\sum_{k=\underline{k}}^{K}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}})\leq\frac{\sqrt{3}\gamma}{2a}\Big{(}\psi(\mathcal{E}_{\underline{k}})-\psi(\mathcal{E}_{K+1})\Big{)}$
$+\frac{\sqrt{3}\tilde{\rho}}{2\gamma}\Big{(}\|\Delta{\mathbf{\dot{X}}}_{\underline{k}}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{W}}_{\underline{k}}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{\underline{k}}\|_{\mathrm{F}}\Big{)}-\frac{\sqrt{3}\tilde{\rho}}{2\gamma}\Big{(}\|\Delta{\mathbf{\dot{X}}}_{K}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{W}}_{K}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{K}\|_{\mathrm{F}}\Big{)}.$
Recall that $\mathcal{E}_{k}$ is monotonically decreasing and $\psi(\mathcal{E}_{k})\geq\psi(\mathcal{E}_{k+1})>0$, it follows that
$\sum_{k=\underline{k}}^{K}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}\leq\frac{\sqrt{3}\gamma}{2a\delta_{0}}\psi(\mathcal{E}_{\underline{k}})+\frac{\sqrt{3}\tilde{\rho}}{2\gamma\delta_{0}}\big{(}\|\Delta{\mathbf{\dot{X}}}_{\underline{k}}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{W}}_{\underline{k}}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{\underline{k}}\|_{\mathrm{F}}\big{)}.$
The right hand side of this inequality is bounded for any $K\geq\underline{k}$. Let $K\to\infty$ and we obtain
$\sum_{k\geq\underline{k}}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}\leq\frac{\sqrt{3}\gamma}{2a\delta_{0}}\psi(\mathcal{E}_{\underline{k}})+\frac{\sqrt{3}\tilde{\rho}}{2\gamma\delta_{0}}\Big{(}\|\Delta{\mathbf{\dot{X}}}_{\underline{k}}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{W}}_{\underline{k}}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{\underline{k}}\|_{\mathrm{F}}\Big{)}.$ (51)
Since $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ is a bounded sequence, for any $\underline{k}\in\mathbb{N}_{+}$, it follows that
$\lambda({\underline{k}}):=\sum_{k=1}^{\underline{k}}\|\Delta{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k}\|_{\mathrm{F}}<+\infty.$ (52)
By combining (51) and (52), we conclude that ${\sum_{k\geq 0}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}}$ is finite.
Note that for any $p,q,K\in\mathbb{N}_{+}$ where $q\geq p>0$, it follows that:
$|||\mathbf{\dot{T}}_{q}-\mathbf{\dot{T}}_{p}|||_{\mathrm{F}}$ $=$ $|||\sum_{k=p}^{q-1}\Delta\mathbf{\dot{T}}_{k+1}|||_{\mathrm{F}}\leq\sum_{k=p}^{q-1}|||\Delta\mathbf{\dot{T}}_{k+1}|||_{\mathrm{F}}$
$\leq$ $\sum_{k=p}^{q-1}\Big{(}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}\Big{)}$
$\leq$ $\sum_{k\geq 0}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}<\infty.$
This implies that $\{\mathbf{\dot{T}}_{k}\}_{k\geq 0}=\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ is a Cauchy sequence and hence converges. Moreover, by Theorem 4.3, it converges to a stationary point. ∎
Corollary 4.1.
Suppose the penalty parameter $\beta$ is large enough, and $L_{1,k},L_{2,k}$ are set as given constants for all $k$ such that (20) and Assumption (A5) hold, then the sequence $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ generated by Algorithm 1 converges to a stationary point of SLRQA (2).
Proof.
It follows directly from the definition of SLRQA (2) that Assumption A1,A2 and A3 are all satisfied. Since (20) and Assumption (A5) hold, Assumption A4, A5 are also satisfied. Furthermore, the Huber function and Frobenius norm are semi-algebraic [3, Example 2]. From [28, Appendix, Lemma 4], it follows that the nuclear norm is also a semi-algebraic function. Since $\phi$ is a semi-algebraic function and the fact that the composition and sum of semi-algebraic functions are also semi-algebraic, we can conclude that $\mathcal{R}$ is semi-algebraic and the KL-inequality 46 holds since any proper closed semi-algebraic function satisfies the KL-inequality [3, Theorem 3]. It follows from Theorem 4.4 that the sequence $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ generated by Algorithm 1 converges to a stationary point of SLRQA (2), which completes the proof. ∎
5 An Algorithm for SLRQA-NF model
In this section, we propose the PL-ADMM-NF algorithm to solve SLRQA-NF (4) which is summarized in Algorithm 3. It is noted that SLRQA-NF is a model of the limiting case of SLRQA. The linearization technique is applied to the update of $\mathbf{\dot{W}}$. Each iteration of PL-ADMM-NF consists of minimizing the augmented Lagrange function with respect to $\mathbf{\dot{W}}$ variable after linearization, minimizing the augmented Lagrange function with respect to ${\mathbf{\dot{X}}}$ variable, and updating the multiplier.
Algorithm 3 PL-ADMM-NF algorithm for SLRQA-NF in (4)
1:Corrupted images $\mathbf{\dot{Y}}$; two parameters $\mu\in(0,2)$ and $\beta>0$; a tolerance $\eta>0$; a sequence of regularization parameters $\{L_{1,k}\}$ satisfying $\sup_{k\geq 0}{L_{1,k}}<\infty$ and $\inf_{k>0}{L_{1,k}}>0$.
2:Recovered $\mathbf{\dot{Z}}_{*}$.
3:Set $\mathbf{\dot{Z}}_{0}=\bm{0}$, $\mathbf{\dot{W}}_{0}=\bm{0}$, $\mathbf{\dot{\Lambda}}_{1,0}=\bm{0}$, $\mathbf{\dot{\Lambda}}_{2,0}=\bm{0}$.
4:for $k=0,1,2,\ldots$ do
5:     $\mathbf{\dot{Z}}$ subproblem:
$\mathbf{\dot{Z}}_{k+1}\in\mathop{\mathrm{arg\,min}}_{\mathbf{\dot{Z}}}\sum_{i}\phi(\sigma_{i}(\mathbf{\dot{Z}}),\gamma)+\left\langle\mathbf{\dot{\Lambda}}_{1,k},\mathbf{\dot{Z}}-\mathbf{\dot{W}}_{k}\right\rangle+\frac{\beta}{2}\|\mathbf{\dot{Z}}-\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}^{2}+\frac{L_{1,k}}{2}\|\mathbf{\dot{Z}}-\mathbf{\dot{Z}}_{k}\|_{\mathrm{F}}^{2};$
6:     $\mathbf{\dot{W}}$ subproblem:
$\mathbf{\dot{W}}_{k+1}$ $=\mathop{\mathrm{arg\,min}}_{\mathbf{\dot{W}}}\lambda\left\langle\nabla p(\mathbf{\dot{W}}_{k}),\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\right\rangle+\beta\left\langle\mathbf{\dot{W}}_{k}-\mathbf{\dot{Z}}_{k+1}+\frac{1}{\beta}\mathbf{\dot{\Lambda}}_{1,k},\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\right\rangle$
$+\beta\left\langle\mathcal{W}\mathcal{A}^{\#}(\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{Y}})+\frac{1}{\beta}\mathbf{\dot{\Lambda}}_{2,k},\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\right\rangle+\frac{\beta\eta}{2}\|\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}^{2};$
7:     $\mathbf{\dot{\Lambda}}_{1,k+1}=\mathbf{\dot{\Lambda}}_{1,k}+\mu\beta\cdot(\mathbf{\dot{Z}}_{k+1}-\mathbf{\dot{W}}_{k+1})$;
8:     $\mathbf{\dot{\Lambda}}_{2,k+1}=\mathbf{\dot{\Lambda}}_{2,k}+\mu\beta\cdot\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{Y}})$;
9:     if $\|\mathbf{\dot{Z}}_{k+1}-\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+\|\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}_{k+1}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}<\eta$ then
10:         $\mathbf{\dot{Z}}_{*}=\mathbf{\dot{Z}}_{k+1}$; break
11:     end if
12:end for
The augmented Lagrangian function of SLRQA-NF (4) is given by:
~$\tilde{\mathcal{L}}_{\beta}(\mathbf{\dot{Z}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}}_{1},\mathbf{\dot{\Lambda}}_{2})=\sum_{i}\phi(\sigma_{i}(\mathbf{\dot{Z}}),\gamma)+\lambda p(\mathbf{\dot{W}})+\left\langle\mathbf{\dot{\Lambda}}_{1},\mathbf{\dot{Z}}-\mathbf{\dot{W}}\right\rangle$ (53)
$+\frac{\beta}{2}\|\mathbf{\dot{Z}}-\mathbf{\dot{W}}\|^{2}_{\mathrm{F}}+\left\langle\mathbf{\dot{\Lambda}}_{2},\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\right\rangle+\frac{\beta}{2}\|\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2}$
$=\sum_{i}\phi(\sigma_{i}(\mathbf{\dot{Z}}),\gamma)+\lambda p(\mathbf{\dot{W}})+\frac{\beta}{2}\|\mathbf{\dot{Z}}-\mathbf{\dot{W}}+\frac{1}{\beta}\mathbf{\dot{\Lambda}}_{1}\|_{\mathrm{F}}^{2}$
$+\frac{\beta}{2}\|\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})+\frac{1}{\beta}\mathbf{\dot{\Lambda}}_{2}\|_{\mathrm{F}}^{2}-\frac{1}{2\beta}\|\mathbf{\dot{\Lambda}}_{1}\|_{\mathrm{F}}^{2}-\frac{1}{2\beta}\|\mathbf{\dot{\Lambda}}_{2}\|^{2}_{\mathrm{F}},$
where $\beta>0$ is the penalty parameter, $\mathbf{\dot{\Lambda}}_{1},\mathbf{\dot{\Lambda}}_{2}\in\mathbb{H}^{m\times n}$ are the Lagrange multipliers.
In Algorithm 3, the subproblem of $\mathbf{\dot{Z}}$ in Step 5 is the proximal ADMM step i.e. the minimization of ~$\tilde{\mathcal{L}}_{\beta}(\mathbf{\dot{Z}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}}_{1},\mathbf{\dot{\Lambda}}_{2})$ for fixed $\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{1,k},\mathbf{\dot{\Lambda}}_{2,k}$ plus a proximal term $\frac{L_{1,k}}{2}\|\mathbf{\dot{Z}}-\mathbf{\dot{Z}}_{k}\|_{\mathrm{F}}^{2}.$ Since the subproblem is equivalent to
$\min_{\mathbf{\dot{Z}}}\,\sum_{i}\phi(\sigma_{i}(\mathbf{\dot{Z}}),\gamma)+\frac{\beta+L_{1,k}}{2}\left\|\mathbf{\dot{Z}}-\frac{1}{\beta+L_{1,k}}\left(\beta\mathbf{\dot{W}}_{k}+\mathbf{\dot{\Lambda}}_{1,k}+L_{1,k}\mathbf{\dot{Z}}_{k}\right)\right\|_{\mathrm{F}}^{2},$
Algorithm 2 can also be used to obtain $\mathbf{\dot{Z}}_{k+1}$.
In Step 6 of Algorithm 3, $\mathbf{\dot{W}}$, according to the linearization strategy in [21], a proximal gradient method is used in Step 6 to linearize the second order term. The penalty parameter of the second order term is denoted by $\beta\eta.$ To be specific, the quadratic term $\frac{\beta}{2}\|\mathbf{\dot{W}}-\mathbf{\dot{Z}}-\frac{1}{\beta}\mathbf{\dot{\Lambda}}_{1}\|_{\mathrm{F}}^{2}+\frac{\beta}{2}\|\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2}$ in ~$\tilde{\mathcal{L}}_{\beta}(\mathbf{\dot{Z}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}}_{1},\mathbf{\dot{\Lambda}}_{2})$ is replaced by the linearized term plus a regularized term $\beta\left\langle\mathbf{\dot{W}}_{k}-\mathbf{\dot{Z}}_{k+1}-\frac{1}{\beta}\mathbf{\dot{\Lambda}}_{1,k},\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\right\rangle+\beta\left\langle\mathcal{W}\mathcal{A}^{\#}(\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{Y}})+\frac{1}{\beta}\mathbf{\dot{\Lambda}}_{2,k},\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\right\rangle$. Hence, the update of $\mathbf{\dot{W}}_{k+1}$ is given by
$\mathbf{\dot{W}}_{k+1}=\mathbf{\dot{W}}_{k}-\frac{\mathcal{W}\mathcal{A}^{\#}(\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{Y}})+\frac{1}{\beta}\mathbf{\dot{\Lambda}}_{2,k})+\mathbf{\dot{W}}_{k}-\mathbf{\dot{Z}}_{k+1}-\frac{1}{\beta}\mathbf{\dot{\Lambda}}_{1,k}}{\eta}-\frac{\lambda}{\beta\eta}\nabla p(\mathbf{\dot{W}}_{k}).$ (54)
Step 7 and 8 is used to update the Lagrange multipliers. Finally, the recovered image ${\mathbf{\dot{X}}}_{*}$ is given by ${\mathbf{\dot{X}}}_{*}=\mathcal{W}^{\#}(\mathbf{\dot{Z}}_{*})$.
Remark 5.1.
{ The combination of two constraints in (4) with respect to $\mathbf{\dot{W}}$ and $\mathbf{\dot{Z}}$ can be represented by $(\mathbf{0};\mathbf{I})$ and $(\mathbf{T};\mathbf{I})$, where $\mathbf{T}$ is the matrix form of $\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#})$. It follows that Assumption A3 is not satisfied. However, this assumption is essential in all known proofs for nonconvex ADMM and its variants [29, 36, 22] for the global convergence. Hence, existing techniques have difficulties proving the convergence of Algorithm 3. It is challenging to prove the global convergence result of variants of nonconvex ADMM without Assumption A3. We leave it as future works.}
6 Numerical experiment
In this section, color image denoising and image inpainting problems are respectively used to demonstrate the performance of SLRQA in (2) with Algorithm 1 and SLRQA-NF in (4) with Algorithm 3. In Section 6.1, we give the parameter setting for SLRQA and SLRQA-NF. In Section 6.2, we compare SLRQA and SLRQA-NF models based on quaternion and RGB representation to demonstrate the rationality of quaternion representation. Furthermore, in Section 6.3 and 6.4, comparisons of SLRQA and SLRQA-NF with other state-of-the-art methods are also presented to respectively show the superiority of SLRQA and SLRQA-NF. All the experiments are performed using MATLAB R2022b running on a desktop with an Intel Core R7-5800H CPU (3.90 GHz) and 16 GB of RAM. The codes are available at https://github.com/dengzhanwang/SLRQA/tree/main.
6.1 Problem description, parameter setting, and testing environment
The models for color image denoising and image inpainting are respectively given by
$\min_{{\mathbf{\dot{X}}},\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}}$ $\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)+\lambda p(\mathbf{\dot{W}})+\frac{1}{2\tau^{2}}\|{\mathbf{\dot{X}}}-\mathbf{\dot{Y}}\|_{\mathrm{F}}^{2},\quad\mathrm{s.t.}\quad\mathcal{W}({\mathbf{\dot{X}}})=\mathbf{\dot{W}},$ (55)
and
$\min_{\mathbf{\dot{Z}},\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}}$ $\sum_{i}\phi(\sigma_{i}(\mathbf{\dot{Z}}),\gamma)+\lambda p(\mathbf{\dot{W}}),\quad\mathrm{s.t.}\quad\mathcal{P}_{\Omega}(\mathcal{W}^{\#}(\mathbf{\dot{W}}))=\mathcal{P}_{\Omega}(\mathbf{\dot{Y}}),\quad\mathbf{\dot{Z}}=\mathbf{\dot{W}},$ (56)
where $\mathcal{W}$ is the QDCT [12], $p(\mathbf{\dot{W}})$ is chosen as the Huber function, i.e., $(p(\mathbf{\dot{W}}))_{ij}=\frac{1}{2\delta}|\mathbf{\dot{W}}_{ij}|^{2}$ if $|\mathbf{\dot{W}}_{ij}|<\delta$; and $(p(\mathbf{\dot{W}}))_{ij}=|\mathbf{\dot{W}}_{ij}|-\frac{1}{2}\delta$ otherwise, $\delta>0$ is a prescribed constant, and the function $\phi$ are chosen as the Schatten-$\gamma$, Laplace, or Weighted Schatten-$\gamma$ functions. Note that the color image denoising problem (55) is the SLRQA in (2) with $\mathcal{A}$ being identity, $\Omega$ being the entire indices of pixels, and $p$ being the Huber function. The color image inpainting problem (56) is the SLRQA-NF in (4) with $\mathcal{A}$ being identity and $p$ being the Huber function. Problem (55) with the Schatten-$\gamma$, Laplace, and Weighted Schatten-$\gamma$ functions for $\phi$ is respectively denoted by SLRQA-1, SLRQA-2, and SLRQA-3. Likewise, Problem (56) using the three options of $\phi$ is denoted by SLRQA-NF-1, SLRQA-NF-2, and SLRQA-NF-3.
The ten images in Figure 1 are used to generate simulation problems. Specifically, for image denoising, the noisy images $\mathbf{\dot{Y}}$ are obtained by adding additive white Gaussian noise with zero mean and variance $\tau^{2}$ to the color images in Figure 1, and for image inpainting, each pixel of an image is chosen to be in the indices set $\Omega$ with the same probability $\chi\in(0,1)$. Therefore, $\chi$ is also called the missing rate. The parameters $\tau$, $\lambda$, and $\gamma$ are specified later when reporting numerical results. Note that the parameters $\gamma$ and $\lambda$ depend on the choice of the surrogate function $\phi$, the noise level, and whether the non-local self-similarity is used. The parameter $\delta$ is set depending on the accuracy of the current iterate. Specifically, $\delta=1$ if $\epsilon_{k}\geq 10^{-2}$, $\delta=10^{-2}$ if $10^{-3}\leq\epsilon_{k}<10^{-2}$, and $\delta=10^{-4}$ otherwise, where $\epsilon_{k}=\|{\mathbf{\dot{X}}}_{k+1}-{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}+\|\mathbf{\dot{W}}_{k+1}-\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}+\|\mathcal{W}({\mathbf{\dot{X}}}_{k+1})-\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}$ in Algorithm 1 and $\epsilon_{k}=\|\mathbf{\dot{Z}}_{k+1}-\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+\|\mathcal{P}_{\Omega}(\mathcal{W}^{\#}(\mathbf{\dot{W}}_{k+1}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}$ in Algorithm 3.
The parameters in Algorithm 1, 2 and 3 are set as $\mu=1.1$, $\beta=10$, $L_{1,k}=L_{2,k}=1,\forall k$, $\eta_{C}=10^{-10}$, and $\eta=10^{-4}$. Note that $\eta_{C}=10^{-10}$ is for the high accuracy of the ${\mathbf{\dot{X}}}$ or $\mathbf{\dot{Z}}$-subproblem.
The qtfm [1] package is used for the quaternion computations except the QSVD. The QSVD of a quaternion matrix ${\mathbf{\dot{X}}}\in\mathbb{H}^{m\times n}$ is computed by taking the SVD of the complex adjoint form of ${\mathbf{\dot{X}}}$, which needs an SVD of a $2m\times 2n$ complex matrix and is more efficient. For the details, we refer to [10].
Peak Signal-to-Noise Ration (PSNR) and Structure Similarity (SSIM) [30] are chosen to measure the quality of the images recovered by all methods. Specifically, the PSNR of two color image $\mathbf{X}$ and $\mathbf{Y}\in\mathbb{R}^{m\times n\times 3}$ is calculated as: ${\rm PSNR}(\mathbf{X},\mathbf{Y})=10\log_{10}\frac{\textit{peakval}}{\rm{MSE}},$ where $\text{MSE}=\frac{1}{3mn}\sum_{i=1}^{m}\sum_{j=1}^{n}\sum_{k=1}^{3}|\mathbf{X}_{i,j,k}-\mathbf{Y}_{i,j,k}|^{2},$ and peakval is taken from the range of the image data type. Furthermore, the SSIM of $\mathbf{X}$ and $\mathbf{Y}$ is given by
$\text{SSIM}(\mathbf{X}_{ij},\mathbf{Y}_{ij})=\frac{1}{MN}\sum_{i=1}^{M}\sum_{j=1}^{N}\text{SSIM}(\mathbf{X}_{ij},\mathbf{Y}_{ij})$ (57)
$\text{SSIM}(\mathbf{X}_{ij},\mathbf{Y}_{ij})=l(\mathbf{X}_{ij},\mathbf{Y}_{ij})\,c(\mathbf{X}_{ij},\mathbf{Y}_{ij})\,s(\mathbf{X}_{ij},\mathbf{Y}_{ij}),$ (58)
where $\mathbf{X}_{ij}$ and $\mathbf{Y}_{ij}$ denote the local image patches, $M$ and $N$ denote the number of local patches along horizontal and vertical direction respectively, the three functions $l(\mathbf{X}_{ij},\mathbf{Y}_{ij})$, $c(\mathbf{X}_{ij},\mathbf{Y}_{ij})$, and $s(\mathbf{X}_{ij},\mathbf{Y}_{ij})$ respectively denote luminance item, contrast item, and structure item, and they are defined as follows:
$l(\mathbf{X}_{ij},\mathbf{Y}_{ij})=\frac{2\mu_{\mathbf{X}_{ij}}\mu_{\mathbf{Y}_{ij}}+C_{1}}{\mu_{\mathbf{X}_{ij}}^{2}+\mu_{\mathbf{Y}_{ij}}^{2}+C_{1}},\quad c(\mathbf{X}_{ij},\mathbf{Y}_{ij})=\frac{2\sigma_{\mathbf{X}_{ij}}\sigma_{\mathbf{Y}_{ij}}+C_{2}}{\sigma_{\mathbf{X}_{ij}}^{2}+\sigma_{\mathbf{Y}_{ij}}^{2}+C_{2}},\quad s(\mathbf{X}_{ij},\mathbf{Y}_{ij})=\frac{\sigma_{\mathbf{X}_{ij}\mathbf{Y}_{ij}}+C_{3}}{\beta_{\mathbf{X}_{ij}}\sigma_{\mathbf{Y}_{ij}}+C_{3}},$ (59)
with $\mu_{\mathbf{X}_{ij}},\mu_{\mathbf{Y}_{ij}},\sigma_{\mathbf{X}_{ij}},\sigma_{\mathbf{Y}_{ij}},\sigma_{{\mathbf{X}_{ij}}{\mathbf{Y}_{ij}}},C_{1},C_{2}$ and $C_{3}$ being the mean, standard deviation and cross-covariance of image patches $\mathbf{X}_{ij}$ and $\mathbf{Y}_{ij}$, $10^{-4},9\times 10^{-4}$ and $4.5\times 10^{-4}$ respectively. MATLAB commands psnr and ssim are used to compute the PSNR and SSIM for color images. In general, the higher values the PSNR and SSIM are, the better the denoising quality is.
6.2 Quaternion and RGB representation
Simulation problems of image denoising in Section 6.2.1 and image inpainting in Section 6.2.2 are used to compare the performance of quaternion and RGB representations.
6.2.1 Image denoising
The parameter $\gamma$ is set as 0.5 in SLRQA-1, SLRQA-2, and SLRQA-3. The parameter $w$ in SLRQA-3 is chosen depending on the number of iteration $k$ and the index $i$, i.e., $w_{i,k}=20/(\sigma_{i}({\mathbf{\dot{X}}}_{k-1})+10^{-4})$. Note that the definition of $\phi$ for SLRQA-3 depends on the index $i$. Such dependency has been used in e.g., [10]. The multiple values of $(\lambda,\tau)$ are used, i.e., $(\lambda,\tau)=(0.01,10),(0.3,30)$, and $(0.5,50)$.
The PSNR and SSIM values of SLRQA with quaternion and RGB representation are reported in Table 2. Furthermore, the comparison of the two representations for SLRQA-1 with different noise levels is shown in Figure 2. The figure and table show that SLRQA with quaternion representation achieves higher PSNR and SSIM compared with RGB representation in all cases. This observation coincides with the intuition that quaternion representation utilizes the information between channels to achieve better performance.
Figure 1: The 10 color images (512 $\times$ 512 $\times$ 3) for numerical experiments.
Table 2: PSNR/SSIM values of SLRQA using the RGB and quaternion representations. “QR” denotes that the quaternion representation is used. Bold fonts denote the best performance on the same conditions.
Model Image Image1 Image2 Image3 Image4 Image5 Image6 Image7 Image8 Image9 Image10
SLRQA-1 $\tau=10$ QR 32.38/0.82 30.43/0.81 31.89/0.81 33.86/0.87 33.27/0.86 31.21/0.81 31.20/0.78 31.20/0.78 33.53/0.70 31.54/0.79
RGB 31.90/0.80 30.27/0.78 31.50/0.77 33.60/0.79 32.56/0.82 30.88/0.77 30.10/0.75 30.34/0.75 32.89/0.68 31.02/0.72
SLRQA-1 $\tau=30$ QR 24.86/0.70 23.91/0.66 24.97/0.72 26.68/0.67 25.65/0.69 22.85/0.69 24.70/0.68 24.19/0.66 25.30/0.64 24.97/0.66
RGB 23.55/0.67 20.71/0.57 23.65/0.70 25.10/0.63 23.26/0.63 21.29/0.66 22.74/0.67 23.91/0.63 24.02/0.60 23.45/0.62
SLRQA-1 $\tau=50$ QR 23.33/0.50 22.04/0.48 23.01/0.51 22.23/0.46 21.84/0.52 21.24/0.50 23.57/0.48 22.14/0.48 22.78/0.56 22.53/0.57
RGB 22.67/0.49 21.56/0.46 22.66/0.49 21.73/0.44 21.48/0.51 20.94/0.48 23.18/0.46 21.55/0.47 21.54/0.55 21.67/0.54
SLRQA-2 $\tau=10$ QR 31.45/0.78 30.19/0.80 31.79/0.78 32.45/0.75 31.18/0.80 30.92/0.80 30.82/0.75 30.75/0.74 32.87/0.70 31.56/0.76
RGB 31.20/0.77 30.72/0.80 31.46/0.78 31.91/0.72 31.04/0.80 30.57/0.79 30.20/0.74 30.35/0.72 31.40/0.67 30.32/0.72
SLRQA-2 $\tau=30$ QR 24.30/0.61 22.77/0.64 25.21/0.64 22.74/0.57 25.32/0.58 21.72/0.64 24.39/0.68 24.06/0.62 25.45/0.68 25.02/0.70
RGB 24.12/0.60 22.47/0.63 24.98/0.62 22.52/0.56 25.02/0.58 21.59/0.62 24.26/0.68 23.75/0.60 24.82/0.65 24.30/0.64
SLRQA-2 $\tau=50$ QR 23.50/0.51 22.77/0.52 24.91/0.58 22.34/0.47 22.62/0.67 21.02/0.51 23.09/0.45 23.26/0.52 23.02/0.62 22.60/0.70
RGB 23.27/0.51 21.96/0.52 24.76/0.57 22.03/0.46 21.88/0.65 20.94/0.51 22.88/0.45 22.90/0.51 22.32/0.60 22.12/0.65
SLRQA-3 $\tau=10$ QR 32.86/0.81 31.20/0.82 32.03/0.84 30.36/0.75 32.50/0.87 30.77/0.76 31.10/0.80 31.92/0.76 33.89/0.73 31.87/0.82
RGB 32.76/0.80 30.80/0.81 31.30/0.84 30.01/0.74 32.04/0.86 30.17/0.74 30.48/0.78 31.65/0.75 32.20/0.68 32.02/0.78
SLRQA-3 $\tau=30$ QR 26.40/0.67 24.02/0.64 26.22/0.68 24.21/0.64 26.34/0.68 22.50/0.63 26.42/0.68 26.12/0.67 25.80/0.68 25.68/0.70
RGB 25.32/0.66 23.78/0.63 26.18/0.67 23.82/0.64 25.89/0.67 22.13/0.61 24.26/0.68 26.05/0.66 24.79/0.62 25.03/0.67
SLRQA-3 $\tau=50$ QR 23.90/0.51 22.06/0.45 24.99/0.60 23.51/0.49 22.72/0.65 22.00/0.53 23.92/0.53 24.32/0.56 23.04/0.63 23.10/0.65
RGB 23.57/0.50 21.46/0.43 24.76/0.60 22.63/0.50 22.08/0.63 21.94/0.53 23.18/0.52 23.58/0.55 22.23/0.56 22.30/0.62
Figure 2: PSNR and SSIM comparison of SLRQA-1 using quaternion and RGB representation. In the legend, the suffix (Q) means that quaternion representation is used, and (RGB) means that RGB representation is used. The numbers 10, 30, and 50 denote the variance of the noise respectively.
6.2.2 Image inpainting
Table 3: PSNR/SSIM values of SLRQA-NF using the RGB and quaternion representations. “QR” denotes that the quaternion representation is used.
Model Image Image1 Image2 Image3 Image4 Image5 Image6 Image7 Image8 Image9 Image10
SLRQA-NF-1 QR 35.58/0.96 31.37/0.97 34.56/0.98 42.48/0.99 39.38/0.98 32.73/0.99 32.45/0.97 35.08/1.00 34.93/0.99 34.41/0.99
($\chi$=0.3) RGB 34.66/0.93 30.83/0.96 33.66/0.97 40.19/0.98 38.63/0.99 30.10/0.90 30.85/0.92 33.91/0.98 34.11/0.99 33.61/0.99
SLRQA-NF-1 QR 32.03/0.94 28.04/0.94 30.40/0.96 37.56/0.98 34.82/0.98 25.75/0.98 27.88/0.92 31.06/0.97 31.58/0.98 31.28/0.98
($\chi$=0.5) RGB 31.19/0.90 27.59/0.93 29.77/0.93 36.42/0.98 34.28/0.97 25.10/0.77 27.35/0.86 30.33/0.96 30.91/0.98 30.63/0.98
SLRQA-NF-1 QR 28.29/0.88 25.20/0.89 26.35/0.93 33.16/0.99 31.00/0.96 23.78/0.96 26.63/0.81 27.12/0.94 28.90/0.97 28.26/0.97
($\chi$=0.7) RGB 25.49/0.71 22.53/0.79 23.23/0.75 29.27/0.89 28.42/0.86 20.59/0.59 24.48/0.77 24.72/0.86 25.85/0.77 25.53/0.93
SLRQA-NF-1 QR 25.88/0.84 23.14/0.86 24.91/0.90 30.74/0.95 28.63/0.93 21.61/0.93 25.29/0.74 25.30/0.92 25.54/0.96 26.67/0.96
($\chi$=0.8) RGB 22.41/0.54 20.63/0.69 21.12/0.67 26.51/0.81 24.92/0.78 18.14/0.46 22.82/0.71 22.38/0.79 23.97/0.61 22.54/0.91
SLRQA-NF-2 QR 35.33/0.96 30.97/0.97 33.56/0.98 40.86/0.99 38.09/0.99 31.26/0.97 32.12/0.97 34.13/0.99 34.49/0.99 34.18/0.99
($\chi$=0.3) RGB 34.18/0.95 30.83/0.97 33.15/0.98 39.19/0.99 37.81/0.99 26.62/0.99 26.85/0.95 33.54/0.98 33.97/0.99 33.84/0.99
SLRQA-NF-2 QR 31.25/0.94 27.53/0.94 29.65/0.96 36.27/0.99 33.75/0.98 26.62/0.91 28.91/0.94 30.15/0.97 31.08/0.98 30.83/0.98
($\chi$=0.5) RGB 30.65/0.92 27.49/0.94 29.46/0.95 35.60/0.97 33.63/0.97 24.62/0.96 26.91/0.90 29.92/0.96 30.78/0.98 30.63/0.98
SLRQA-NF-2 QR 27.30/0.88 24.05/0.89 25.87/0.93 31.87/0.99 29.95/0.96 22.70/0.80 26.19/0.89 26.41/0.94 27.56/0.97 27.46/0.97
($\chi$=0.7) RGB 25.90/0.86 23.19/0.87 25.63/0.90 31.27/0.96 27.39/0.94 20.26/0.77 24.52/0.85 25.17/0.93 25.53/0.96 25.25/0.96
SLRQA-NF-2 QR 24.54/0.79 22.24/0.84 23.61/0.89 28.74/0.94 27.09/0.91 20.11/0.71 24.52/0.86 23.89/0.90 25.56/0.95 25.33/0.95
($\chi$=0.8) RGB 21.90/0.76 21.84/0.83 21.13/0.86 26.35/0.93 24.92/0.91 17.85/0.81 22.35/0.85 21.23/0.89 23.07/0.94 22.75/0.95
SLRQA-NF-3 QR 35.91/0.96 31.41/0.97 34.46/0.98 41.86/0.99 39.00/0.99 32.80/0.97 32.52/0.97 34.61/1.00 35.77/0.99 34.63/0.99
($\chi$=0.3) RGB 34.66/0.93 30.83/0.96 33.66/0.97 40.19/0.99 38.63/0.99 28.79/0.95 30.33/0.95 33.91/0.98 34.11/0.99 33.61/0.99
SLRQA-NF-3 QR 32.02/0.93 28.21/0.94 30.52/0.96 37.69/0.99 34.80/0.98 27.97/0.92 29.33/0.94 30.92/0.97 32.14/0.98 31.60/0.98
($\chi$=0.5) RGB 31.23/0.90 27.59/0.93 29.77/0.93 36.42/0.98 34.28/0.97 25.46/0.90 27.37/0.91 30.33/0.96 30.91/0.98 30.63/0.98
SLRQA-NF-3 QR 28.40/0.88 25.05/0.89 26.87/0.93 33.37/0.97 30.95/0.96 24.00/0.83 26.77/0.89 27.41/0.94 28.56/0.97 28.46/0.97
($\chi$=0.7) RGB 25.59/0.72 22.53/0.80 23.23/0.76 29.27/0.90 28.42/0.89 22.10/0.80 24.86/0.85 24.72/0.87 25.85/0.78 25.53/0.94
SLRQA-NF-3 QR 26.08/0.84 23.36/0.86 25.11/0.91 30.86/0.95 28.70/0.93 21.83/0.76 25.43/0.86 25.58/0.92 26.74/0.96 26.75/0.96
($\chi$=0.8) RGB 22.60/0.56 20.84/0.70 21.24/0.67 26.72/0.81 25.12/0.79 21.83/0.74 25.43/0.82 22.60/0.80 23.97/0.61 22.89/0.91
The parameter $\gamma$ is set as 0.7, 1, and 0.7 respectively for SLRQA-NF-1, SLRQA-NF-2, and SLRQA-NF-3. The parameter $w$ in SLRQA-NF-3 is set as $w_{i,k}=10/(\sigma_{i}({\mathbf{\dot{X}}}_{k-1})+10^{-4})$.
It is shown in Table 3 that the PSNR and SSIM values of the SLRQA-NF with quaternion representation are higher than SLRQA-NF with RGB representation for all different missing rate $\chi$. Furthermore, the PSNR and SSIM values of different $\chi$ and different images are given in Figure 3 which shows that SLRQA-NF-1 with quaternion representation method outperforms SLRQA-NF-1 with RGB representation by a large margin. Moreover, the improvements in PSNR and SSIM increase with the rate of corruption.
Figure 3: PSNR and SSIM comparisons of quaternion and RGB based SLRQA-NF-1. In the legend, the suffix (Q) means that quaternion representation is used, and (RGB) means that RGB representation is used. The numbers 0.3, 0.5, and 0.7 denote the missing rate of the images respectively.
It is concluded that the quaternion-based model achieves better denoising and inpainting performance. Hence, it is preferable to use the quaternion model in color image processing.
6.3 Color image denoising
NSS is a significant technique in color image denoising problems. The comparison without NSS directly reflects the superiority of the compared models, and the comparison with NSS better reflects the robustness of the models in practice. To fully compare the SLRQA with other state-of-the-art color image denoising methods, we tested their performance without and with the usage of non-local self-similarity (NSS) respectively in Section 6.3.1 and Section 6.3.2.
6.3.1 Color image denoising without NSS
The compared methods include SRRC[39], WNNM[14], LRQA[10], and DWT[5]333 The implementations that we use are available from:
SRRC: https://github.com/zt9877/SRRC.git
WNNM: http://www4.comp.polyu.edu.hk/ cslzhang/papers.htm
LRQA: https://www.fst.um.edu.mo/personal/wp-content/uploads/2021/05/LRQA.zip
. SRRC exploits the weighted Schatten $p$-norm as the regularizer for the rank residual to obtain a new rank minimization model. WNNM uses the weighted nuclear norm as an alternative function to the rank function. LRQA is a quaternion-based model that makes use of the low-rankness property by a nonconvex surrogate function. Compared with SLRQA, none of SRRC, WNNM, and LRQA exploits the sparsity feature of the image. DWT [5] is a vanilla analysis-based approach (5) where low-rankness prior is not used and is implemented on our own. It is noted that SRRC, WNNM, and DWT are RGB-based methods, and LRQA and SLRQA are quaternion-based models. LRQA-1 and LRQA-2 denote the LRQA models using Schatten-$\gamma$ and Laplace function, respectively. For the compared methods, we use the default parameters stated in the corresponding papers.
The PSNR and SSIM results of the compared methods are shown in Table 4. It can be seen that SLRQA has better results than the other methods in terms of PSNR and SSIM in most cases. Furthermore, The PSNR value of SRRC, WNNM, and DWT are much lower than that of SLRQA when $\tau$ is large. Note that since NSS is not utilized in these methods, the results in Table 4 underscore the effectiveness and robustness of the SLRQA model and further validate the advanced nature and adaptability of its intrinsic architecture. We also give the visual comparison between all competing methods and SLRQA as shown in Figure 4. It can be seen in the highlighted red rectangles that when compared with other denoising methods, SLRQA can preserve more details of images. In particular, in terms of hair details, the images denoised by the DWT and LRQA-1 methods are still noisy. The images recovered by LRQA and SRRC methods contain blurry parts and hence the detail features of the images are lost. In contrast, the images recovered by SLRQA models have better results both numerically and visually.
To further analyze the details of PL-ADMM, we plot the empirical convergence of PL-ADMM for Problem (55) on ten tested images with $\tau=30$. The results are shown in Figure 5. It can be observed that the error curve has a downward trend, illustrating the empirical convergence of Algorithm 1. We can observe that $\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}\rightarrow 0$. Hence, it follows from (42) that $|\|\widetilde{\mathbf{\dot{D}}}_{k+1}\||\rightarrow 0$. According to the update of $\mathbf{\dot{\Lambda}}_{k+1}$ and Lemma 4.7, it follows that $({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})$ converges to a stationary point of $\mathcal{L}_{\beta}$ and hence a KKT point of the original problem in (55).
Table 4: PSNR/SSIM results on color image denoising problem of different methods without NSS.
Image Algorithm SRRC[39] WNNM[14] DWT[5] LRQA-1[10] LRQA-2[10] SLRQA-1 SLRQA-2 SLRQA-3
$\tau=10$ Image1 29.23/0.70 28.06/0.73 27.67/0.61 28.48/0.61 29.21/0.65 31.55/0.76 31.45/0.78 32.86/0.81
Image2 28.56/0.73 28.01/0.70 27.98/0.71 28.44/0.72 28.42/0.72 30.36/0.80 30.19/0.80 31.20/0.82
Image3 28.03/0.60 27.81/0.61 28.21/0.61 28.49/0.64 28.56/0.65 31.33/0.77 31.79/0.78 32.03/0.84
Image4 28.12/0.60 27.90/0.65 28.29/0.66 28.53/0.64 28.55/0.68 30.09/0.74 31.45/0.75 30.36/0.75
Image5 28.82/0.66 28.72/0.64 28.67/0.64 28.85/0.65 28.71/0.67 32.13/0.85 31.18/0.80 32.50/0.87
Image6 28.32/0.60 27.89/0.56 28.02/0.61 28.72/0.65 28.66/0.65 30.75/0.76 30.92/0.80 30.77/0.76
Image7 28.45/0.62 27.32/0.58 28.02/0.59 28.65/0.60 29.04/0.94 30.95/0.73 30.82/0.75 31.10/0.80
Image8 28.01/0.60 28.11/0.59 28.46/0.58 28.56/0.63 28.24/ 0.63 31.23/0.76 30.75/0.74 31.92/0.76
Image9 28.02/0.74 27.98/0.73 27.62/0.71 28.82/0.76 29.10/0.76 30.14/0.70 29.66/0.78 30.80/0.79
Image10 30.52/0.74 29.72/0.72 29.62/0.70 30.90/0.76 31.10/0.79 31.14/0.78 31.26/0.80 31.32/0.81
$\tau=30$ Image1 20.32/0.30 20.41/0.25 20.36/0.28 20.49/0.27 20.80/0.30 25.47/0.66 24.30/0.61 26.40/0.67
Image2 20.01/0.32 20.73/0.32 20.09/0.33 20.44/0.37 20.40/0.37 23.90/0.61 22.77/0.64 24.02/0.64
Image3 00.12/0.21 20.24/0.22 20.55/0.25 20.50/0.27 21.10/0.31 26.15/0.65 25.21/0.64 26.22/0.68
Image4 20.22/0.29 20.10/0.26 20.01/0.29 20.68/0.29 20.20/0.31 23.81/0.59 22.74/0.57 24.21/0.64
Image5 21.02/0.32 20.58/0.33 20.65/0.37 21.23/0.30 21.30/0.33 26.60/0.68 25.32/0.58 26.34/0.68
Image6 20.87/0.33 21.01/0.37 20.37/0.37 20.76/0.36 20.61/0.35 23.95/0.66 21.72/0.64 22.50/0.63
Image7 20.33/0.29 20.71/0.29 20.94/0.22 20.93/0.26 20.21/0.21 25.12/0.63 24.39/0.68 26.42/0.68
Image8 20.56/0.22 20.42/0.24 21.08/0.28 20.77/0.29 20.90/0.29 25.14/0.64 24.06/0.62 26.12/0.67
Image9 24.82/0.58 24.22/0.58 23.62/0.55 24.90/0.58 25.10/0.61 25.30/0.64 25.45/0.68 25.80/0.68
Image10 24.32/0.64 23.82/0.62 23.62/0.61 24.34/0.66 24.90/0.68 24.97/0.66 25.02/0.70 25.68/0.70
$\tau=50$ Image1 17.92/0.17 17.01/0.14 18.01/0.18 17.66/0.17 17.30/0.18 23.33/0.50 23.50/0.51 23.90/0.51
Image2 16.02/0.21 17.03/0.20 17.94/0.25 17.53/0.25 17.10/0.26 22.04/0.48 22.77/0.52 22.06/0.45
Image3 17.75/0.15 17.40/0.21 17.24/0.19 17.56/0.17 17.90/0.18 24.21/0.51 24.91/0.58 24.99/0.60
Image4 17.08/0.14 16.60/0.16 16.87/0.17 17.82/0.19 17.20/0.14 22.23/0.46 22.34/0.47 23.51/0.49
Image5 18.04/0.17 18.08/0.17 18.31/0.18 18.39/0.19 18.30/0.17 21.84/0.52 21.62/0.67 22.72/0.65
Image6 17.53/0.25 18.01/0.26 17.11/0.22 17.75/0.26 18.61/0.29 21.24/0.50 21.02/0.51 22.00/0.53
Image7 17.60/0.17 18.21/0.18 17.11/0.13 18.08/0.17 18.71/0.17 23.57/0.48 23.09/0.45 23.92/0.53
Image8 17.02/0.14 17.42/0.18 17.62/0.19 17.90/0.18 18.10/0.29 23.14/0.48 23.26/0.52 24.32/0.56
Image9 22.02/0.51 21..92/0.49 21.62/0.46 22.34/0.52 22.67/0.55 22.78/0.56 23.02/0.62 23.04/0.63
Image10 21.82/0.50 21.42/0.48 21.32/0.45 22.32/0.51 22.10/0.51 22.74/0.54 22.60/0.52 23.10/0.58
(a) Image5
(b) Noisy Image
(c) SRRC
(d) WNNM
(e) DWT
(f) LRQA-1
(g) LRQA-2
(h) SLRQA-1
(i) SLRQA-2
(j) SLRQA-3
Figure 4: Color image denoising results of different methods on “Image5” with $\tau=10$.
Figure 5: The error convergence of SLRQA-1 for image denoising problems.
6.3.2 Color image denoising with NSS
For color image denoising problems with NSS, in addition to the methods compared in Section 6.3.1, we also compare SLRQA with CBM3D [11] [^4] The implementation is available from http://www.cs.tut.fi/~foi/GCF-BM3D.. Instead of constructing a denoising model, CBM3D employs block matching and sparse 3D collaborative filtering techniques to denoise the image, a process that cannot be achieved without using NSS. Hence we do not conduct a comparison with the CBM3D algorithm in the experiments of Section 6.3.1.
The NSS prior is based on the fact that for a given local patch in a natural image, one can find many similar patches across the whole image. This technique has been widely used for image processing problems, such as image denoising, image deblurring, and image repairing. Here, the NSS process in [10] is used. Specifically, a noisy image $\mathbf{\dot{Y}}$ is divided into $n$ overlapping patches of size $\sqrt{d}\times\sqrt{d}$. Each patch is transformed into a vector $\mathbf{\dot{y}}^{i}\in\mathbb{H}^{d},i=1,2,...,n$. Then for each patch, $s$ nearest neighbor patches are selected from a searching window with $L\times L$ pixels to form a set $K_{i}$. All patches in $K_{i}$ are stacked into a matrix $\mathbf{\dot{Y}}^{i}\in\mathbb{H}^{d\times s}$. We denoise $\mathbf{\dot{Y}}^{i}$ and aggregate all denoised patches together to form the clean color image $\mathbf{\dot{X}}$. Since all patches in each data matrix have similar structures, the constructed data matrix $\mathbf{\dot{Y}}^{i}$ is low rank. We also adopt the iterative regularization scheme in [10] and the iterative relaxation parameter is fixed to 0.1. For the parameter of NSS, we set patch size to 10$\times$10, 12$\times$12, and 14$\times$14 for $\tau$ = 10, 30, and 50, respectively. The number of selected non-local similar patches $n$ is set as 70, 80, and 90, respectively. The searching window $L$ is fixed at 30.
The parameter $\gamma$ is set as 0.3, 0.8, and 0.3 respectively for SLRQA-1, SLRQA-2, and SLRQA-3. The parameter $w$ in SLRQA-3 is set as $w_{i,k}=10/(\sigma_{i}({\mathbf{\dot{X}}}_{k-1})+10^{-4})$. The multiple values of $(\lambda,\tau)$ are used, i.e., $(\lambda,\tau)=(0.001,10),(0.01,30)$, and $(0.01,50)$. Intuitively, NSS enhances the low-rank property of the image. Hence, the parameter $\lambda$ is chosen to be smaller when compared to that without using NSS.
The PSNR and SSIM values of each image between the competing methods under different noise levels are reported in Table 5. Most proposed SLRQA methods achieve higher PSNR and SSIM values than other methods. Moreover, SLRQA-3 achieves the best results in most cases. The visual comparisons of Image5 with $\tau=30$ are shown in Figure 6 as a typical example. It is shown in Figure 6 that the images recovered by CBM3D and WNNM methods are still noisy. In addition, the images recovered by DWT, LRQA-1, and LRQA-2 are overly smooth, resulting in the loss of the features of the original images. In contrast, the image recovered by SLRQA-3 can not only remove the noise but also keep the features of the original images.
From the above numerical experiments, we can conclude that the SLRQA has better denoising performance compared to other methods with or without using NSS in the sense of PSNR and SSIM.
Table 5: PSNR/SSIM results on color image denoising problem of different methods with NSS.
Image Algorithm CBM3D[11] SRRC[39] WNNM[14] DWT[5] LRQA-1[10] LRQA-2[10] SLRQA-1 SLRQA-2 SLRQA-3
$\tau=10$ Image1 35.71/0.92 35.67/0.92 34.85/0.91 34.62/0.91 35.19/0.92 35.34/0.92 35.69/0.92 35.57/0.92 35.86/0.93
Image2 32.22/0.86 32.02/0.85 31.44/0.86 30.89/0.85 31.94/0.87 31.88/0.87 32.20/0.87 32.19/0.87 32.40/0.87
Image3 34.67/0.88 34.72/0.88 34.37/0.88 34.22/0.88 34.63/0.88 34.72/0.88 34.75/0.88 34.79/0.88 35.03/0.88
Image4 34.36/0.91 34.12/0.91 33.95/0.91 33.89/0.91 34.01/0.91 34.05/0.91 34.43/0.91 34.45/0.91 34.36/0.91
Image5 37.08/0.95 36.82/0.94 35.94/0.95 35.87/0.94 36.48/0.95 36.71/0.95 37.10/0.95 37.18/0.95 37.50/0.95
Image6 37.02/0.97 36.32/0.96 36.04/0.96 36.02/0.96 36.65/0.97 36.66/0.97 37.39/0.97 37.22/0.97 37.27/0.97
Image7 34.25/0.91 34.01/0.89 33.80/0.90 33.92/0.90 34.25/0.90 34.24/0.90 34.74/0.91 34.82/0.91 35.00/0.91
Image8 35.94/0.94 34.82/0.93 34.71/0.93 34.66/0.93 35.14/0.94 35.24/ 0.94 35.93/0.94 35.75/0.94 35.92/0.94
Image9 33.25/0.81 32.81/0.79 32.60/0.79 32.32/0.75 33.75/0.80 34.14/0.88 34.24/0.90 34.92/0.89 35.01/0.90
Image10 32.94/0.78 33.52/0.80 32.33/0.83 32.85/0.82 32.84/0.89 32.24/ 0.80 33.96/0.84 35.45/0.82 34.08/0.84
$\tau=30$ Image1 30.82/0.85 30.76/0.84 30.41/0.85 30.36/0.85 30.66/0.86 31.00/0.86 31.30/0.87 31.30/0.87 31.40/0.88
Image2 27.76/0.74 27.01/0.73 26.73/0.73 26.99/0.74 27.30/0.73 27.40/0.73 27.67/0.74 27.77/0.74 28.00/0.74
Image3 31.02/0.81 30.82/0.80 30.74/0.80 30.55/0.80 30.93/0.81 31.10/0.81 31.15/0.81 31.21/0.81 31.22/0.81
Image4 29.45/0.80 29.22/0.79 29.10/0.81 29.00/0.80 29.37/0.81 29.20/0.81 29.57/0.81 29.74/0.81 29.81/0.81
Image5 32.21/0.85 31.02/0.85 31.08/0.85 30.65/0.84 31.34/0.85 31.30/0.85 31.50/0.86 31.32/0.85 31.34/0.85
Image6 31.55/0.93 30.87/0.92 31.01/0.92 30.37/0.93 31.56/0.94 31.61/0.94 31.68/0.94 31.72/0.94 31.50/0.94
Image7 29.23/0.80 28.81/0.79 28.71/0.79 28.94/0.79 29.09/0.80 29.21/0.80 29.32/0.80 29.39/0.80 29.42/0.80
Image8 29.92/0.82 29.56/0.82 29.42/0.81 28.68/0.82 29.82/0.82 29.90/0.82 29.80/0.82 30.06/0.82 30.12/0.83
Image9 29.25/0.80 29.01/0.79 28.80/0.74 28.12/0.69 29.05/0.80 29.24/0.80 29.74/0.81 29.82/0.81 30.02/0.82
Image10 30.94/0.94 30.82/0.93 30.71/0.93 30.66/0.93 31.14/0.94 31.34/ 0.94 31.93/0.94 31.75/0.94 31.92/0.94
$\tau=50$ Image1 28.32/0.82 27.92/0.81 28.01/0.81 28.15/0.82 28.31/0.82 28.30/0.82 28.60/0.82 28.50/0.82 28.60/0.82
Image2 25.23/0.67 25.02/0.66 25.03/0.66 24.94/0.65 25.22/0.67 25.10/0.67 25.47/0.67 25.77/0.67 25.60/0.67
Image3 28.82/0.28 28.75/0.77 28.74/0.76 28.24/0.75 28.85/0.77 28.90/0.77 28.95/0.78 28.91/0.78 28.99/0.78
Image4 27.12/0.74 27.08/0.74 26.60/0.74 26.87/0.74 27.06/0.74 27.20/0.74 27.27/0.75 27.34/0.75 27.51/0.75
Image5 28.56/0.76 28.34/0.76 28.08/0.76 28.01/0.76 28.41/0.77 28.30/0.77 28.50/0.77 28.62/0.78 28.72/0.78
Image6 28.77/0.90 28.53/0.90 28.01/0.89 28.11/0.90 28.52/0.92 28.61/0.92 28.98/0.93 29.02/0.93 29.00/0.93
Image7 26.53/0.75 26.43/0.75 26.21/0.74 26.11/0.74 26.74/0.75 26.71/0.74 26.82/0.75 26.89/0.75 26.92/0.75
Image8 28.12/0.72 27.90/0.71 27.42/0.71 27.62/0.72 28.02/0.72 28.10/0.73 28.10/0.73 28.26/0.73 28.32/0.73
Image9 26.65/0.66 26.71/0.66 26.80/0.68 26.02/0.63 27.05/0.69 27.24/0.70 27.34/0.70 27.62/0.71 28.70/0.74
Image10 27.34/0.70 27.02/0.68 26.81/0.64 26.67/0.62 27.34/0.65 27.64/ 0.70 27.73/0.69 27.75/0.70 28.02/0.72
(a) Image5
(b) Noisy Image
(c) CBM3D
(d) SRRC
(e) WNNM
(f) DWT
(g) LRQA-1
(h) LRQA-2
(i) SLRQA-1
(j) SLRQA-2
(k) SLRQA-3
Figure 6: Color image denoising result of tested methods on “Image5” with $\tau=30$.
6.4 Color Image Inpainting
For color image inpainting problems, the compared methods include Bilinear Factorization (BF) methods such as LMafit [31], MCNMF [33], and Rank Minimization (RM) methods such as LADM [21], LRQA [10], and WNNM [13]555 The implementations that we use are available from:
LMafit: https://github.com/optsuite/LMaFit
MCNMF: https://xu-yangyang.github.io/codes/MCNF.zip
LRQA: https://www.fst.um.edu.mo/personal/wp-content/uploads/2021/05/LRQA.zip
WNNM: https://www4.comp.polyu.edu.hk/cslzhang/code/WNNM_MC_code.zip
. We note that the BF methods require a rank estimate as a prior. For fairness, the LADM is extended to quaternion representation, which is referred to as QLADM and performs better than the LADM method. To fully demonstrate the performance of LMaFit and MCNMF, we choose the best PSNR and SSIM from the initial rank range [50, 80, 100, 120, 150].5080100120150[50,\,80,\,100,\,120,\,150].[ 50 , 80 , 100 , 120 , 150 ] .
The results of all tested algorithms are shown in Table 6. It is shown therein that SLRQA-NF-3 always enjoys the best performance compared with other methods in terms of PSNR and SSIM. The visual comparisons between SLRQA-NF and all competing inpainting methods on the Image2 with $\chi=0.7$ are shown in Figure 7 as a typical example. Although the LMaFit and MCNMF methods complete the image, the details of the recovered image are lost, and the recovered images are blurry. The images recovered by LRQA-1 and LRQA-2 methods have a relatively good effect, but the images are still slightly noisy or blurry. In addition, the images recovered by WNNM are noisy. In contrast, SLRQA-NF methods all demonstrate considerable inpainting results and have the highest PSNR and SSIM values.
(a) Image2
(b) Corrupted image
(c) WNNM
(d)LMaFit
(e) MCNMF
(f) QLADM
(g) LRQA-1
(h) LRQA-2
(i) SLRQA-NF-1
(j) SLRQA-NF-2
(k) SLRQA-NF-3
Figure 7: Color image inpainting result of tested methods on “Image2” with $\chi=0.7$.
Table 6: PSNR/SSIM results on image inpainting problem of different methods.
Image Algorithm LMaFit[31] WNNM [13] MCNMF [33] QLADM[21] LRQA-1 [10] LRQA-2 [10] SLRQA-NF-1 SLRQA-NF-2 SLRQA-NF-3
$\chi=0.5$ Image1 24.69/0.74 25.25/0.60 29.35/0.83 31.16/0.95 30.87/0.89 30.25/0.87 32.32/0.94 31.26/0.94 32.47/0.94
Image2 22.14/0.80 23.55/0.81 24.22/0.86 30.89/0.85 26.63/0.91 26.20/0.90 28.04/0.94 27.54/0.94 28.16/0.94
Image3 23.43/0.80 25.23/0.75 27.05/0.88 29.65/0.97 28.37/0.90 27.94/0.89 30.40/0.96 29.65/0.97 30.60/0.97
Image4 29.09/0.91 31.16/0.85 34.73/0.96 36.39/0.99 35.99/0.97 35.51/0.97 37.69/0.99 36.27/0.98 37.90/0.99
Image5 26.79/0.85 30.09/0.87 30.02/0.92 33.73/0.98 33.27/0.96 32.72/0.95 34.82/0.98 33.75/0.98 34.98/0.98
Image6 19.64/0.61 22.01/0.61 22.04/0.87 26.56/0.92 25.27/0.79 25.72/0.80 28.25/0.93 26.62/0.91 27.97/0.92
Image7 25.00/0.81 24.96/0.75 25.14/0.82 28.84/0.94 27.59/0.87 27.86/0.88 27.82/0.93 28.92/0.94 29.33/0.94
Image8 24.18/0.87 25.17/0.86 27.21/0.92 30.03/0.97 29.62/0.95 29.06/0.94 31.06/0.97 30.16/0.97 31.22/0.97
Image9 24.93/0.94 26.02/0.94 27.04/0.95 30.99/0.98 30.12/0.98 29.56/0.97 31.58/0.98 31.09/0.98 31.71/0.98
Image10 24.36/0.93 26.18/0.94 23.73/0.92 30.69/0.98 29.86/0.98 29.20/0.97 31.28/0.98 30.84/0.98 31.43/0.98
$\chi=0.7$ Image1 22.32/0.66 23.05/0.52 24.40/0.67 27.29/0.89 25.97/0.74 25.47/0.72 28.40/0.88 27.30/0.88 28.62/0.89
Image2 20.19/0.73 20.13/0.67 21.52/0.77 24.19/0.89 22.51/0.79 22.44/0.79 25.05/0.90 24.30/0.89 25.21/0.90
Image3 21.20/0.72 21.50/0.63 22.78/0.75 26.09/0.93 23.69/0.77 23.67/0.77 26.87/0.93 26.00/0.92 27.10/0.94
Image4 25.87/0.85 28.34/0.80 28.50/0.89 31.89/0.97 31.07/0.93 30.44/0.92 33.37/0.97 31.80/0.97 33.63/0.98
Image5 24.25/0.78 25.75/0.77 26.19/0.83 29.58/0.95 28.53/0.88 27.85/0.86 30.95/0.96 29.68/0.95 31.14/0.96
Image6 17.49/0.52 18.13/0.44 19.66/0.67 22.78/0.82 20.28/0.57 19.86/0.55 23.78/0.81 22.70/0.80 24.00/0.83
Image7 22.59/0.74 22.40/0.67 21.17/0.82 26.24/0.90 24.62/0.78 24.58/0.78 26.63/0.88 26.19/0.89 26.57/0.88
Image8 21.79/0.81 22.34/0.79 23.17/0.84 26.34/0.94 24.87/0.87 24.54/0.87 27.41/0.94 26.42/0.94 27.60/0.95
Image9 22.66/0.91 23.33/0.90 24.77/0.93 27.83/0.97 26.13/0.94 25.69/0.94 28.56/0.97 27.81/0.97 28.74/0.97
Image10 21.55/0.89 23.22/0.90 24.18/0.92 27.54/0.96 25.95/0.95 25.34/0.94 28.46/0.97 27.58/0.96 28.64/0.97
$\chi=0.8$ Image1 21.13/0.59 20.77/0.43 19.95/0.44 25.05/0.85 22.87/0.61 22.76/0.60 26.08/0.84 24.54/0.79 26.30/0.85
Image2 19.32/0.68 17.72/0.54 17.87/0.58 22.34/0.84 19.44/0.65 20.27/0.69 23.36/0.86 22.25/0.84 23.53/0.86
Image3 20.26/0.65 18.71/0.53 18.93/0.59 23.98/0.90 20.52/0.65 21.35/0.69 24.91/0.90 23.61/0.88 25.13/0.91
Image4 24.62/0.79 25.48/0.72 23.53/0.75 29.09/0.95 27.75/0.87 27.20/0.85 30.74/0.96 28.73/0.94 31.04/0.96
Image5 23.37/0.74 22.52/0.65 21.25/0.66 27.10/0.93 25.53/0.80 25.05/0.78 28.70/0.93 27.09/0.92 28.97/0.94
Image6 16.86/0.46 15.41/0.30 15.94/0.34 20.82/0.74 17.93/0.45 14.65/0.31 21.61/0.74 20.10/0.70 21.82/0.76
Image7 21.53/0.69 20.67/0.60 19.35/0.57 24.79/0.87 22.81/0.71 22.64/0.71 25.59/0.85 24.52/0.86 25.43/0.86
Image8 20.53/0.77 19.88/0.71 19.22/0.71 24.26/0.91 21.69/0.79 21.93/0.80 25.30/0.92 23.90/0.90 25.50/0.92
Image9 21.93/0.90 20.78/0.84 21.23/0.85 25.89/0.95 23.43/0.91 23.24/0.91 26.74/0.96 25.56/0.95 26.98/0.96
Image10 20.47/0.87 20.48/0.84 20.34/0.85 25.52/0.95 23.10/0.90 22.85/0.90 26.67/0.96 25.32/0.95 26.90/0.96
7 Conclusion
In this paper, we propose a novel SLRQA model for the color image process problems. Different from most existing models which only consider one or two properties from low-rankness, sparsity, and quaternion representation, SLRQA utilizes these properties all. Furthermore, SLRQA does not need an initial rank esamite. A PL-ADMM algorithm is proposed to solve the SLRQA model. Furthermore, we prove that any limit point of PL-ADMM for SLRQA is a stationary point under mild assumptions. In addition, the global convergence analysis of the sequence generated by PL-ADMM under the KL-inequality assumption is presented. When the observation is noise-free, a SLRQA-NF model of the limiting case of SLRQA model is proposed. Subsequently, a nonconvex PL-ADMM-NF algorithm is also proposed to solve the SLRQA-NF model. Extensive experiments for color image denoising and inpainting problems have demonstrated the robustness and effectiveness of SLRQA and SLRQA-NF.

### v2

@@
 In this paper, we propose a Sparse Low-rank Quaternion Approximation (SLRQA) model for color image processing problems with noisy observations.
-The proposed SLRQA is a quaternion model that combines low-rankness and sparsity priors.
-Furthermore, it does not need an initial rank estimate.
-To solve the SLRQA model, a nonconvex proximal linearized ADMM (PL-ADMM) algorithm is proposed.
-Furthermore, the global convergence analysis of the PL-ADMM under mild assumptions is presented.
-When the observation is noise-free, an SLRQA-NF model of the limiting case of the SLRQA is proposed.
-Subsequently, a nonconvex proximal linearized ADMM (PL-ADMM-NF) algorithm for the SLRQA-NF is given.
+The proposed SLRQA is a quaternion model that combines low-rankness and sparsity priors without an initial rank estimation.
+A proximal linearized ADMM (PL-ADMM) algorithm is proposed to solve SLRQA and the global convergence is guaranteed under standard assumptions.
+When the observation is noise-free, a limiting case of the SLRQA, called SLRQA-NF, is proposed.
+Subsequently, a proximal linearized ADMM (PL-ADMM-NF) algorithm for SLRQA-NF is given.
+Since SLRQA-NF does not satisfy a widely-used assumption for global convergence of ADMM-type algorithms, we propose a novel assumption, under which the global convergence of PL-ADMM-NF is established.
 In numerical experiments, we verify the effectiveness of quaternion representation.
-Furthermore, for color image denoising and color image inpainting problems, SLRQA and SLRQA-NF models demonstrate superior performance both quantitatively and visually when compared with some state-of-the-art methods.
+Furthermore, for color image denoising and color image inpainting problems, SLRQA and SLRQA-NF demonstrate superior performance both quantitatively and visually when compared with some state-of-the-art methods.
 Key words: Color image denoising, Quaternion matrix completion, Nonconvex linearized ADMM.
 1 Introduction
-Low-rankness plays an important role in image processing and has been used in image denoising [10], image inpainting [15], image deblurring [26], and image filtering [7].
+Low-rankness plays an important role in image processing and has been used in image denoising [12], image inpainting [16], image deblurring [31], and image filtering problems [9].
 A commonly encountered approach for exploiting the low-rankness is to find a matrix with the lowest rank such that certain matching errors are minimized.
@@
 To overcome this difficulty, Candès et al.
-[6] have proved that the nuclear norm is the tightest convex relaxation of the NP-hard rank minimization function.
+have proved that the nuclear norm is the tightest convex relaxation of the NP-hard rank minimization function [8].
 In addition to the nuclear norm, other surrogates have also been proposed, such as weighted nuclear norm (WNNM)[15], Schatten $p$-norm [29], weighted Schatten $p$-norm [38], and log-determinant penalty [21], which have shown competitive effectiveness in various applications.
@@
 A color image can be encoded as a pure quaternion matrix.
-As claimed in [10], the main advantage of quaternion representation is that a color image is treated as a pure quaternion matrix naturally and holistically, meaning that the correlation information among the three RGB channels is used.
-It is empirically shown in [10] that methods using quaternion representation outperform methods with monochromatic techniques.
-Moreover, quaternion has been increasingly used for color image processing in the past two decades [10].
-These three priors, i.e.
+The main advantage of quaternion representation is that the three color channels simultaneously are treated by quaternion representation, which can well keep the connection between image color channels [39].
+Moreover, compared to tensor-based methods, quaternion-based methods not only use the information among channels but also preserve the orthogonality of the coefficient matrix for the three channels, see detailed discussions under the framework of dictionary learning in [40, Section III].
+Furthermore, it is empirically shown in [40, 12] that quaternion representation consistently provides a more accurate approximation than monochromatic representation using the reconstructed matrix of the same rank.
+The three priors, i.e.
 low-rankness, sparsity, and quaternion representation, all play important roles in image processing.
@@
 It follows that $E(\mathbf{\dot{Y}};{\mathbf{\dot{X}}})=\frac{1}{2\tau^{2}}\|\mathcal{A}({\mathbf{\dot{X}}})-\mathbf{\dot{Y}}\|_{\mathrm{F}}^{2}$ if all the entries of $\mathbf{\dot{Y}}$ are observed, and $E(\mathbf{\dot{Y}};{\mathbf{\dot{X}}})=\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2}$ if only partial entries of $\mathbf{\dot{Y}}$ are observed, where $\Omega$ denotes the indices of the observed entries, that is, $(\mathcal{P}_{\Omega}(\mathbf{\dot{Y}}))_{ij}=\mathbf{\dot{Y}}_{ij}$ if $(i,j)\in\Omega$ and $(\mathcal{P}_{\Omega}(\mathbf{\dot{Y}}))_{ij}=0$ if $(i,j)\notin\Omega$.
-By integrating the low-rankness and sparsity into the prior term, i.e., $R({\mathbf{\dot{X}}})=\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)+\lambda p(\mathcal{W}({\mathbf{\dot{X}}}))$, and letting $\mathbf{\dot{W}}=\mathcal{W}({\mathbf{\dot{X}}})$, we propose the following Sparse Low-rank Quaternion Approximation (SLRQA) model for color image processing:
-$\min_{{\mathbf{\dot{X}}},\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}}$ $\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)+\lambda p(\mathbf{\dot{W}})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2},$ (2)
+By integrating the low-rankness and sparsity into the prior term, i.e., $R({\mathbf{\dot{X}}})=\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma)+\lambda p(\mathcal{W}({\mathbf{\dot{X}}}))$, and letting $\mathbf{\dot{W}}=\mathcal{W}({\mathbf{\dot{X}}})$, we propose the following Sparse Low-rank Quaternion Approximation (SLRQA) model for color image processing:
+$\min_{{\mathbf{\dot{X}}},\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}}$ $\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma)+\lambda p(\mathbf{\dot{W}})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2},$ (2)
 $\mathrm{s.t.}$ $\mathcal{W}({\mathbf{\dot{X}}})=\mathbf{\dot{W}},$
-where ${\mathbf{\dot{X}}}\in\mathbb{H}^{m\times n}$ denotes a color image, $\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}$ is the coefficient matrix under an orthogonal transform $\mathcal{W}$ such as quaternion wavelet transform or quaternion discrete cosine transform (QDCT), $\mathcal{A}$ is a known operator such as a total variation operator, a blurring operator or the Radon transform, $p(\mathbf{\dot{W}})$ is a function that has Lipschitz continuously differentiable and promotes the sparsity of $\mathbf{\dot{W}}$, $\phi$ is a nonconvex surrogate function that substitutes the rank function, $\sigma_{i}({\mathbf{\dot{X}}})$ is the $i$-th singular value of ${\mathbf{\dot{X}}}$ defined in Section 2, $\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)$ promotes the low-rankness of ${\mathbf{\dot{X}}}$.
+where ${\mathbf{\dot{X}}}\in\mathbb{H}^{m\times n}$ denotes a color image, $\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}$ is the coefficient matrix under an orthogonal transform $\mathcal{W}$ such as quaternion wavelet transform or quaternion discrete cosine transform (QDCT), $\mathcal{A}$ is a known operator such as a blurring operator or the Radon transform, $p(\mathbf{\dot{W}})$ is a function that promotes sparsity and has Lipschitz continuously differentiable such as a Huber function, i.e., $(p(\mathbf{\dot{W}}))_{ij}=\frac{1}{2\delta}|\mathbf{\dot{W}}_{ij}|^{2}$ if $|\mathbf{\dot{W}}_{ij}|<\delta$; and $(p(\mathbf{\dot{W}}))_{ij}=|\mathbf{\dot{W}}_{ij}|-\frac{1}{2}\delta$ otherwise, $\delta>0$ is a prescribed constant, $\varepsilon>0$ is a smoothing parameter, $\phi$ is a nonconvex surrogate function that substitutes the rank function, $\sigma_{i}({\mathbf{\dot{X}}})$ is the $i$-th singular value of ${\mathbf{\dot{X}}}$ defined in Section 2.
+The function $\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma)$ promotes the low-rankness of ${\mathbf{\dot{X}}}$.
+We note that if $\varepsilon=0$, then $\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma)$ reduces to $\sum_{i}\phi(\sigma_{i}({{\mathbf{\dot{X}}}}),\gamma)$.
+If we futher assumes that $\phi(x,\gamma)=x$, then $\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma)$ reduces to the nuclear norm.
 Some nonconvex surrogate functions for $\phi$ have been given in [12, 28] and listed in Table 1.
-Table 1: nonconvex surrogate functions $\phi(x,\gamma)$
+Table 1: Some surrogate functions $\phi(x,\gamma)$
 Nonconvex function Log-determinant Schatten-$\gamma$ Logarithm Laplace [11] Weighted Schatten-$\gamma$ [38] ETP
 $\phi(x,\gamma)$ log$(1+x^{2})$ $x^{\gamma}$ log($\gamma$+$x$) $1-e^{-{\frac{x}{\gamma}}}$ $wx^{\gamma}$ $\frac{\lambda(1-\exp(-\gamma x))}{1-\exp(-\gamma)}$
-The $\ell_{1}$ norm has been widely used to promote sparsity.
+The $\ell_{1}$ norm and nuclear norm have been widely used to promote sparsity and low rankness respectively.
 However, the Lipschitz continuous differentiability of $p(\mathbf{\dot{W}})$ is assumed in our convergence analysis [^2] We use a variant of the ADMM method to solve SLRQA.
-Note that such an assumption has been made for all the ADMM methods with global convergence guaranteed for nonconvex problems as far as we know, see, e.g., a review in [22].
+Note that such an assumption has been made for all the ADMM methods with global convergence guaranteed for nonconvex problems as far as we know, see, e.g., a review in [27]..
 Therefore, a smoothing function such as the Huber function can be used as a substitution for $p(\mathbf{\dot{W}})$.
+For more surrogate function of $\|\cdot\|_{1}$, we refer the readers to [28].
+For the subsequent convergence analysis for Model (4), it is required that the term that promotes the low-rankness is also differentiable with respect to ${\mathbf{\dot{X}}}$.
+To satisfy this condition and ensure numerical stability, we also use a smoothed version of $\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)$ by adding a small positive perturbation $\varepsilon$, i.e., replacing $\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)$ with $\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma),$ as discussed in [46, 47].
 It is empirically shown in Section 6 that such a substitution yields satisfactory performance.
-This strategy has been used in [38] to guarantee convergence theoretically.
+This strategy has also been used in [46, 47] to guarantee convergence theoretically.
 When the observed color image is noise-free, or the variance $\tau^{2}$ in the noise $\bm{\varepsilon}$ is zero, Model (2) is undefined due to the existence of a zero in the denominator of the data fitting term.
@@
 It follows that a variant of SLRQA for the noise-free problem is given by
-$\min_{{\mathbf{\dot{X}}},\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}}$ $\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)+\lambda p(\mathbf{\dot{W}}),$ (3)
+$\min_{{\mathbf{\dot{X}}},\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}}$ $\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma)+\lambda p(\mathbf{\dot{W}}),$ (3)
 $\mathrm{s.t.}$ $\mathcal{W}({\mathbf{\dot{X}}})=\mathbf{\dot{W}},\quad\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}))=\mathcal{P}_{\Omega}(\mathbf{\dot{Y}}).$
-Since $\mathcal{W}$ is an orthogonal transform such as quaternion wavelet transform or quaternion discrete cosine transform (QDCT) [12], we have $\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)=\sum_{i}\phi(\sigma_{i}(\mathbf{\dot{Q}}_{1}\mathbf{\dot{W}}\mathbf{\dot{Q}}_{2}^{*}),\gamma)=\sum_{i}\phi(\sigma_{i}(\mathbf{\dot{W}}),\gamma)$, the superscript * denotes the conjugate transpose operator.
+Since $\mathcal{W}$ is an orthogonal transform such as quaternion wavelet transform or quaternion discrete cosine transform (QDCT) [14], we have $\sum_{i}\phi(\sigma_{i}(\sqrt{{\mathbf{\dot{X}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma)=\sum_{i}\phi(\sigma_{i}(\sqrt{\mathbf{\dot{W}}^{*}\mathbf{\dot{W}}+\varepsilon^{2}\mathbf{I}}),\gamma)$, where $\mathbf{\dot{Q}}_{1},\mathbf{\dot{Q}}_{2}$ are unitary matrix that defined later, the superscript * denotes the conjugate transpose operator.
 Let $\mathbf{\dot{Z}}=\mathcal{W}({\mathbf{\dot{X}}})$.
 We propose a model named SLRQA-NF which is equivalent to (3) for noise-free problems:
-$\min_{\mathbf{\dot{Z}},\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}}$ $\sum_{i}\phi(\sigma_{i}(\mathbf{\dot{Z}}),\gamma)+\lambda p(\mathbf{\dot{W}}),$ (4)
+$\min_{\mathbf{\dot{Z}},\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}}$ $\sum_{i}\phi(\sigma_{i}(\sqrt{{\mathbf{\dot{Z}}}^{*}\mathbf{\dot{Z}}+\varepsilon^{2}\mathbf{I}}),\gamma)+\lambda p(\mathbf{\dot{W}}),$ (4)
 $\mathrm{s.t.}$ $\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}))=\mathcal{P}_{\Omega}(\mathbf{\dot{Y}}),\quad\mathbf{\dot{Z}}=\mathbf{\dot{W}},$
@@
 1.2 Related work
-The derivations of SLRQA and SLRQA-NF models are motivated by LADM [21] and LRQA [10] models.
-To be specific, LADM is a convex model that utilizes the sparsity and low-rankness properties while quaternion representation and nonconvex surrogate rank function have not been considered.
+The derivations of SLRQA and SLRQA-NF are motivated by sparse low rank inpainting (SLRI) [26] and Low-Rank Quaternion Approximation (LRQA) [12] models.
+To be specific, SLRI is a convex model that uses the sparsity and low-rankness properties while quaternion representation and nonconvex surrogate rank function have not been considered.
 LRQA is a quaternion-based model that makes use of the low-rankness property by replacing the nuclear norm with nonconvex surrogate functions, but it has not taken advantage of sparsity.
-SLRQA and SLRQA-NF models take advantage of LADM and LRQA by using quaternion representation, low-rankness, and sparsity.
+SLRQA and SLRQA-NF take advantage of SLRI and LRQA by using quaternion representation, low-rankness, and sparsity.
 Next, we discuss some image processing models related to SLRQA and SLRQA-NF.
@@
 where the variable $\mathbf{\dot{W}}$ is eliminated by $\mathbf{\dot{W}}=\mathcal{W}({\mathbf{\dot{X}}})$.
-If the sparse term $p(\mathbf{\dot{W}})$ is not used, then SLRQA becomes the LRQA denoising model in [10, Equation 20]
+If the sparse term $p(\mathbf{\dot{W}})$ is not used, the low-rankness prior is used and $\varepsilon$ is set to 0, then SLRQA becomes LRQA denoising model in [12, Equation (20)]
 $\min_{{\mathbf{\dot{X}}}\in\mathbb{H}^{m\times n}}\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)+\frac{1}{2\tau^{2}}\|{\mathbf{\dot{X}}}-\mathbf{\dot{Y}}\|_{\mathrm{F}}^{2}.$
-As noted in [10], LRQA is a generalization of other models such as the WNNM model.
+As noted in [12], LRQA is a generalization of other models such as the WNNM [15].
 Image deblurring (noisy observation and blurring kernel operator $\mathcal{A}$):
@@
 $\min_{{\mathbf{\dot{X}}}\in\mathbb{H}^{m\times n}}\lambda\|\mathcal{W}({\mathbf{\dot{X}}})\|_{1}+\frac{1}{2\tau^{2}}\|\mathcal{A}({\mathbf{\dot{X}}})-\mathbf{\dot{Y}}\|_{\mathrm{F}}^{2}.$
-Image inpainting (partially observed images): If the observed image is noise-free, the linear operator $\mathcal{A}$ is the identity, $p(\mathbf{\dot{W}})=\|\mathbf{\dot{W}}\|_{1}$, and $\phi$ is the nuclear norm, then the SLRQA-NF model becomes the LADM model in [21] with an extension to quaternion representation:
+Image inpainting (partially observed images): If the observed image is noise-free, the linear operator $\mathcal{A}$ is the identity, $p(\mathbf{\dot{W}})=\|\mathbf{\dot{W}}\|_{1}$, $\phi$ is the nuclear norm and $\varepsilon=0$, then SLRQA-NF becomes SLRI in [26] with an extension to quaternion representation:
 $\min_{\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}}$ $\|\mathbf{\dot{W}}\|_{*}+\lambda\|\mathbf{\dot{W}}\|_{1},\quad\mathrm{s.t.}\,\mathcal{P}_{\Omega}(\mathcal{W}^{\#}(\mathcal{\mathbf{\dot{W}}}))=\mathcal{P}_{\Omega}(\mathbf{\dot{Y}}).$ (6)
-Furthermore, if the sparsity term $\phi$ is dropped and a nonconvex surrogate is used, then SLRQA-NF model becomes the LRQA [10, Equation 22] given by
+Furthermore, if the sparsity term is dropped, a nonconvex surrogate $\phi$ is used and $\varepsilon$ is set to 0, then SLRQA-NF becomes LRQA [12, Equation (22)] given by
 $\min_{{\mathbf{\dot{X}}}\in\mathbb{H}^{m\times n}}$ $\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma),\quad\mathrm{s.t.}\,\mathcal{P}_{\Omega}(\mathcal{\mathcal{{\mathbf{\dot{X}}}}})=\mathcal{P}_{\Omega}(\mathbf{\dot{Y}}).$ (7)
@@
 $\mathrm{s.t.}$ $\mathbf{\dot{L}}^{*}\mathbf{\dot{L}}=\mathbf{I}_{r},\,\mathbf{\dot{R}}\mathbf{\dot{R}}^{*}=\mathbf{I}_{r},~{}~{}\mathcal{W}({\mathbf{\dot{X}}})=\mathbf{\dot{W}},~{}~{}{\mathbf{\dot{X}}}=\mathbf{\dot{L}}\mathbf{\dot{D}}\mathbf{\dot{R}},~{}~{}\mathcal{P}_{\Omega}(\mathbf{\dot{Y}}-\mathbf{\dot{L}}\mathbf{\dot{D}}\mathbf{\dot{R}})=\mathbf{0}.$
-where $r$ is a prescribed number and the superscript * denotes the conjugate transpose operator, see its definition in Section 2.
-Different from SLRQA models, Model (8) uses the nuclear norm to promote low-rank instead of nonconvex surrogate functions $\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)$.
+where $r$ is a prescribed integer and the superscript * denotes the conjugate transpose operator, see its definition in Section 2.
+Different from SLRQAs, Model (8) uses the nuclear norm to promote low-rank instead of nonconvex surrogate functions $\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)$.
 Furthermore, it needs to estimate the rank $r$ in the initialization phase while SLRQA does not.
@@
 The model is given by
-$\min_{{\mathbf{\dot{X}}}}$ $\|{\mathbf{\dot{X}}}\|_{*}-\sum_{i=r}^{\min(m,n)}\sigma_{i}({\mathbf{\dot{X}}})+\lambda\|\mathbf{\dot{W}}\|_{1},$ (9)
+$\min_{{\mathbf{\dot{X}}}\in\mathbb{R}^{m\times n}}$ $\|{\mathbf{\dot{X}}}\|_{*}-\sum_{i=r}^{\min(m,n)}\sigma_{i}({\mathbf{\dot{X}}})+\lambda\|\mathbf{\dot{W}}\|_{1},$ (9)
 $\mathrm{s.t.}$ $\mathcal{P}_{\Omega}({\mathbf{\dot{X}}}-\mathbf{\dot{Y}})=0,\,\mathcal{W}({\mathbf{\dot{X}}})=\mathbf{\dot{W}}.$
-Note that Model (9) also needs to estimate the rank $r$ and the low-rankness is promoted by penalizing a truncated nuclear norm, which is convex.
-Moreover, convergence analyses of the proposed algorithms for solving the existing models in [17, 34, 35] with three priors are not given.
-In this paper, proximal linearized ADMM algorithms are proposed to solve Models (2) and (4).
-Note that the convergence of nonconvex ADMM and its variants has been studied widely.
-However, as far as we know, the existing ADMM and its variants either consider different optimization problems or use different optimization techniques.
-A nonexhaustive review is given as follows.
-Wang et al.
-[29] comprehensively study the convergence of ADMM for nonconvex and nonsmooth objective functions subject to coupled linear equality constraints.
-Their assumptions require that the solution of the subproblem is Lipschitz continuous with respect to the input, which is often overly stringent and hard to verify for practical applications.
-Hong et al.
-[18] study the convergence of ADMM for solving nonconvex consensus problems, but it is under the case where every constraint matrix has full column rank.
-Compared with the above analyses, our convergence analysis only needs one constraint matrix that has full column rank.
-In [36], Yashtini proposes a proximal linearized ADMM for solving linearly constrained nonconvex and possibly nonsmooth optimization problems.
-The proposed method employs a linearization strategy that targets both the quadratic penalty term and the objective function.
-Above all, all the nonconvex ADMM and their variant algorithms for nonconvex situations are different from PL-ADMM where only one variable is linearized in the situation.
-For the reviewed algorithms, the first two algorithms are traditional ADMM algorithms for nonconvex problems, while the ADMM of the last algorithm uses the technique of linearization for both two variables.
-For a detailed review of other non-convex ADMM algorithms, we refer the readers to [22].
+Note that Model (9) also needs to estimate the rank $r$ and the low-rankness is promoted by penalizing a truncated nuclear norm, which is nonconvex.
+Moreover, convergence analysis of the proposed algorithms for solving the existing models in [18, 41, 42] with three priors are not given.
+In this paper, proximal linearized ADMM algorithms, i.e., PL-ADMM and PL-ADMM-NF are proposed to solve Models (2) and (4) and the global convergence of these algorithms is established.
+A generalization of SLRQA is formulated as
+$\min_{{\mathbf{\dot{X}}},\mathbf{\dot{W}}}$ $f({\mathbf{\dot{X}}})+g(\mathbf{\dot{W}})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2},$ (10)
+$\mathrm{s.t.}$ $\mathcal{C}({\mathbf{\dot{X}}})+\mathcal{B}(\mathbf{\dot{W}})=\mathbf{\dot{B}},$
+where $f$ is semi-lowercontinuous, $g$ is continuously differentiable, $\mathcal{C},\mathcal{B}:\mathbb{H}^{m\times n}\rightarrow\mathbb{H}^{p\times q}$ are the constraint matrices and $\mathbf{\dot{B}}\in\mathbb{H}^{p\times q}$.
+Specifically, if $f({\mathbf{\dot{X}}})=\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma),g(\mathbf{\dot{W}})=p(\mathbf{\dot{W}})$, $\mathcal{B}=-\mathcal{I}$, $\mathcal{C}=\mathcal{W}$, and $\mathbf{\dot{B}}=0$, then Problem (21) reduces to the SLRQA in (2), where $\mathcal{I}$ denotes the identity operator.
+The convergence of nonconvex ADMM and its variants has been extensively studied in literatures and the global convergence of PL-ADMM is established.
+However, SLRQA-NF does not satisfy a widely-used assumption in existing ADMM-type algorithms. Specifically, SLRQA-NF can be formulated as
+$\min_{{\mathbf{\dot{X}}},\mathbf{\dot{W}}}$ $f({\mathbf{\dot{X}}})+g(\mathbf{\dot{W}}),$ (11)
+$\mathrm{s.t.}$ $\mathcal{C}_{1}({\mathbf{\dot{X}}})+\mathcal{B}_{1}(\mathbf{\dot{W}})=\mathbf{\dot{B}}_{1},$
+$\qquad\quad~{}\mathcal{B}_{2}(\mathbf{\dot{W}})=\mathbf{\dot{B}}_{2},$
+where $f$ and $g$ are both continuously differentiable, $\mathcal{B}_{1},\mathcal{C}_{1}:\mathbb{H}^{m\times n}\rightarrow\mathbb{H}^{p_{1}\times q_{1}}$, $\mathcal{B}_{2}:\mathbb{R}^{m\times n}\rightarrow\mathbb{R}^{p_{2}\times q_{2}}$ are given linear maps, $\mathbf{\dot{B}}_{1}\in\mathbb{H}^{p_{1}\times q_{1}}$, and $\mathbf{\dot{B}}_{2}\in\mathbb{H}^{p_{2}\times q_{2}}$.
+Specifically, if $f({\mathbf{\dot{X}}})=\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma),$ $g(\mathbf{\dot{W}})=p(\mathbf{\dot{W}})$, $\mathcal{B}_{1}=-\mathcal{I}$, $\mathcal{C}_{1}=\mathcal{I}$, $\mathcal{B}_{2}=\mathcal{P}_{\Omega}\mathcal{A}\mathcal{W}^{\#}$, $\mathbf{\dot{B}}_{1}=0$, and $\mathbf{\dot{B}}_{2}=\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})$, then Problem (11) reduces to the SLRQA-NF in (4).
+Note that the linear constraints in Problem (11) can be viewed as those in (21) if the two constraints are considered as one, that is $\mathcal{C}=\begin{pmatrix}\mathcal{C}_{1}\\ 0\end{pmatrix}$, $\mathcal{B}=\begin{pmatrix}\mathcal{B}_{1}\\ \mathcal{B}_{2}\end{pmatrix}$, and $\mathbf{\dot{B}}=\begin{pmatrix}\mathbf{\dot{B}}_{1}\\ \mathbf{\dot{B}}_{2}\end{pmatrix}$.
+In the convergence analysis of existing ADMM-type algorithms for problems with constraints as in (21), the assumption
+[…]
+which we refer to as “range assumption”, are required [35, 43, 27, 19, 24].
+The authors in [35] show the tightness of the “range assumption” in the sense that there exists a problem such that the “range assumption” does not hold while the ADMM algorithm for this problem diverges.
+This tightness does not prevent a weaker version of the “range assumption” that guarantees convergence of ADMM.
+Since the constraints in (11) do not satisfy the “range assumption”, we propose an alternative weaker assumption that guarantees global convergence of an ADMM-type algorithm.
 1.3 Contribution
 The main contributions of this paper are listed as follows:
-  • We propose a novel SLRQA model to address image processing problems with noise by maximizing the a posteriori probability estimate.
-The proposed SLRQA takes advantage of quaternion representation, low-rankness, and sparsity without requiring an estimation of the rank.
-  • We give a proximal linearized alternating direction method of multipliers (PL-ADMM) algorithm.
-Under mild conditions, we prove that the sequence generated by the algorithm converges to a stationary point.
-To the best of our knowledge, this is the first convergence analysis result of an algorithm for a quaternion-based image processing model.
-  • For noise-free observations, we propose a variant of SLRQA called SLRQA-NF. Inspired by PL-ADMM, an algorithm named PL-ADMM-NF is presented to solve the SLRQA-NF model.
+  • We propose a novel model called SLRQA to address image processing problems with noise by maximizing the a posteriori probability estimate.
+Unlike the existing image processing models, the proposed SLRQA (2) is a quaternion model combining low rank and sparsity, providing flexibility to choose surrogate functions and not needing an initial rank estimate.
+We propose a PL-ADMM algorithm to solve SLRQA and its global convergence is guaranteed.
+  • When the observations are noise-free, we propose a model named SLRQA-NF (4) and a corresponding algorithm named PL-ADMM-NF.
+The global convergence of PL-ADMM-NF is established under a newly proposed assumption.
+To the best of our knowledge, this is the first nonconvex ADMM-type algorithm without the “range assumption” in (12) and still guarantees global convergence.
   • In the numerical experiments, we verify that the quaternion representation outperforms the RGB monochromatic representation.
@@
 In Section 2, the preliminaries and notations are introduced.
-The PL-ADMM algorithm framework for solving SLRQA model is given in Section 3.
-The global convergence is established in Section 4.
-In Section 5, the PL-ADMM-NF algorithm for solving SLRQA-NF model is presented.
+The PL-ADMM algorithm for solving SLRQA and PL-ADMM-NF algorithm for solving SLRQA-NF are presented in Section 3.
+The global convergence of these two algorithms are established in Section 4.
 In Section 5, extensive numerical experiments are conducted to demonstrate the performance of the models and algorithms by comparing them with other state-of-the-art methods.
@@
 2 Preliminaries
-Throughout this paper, scalars, real vectors, real matrices, and real tensor are respectively denoted by lowercase letters $x$, boldface lowercase letters $\mathbf{x}$, boldface capital letters $\mathbf{X}$, and boldface capital letters with a tilde $\widetilde{\mathbf{X}}$ respectively.
+Throughout this paper, scalars, real vectors, real matrices, and real tensors are respectively denoted by lowercase letters $x$, boldface lowercase letters $\mathbf{x}$, boldface capital letters $\mathbf{X}$, and boldface capital letters with a tilde $\widetilde{\mathbf{X}}$ respectively.
 $\mathbb{H}$ denotes the quaternion algebra proposed in [17].
 A dot above a variable $\dot{q}$ represents a quaternion number.
-A quaternion number consists of one real part and three imaginary parts:
-$\dot{q}=x_{s}+x_{i}\mathbf{i}+x_{j}\mathbf{j}+x_{k}\mathbf{k}.$
-where $x_{s},x_{i},x_{j},x_{k}\in\mathbb{R}$, and $\mathbf{i},\mathbf{j},\mathbf{k}$ are three imaginary units.
+A quaternion number consists of one real part and three imaginary parts: $\dot{q}=x_{s}+x_{i}\mathbf{i}+x_{j}\mathbf{j}+x_{k}\mathbf{k}.$ where $x_{s},x_{i},x_{j},x_{k}\in\mathbb{R}$, and $\mathbf{i},\mathbf{j},\mathbf{k}$ are three imaginary units.
 The quaternion $\dot{q}$ is called a pure quaternion if and only if it has a zero real part, $\mathbf{\;i.e.\;}x_{s}=0$.
@@
 For $\widetilde{\mathbf{\dot{G}}}:=(\mathbf{\dot{G}}_{1},\cdots,\mathbf{\dot{G}}_{p})\in\mathbb{H}^{n_{1}\times m_{1}}\times\cdots\mathbb{H}^{n_{p}\times m_{p}},$ $|||\widetilde{\mathbf{\dot{G}}}\||_{\mathrm{F}}:=\sqrt{\sum_{i=1}^{p}\|\mathbf{\dot{G}}_{i}\|_{\mathrm{F}}^{2}}$.
-Hence we have $\sum_{i=1}^{p}\frac{1}{\sqrt{p}}\|\mathbf{\dot{G}}_{i}\|_{\mathrm{F}}\leq\||\widetilde{\mathbf{\dot{G}}}\||_{\mathrm{F}}\leq\sum_{i=1}^{p}\|\mathbf{\dot{G}}_{i}\|_{\mathrm{F}},$ where the first equality follows from Cauchy-Schwarz inequality.
+Hence we have $\sum_{i=1}^{p}\frac{1}{\sqrt{p}}\|\mathbf{\dot{G}}_{i}\|_{\mathrm{F}}\leq\||\widetilde{\mathbf{\dot{G}}}\||_{\mathrm{F}}\leq\sum_{i=1}^{p}\|\mathbf{\dot{G}}_{i}\|_{\mathrm{F}},$ where the first equality follows from the Cauchy-Schwarz inequality.
 The quaternion singular value decomposition theorem will be used in the solving process of our algorithm. We state it as follows.
@@
 According to the QSVD, the rank of a quaternion matrix is defined as the number of nonzero singular values while the nuclear norm $\|\mathbf{\dot{A}}\|_{*}$ is defined as the sum of all nonzero singular values.
-3 An Algorithm for SLRQA model
+3 Algorithm framework
+In this section, two algorithms, i.e.
+PL-ADMM and PL-ADMM-NF, are proposed to solve Model (2) and (4) respectively.
+The details of PL-ADMM are clarified in Section 3.1 and the PL-ADMM-NF algorithm is presented in Section 3.2.
+3.1 An Algorithm for SLRQA
 The proposed proximal linearized alternating direction method of multipliers (PL-ADMM) is stated in Algorithm 1.
 Algorithm 1 Proixmal linearized-ADMM (PL-ADMM) for SLRQA in (2)
-1:A noisy observed image $\mathbf{\dot{Y}}$; an initial iterate ${\mathbf{\dot{X}}}_{0}:=\mathbf{\dot{Y}}$; two parameters $\mu\in(0,2)$ and $\beta>0$; a tolerance $\eta>0$; two sequences of regularization parameters $\{L_{1,k}\}$ and $\{L_{2,k}\}$ satisfying $\sup_{k\geq 0}{L_{i,k}}<\infty$ and $\inf_{k>0}{L_{i,k}}>0$ for $i=1,2$;
-2:A recovered image ${\mathbf{\dot{X}}}_{*}$;
+1:A noisy observed image $\mathbf{\dot{Y}}$, an initial iterate ${\mathbf{\dot{X}}}_{0}:=\mathbf{\dot{Y}}$, two parameters $\mu\in(0,2)$ and $\beta>0$, a tolerance $\eta>0$, two sequences of regularization parameters $\{L_{k,1}\}$ and $\{L_{k,2}\}$ satisfying $\sup_{k\geq 0}{L_{i,k}}<\infty$ and $\inf_{k>0}{L_{i,k}}>0$ for $i=1,2$.
+2:A recovered image ${\mathbf{\dot{X}}}_{*}$.
 3:Set $\mathbf{\dot{W}}_{0}=\bm{0}$ and $\mathbf{\dot{\Lambda}}_{0}=\bm{0}$;
@@
 5:      ${\mathbf{\dot{X}}}$-subproblem:
-${\mathbf{\dot{X}}}_{k+1}\in$ $\mathop{\mathrm{arg\,min}}_{{\mathbf{\dot{X}}}}\;\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)$
-$+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})-\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}))\|_{\mathrm{F}}^{2}+\frac{\beta}{2}\|\mathcal{W}({\mathbf{\dot{X}}})-\mathbf{\dot{W}}_{k}+\frac{\mathbf{\dot{\Lambda}}_{k}}{\beta}\|_{\mathrm{F}}^{2}+\frac{L_{1,k}}{2}\|{\mathbf{\dot{X}}}-{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}^{2};$
+${\mathbf{\dot{X}}}_{k+1}\in$ $\mathop{\mathrm{arg\,min}}_{{\mathbf{\dot{X}}}}\;\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma)$
+$+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})-\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}))\|_{\mathrm{F}}^{2}+\frac{\beta}{2}\|\mathcal{W}({\mathbf{\dot{X}}})-\mathbf{\dot{W}}_{k}+\frac{\mathbf{\dot{\Lambda}}_{k}}{\beta}\|_{\mathrm{F}}^{2}+\frac{L_{k,1}}{2}\|{\mathbf{\dot{X}}}-{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}^{2};$
 6:      $\mathbf{\dot{W}}$-subproblem:
-$\mathbf{\dot{W}}_{k+1}=\mathop{\mathrm{arg\,min}}_{\mathbf{\dot{W}}}$ $\;\lambda\left\langle\nabla p(\mathbf{\dot{W}}_{k}),\mathbf{\dot{W}}\right\rangle+\frac{\beta}{2}\|\mathbf{\dot{W}}-(\mathcal{W}({\mathbf{\dot{X}}}_{k+1})+\frac{\mathbf{\dot{\Lambda}}_{k}}{\beta})\|_{\mathrm{F}}^{2}+\frac{L_{2,k}}{2}\|\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}^{2};$
+$\mathbf{\dot{W}}_{k+1}=\mathop{\mathrm{arg\,min}}_{\mathbf{\dot{W}}}$ $\;\lambda\left\langle\nabla p(\mathbf{\dot{W}}_{k}),\mathbf{\dot{W}}\right\rangle+\frac{\beta}{2}\|\mathbf{\dot{W}}-(\mathcal{W}({\mathbf{\dot{X}}}_{k+1})+\frac{\mathbf{\dot{\Lambda}}_{k}}{\beta})\|_{\mathrm{F}}^{2}+\frac{L_{k,2}}{2}\|\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}^{2};$
 7:      Update Lagrange multipliers: $\mathbf{\dot{\Lambda}}_{k+1}=\mathbf{\dot{\Lambda}}_{k}+\mu\beta(\mathcal{W}({\mathbf{\dot{X}}}_{k+1})-\mathbf{\dot{W}}_{k+1});$
@@
 Each iteration of PL-ADMM minimizes the augmented Lagrangian function by using the alternating direction method. Specifically, the augmented Lagrangian function of SLRQA in (2) is given by
-$\mathcal{L}_{\beta}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})=\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)+$ $\lambda p(\mathbf{\dot{W}})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})-\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}))\|_{\mathrm{F}}^{2}$
+$\mathcal{L}_{\beta}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})=\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma)+$ $\lambda p(\mathbf{\dot{W}})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})-\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}))\|_{\mathrm{F}}^{2}$
 $+\left\langle\mathbf{\dot{\Lambda}},\mathcal{W}({\mathbf{\dot{X}}})-\mathbf{\dot{W}}\right\rangle+\frac{\beta}{2}\|\mathcal{W}({\mathbf{\dot{X}}})-\mathbf{\dot{W}}\|^{2}_{\mathrm{F}},$
 where $\mathbf{\dot{\Lambda}}\in\mathbb{H}^{m\times n}$ denotes the Lagrange multipliers and $\beta>0$.
-In Step 5, ${\mathbf{\dot{X}}}_{k+1}$ is obtained by minimizing $\mathcal{L}_{\beta}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})+\frac{L_{1,k}}{2}\|{\mathbf{\dot{X}}}-{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}^{2}$ for a fixed $\mathbf{\dot{W}}_{k}$ and $\mathbf{\dot{\Lambda}}_{k}$.
+In Step 5, ${\mathbf{\dot{X}}}_{k+1}$ is obtained by minimizing $\mathcal{L}_{\beta}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})+\frac{L_{k,1}}{2}\|{\mathbf{\dot{X}}}-{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}^{2}$ for a fixed $\mathbf{\dot{W}}_{k}$ and $\mathbf{\dot{\Lambda}}_{k}$.
 This subproblem does not yield a closed-form solution in general and therefore may be solved by nonsmooth algorithms such as a subgradient method or the linearized Bregman algorithm in [6].
@@
 Specifically, the subproblem can be reformulated as
-${\mathbf{\dot{X}}}_{k+1}=$ $\mathop{\mathrm{arg\,min}}_{{\mathbf{\dot{X}}}}\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)+$ (10)
-$\qquad\qquad\frac{\tau^{2}(\beta+L_{1,k})+1}{2\tau^{2}}\left\|{\mathbf{\dot{X}}}-\frac{\tau^{2}}{\tau^{2}(\beta+L_{1,k})+1}\left(\frac{1}{\tau^{2}}\mathbf{\dot{Y}}+\beta\mathcal{W}^{\#}(\mathbf{\dot{W}}_{k}-\frac{\mathbf{\dot{\Lambda}}_{k}}{\beta})+L_{1,k}{\mathbf{\dot{X}}}_{k}\right)\right\|_{\mathrm{F}}^{2},$
+${\mathbf{\dot{X}}}_{k+1}=$ $\mathop{\mathrm{arg\,min}}_{{\mathbf{\dot{X}}}}\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma)+$ (13)
+$\frac{\tau^{2}(\beta+L_{k,1})+1}{2\tau^{2}}\left\|{\mathbf{\dot{X}}}-\frac{\tau^{2}}{\tau^{2}(\beta+L_{k,1})+1}\left(\frac{1}{\tau^{2}}\mathbf{\dot{Y}}+\beta\mathcal{W}^{\#}(\mathbf{\dot{W}}_{k}-\frac{\mathbf{\dot{\Lambda}}_{k}}{\beta})+L_{k,1}{\mathbf{\dot{X}}}_{k}\right)\right\|_{\mathrm{F}}^{2},$
 If $\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)=\sigma_{i}({\mathbf{\dot{X}}}),$ the ${\mathbf{\dot{X}}}$-subproblem has a closed-form solution by using the soft thresholding method on the singular values [6].
 When a nonconvex surrogate function $\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)$ is used, the subproblem of ${\mathbf{\dot{X}}}$ can be solved via a convex-concave procedure (CCP) in [22].
-For completeness, CCP is stated in Algorithm 2.
-Lemma 3.1 is crucial in CCP and it transforms the matrix optimization problem of ${\mathbf{\dot{X}}}$ into an optimization problem of singular values.
-Lemma 3.1 ([10]).
+For completeness, we summarize CCP in Algorithm 2.
+Lemma 3.1 is crucial since it transforms the matrix optimization problem of ${\mathbf{\dot{X}}}$ into an optimization problem of singular values.
+We note that the case of $\varepsilon=0$ has been proven in [12, Theorem 3].
+We extend the theorem into the case that $\varepsilon\geq 0$.
+Lemma 3.1.
 Let $\mathbf{\hat{{\mathbf{\dot{X}}}}}=\mathbf{\dot{U}}\bm{\Sigma}\mathbf{\dot{V}}^{*}\in\mathbb{H}^{m\times n}$ be a singular value decomposition of $\mathbf{\hat{{\mathbf{\dot{X}}}}}$.
 The solution of
-$\mathop{\mathrm{arg\,min}}_{{\mathbf{\dot{X}}}\in\mathbb{H}^{m\times n}}\,\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)+\frac{\mu}{2}\|\mathbf{\hat{{\mathbf{\dot{X}}}}}-{\mathbf{\dot{X}}}\|_{\mathrm{F}}^{2}$ (11)
+$\mathop{\mathrm{arg\,min}}_{{\mathbf{\dot{X}}}\in\mathbb{H}^{m\times n}}\,\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma)+\frac{\mu}{2}\|\mathbf{\hat{{\mathbf{\dot{X}}}}}-{\mathbf{\dot{X}}}\|_{\mathrm{F}}^{2}$ (14)
 can be represented by ${\mathbf{\dot{X}}}_{*}=\mathbf{\dot{U}}\bm{\Sigma_{*}}\mathbf{\dot{V}}^{*},$ where $\bm{(\Sigma_{*})}_{i,i}=(\bm{\sigma_{*}})_{i},$ for $i=1,\cdots,\min\{m,n\}$ where $\bm{\sigma_{*}}$ is obtained by
-$\bm{\sigma_{*}}=\mathop{\mathrm{arg\,min}}_{\bm{\sigma}\,\geq 0}\phi(\bm{\sigma},\gamma)+\frac{\mu}{2}\|\bm{\sigma}-\bm{\sigma_{\bm{\hat{{\mathbf{\dot{X}}}}}}}\|_{2}^{2}.$ (12)
-In the $k$-th iteration of Algorithm 2, the function $\phi(\bm{\sigma},\gamma)$ is approximated by a linearized function $\phi(\bm{\sigma}_{k},\gamma)+\nabla_{\bm{\sigma}}\phi(\bm{\sigma}_{k},\gamma)^{\mathrm{T}}\bm{\sigma}$.
-It follows that the update of $\bm{\sigma}$ in Problem (12) at $\bm{\sigma}_{k}$ is given by
-$\bm{\sigma}_{k+1}=\max\left\{\bm{\sigma}_{\bm{\hat{{\mathbf{\dot{X}}}}}}-\frac{\nabla_{\bm{\sigma}}\phi(\bm{\sigma}_{k},\gamma)}{\mu},\bm{0}\right\}.$ (13)
-It has been shown in [20, Theorem 4] that any limit points of $\{\bm{\sigma}_{k}\}_{k=1}^{\infty}$ are stationary points of (12).
+$\bm{\sigma_{*}}=\mathop{\mathrm{arg\,min}}_{\bm{\sigma}\,\geq 0}\varphi(\bm{\sigma},\varepsilon;\gamma)+\frac{\mu}{2}\|\bm{\sigma}-\bm{\sigma_{\bm{\hat{{\mathbf{\dot{X}}}}}}}\|_{2}^{2},$ (15)
+where $\varphi(\bm{\sigma},\varepsilon;\gamma):=\sum_{i}\phi(\sqrt{\bm{\sigma}_{i}^{2}+\varepsilon^{2}},\gamma)$.
+Proof.
+According to the definition of QSVD, the following inequality holds.
+$\frac{\mu}{2}\|\mathbf{\hat{{\mathbf{\dot{X}}}}}-{\mathbf{\dot{X}}}\|_{\mathrm{F}}^{2}+\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma)=\frac{\mu}{2}\|\bm{\Sigma}-\mathbf{\dot{U}}^{*}{\mathbf{\dot{X}}}\mathbf{\dot{V}}\|_{\mathrm{F}}^{2}+\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma),$ (16)
+where the equality is based on the fact that the Frobenius norm is unitarily variant [12, Equation (37)].
+Suppose $\mathbf{\dot{A}}=\mathbf{\dot{U}}^{*}{\mathbf{\dot{X}}}\mathbf{\dot{V}}$, thus (16) is equal to
+$=\frac{\mu}{2}\|\bm{\Sigma}-\mathbf{\dot{U}}^{*}{\mathbf{\dot{X}}}\mathbf{\dot{V}}\|_{\mathrm{F}}^{2}+\sum_{i}\phi(\sigma_{i}(\sqrt{{\mathbf{\dot{A}}}^{*}\mathbf{\dot{A}}+\varepsilon^{2}\mathbf{I}}),\gamma)$ (17)
+$\geq\frac{\mu}{2}\left\|\bm{\Sigma}-\bm{\Sigma}_{\mathbf{\dot{A}}}\right\|_{\mathrm{F}}^{2}+\sum_{i}\phi(\sigma_{i}(\sqrt{{\mathbf{\dot{A}}}^{*}\mathbf{\dot{A}}+\varepsilon^{2}\mathbf{I}}),\gamma)$
+$=\frac{\mu}{2}\left\|\bm{\Sigma}-\bm{\Sigma}_{{\mathbf{\dot{X}}}}\right\|_{\mathrm{F}}^{2}+\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma)$
+$=\frac{\mu}{2}\left\|\bm{\sigma}-\bm{\sigma}({\mathbf{\dot{X}}})\right\|_{2}^{2}+\varphi(\bm{\sigma}({\mathbf{\dot{X}}}),\varepsilon;\gamma)$
+$\geq\frac{\mu}{2}\left\|\bm{\sigma}-\bm{\sigma}_{*}\right\|_{2}^{2}+\varphi(\bm{\sigma}_{*},\varepsilon;\gamma),$
+where $\bm{\sigma}({\mathbf{\dot{X}}})$ is the singular value vectors of ${\mathbf{\dot{X}}}$, the first inequality follows from Hoffman-Wielandt inequality.
+According to the definition of ${\mathbf{\dot{X}}}_{*}$, $\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}_{*}}^{*}{\mathbf{\dot{X}}}_{*}+\varepsilon^{2}\mathbf{I}}),\gamma)+\frac{\mu}{2}\|\mathbf{\hat{{\mathbf{\dot{X}}}}}-{\mathbf{\dot{X}}}_{*}\|_{\mathrm{F}}^{2}=\frac{\mu}{2}\left\|\bm{\sigma}-\bm{\sigma}_{*}\right\|_{2}^{2}+\varphi(\bm{\sigma}_{*},\varepsilon;\gamma)$.
+The proof is completed.
+∎
+In the $k$-th iteration of Algorithm 2, the function $\varphi(\bm{\sigma},\varepsilon;\gamma)$ is approximated by a linearized function $\varphi(\bm{\sigma}_{k},\varepsilon;\gamma)+\nabla_{\bm{\sigma}}\varphi(\bm{\sigma}_{k},\varepsilon;\gamma)^{\mathrm{T}}\bm{\sigma}$.
+It follows that the update of $\bm{\sigma}$ in Problem (15) is given by
+$\bm{\sigma}_{k+1}=\max\left\{\bm{\sigma}_{\bm{\hat{{\mathbf{\dot{X}}}}}}-\frac{\nabla_{\bm{\sigma}}\varphi(\bm{\sigma}_{k},\varepsilon;\gamma)}{\mu},\bm{0}\right\}.$ (18)
+It has been shown in [22, Theorem 4] that any limit points of $\{\bm{\sigma}_{k}\}_{k=1}^{\infty}$ are stationary points of (15).
 Algorithm 2 CCP to solve Problem (14)
-1:A quaternion matrix $\bm{\hat{{\mathbf{\dot{X}}}}}\in\mathbb{H}^{m\times n}$; a nonconvex function $\phi$; a constant $\mu>0$; a tolerance $\eta_{C}$; and an initial iterate $\bm{\sigma}_{0}=\mathbf{0}$;
-2:$\bm{{\mathbf{\dot{X}}}}_{*}$;
+1:A quaternion matrix $\bm{\hat{{\mathbf{\dot{X}}}}}\in\mathbb{H}^{m\times n}$, a nonconvex function $\phi$, a constant $\mu>0$, a tolerance $\eta_{C}$, and an initial iterate $\bm{\sigma}_{0}=\mathbf{0}$.
+2:$\bm{{\mathbf{\dot{X}}}}_{*}$.
 3:Compute QSVD $\mathbf{\hat{{\mathbf{\dot{X}}}}}=\mathbf{\dot{U}}\bm{\Sigma}\mathbf{\dot{V}}^{*}$ and let $\bm{\sigma_{\hat{{\mathbf{\dot{X}}}}}}$ satisfy $\bm{\Sigma}=\mathrm{diag}(\bm{\sigma_{\hat{{\mathbf{\dot{X}}}}}})$;
@@
 9:end for
-In Step 6, the $\mathbf{\dot{W}}$-subproblem approximately minimizes $\mathcal{L}_{\beta}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})+\frac{L_{2,k}}{2}\|\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}^{2}$ for a fixed ${\mathbf{\dot{X}}}_{k+1}$ and $\mathbf{\dot{\Lambda}}_{k}$.
-Specifically, the objective function of the $\mathbf{\dot{W}}$-subproblem is obtained by using a linearized strategy for $p(\mathbf{\dot{W}})$ in $\mathcal{L}_{\beta}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})+\frac{L_{2,k}}{2}\|\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}^{2}$.
+In Step 6, the $\mathbf{\dot{W}}$-subproblem approximately minimizes $\mathcal{L}_{\beta}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})+\frac{L_{k,2}}{2}\|\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}^{2}$ for a fixed ${\mathbf{\dot{X}}}_{k+1}$ and $\mathbf{\dot{\Lambda}}_{k}$.
+Specifically, the objective function of the $\mathbf{\dot{W}}$-subproblem is obtained by using a linearized strategy for $p(\mathbf{\dot{W}})$ in $\mathcal{L}_{\beta}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})+\frac{L_{k,2}}{2}\|\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}^{2}$.
 It follows that $\mathbf{\dot{W}}_{k+1}$ has a closed form solution
-$\mathbf{\dot{W}}_{k+1}=\frac{1}{\beta+L_{2,k}}\left(\beta\mathcal{W}({\mathbf{\dot{X}}}_{k+1})+\mathbf{\dot{\Lambda}}_{k}+L_{2,k}\mathbf{\dot{W}}_{k}\right)-\frac{\lambda}{\beta+L_{2,k}}\nabla p(\mathbf{\dot{W}}_{k}).$ (14)
-Step 7 is used to update the Lagrange multipliers.
-The difference between the classic ADMM for convex function and PL-ADMM is that $\mu\in(0,2)$ in PL-ADMM rather than $\mu\in(0,\frac{\sqrt{5}+1}{2})$ [8].
+$\mathbf{\dot{W}}_{k+1}=\frac{1}{\beta+L_{k,2}}\left(\beta\mathcal{W}({\mathbf{\dot{X}}}_{k+1})+\mathbf{\dot{\Lambda}}_{k}+L_{k,2}\mathbf{\dot{W}}_{k}\right)-\frac{\lambda}{\beta+L_{k,2}}\nabla p(\mathbf{\dot{W}}_{k}).$ (19)
+Step 7 is used to update the Lagrange multiplier.
+The difference between the classic ADMM for convex function and PL-ADMM is that $\mu\in(0,2)$ in PL-ADMM rather than $\mu\in(0,\frac{\sqrt{5}+1}{2})$ [10], which means that a larger stepsize can be chosen with convergence guarantee.
+3.2 An Algorithm for SLRQA-NF
+The proposed algorithm PL-ADMM-NF for solving SLRQA-NF is stated in Algorithm 3.
+Algorithm 3 PL-ADMM-NF algorithm for SLRQA-NF in (4)
+1:Corrupted images $\mathbf{\dot{Y}}$, two parameters $\mu\in(0,2)$ and $\beta>0$, a tolerance $\eta>0$, a sequence of regularization parameters $\{L_{k,1}\},\{L_{k,2}\}$ satisfying $\sup_{k\geq 0}{L_{k,1}}<\infty$, $\sup_{k\geq 0}{L_{k,2}}<\infty$ and $\inf_{k>0}{L_{k,1}}>0$, $\inf_{k>0}{L_{k,2}}>0$.
+2:Recovered $\mathbf{\dot{Z}}_{*}$.
+3:Set $\mathbf{\dot{Z}}_{0}=\bm{0}$, $\mathbf{\dot{W}}_{0}=\bm{0}$, $\mathbf{\dot{\Lambda}}_{1,0}=\bm{0}$, $\mathbf{\dot{\Lambda}}_{2,0}=\bm{0}$.
+4:for $k=0,1,2,\ldots$ do
+5:     $\mathbf{\dot{Z}}$ subproblem:
+$\mathbf{\dot{Z}}_{k+1}\in\mathop{\mathrm{arg\,min}}_{\mathbf{\dot{Z}}}\sum_{i}\phi(\sigma_{i}(\sqrt{{\mathbf{\dot{Z}}}^{*}\mathbf{\dot{Z}}+\varepsilon^{2}\mathbf{I}}),\gamma)+\left\langle\mathbf{\dot{\Lambda}}_{k,1},\mathbf{\dot{Z}}-\mathbf{\dot{W}}_{k}\right\rangle+\frac{\beta_{1}}{2}\|\mathbf{\dot{Z}}-\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}^{2}+\frac{L_{k,1}}{2}\|\mathbf{\dot{Z}}-\mathbf{\dot{Z}}_{k}\|_{\mathrm{F}}^{2};$
+6:     $\mathbf{\dot{W}}$ subproblem:
+$\mathbf{\dot{W}}_{k+1}$ $=\mathop{\mathrm{arg\,min}}_{\mathbf{\dot{W}}}\lambda\left\langle\nabla p(\mathbf{\dot{W}}_{k}),\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\right\rangle+\frac{\beta_{1}}{2}\|\mathbf{\dot{W}}-\mathbf{\dot{Z}}_{k+1}+\frac{1}{\beta_{1}}\mathbf{\dot{\Lambda}}_{k,1}\|_{\mathrm{F}}^{2}$
+$+\frac{\beta_{2}}{2}\left\|(\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}})-\mathbf{\dot{Y}})+\frac{1}{\beta_{2}}\mathbf{\dot{\Lambda}}_{k,2}\right\|_{\mathrm{F}}^{2}+\frac{L_{k,2}}{2}\|\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}^{2};$
+7:     $\mathbf{\dot{\Lambda}}_{k+1,1}=\mathbf{\dot{\Lambda}}_{k,1}+\mu\beta(\mathbf{\dot{Z}}_{k+1}-\mathbf{\dot{W}}_{k+1})$;
+8:     $\mathbf{\dot{\Lambda}}_{k+1,2}=\mathbf{\dot{\Lambda}}_{k,2}+\mu\beta\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{Y}})$;
+9:     if $\|\mathbf{\dot{Z}}_{k+1}-\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+\|\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}_{k+1}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}<\eta$ then
+10:         Return $\mathbf{\dot{Z}}_{*}=\mathbf{\dot{Z}}_{k+1}$; break
+11:     end if
+12:end for
+The augmented Lagrangian function of SLRQA-NF (4) is given by:
+~$\tilde{\mathcal{L}}_{\beta_{1},\beta_{2}}(\mathbf{\dot{Z}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}}_{1},\mathbf{\dot{\Lambda}}_{2})=\sum_{i}\phi(\sigma_{i}(\sqrt{{\mathbf{\dot{Z}}}^{*}\mathbf{\dot{Z}}+\varepsilon^{2}\mathbf{I}}),\gamma)+\lambda p(\mathbf{\dot{W}})+\left\langle\mathbf{\dot{\Lambda}}_{1},\mathbf{\dot{Z}}-\mathbf{\dot{W}}\right\rangle$ (20)
+$+\frac{\beta_{1}}{2}\|\mathbf{\dot{Z}}-\mathbf{\dot{W}}\|^{2}_{\mathrm{F}}+\left\langle\mathbf{\dot{\Lambda}}_{2},\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\right\rangle+\frac{\beta_{2}}{2}\|\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2}$
+$=\sum_{i}\phi(\sigma_{i}(\sqrt{{\mathbf{\dot{Z}}}^{*}\mathbf{\dot{Z}}+\varepsilon^{2}\mathbf{I}}),\gamma)+\lambda p(\mathbf{\dot{W}})+\frac{\beta_{1}}{2}\|\mathbf{\dot{Z}}-\mathbf{\dot{W}}+\frac{1}{\beta_{1}}\mathbf{\dot{\Lambda}}_{1}\|_{\mathrm{F}}^{2}$
+$+\frac{\beta_{2}}{2}\|\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})+\frac{1}{\beta_{2}}\mathbf{\dot{\Lambda}}_{2}\|_{\mathrm{F}}^{2}-\frac{1}{2\beta_{1}}\|\mathbf{\dot{\Lambda}}_{1}\|_{\mathrm{F}}^{2}-\frac{1}{2\beta_{2}}\|\mathbf{\dot{\Lambda}}_{2}\|^{2}_{\mathrm{F}},$
+where $\beta_{1},\beta_{2}>0$ are the penalty parameters, $\mathbf{\dot{\Lambda}}_{1},\mathbf{\dot{\Lambda}}_{2}\in\mathbb{H}^{m\times n}$ are the Lagrange multipliers.
+In Step 3 of Algorithm 3, since the subproblem of $\mathbf{\dot{Z}}$ is equivalent to
+$\min_{\mathbf{\dot{Z}}}\,\sum_{i}\phi(\sigma_{i}(\sqrt{{\mathbf{\dot{Z}}}^{*}\mathbf{\dot{Z}}+\varepsilon^{2}\mathbf{I}}),\gamma)+\frac{\beta+L_{k,1}}{2}\left\|\mathbf{\dot{Z}}-\frac{1}{\beta+L_{k,1}}\left(\beta\mathbf{\dot{W}}_{k}+\mathbf{\dot{\Lambda}}_{k,1}+L_{k,1}\mathbf{\dot{Z}}_{k}\right)\right\|_{\mathrm{F}}^{2},$
+Algorithm 2 can be used to obtain $\mathbf{\dot{Z}}_{k+1}$. In Step 6 of Algorithm 3, the solution of $\mathbf{\dot{W}}$ has an explicit solution:
+$\mathbf{\dot{W}}_{k+1}=\mathcal{G}^{-1}\left(\beta_{1}\mathbf{\dot{Z}}_{k+1}-\mathbf{\dot{\Lambda}}_{k,1}+\mathcal{W}\mathcal{A}^{\#}\mathcal{P}_{\Omega}^{\#}\left(\beta_{2}\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})-\mathbf{\dot{\Lambda}}_{k,2}\right)+L_{k,2}\mathbf{\dot{W}}_{k}-\lambda\nabla p(\mathbf{\dot{W}}_{k})\right),$
+where $\mathcal{G}=(\beta_{1}+L_{k,2})\mathcal{I}+\beta_{2}\mathcal{W}\mathcal{A}^{\#}\mathcal{P}_{\Omega}^{\#}\mathcal{P}_{\Omega}\mathcal{A}\mathcal{W}^{\#}$.
+Steps 7 and 8 are used to update the Lagrange multipliers.
+Finally, the recovered image ${\mathbf{\dot{X}}}_{*}$ is obtained by ${\mathbf{\dot{X}}}_{*}=\mathcal{W}^{\#}(\mathbf{\dot{Z}}_{*})$.
 4 Convergence analysis
+In this section, we present the global convergence analysis for Algorithm 1 and 3 respectively.
+The convergence analysis for 1 is similar to that in [43].
+However, the algorithm we analysis is slightly different from [43].
+Furthermore, we present more flexible assumptions.
+The convergence analysis for Algorithm 3 is different from the previous works since two linear constraints are considered.
+To the best of our knowledge, this is the first convergence analysis for ADMM with two linear constrants.
+4.1 Global Convergence for Algorithm 1
 The global convergence is established for a generalization of SLRQA in (2):
@@
 $\mathrm{s.t.}$ $\mathcal{C}({\mathbf{\dot{X}}})+\mathcal{B}(\mathbf{\dot{W}})=\mathbf{\dot{B}},$
-where $f$ is continuous and $g$ is continuously differentiable, $\mathcal{C},\mathcal{B}:\mathbb{H}^{m\times n}\rightarrow\mathbb{H}^{m\times n}$ are given linear maps and $\mathbf{\dot{B}}\in\mathbb{H}^{m\times n}$.
+where $f$ is semi-lowercontinuous and $g$ is continuously differentiable, $\mathcal{C},\mathcal{B}:\mathbb{H}^{m\times n}\rightarrow\mathbb{H}^{m\times n}$ are given linear maps and $\mathbf{\dot{B}}\in\mathbb{H}^{m\times n}$.
 Specifically, if $f({\mathbf{\dot{X}}})=\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma),g(\mathbf{\dot{W}})=p(\mathbf{\dot{W}})$, $\mathcal{B}=-\mathbf{I}$, $\mathcal{C}=\mathcal{W}$, and $\mathbf{\dot{B}}=0$, then Problem (21) becomes the SLRQA in (2).
@@
 $=f({\mathbf{\dot{X}}})+g(\mathbf{\dot{W}})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2}+\frac{\beta}{2}\left\|\mathcal{C}({\mathbf{\dot{X}}})+\mathcal{B}(\mathbf{\dot{W}})-\mathbf{\dot{B}}+\frac{\mathbf{\dot{\Lambda}}}{\beta}\right\|_{\mathrm{F}}^{2}-\frac{1}{2\beta}\left\|\mathbf{\dot{\Lambda}}\right\|_{\mathrm{F}}^{2}.$
-The proposed PL-ADMM for (15) is
+The iteration scheme for the proposed PL-ADMM for (21) is
 ${\mathbf{\dot{X}}}_{k+1}$ $\in\mathop{\mathrm{arg\,min}}_{{\mathbf{\dot{X}}}}f({\mathbf{\dot{X}}})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2}$ (23)
-$+\frac{\beta}{2}\left\|\mathcal{C}({\mathbf{\dot{X}}})+\mathcal{B}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}+\frac{\mathbf{\dot{\Lambda}}_{k}}{\beta}\right\|_{\mathrm{F}}^{2}+\frac{L_{1,k}}{2}\|{\mathbf{\dot{X}}}-{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}^{2},$
-$\mathbf{\dot{W}}_{k+1}$ $\in\mathop{\mathrm{arg\,min}}_{\mathbf{\dot{W}}}\left\langle\nabla g(\mathbf{\dot{W}}_{k}),\mathbf{\dot{W}}\right\rangle+\frac{\beta}{2}\left\|\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}})-\mathbf{\dot{B}}+\frac{\mathbf{\dot{\Lambda}}_{k}}{\beta}\right\|_{\mathrm{F}}^{2}+\frac{L_{2,k}}{2}\left\|\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\right\|^{2}_{\mathrm{F}},$
+$+\frac{\beta}{2}\left\|\mathcal{C}({\mathbf{\dot{X}}})+\mathcal{B}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}+\frac{\mathbf{\dot{\Lambda}}_{k}}{\beta}\right\|_{\mathrm{F}}^{2}+\frac{L_{k,1}}{2}\|{\mathbf{\dot{X}}}-{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}^{2},$
+$\mathbf{\dot{W}}_{k+1}$ $\in\mathop{\mathrm{arg\,min}}_{\mathbf{\dot{W}}}\left\langle\nabla g(\mathbf{\dot{W}}_{k}),\mathbf{\dot{W}}\right\rangle+\frac{\beta}{2}\left\|\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}})-\mathbf{\dot{B}}+\frac{\mathbf{\dot{\Lambda}}_{k}}{\beta}\right\|_{\mathrm{F}}^{2}+\frac{L_{k,2}}{2}\left\|\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\right\|^{2}_{\mathrm{F}},$
 $\mathbf{\dot{\Lambda}}_{k+1}$ $=\mathbf{\dot{\Lambda}}_{k}+\mu\beta\cdot(\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}),$
 which is a generalization of Algorithm 1.
-The choices of $\beta,L_{1,k}$ and $L_{2,k}$ are crucial for the convergence of (17).
-Generally speaking, $\beta$ should be chosen large enough, and $L_{1,k}$, $L_{2,k}$ should be chosen in a suitable range.
-In order to give the conditions of $\beta$ and $L_{1,k},L_{2,k}$, we introduce the following notations.
+The choices of $\beta,L_{k,1}$ and $L_{k,2}$ are crucial for the convergence of (59).
+Generally speaking, $\beta$ should be chosen large enough, and $L_{k,1}$, $L_{k,2}$ should be chosen in a suitable range.
+In order to give the conditions of $\beta$ and $L_{k,1},L_{k,2}$, we introduce the following notations.
 Let $q_{i}:=\sup_{k\geq 0}{L_{i,k}}<\infty$ and $q_{i}^{-}:=\inf_{k>0}{L_{i,k}}>0$ for $i=1,2$.
@@
 For any sequence $\{\mathbf{\dot{U}}_{k}\}_{k\geq 0}$, define $\Delta$ operator by $\Delta\mathbf{\dot{U}}_{k}:=\mathbf{\dot{U}}_{k}-\mathbf{\dot{U}}_{k-1}.$
-The assumptions for the convergence analysis are given in Assumption 4.1, including the conditions of $\beta$, $L_{1,k}$, and $L_{2,k}$.
+The assumptions for the convergence analysis are given in Assumption 4.1, including the conditions of $\beta$, $L_{k,1}$, and $L_{k,2}$.
 Assumption 4.1.
@@
     $\|\nabla g(\mathbf{\dot{W}}_{1})-\nabla g(\mathbf{\dot{W}}_{2})\|_{\mathrm{F}}\leq L_{g}\|\mathbf{\dot{W}}_{1}-\mathbf{\dot{W}}_{2}\|_{\mathrm{F}}.$ (24)
-    It has been shown in [2, Lemma 5.7] that (18) implies that for every $\mathbf{\dot{W}}_{1}$ and $\mathbf{\dot{W}}_{2}$, it holds that
+    It has been shown in [4, Lemma 5.7] that (24) implies that for every $\mathbf{\dot{W}}_{1}$ and $\mathbf{\dot{W}}_{2}$, it holds that
     $g(\mathbf{\dot{W}}_{1})\leq g(\mathbf{\dot{W}}_{2})+\left\langle\nabla g(\mathbf{\dot{W}}_{2}),\mathbf{\dot{W}}_{1}-\mathbf{\dot{W}}_{2}\right\rangle+\frac{L_{g}}{2}\|\mathbf{\dot{W}}_{1}-\mathbf{\dot{W}}_{2}\|_{\mathrm{F}}^{2}.$ (25)
   3. A3
-    The matrix $\mathcal{B}^{\#}\mathcal{B}$ is full rank, ${\rm range}(\mathcal{C})\subseteq{\rm range}(\mathcal{B})$, and $\mathbf{\dot{B}}\in{\rm range}(\mathcal{B})$.
+    The matrix $\mathcal{B}^{\#}\mathcal{B}$ is full rank or $g(\mathbf{\dot{W}})$ is coercive, ${\rm range}(\mathcal{C})\subseteq{\rm range}(\mathcal{B})$, and $\mathbf{\dot{B}}\in{\rm range}(\mathcal{B})$.
   4. A4
     The parameters $\beta>0$, $\mu\in(0,2)$, and there exist three constants $a_{1}>0,a_{2}>0$, and $r>1$ such that
-    $q_{1}^{-}\mathbf{I}\succeq a_{1}\mathbf{I}\,\quad\mbox{and}\,\quad q_{2}^{-}\mathbf{I}+\beta\mathcal{B}^{\#}\mathcal{B}-(r\theta_{0}+r\theta_{1}+L_{g})\mathbf{I}\succeq a_{2}\mathbf{I},$ (20)
+    $q_{1}^{-}\mathbf{I}\succeq a_{1}\mathbf{I}\,\quad\mbox{and}\,\quad q_{2}^{-}\mathbf{I}+\beta\mathcal{B}^{\#}\mathcal{B}-(r\varsigma_{0}+r\varsigma_{1}+L_{g})\mathbf{I}\succeq a_{2}\mathbf{I},$ (26)
     where $\mathbf{I}$ is the identity matrix with correct size,
-    $\theta_{0}:=\frac{2\mu(q_{2}+L_{g})^{2}}{\beta\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}(\rho(\mu))^{2}},\,\hbox{ and }\theta_{1}:=\frac{2\mu q_{2}^{2}}{\beta\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}(\rho(\mu))^{2}}.$ (21)
+    $\varsigma_{0}:=\frac{2\mu(q_{2}+L_{g})^{2}}{\beta\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}(\rho(\mu))^{2}},\,\hbox{ and }\varsigma_{1}:=\frac{2\mu q_{2}^{2}}{\beta\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}(\rho(\mu))^{2}}.$ (27)
   5. A5
-    The parameter $\beta$ satisfies $\beta>\frac{2L_{g}}{\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}\rho(\mu)}.$
+    The parameter $\beta$ satisfies $\beta>\frac{2L_{g}}{\kappa\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}\rho(\mu)}$, where $\kappa\in(0,1)$ is a given constant.
 One can verify that A1, A2, and A3 in Assumption 4.1 are satisfied for SLRQA in (2).
-If $\beta$ is chosen sufficiently large and the sequences $\{L_{1,k}\}$ and $\{L_{2,k}\}$ are chosen such that $q_{i}^{-},q_{i}$, $i=1,2$ in suitable range, then A4 and A5 in Assumption 4.1 are also satisfied.
-For example, if set $L_{1,k}=L_{2,k}=1$ for all $k$ and let
+If $\beta$ is chosen sufficiently large and the sequences $\{L_{k,1}\}$ and $\{L_{k,2}\}$ are chosen such that $q_{i}^{-},q_{i}$, $i=1,2$ in suitable range, then A4 and A5 in Assumption 4.1 are also satisfied.
+For example, if set $L_{k,1}=L_{k,2}=1$ for all $k$ and let
 $\beta>\max\left\{\frac{2L_{g}}{\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}\rho(\mu)},\frac{r(2\mu(1+L_{g})^{2}+2\mu)}{(\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}})^{2}(\rho(\mu))^{2}}+\frac{L_{g}}{\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}},1\right\},$
@@
 The sequence $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ generated by (59) satisfies:
-$\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})-\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\geq\frac{L_{1,k}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2}_{\mathrm{F}}.$
+$\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})-\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\geq\frac{L_{k,1}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2}_{\mathrm{F}}.$
 Proof.
 According to the ${\mathbf{\dot{X}}}_{k+1}$ update in (59), the following inequality holds:
-$f({\mathbf{\dot{X}}}_{k+1})-f({\mathbf{\dot{X}}}_{k})+\frac{\beta}{2}\|\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta\|_{\mathrm{F}}^{2}+\frac{L_{1,k}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}$ (22)
+$f({\mathbf{\dot{X}}}_{k+1})-f({\mathbf{\dot{X}}}_{k})+\frac{\beta}{2}\|\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta\|_{\mathrm{F}}^{2}+\frac{L_{k,1}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}$ (28)
 $+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}((\mathcal{A}({\mathbf{\dot{X}}}_{k+1})-\mathbf{\dot{Y}}))\|_{\mathrm{F}}^{2}\leq\frac{\beta}{2}\left\|\mathcal{C}({\mathbf{\dot{X}}}_{k})+\mathcal{B}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta\right\|_{\mathrm{F}}^{2}+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}((\mathcal{A}({\mathbf{\dot{X}}}_{k})-\mathbf{\dot{Y}}))\|_{\mathrm{F}}^{2}.$
@@
 $=f({\mathbf{\dot{X}}}_{k})-f({\mathbf{\dot{X}}}_{k+1})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}((\mathcal{A}({\mathbf{\dot{X}}}_{k})-\mathbf{\dot{Y}}))\|_{\mathrm{F}}^{2}-\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}((\mathcal{A}({\mathbf{\dot{X}}}_{k+1})-\mathbf{\dot{Y}}))\|_{\mathrm{F}}^{2}$
-$-\frac{\beta}{2}\|\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta\|_{\mathrm{F}}^{2}+\frac{\beta}{2}\|\mathcal{C}({\mathbf{\dot{X}}}_{k})+\mathcal{B}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta\|_{\mathrm{F}}^{2}\geq\frac{L_{1,k}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2},$
+$-\frac{\beta}{2}\|\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta\|_{\mathrm{F}}^{2}+\frac{\beta}{2}\|\mathcal{C}({\mathbf{\dot{X}}}_{k})+\mathcal{B}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta\|_{\mathrm{F}}^{2}\geq\frac{L_{k,1}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2},$
 which completes the proof. ∎
@@
 $\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})-\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k})\geq\|\Delta\mathbf{\dot{W}}_{k+1}\|^{2}_{B_{k}},$
-where $B_{k}:=L_{2,k}\mathbf{I}-L_{g}\mathbf{I}+\frac{\beta}{2}\mathcal{B}\mathcal{B}^{\#}$.
+where $B_{k}:=L_{k,2}\mathbf{I}-L_{g}\mathbf{I}+\frac{\beta}{2}\mathcal{B}\mathcal{B}^{\#}$.
 Proof.
 According to the optimality condition of $\mathbf{\dot{W}}_{k+1},$ the following equality holds:
-$\nabla g(\mathbf{\dot{W}}_{k})+\beta\mathcal{B}^{\#}(\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta)+L_{2,k}\Delta\mathbf{\dot{W}}_{k+1}=0.$ (23)
+$\nabla g(\mathbf{\dot{W}}_{k})+\beta\mathcal{B}^{\#}(\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta)+L_{k,2}\Delta\mathbf{\dot{W}}_{k+1}=0.$ (29)
 It follows that
@@
 Suppose A2 in Assumption 4.1 holds. The sequence $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ generated by (59) satisfies:
-$\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1})\leq\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})-\|\Delta\mathbf{\dot{W}}_{k+1}\|^{2}_{B_{k}}-\frac{L_{1,k}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2}_{\mathrm{F}}+\frac{1}{\beta\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}.$
+$\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1})\leq\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})-\|\Delta\mathbf{\dot{W}}_{k+1}\|^{2}_{B_{k}}-\frac{L_{k,1}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2}_{\mathrm{F}}+\frac{1}{\beta\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}.$
 Proof.
@@
 $\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1})$ $=\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k})+\frac{1}{\beta\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}$
-$\leq\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})-\|\Delta\mathbf{\dot{W}}_{k+1}\|_{B_{k}}^{2}-\frac{L_{1,k}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}+\frac{1}{\beta\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}.$
+$\leq\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})-\|\Delta\mathbf{\dot{W}}_{k+1}\|_{B_{k}}^{2}-\frac{L_{k,1}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}+\frac{1}{\beta\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}.$
 Lemma 4.4.
-Suppose A2,A3 in Assumption 4.1 hold, the sequence $\{({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1})\}_{k\geq 0}$ generated by (17) satisfies:
-$\frac{1}{\beta\mu}\|\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}\leq\theta_{0}\|\Delta\mathbf{\dot{W}}_{k}\|^{2}_{\mathrm{F}}+\theta_{1}\|\Delta\mathbf{\dot{W}}_{k+1}\|^{2}_{\mathrm{F}}+\theta_{2}\|\mathcal{B}^{\#}(\Delta\mathbf{\dot{\Lambda}}_{k})\|^{2}_{\mathrm{F}}-\theta_{2}\|\mathcal{B}^{\#}(\Delta\mathbf{\dot{\Lambda}}_{k+1})\|^{2}_{\mathrm{F}},$ (24)
-$\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1})+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{B_{k}-r\theta_{1}\mathbf{I}}^{2}+\frac{L_{1,k}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2}_{\mathrm{F}}+\frac{r-1}{\beta\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}$
-$+r\theta_{2}\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}\leq\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})+r\theta_{0}\|\Delta\mathbf{\dot{W}}_{k}\|^{2}_{\mathrm{F}}+r\theta_{2}\|\mathcal{B}^{\#}(\Delta\mathbf{\dot{\Lambda}}_{k})\|^{2}_{\mathrm{F}},$ (25)
-where $r>1,\theta_{0},\theta_{1}$ are defined in (21), and $\theta_{2}:=\frac{|1-\mu|}{\beta\mu\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}\rho(\mu)}.$
+Suppose A2,A3 in Assumption 4.1 hold, the sequence $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ generated by (59) satisfies:
+$\frac{1}{\beta\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}\leq\varsigma_{0}\|\Delta\mathbf{\dot{W}}_{k}\|^{2}_{\mathrm{F}}+\varsigma_{1}\|\Delta\mathbf{\dot{W}}_{k+1}\|^{2}_{\mathrm{F}}+\varsigma_{2}\|\mathcal{B}^{\#}(\Delta\mathbf{\dot{\Lambda}}_{k})\|^{2}_{\mathrm{F}}-\varsigma_{2}\|\mathcal{B}^{\#}(\Delta\mathbf{\dot{\Lambda}}_{k+1})\|^{2}_{\mathrm{F}},$ (30)
+$\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1})+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{B_{k}-r\varsigma_{1}\mathbf{I}}^{2}+\frac{L_{k,1}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2}_{\mathrm{F}}+\frac{r-1}{\beta\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}$
+$+r\varsigma_{2}\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}\leq\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})+r\varsigma_{0}\|\Delta\mathbf{\dot{W}}_{k}\|^{2}_{\mathrm{F}}+r\varsigma_{2}\|\mathcal{B}^{\#}(\Delta\mathbf{\dot{\Lambda}}_{k})\|^{2}_{\mathrm{F}},$ (31)
+where $r>1,\varsigma_{0},\varsigma_{1}$ are defined in (27), and $\varsigma_{2}:=\frac{|1-\mu|}{\beta\mu\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}\rho(\mu)}.$
 Proof.
 Let $k\geq 1$ be fixed and define the matrix
-$\mathbf{\dot{G}}_{k+1}:=-L_{2,k}\Delta\mathbf{\dot{W}}_{k+1}-\nabla g(\mathbf{\dot{W}}_{k}).$ (26)
-It follows that $\Delta\mathbf{\dot{G}}_{k+1}=L_{2,k-1}\Delta\mathbf{\dot{W}}_{k}-L_{2,k}\Delta\mathbf{\dot{W}}_{k+1}+\nabla g(\mathbf{\dot{W}}_{k-1})-\nabla g(\mathbf{\dot{W}}_{k}).$ By the triangle inequality, the following inequality holds:
-$\|\Delta\mathbf{\dot{G}}_{k+1}\|_{\mathrm{F}}\leq\|\nabla g(\mathbf{\dot{W}}_{k-1})-\nabla g(\mathbf{\dot{W}}_{k})\|_{\mathrm{F}}+L_{2,k}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+L_{2,k-1}\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}.$
-By A2 in Assumption 4.1, $\nabla g(\mathbf{\dot{W}})$ is $L_{g}$ Lipschitz continuous and $q_{2}=\sup_{k\geq 0}L_{2,k}<\infty,$ we have $\|\Delta\mathbf{\dot{G}}_{k+1}\|_{\mathrm{F}}\leq(L_{g}+q_{2})\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}+q_{2}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}.$ Hence it follows that:
+$\mathbf{\dot{G}}_{k+1}:=-L_{k,2}\Delta\mathbf{\dot{W}}_{k+1}-\nabla g(\mathbf{\dot{W}}_{k}).$ (32)
+It follows that $\Delta\mathbf{\dot{G}}_{k+1}=L_{2,k-1}\Delta\mathbf{\dot{W}}_{k}-L_{k,2}\Delta\mathbf{\dot{W}}_{k+1}+\nabla g(\mathbf{\dot{W}}_{k-1})-\nabla g(\mathbf{\dot{W}}_{k}).$ By the triangle inequality, the following inequality holds:
+$\|\Delta\mathbf{\dot{G}}_{k+1}\|_{\mathrm{F}}\leq\|\nabla g(\mathbf{\dot{W}}_{k-1})-\nabla g(\mathbf{\dot{W}}_{k})\|_{\mathrm{F}}+L_{k,2}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+L_{2,k-1}\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}.$
+By A2 in Assumption 4.1, $\nabla g(\mathbf{\dot{W}})$ is $L_{g}$ Lipschitz continuous and $q_{2}=\sup_{k\geq 0}L_{k,2}<\infty,$ we have $\|\Delta\mathbf{\dot{G}}_{k+1}\|_{\mathrm{F}}\leq(L_{g}+q_{2})\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}+q_{2}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}.$ Hence it follows that:
 $\|\Delta\mathbf{\dot{G}}_{k+1}\|^{2}_{\mathrm{F}}\leq 2(L_{g}+q_{2})^{2}\|\Delta\mathbf{\dot{W}}_{k}\|^{2}_{\mathrm{F}}+2q_{2}^{2}\|\Delta\mathbf{\dot{W}}_{k+1}\|^{2}_{\mathrm{F}}.$ (33)
@@
 ∎
-Let $\mathcal{R}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}},\mathbf{\dot{W}}^{\prime},\mathbf{\dot{\Lambda}}^{\prime}):=\mathcal{L}_{\beta}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})+r\theta_{0}\|\mathbf{\dot{W}}-\mathbf{\dot{W}}^{\prime}\|^{2}_{\mathrm{F}}+r\theta_{2}\|\mathcal{B}^{\#}(\mathbf{\dot{\Lambda}}-\mathbf{\dot{\Lambda}}^{\prime})\|^{2}_{\mathrm{F}}$.
+Let $\mathcal{R}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}},\mathbf{\dot{W}}^{\prime},\mathbf{\dot{\Lambda}}^{\prime}):=\mathcal{L}_{\beta}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})+r\varsigma_{0}\|\mathbf{\dot{W}}-\mathbf{\dot{W}}^{\prime}\|^{2}_{\mathrm{F}}+r\varsigma_{2}\|\mathcal{B}^{\#}(\mathbf{\dot{\Lambda}}-\mathbf{\dot{\Lambda}}^{\prime})\|^{2}_{\mathrm{F}}$.
 The merit function $\mathcal{R}_{k}$ is defined by
-$\mathcal{R}_{k}:=\mathcal{R}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},\mathbf{\dot{\Lambda}}_{k-1})=\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})+r\theta_{0}\|\Delta\mathbf{\dot{W}}_{k}\|^{2}_{\mathrm{F}}+r\theta_{2}\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k}\|^{2}_{\mathrm{F}}.$ (30)
+$\mathcal{R}_{k}:=\mathcal{R}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},\mathbf{\dot{\Lambda}}_{k-1})=\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})+r\varsigma_{0}\|\Delta\mathbf{\dot{W}}_{k}\|^{2}_{\mathrm{F}}+r\varsigma_{2}\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k}\|^{2}_{\mathrm{F}}.$ (36)
 If Assumption 4.1 holds, then it can be shown that the merit function is descent. Specifically, we have
 $\mathcal{R}_{k+1}+a\left(\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}\right)$ (37)
-$\leq\mathcal{R}_{k+1}+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{B_{k}-r(\theta_{1}+\theta_{0})\mathbf{I}}^{2}+\frac{L_{1,k}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2}_{\mathrm{F}}+\frac{r-1}{\beta\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}\leq\mathcal{R}_{k}\leq\mathcal{R}_{k_{0}},$
+$\leq\mathcal{R}_{k+1}+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{B_{k}-r(\varsigma_{1}+\varsigma_{0})\mathbf{I}}^{2}+\frac{L_{k,1}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2}_{\mathrm{F}}+\frac{r-1}{\beta\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}\leq\mathcal{R}_{k}\leq\mathcal{R}_{k_{0}},$
 where $a=\min\{a_{1},a_{2},\frac{r-1}{\beta\mu}\}$ and $a_{1},a_{2}$ are defined in Assumption 4.1, the first inequality follows from A4, A5 in Assumption 4.1, the second inequality follows from Lemma 4.4, and the third inequality is due to induction of $\mathcal{R}_{k}\leq\mathcal{R}_{k+1}$ for any $k\geq k_{0}$.
@@
 $f({\mathbf{\dot{X}}}_{k+1})+g(\mathbf{\dot{W}}_{k+1})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}_{k}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2}+\frac{\beta}{2}\left\|\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}+\frac{\mathbf{\dot{\Lambda}}_{k+1}}{\beta}\right\|_{\mathrm{F}}^{2}$ (38)
-$-\frac{1}{2\beta}\left\|\mathbf{\dot{\Lambda}}_{k+1}\right\|_{\mathrm{F}}^{2}+(r\theta_{0}+a)\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}+a(\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2})+r\theta_{2}\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2}\leq\mathcal{R}_{k_{0}}.$
+$-\frac{1}{2\beta}\left\|\mathbf{\dot{\Lambda}}_{k+1}\right\|_{\mathrm{F}}^{2}+(r\varsigma_{0}+a)\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}+a(\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2})+r\varsigma_{2}\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2}\leq\mathcal{R}_{k_{0}}.$
 According to (34), we have
@@
 $\mu\mathcal{B}^{\#}\mathbf{\dot{\Lambda}}_{k+1}=\rho(\mu)\frac{\mu\mathbf{\dot{G}}_{k+1}}{\rho(\mu)}+|1-\mu|\left(\text{sign}(1-\mu)\mathcal{B}^{\#}(\mathbf{\dot{\Lambda}}_{k}-\mathbf{\dot{\Lambda}}_{k+1})\right).$
-By the convexity of $\|\cdot\|^{2}$, we have $\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}\mu^{2}\|\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2}\leq\frac{\mu^{2}}{\rho(\mu)}\|\mathbf{\dot{G}}_{k+1}\|_{\mathrm{F}}^{2}+|1-\mu|\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2}.$ According to definition (26), we have $\|\mathbf{\dot{G}}_{k+1}\|_{\mathrm{F}}^{2}\leq 2q_{2}^{2}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}+2\|\nabla g(\mathbf{\dot{W}}_{k})\|_{\mathrm{F}}^{2}\leq 2(q_{2}+L_{g})^{2}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}+2\|\nabla g(\mathbf{\dot{W}}_{k+1})\|_{\mathrm{F}}^{2}.$ It follows that:
-$-\frac{1}{2\beta}\|\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2}\geq-\theta_{3}\|\nabla g(\mathbf{\dot{W}}_{k+1})\|_{\mathrm{F}}^{2}-\theta_{4}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}-\theta_{5}\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2},$ (34)
+It follows from the convexity of $\|\cdot\|^{2}$ that $\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}\mu^{2}\|\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2}\leq\frac{\mu^{2}}{\rho(\mu)}\|\mathbf{\dot{G}}_{k+1}\|_{\mathrm{F}}^{2}+|1-\mu|\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2}.$ According to definition (32), we have $\|\mathbf{\dot{G}}_{k+1}\|_{\mathrm{F}}^{2}\leq 2q_{2}^{2}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}+2\|\nabla g(\mathbf{\dot{W}}_{k})\|_{\mathrm{F}}^{2}\leq 2(q_{2}+L_{g})^{2}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}+2\|\nabla g(\mathbf{\dot{W}}_{k+1})\|_{\mathrm{F}}^{2}.$ Hence it follows that:
+$-\frac{1}{2\beta}\|\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2}\geq-\varsigma_{3}\|\nabla g(\mathbf{\dot{W}}_{k+1})\|_{\mathrm{F}}^{2}-\varsigma_{4}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}-\varsigma_{5}\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2},$ (40)
 where
-$\theta_{3}:=\frac{1}{\beta\rho(\mu)\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}},\quad\theta_{4}:=\frac{(q_{2}+L_{g})^{2}}{\beta\rho(\mu)\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}},\quad\theta_{5}:=\frac{|1-\mu|}{2\beta\mu^{2}\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}}.$
+$\varsigma_{3}:=\frac{1}{\beta\rho(\mu)\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}},\quad\varsigma_{4}:=\frac{(q_{2}+L_{g})^{2}}{\beta\rho(\mu)\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}},\quad\varsigma_{5}:=\frac{|1-\mu|}{2\beta\mu^{2}\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}}.$
 Using Inequalities (40) and (38), we obtain
-$f({\mathbf{\dot{X}}}_{k+1})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}_{k}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2}+\frac{\beta}{2}\left\|\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}+\frac{\mathbf{\dot{\Lambda}}_{k+1}}{\beta}\right\|_{\mathrm{F}}^{2}$ (35)
-$+a(\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2})+(r\theta_{0}+a-\theta_{4})\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}+(r\theta_{2}-\theta_{5})\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2}$
-$\leq\mathcal{R}_{k_{0}}-\inf_{\mathbf{\dot{W}}}\left\{g(\mathbf{\dot{W}})-\theta_{3}\|\nabla g(\mathbf{\dot{W}})\|_{\mathrm{F}}^{2}\right\}.$
-According to (19) in Assumption A2, setting $\mathbf{\dot{W}}_{1}=\mathbf{\dot{W}}-\delta\nabla g(\mathbf{\dot{W}})$ and $\mathbf{\dot{W}}_{2}=\mathbf{\dot{W}}$, it follows that $g(\mathbf{\dot{W}}_{k}-\delta\nabla g(\mathbf{\dot{W}}_{k}))\leq g(\mathbf{\dot{W}}_{k})-(\delta-\frac{L_{g}\delta^{2}}{2})\|\nabla g(\mathbf{\dot{W}}_{k})\|_{\mathrm{F}}^{2}.$ Since $g$ is bounded from below, there exist $M$ such that
-$-M<\inf\{g(\mathbf{\dot{W}})-(\delta-\frac{L_{g}\delta^{2}}{2})\|\nabla g(\mathbf{\dot{W}})\|_{\mathrm{F}}^{2}:\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}\}.$ (36)
+$f({\mathbf{\dot{X}}}_{k+1})+(1-\kappa)g(\mathbf{\dot{W}}_{k+1})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}_{k}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2}+\frac{\beta}{2}\left\|\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}+\frac{\mathbf{\dot{\Lambda}}_{k+1}}{\beta}\right\|_{\mathrm{F}}^{2}$ (41)
+$+a(\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2})+(r\varsigma_{0}+a-\varsigma_{4})\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}+(r\varsigma_{2}-\varsigma_{5})\|\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2}$
+$\leq\mathcal{R}_{k_{0}}-\inf_{\mathbf{\dot{W}}}\left\{\kappa g(\mathbf{\dot{W}})-\varsigma_{3}\|\nabla g(\mathbf{\dot{W}})\|_{\mathrm{F}}^{2}\right\},$
+where $\kappa\in(0,1).$ According to (25) in Assumption A2, setting $\mathbf{\dot{W}}_{1}=\mathbf{\dot{W}}-\delta\nabla g(\mathbf{\dot{W}})$ and $\mathbf{\dot{W}}_{2}=\mathbf{\dot{W}}$, it follows that $\kappa g(\mathbf{\dot{W}}_{k}-\delta\nabla g(\mathbf{\dot{W}}_{k}))\leq\kappa g(\mathbf{\dot{W}}_{k})-\kappa(\delta-\frac{L_{g}\delta^{2}}{2})\|\nabla g(\mathbf{\dot{W}}_{k})\|_{\mathrm{F}}^{2}.$ Since $g$ is bounded from below, there exist $M$ such that
+$-M<\inf\{\kappa g(\mathbf{\dot{W}})-\kappa(\delta-\frac{L_{g}\delta^{2}}{2})\|\nabla g(\mathbf{\dot{W}})\|_{\mathrm{F}}^{2}:\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}\}.$ (42)
 We choose $\delta=\frac{1}{L_{g}}$.
-According to Assumption A5, we have $\theta_{3}<\frac{1}{2L_{g}}=\delta-\frac{L_{g}\delta^{2}}{2}.$ Since $r>1$ and $\mu\in(0,2),$ according to the denifition of $\theta_{0},\theta_{2},\theta_{4}$ and $\theta_{5}$, it holds that $r\theta_{2}-\theta_{5}>0$ and $r\theta_{0}+a-\theta_{4}>0$.
+According to Assumption A5, we have $\varsigma_{3}<\frac{\kappa}{2L_{g}}=\kappa(\delta-\frac{L_{g}\delta^{2}}{2}).$ Since $r>1$ and $\mu\in(0,2),$ according to the denifition of $\varsigma_{0},\varsigma_{2},\varsigma_{4}$ and $\varsigma_{5}$, it holds that $r\varsigma_{2}-\varsigma_{5}>0$ and $r\varsigma_{0}+a-\varsigma_{4}>0$.
 It follows from (41) that
-$f({\mathbf{\dot{X}}}_{k+1})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}_{k}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2}$ (37)
+$f({\mathbf{\dot{X}}}_{k+1})+(1-\kappa)g(\mathbf{\dot{W}}_{k+1})+\frac{1}{2\tau^{2}}\|\mathcal{P}_{\Omega}(\mathcal{A}({\mathbf{\dot{X}}}_{k}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2}$ (43)
 $+\frac{\beta}{2}\left\|\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}+\frac{\mathbf{\dot{\Lambda}}_{k+1}}{\beta}\right\|_{\mathrm{F}}^{2}+a(\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}^{2})<\mathcal{R}_{k_{0}}+M.$
@@
 According to (43), $\Delta\mathbf{\dot{\Lambda}}_{k+1}$ is bounded.
-By the $\mathbf{\dot{\Lambda}}$ update in (17), we have $\mathcal{B}\mathbf{\dot{W}}_{k+1}=\frac{1}{\beta\mu}\Delta\mathbf{\dot{\Lambda}}_{k+1}-\mathcal{C}{\mathbf{\dot{X}}}_{k+1}+\mathbf{\dot{B}}.$ Since $\{\mathcal{C}{\mathbf{\dot{X}}}_{k}\}_{k\geq k_{0}}$ is bounded and $\mathcal{B}^{\#}\mathcal{B}$ is full rank, it follows that $\{\mathbf{\dot{W}}_{k}\}_{k\geq k_{0}}$ is bounded.
+By the $\mathbf{\dot{\Lambda}}$ update in (59), we have $\mathcal{B}\mathbf{\dot{W}}_{k+1}=\frac{1}{\beta\mu}\Delta\mathbf{\dot{\Lambda}}_{k+1}-\mathcal{C}{\mathbf{\dot{X}}}_{k+1}+\mathbf{\dot{B}}.$ Since $\{\mathcal{C}{\mathbf{\dot{X}}}_{k}\}_{k\geq k_{0}}$ is bounded, if $\mathcal{B}^{\#}\mathcal{B}$ is full rank or $g$ is coercive, $\{\mathbf{\dot{W}}_{k}\}_{k\geq k_{0}}$ is bounded.
 From the fact that $\frac{\beta}{2}\left\|\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}+\frac{\mathbf{\dot{\Lambda}}_{k+1}}{\beta}\right\|_{\mathrm{F}}^{2}$ is bounded, it follows that $\{\mathbf{\dot{\Lambda}}_{k}\}_{k\geq k_{0}}$ is bounded.
@@
 Then $\widetilde{\mathbf{\dot{D}}}_{k+1}:=(\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k+1}},\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k+1}},\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k+1}})\in\partial\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1})$, where
-$\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k+1}}$ $:=\mathcal{C}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}-L_{2,k}\Delta{\mathbf{\dot{X}}}_{k+1}-\beta\mathcal{C}^{\#}\mathcal{B}\Delta\mathbf{\dot{W}}_{k+1},\quad\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k+1}}:=\frac{1}{\beta\mu}\Delta\mathbf{\dot{\Lambda}}_{k+1},\,$
-$\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k+1}}$ $:=\nabla g(\mathbf{\dot{W}}_{k+1})-\nabla g(\mathbf{\dot{W}}_{k})+\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}-L_{1,k}\Delta\mathbf{\dot{W}}_{k+1}.$
-Furthermore, there exists $\rho>0$ such that
+$\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k+1}}$ $:=\mathcal{C}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}-L_{k,1}\Delta{\mathbf{\dot{X}}}_{k+1}-\beta\mathcal{C}^{\#}\mathcal{B}\Delta\mathbf{\dot{W}}_{k+1},\quad\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k+1}}:=\frac{1}{\beta\mu}\Delta\mathbf{\dot{\Lambda}}_{k+1},\,$
+$\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k+1}}$ $:=\nabla g(\mathbf{\dot{W}}_{k+1})-\nabla g(\mathbf{\dot{W}}_{k})+\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}-L_{k,2}\Delta\mathbf{\dot{W}}_{k+1}.$
+Furthermore, there exists $\pi>0$ such that
 $|\|\widetilde{\mathbf{\dot{D}}}_{k+1}\||\leq\pi\left(\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}\right),$
@@
 According to the optimality condition of the update of ${\mathbf{\dot{X}}}_{k+1},$ it follows that
-$-\frac{1}{\tau^{2}}\mathcal{A}^{\#}\mathcal{P}^{\#}_{\Omega}(\mathcal{P}_{\Omega}(\mathcal{A}{\mathbf{\dot{X}}}_{k+1}-\mathbf{\dot{Y}}))-\beta\mathcal{C}^{\#}(\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta)-L_{1,k}\Delta{\mathbf{\dot{X}}}_{k+1}\in\partial f({\mathbf{\dot{X}}}_{k+1}),$
+$-\frac{1}{\tau^{2}}\mathcal{A}^{\#}\mathcal{P}^{\#}_{\Omega}(\mathcal{P}_{\Omega}(\mathcal{A}{\mathbf{\dot{X}}}_{k+1}-\mathbf{\dot{Y}}))-\beta\mathcal{C}^{\#}(\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta)-L_{k,1}\Delta{\mathbf{\dot{X}}}_{k+1}\in\partial f({\mathbf{\dot{X}}}_{k+1}),$
 which yields
-$\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k+1}}=-\mathcal{C}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}-L_{1,k}\Delta{\mathbf{\dot{X}}}_{k+1}+\beta\mathcal{C}^{\#}\mathcal{B}\Delta\mathbf{\dot{W}}_{k+1}\in\partial_{{\mathbf{\dot{X}}}}\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1}).$ (39)
+$\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k+1}}=-\mathcal{C}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}-L_{k,1}\Delta{\mathbf{\dot{X}}}_{k+1}+\beta\mathcal{C}^{\#}\mathcal{B}\Delta\mathbf{\dot{W}}_{k+1}\in\partial_{{\mathbf{\dot{X}}}}\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1}).$ (45)
 According to the optimality condition of the update of $\mathbf{\dot{W}}_{k+1},$ the following equality holds:
-$\nabla g(\mathbf{\dot{W}}_{k})+\beta\mathcal{B}^{\#}(\mathcal{C}{\mathbf{\dot{X}}}_{k+1}-\mathcal{B}\mathbf{\dot{W}}_{k+1}-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta)+L_{2,k}\Delta\mathbf{\dot{W}}_{k+1}=0.$ (40)
+$\nabla g(\mathbf{\dot{W}}_{k})+\beta\mathcal{B}^{\#}(\mathcal{C}{\mathbf{\dot{X}}}_{k+1}-\mathcal{B}\mathbf{\dot{W}}_{k+1}-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta)+L_{k,2}\Delta\mathbf{\dot{W}}_{k+1}=0.$ (46)
 It follows from (46) and the update of $\mathbf{\dot{\Lambda}}_{k+1}$ that
-$\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k+1}}$ $=\nabla g(\mathbf{\dot{W}}_{k+1})-\nabla g(\mathbf{\dot{W}}_{k})+\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}-L_{2,k}\Delta\mathbf{\dot{W}}_{k+1}\in\partial_{\mathbf{\dot{W}}}\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1}).$ (41)
+$\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k+1}}$ $=\nabla g(\mathbf{\dot{W}}_{k+1})-\nabla g(\mathbf{\dot{W}}_{k})+\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}-L_{k,2}\Delta\mathbf{\dot{W}}_{k+1}\in\partial_{\mathbf{\dot{W}}}\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1}).$ (47)
 $\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k+1}}$ $=\frac{1}{\beta\mu}\Delta\mathbf{\dot{\Lambda}}_{k+1}\in\partial_{\mathbf{\dot{\Lambda}}}\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1}).$
@@
 There exists at least one limit point of $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ and $\omega\left(\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}\right)$ is bounded.
-According to [27, Chapter 2, Exercise 6], it follows that $\omega\left(\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}\right)$ is closed.
+According to [32, Chapter 2, Exercise 6], it follows that $\omega\left(\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}\right)$ is closed.
 Hence $\omega\left(\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}\right)$ is nonempty and compact.
@@
 Suppose that the Assumption 4.1 holds. Let $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ be a sequence generated by (59). Define
-$\mathbf{\dot{S}}_{{\mathbf{\dot{X}}}_{k}}:=\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k}},\quad\mathbf{\dot{S}}_{\mathbf{\dot{W}}_{k}}:=\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k}}+2r\theta_{0}\Delta\mathbf{\dot{W}}_{k},\quad\mathbf{\dot{S}}_{\mathbf{\dot{\Lambda}}_{k}}:=\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k}}+2r\theta_{2}\mathcal{B}\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k},$
-$\mathbf{\dot{S}}_{{\mathbf{\dot{W}}^{\prime}}_{k}}:=-2r\theta_{0}\Delta\mathbf{\dot{W}}_{k},\quad\mathbf{\dot{S}}_{{\mathbf{\dot{\Lambda}}^{\prime}}_{k}}:=-2\theta_{2}\mathcal{B}\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k}$
+$\mathbf{\dot{S}}_{{\mathbf{\dot{X}}}_{k}}:=\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k}},\quad\mathbf{\dot{S}}_{\mathbf{\dot{W}}_{k}}:=\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k}}+2r\varsigma_{0}\Delta\mathbf{\dot{W}}_{k},\quad\mathbf{\dot{S}}_{\mathbf{\dot{\Lambda}}_{k}}:=\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k}}+2r\varsigma_{2}\mathcal{B}\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k},$
+$\mathbf{\dot{S}}_{{\mathbf{\dot{W}}^{\prime}}_{k}}:=-2r\varsigma_{0}\Delta\mathbf{\dot{W}}_{k},\quad\mathbf{\dot{S}}_{{\mathbf{\dot{\Lambda}}^{\prime}}_{k}}:=-2\varsigma_{2}\mathcal{B}\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k}$
 where $(\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k}},\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k}},\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k}})\in\partial\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})$.
@@
 where
-$\tilde{\pi}=\sqrt{3}\pi+4r\max\{\theta_{0},\theta_{2}\|\mathcal{B}\|^{2}\},$ (44)
+$\tilde{\pi}=\sqrt{3}\pi+4r\max\{\varsigma_{0},\varsigma_{2}\|\mathcal{B}\|^{2}\},$ (50)
 $r>1$, $\pi$ is given in (44).
@@
 $\mathbf{\dot{S}}_{{\mathbf{\dot{X}}}_{k}}$ ::=:= $\partial_{\mathbf{\dot{X}}}\mathcal{R}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},\mathbf{\dot{\Lambda}}_{k-1})=\partial_{\mathbf{\dot{X}}}\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})=\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k}},$
-$\mathbf{\dot{S}}_{\mathbf{\dot{W}}_{k}}$ ::=:= $\nabla_{\mathbf{\dot{W}}}\mathcal{R}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},\mathbf{\dot{\Lambda}}_{k-1})=\nabla_{\mathbf{\dot{W}}}\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})+2r\theta_{0}\Delta\mathbf{\dot{W}}_{k}=\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k}}+2r\theta_{0}\Delta\mathbf{\dot{W}}_{k},$
-$\mathbf{\dot{S}}_{\mathbf{\dot{\Lambda}}_{k}}$ ::=:= $\nabla_{\mathbf{\dot{\Lambda}}}\mathcal{R}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},\mathbf{\dot{\Lambda}}_{k-1})=\nabla_{\mathbf{\dot{\Lambda}}}\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})+2r\theta_{2}\mathcal{B}\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k}=\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k}}+2r\theta_{2}\mathcal{B}\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k},$
-$\mathbf{\dot{S}}_{{\mathbf{\dot{W}}^{\prime}}_{k}}$ ::=:= $\nabla_{\mathbf{\dot{W}}^{\prime}}\mathcal{R}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},\mathbf{\dot{\Lambda}}_{k-1})=-2r\theta_{0}\Delta\mathbf{\dot{W}}_{k},$
-$\mathbf{\dot{S}}_{{\mathbf{\dot{\Lambda}}^{\prime}}_{k}}$ ::=:= $\nabla_{\mathbf{\dot{\Lambda}}^{\prime}}\mathcal{R}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},\mathbf{\dot{\Lambda}}_{k-1})=-2r\theta_{2}\mathcal{B}\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k}.$
+$\mathbf{\dot{S}}_{\mathbf{\dot{W}}_{k}}$ ::=:= $\nabla_{\mathbf{\dot{W}}}\mathcal{R}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},\mathbf{\dot{\Lambda}}_{k-1})=\nabla_{\mathbf{\dot{W}}}\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})+2r\varsigma_{0}\Delta\mathbf{\dot{W}}_{k}=\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k}}+2r\varsigma_{0}\Delta\mathbf{\dot{W}}_{k},$
+$\mathbf{\dot{S}}_{\mathbf{\dot{\Lambda}}_{k}}$ ::=:= $\nabla_{\mathbf{\dot{\Lambda}}}\mathcal{R}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},\mathbf{\dot{\Lambda}}_{k-1})=\nabla_{\mathbf{\dot{\Lambda}}}\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})+2r\varsigma_{2}\mathcal{B}\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k}=\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k}}+2r\varsigma_{2}\mathcal{B}\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k},$
+$\mathbf{\dot{S}}_{{\mathbf{\dot{W}}^{\prime}}_{k}}$ ::=:= $\nabla_{\mathbf{\dot{W}}^{\prime}}\mathcal{R}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},\mathbf{\dot{\Lambda}}_{k-1})=-2r\varsigma_{0}\Delta\mathbf{\dot{W}}_{k},$
+$\mathbf{\dot{S}}_{{\mathbf{\dot{\Lambda}}^{\prime}}_{k}}$ ::=:= $\nabla_{\mathbf{\dot{\Lambda}}^{\prime}}\mathcal{R}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k},\mathbf{\dot{W}}_{k-1},\mathbf{\dot{\Lambda}}_{k-1})=-2r\varsigma_{2}\mathcal{B}\mathcal{B}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k}.$
 By the triangle inequality, we obtain
-$\|\mathbf{\dot{S}}_{{\mathbf{\dot{X}}}_{k}}\|_{\mathrm{F}}=\|\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k}}\|_{\mathrm{F}},\quad\|\mathbf{\dot{S}}_{\mathbf{\dot{W}}_{k}}\|_{\mathrm{F}}\leq\|\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k}}\|_{\mathrm{F}}+2r\theta_{0}\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}},$
-$\|\mathbf{\dot{S}}_{\mathbf{\dot{\Lambda}}_{k}}\|_{\mathrm{F}}\leq\|\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k}}\|_{\mathrm{F}}+2r\theta_{2}\|\mathcal{B}\|^{2}\|\Delta\mathbf{\dot{\Lambda}}_{k}\|_{\mathrm{F}},\quad\|\mathbf{\dot{S}}_{{\mathbf{\dot{W}}^{\prime}}_{k}}\|_{\mathrm{F}}=2r\theta_{0}\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}},\quad\|\mathbf{\dot{S}}_{{\mathbf{\dot{\Lambda}}^{\prime}}_{k}}\|_{\mathrm{F}}=2r\theta_{2}\|\mathcal{B}\|^{2}\|\Delta\mathbf{\dot{\Lambda}}_{k}\|_{\mathrm{F}}.$
+$\|\mathbf{\dot{S}}_{{\mathbf{\dot{X}}}_{k}}\|_{\mathrm{F}}=\|\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k}}\|_{\mathrm{F}},\quad\|\mathbf{\dot{S}}_{\mathbf{\dot{W}}_{k}}\|_{\mathrm{F}}\leq\|\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k}}\|_{\mathrm{F}}+2r\varsigma_{0}\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}},$
+$\|\mathbf{\dot{S}}_{\mathbf{\dot{\Lambda}}_{k}}\|_{\mathrm{F}}\leq\|\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k}}\|_{\mathrm{F}}+2r\varsigma_{2}\|\mathcal{B}\|^{2}\|\Delta\mathbf{\dot{\Lambda}}_{k}\|_{\mathrm{F}},\quad\|\mathbf{\dot{S}}_{{\mathbf{\dot{W}}^{\prime}}_{k}}\|_{\mathrm{F}}=2r\varsigma_{0}\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}},\quad\|\mathbf{\dot{S}}_{{\mathbf{\dot{\Lambda}}^{\prime}}_{k}}\|_{\mathrm{F}}=2r\varsigma_{2}\|\mathcal{B}\|^{2}\|\Delta\mathbf{\dot{\Lambda}}_{k}\|_{\mathrm{F}}.$
 By Lemma 4.6, it follows that
 $|||\widetilde{\mathbf{\dot{S}}}_{k}|||$$\leq$ $\|\mathbf{\dot{S}}_{{\mathbf{\dot{X}}}_{k}}\|_{\mathrm{F}}+\|\mathbf{\dot{S}}_{\mathbf{\dot{W}}_{k}}\|_{\mathrm{F}}+\|\mathbf{\dot{S}}_{\mathbf{\dot{\Lambda}}_{k}}\|_{\mathrm{F}}+\|\mathbf{\dot{S}}_{{\mathbf{\dot{W}}^{\prime}}_{k}}\|_{\mathrm{F}}+\|\mathbf{\dot{S}}_{{\mathbf{\dot{\Lambda}}^{\prime}}_{k}}\|_{\mathrm{F}}$
-$\leq$ $\|\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k}}\|_{\mathrm{F}}+\|\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k}}\|_{\mathrm{F}}+\|\mathbf{\dot{D}}^{k}_{\mathbf{\dot{\Lambda}}}\|_{\mathrm{F}}+4r\theta_{0}\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}+4r\theta_{2}\|\mathcal{B}\|^{2}\|\Delta\mathbf{\dot{\Lambda}}_{k}\|_{\mathrm{F}}$
-$\leq$ $\sqrt{3}|||\mathbf{\dot{D}}_{k}|||+4r\theta_{0}\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}+4r\theta_{2}\|\mathcal{B}\|^{2}\|\Delta\mathbf{\dot{\Lambda}}_{k}\|_{\mathrm{F}}$
-$\leq$ $\sqrt{3}\rho\|\Delta{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}+(\sqrt{3}\rho+4r\theta_{0})\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}+(\sqrt{3}\rho+4r\theta_{2}\|\mathcal{B}\|^{2})\|\Delta\mathbf{\dot{\Lambda}}_{k}\|_{\mathrm{F}},$
+$\leq$ $\|\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k}}\|_{\mathrm{F}}+\|\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k}}\|_{\mathrm{F}}+\|\mathbf{\dot{D}}^{k}_{\mathbf{\dot{\Lambda}}}\|_{\mathrm{F}}+4r\varsigma_{0}\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}+4r\varsigma_{2}\|\mathcal{B}\|^{2}\|\Delta\mathbf{\dot{\Lambda}}_{k}\|_{\mathrm{F}}$
+$\leq$ $\sqrt{3}|||\mathbf{\dot{D}}_{k}|||+4r\varsigma_{0}\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}+4r\varsigma_{2}\|\mathcal{B}\|^{2}\|\Delta\mathbf{\dot{\Lambda}}_{k}\|_{\mathrm{F}}$
+$\leq$ $\sqrt{3}\rho\|\Delta{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}+(\sqrt{3}\rho+4r\varsigma_{0})\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}+(\sqrt{3}\rho+4r\varsigma_{2}\|\mathcal{B}\|^{2})\|\Delta\mathbf{\dot{\Lambda}}_{k}\|_{\mathrm{F}},$
 which completes the proof. ∎
@@
 The results follow immediately from Theorem 4.1, 4.2 and Lemma 4.9. ∎
-The following convergence result of the whole sequence, established under the Kurdyka-Łojasiewicz property, follows from [36, Theorem 2].
+The following convergence result of the whole sequence, established under the Kurdyka-Łojasiewicz(KL) property, follows from [43, Theorem 2].
 The difference lies in the construction of $\mathcal{R}_{k}$.
 The proofs are given for completeness.
+Furthermore, the rationality of the KL property (52) assumption for Model (2) is clarified in Corollary 4.1.
 Theorem 4.4 (Global Convergence).
@@
 Corollary 4.1.
-Suppose the penalty parameter $\beta$ is large enough, and $L_{1,k},L_{2,k}$ are set as given constants for all $k$ such that (20) and Assumption (A5) hold, then the sequence $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ generated by Algorithm 1 converges to a stationary point of SLRQA (2).
+Suppose the penalty parameter $\beta$ is large enough, and $L_{k,1},L_{k,2}$ are set as given constants for all $k$ such that (26) and Assumption (A5) hold, then the sequence $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ generated by Algorithm 1 converges to a stationary point of SLRQA (2).
 Proof.
@@
 Since (26) and Assumption (A5) hold, Assumption A4, A5 are also satisfied.
-Furthermore, the Huber function and Frobenius norm are semi-algebraic [3, Example 2].
-From [28, Appendix, Lemma 4], it follows that the nuclear norm is also a semi-algebraic function.
-Since $\phi$ is a semi-algebraic function and the fact that the composition and sum of semi-algebraic functions are also semi-algebraic, we can conclude that $\mathcal{R}$ is semi-algebraic and the KL-inequality 46 holds since any proper closed semi-algebraic function satisfies the KL-inequality [3, Theorem 3].
+Furthermore, the Huber function and Frobenius norm are semi-algebraic [5, Example 2].
+From [34, Appendix, Lemma 4], it follows that the nuclear norm is also a semi-algebraic function.
+Since $\phi$ is a semi-algebraic function and the fact that the composition and sum of semi-algebraic functions are also semi-algebraic, we can conclude that $\mathcal{R}$ is semi-algebraic and the KL-inequality 52 holds since any proper closed semi-algebraic function satisfies the KL-inequality [5, Theorem 3].
 It follows from Theorem 4.4 that the sequence $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\}_{k\geq 0}$ generated by Algorithm 1 converges to a stationary point of SLRQA (2), which completes the proof.
 ∎
-5 An Algorithm for SLRQA-NF model
-In this section, we propose the PL-ADMM-NF algorithm to solve SLRQA-NF (4) which is summarized in Algorithm 3.
-It is noted that SLRQA-NF is a model of the limiting case of SLRQA.
-The linearization technique is applied to the update of $\mathbf{\dot{W}}$.
-Each iteration of PL-ADMM-NF consists of minimizing the augmented Lagrange function with respect to $\mathbf{\dot{W}}$ variable after linearization, minimizing the augmented Lagrange function with respect to ${\mathbf{\dot{X}}}$ variable, and updating the multiplier.
-Algorithm 3 PL-ADMM-NF algorithm for SLRQA-NF in (4)
-1:Corrupted images $\mathbf{\dot{Y}}$; two parameters $\mu\in(0,2)$ and $\beta>0$; a tolerance $\eta>0$; a sequence of regularization parameters $\{L_{1,k}\}$ satisfying $\sup_{k\geq 0}{L_{1,k}}<\infty$ and $\inf_{k>0}{L_{1,k}}>0$.
-2:Recovered $\mathbf{\dot{Z}}_{*}$.
-3:Set $\mathbf{\dot{Z}}_{0}=\bm{0}$, $\mathbf{\dot{W}}_{0}=\bm{0}$, $\mathbf{\dot{\Lambda}}_{1,0}=\bm{0}$, $\mathbf{\dot{\Lambda}}_{2,0}=\bm{0}$.
-4:for $k=0,1,2,\ldots$ do
-5:     $\mathbf{\dot{Z}}$ subproblem:
-$\mathbf{\dot{Z}}_{k+1}\in\mathop{\mathrm{arg\,min}}_{\mathbf{\dot{Z}}}\sum_{i}\phi(\sigma_{i}(\mathbf{\dot{Z}}),\gamma)+\left\langle\mathbf{\dot{\Lambda}}_{1,k},\mathbf{\dot{Z}}-\mathbf{\dot{W}}_{k}\right\rangle+\frac{\beta}{2}\|\mathbf{\dot{Z}}-\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}^{2}+\frac{L_{1,k}}{2}\|\mathbf{\dot{Z}}-\mathbf{\dot{Z}}_{k}\|_{\mathrm{F}}^{2};$
-6:     $\mathbf{\dot{W}}$ subproblem:
-$\mathbf{\dot{W}}_{k+1}$ $=\mathop{\mathrm{arg\,min}}_{\mathbf{\dot{W}}}\lambda\left\langle\nabla p(\mathbf{\dot{W}}_{k}),\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\right\rangle+\beta\left\langle\mathbf{\dot{W}}_{k}-\mathbf{\dot{Z}}_{k+1}+\frac{1}{\beta}\mathbf{\dot{\Lambda}}_{1,k},\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\right\rangle$
-$+\beta\left\langle\mathcal{W}\mathcal{A}^{\#}(\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{Y}})+\frac{1}{\beta}\mathbf{\dot{\Lambda}}_{2,k},\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\right\rangle+\frac{\beta\eta}{2}\|\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}^{2};$
-7:     $\mathbf{\dot{\Lambda}}_{1,k+1}=\mathbf{\dot{\Lambda}}_{1,k}+\mu\beta\cdot(\mathbf{\dot{Z}}_{k+1}-\mathbf{\dot{W}}_{k+1})$;
-8:     $\mathbf{\dot{\Lambda}}_{2,k+1}=\mathbf{\dot{\Lambda}}_{2,k}+\mu\beta\cdot\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{Y}})$;
-9:     if $\|\mathbf{\dot{Z}}_{k+1}-\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+\|\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}_{k+1}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}<\eta$ then
-10:         $\mathbf{\dot{Z}}_{*}=\mathbf{\dot{Z}}_{k+1}$; break
-11:     end if
-12:end for
-The augmented Lagrangian function of SLRQA-NF (4) is given by:
-~$\tilde{\mathcal{L}}_{\beta}(\mathbf{\dot{Z}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}}_{1},\mathbf{\dot{\Lambda}}_{2})=\sum_{i}\phi(\sigma_{i}(\mathbf{\dot{Z}}),\gamma)+\lambda p(\mathbf{\dot{W}})+\left\langle\mathbf{\dot{\Lambda}}_{1},\mathbf{\dot{Z}}-\mathbf{\dot{W}}\right\rangle$ (53)
-$+\frac{\beta}{2}\|\mathbf{\dot{Z}}-\mathbf{\dot{W}}\|^{2}_{\mathrm{F}}+\left\langle\mathbf{\dot{\Lambda}}_{2},\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\right\rangle+\frac{\beta}{2}\|\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2}$
-$=\sum_{i}\phi(\sigma_{i}(\mathbf{\dot{Z}}),\gamma)+\lambda p(\mathbf{\dot{W}})+\frac{\beta}{2}\|\mathbf{\dot{Z}}-\mathbf{\dot{W}}+\frac{1}{\beta}\mathbf{\dot{\Lambda}}_{1}\|_{\mathrm{F}}^{2}$
-$+\frac{\beta}{2}\|\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})+\frac{1}{\beta}\mathbf{\dot{\Lambda}}_{2}\|_{\mathrm{F}}^{2}-\frac{1}{2\beta}\|\mathbf{\dot{\Lambda}}_{1}\|_{\mathrm{F}}^{2}-\frac{1}{2\beta}\|\mathbf{\dot{\Lambda}}_{2}\|^{2}_{\mathrm{F}},$
-where $\beta>0$ is the penalty parameter, $\mathbf{\dot{\Lambda}}_{1},\mathbf{\dot{\Lambda}}_{2}\in\mathbb{H}^{m\times n}$ are the Lagrange multipliers.
-In Algorithm 3, the subproblem of $\mathbf{\dot{Z}}$ in Step 5 is the proximal ADMM step i.e.
-the minimization of ~$\tilde{\mathcal{L}}_{\beta}(\mathbf{\dot{Z}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}}_{1},\mathbf{\dot{\Lambda}}_{2})$ for fixed $\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{1,k},\mathbf{\dot{\Lambda}}_{2,k}$ plus a proximal term $\frac{L_{1,k}}{2}\|\mathbf{\dot{Z}}-\mathbf{\dot{Z}}_{k}\|_{\mathrm{F}}^{2}.$ Since the subproblem is equivalent to
-$\min_{\mathbf{\dot{Z}}}\,\sum_{i}\phi(\sigma_{i}(\mathbf{\dot{Z}}),\gamma)+\frac{\beta+L_{1,k}}{2}\left\|\mathbf{\dot{Z}}-\frac{1}{\beta+L_{1,k}}\left(\beta\mathbf{\dot{W}}_{k}+\mathbf{\dot{\Lambda}}_{1,k}+L_{1,k}\mathbf{\dot{Z}}_{k}\right)\right\|_{\mathrm{F}}^{2},$
-Algorithm 2 can also be used to obtain $\mathbf{\dot{Z}}_{k+1}$.
-In Step 6 of Algorithm 3, $\mathbf{\dot{W}}$, according to the linearization strategy in [21], a proximal gradient method is used in Step 6 to linearize the second order term.
-The penalty parameter of the second order term is denoted by $\beta\eta.$ To be specific, the quadratic term $\frac{\beta}{2}\|\mathbf{\dot{W}}-\mathbf{\dot{Z}}-\frac{1}{\beta}\mathbf{\dot{\Lambda}}_{1}\|_{\mathrm{F}}^{2}+\frac{\beta}{2}\|\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2}$ in ~$\tilde{\mathcal{L}}_{\beta}(\mathbf{\dot{Z}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}}_{1},\mathbf{\dot{\Lambda}}_{2})$ is replaced by the linearized term plus a regularized term $\beta\left\langle\mathbf{\dot{W}}_{k}-\mathbf{\dot{Z}}_{k+1}-\frac{1}{\beta}\mathbf{\dot{\Lambda}}_{1,k},\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\right\rangle+\beta\left\langle\mathcal{W}\mathcal{A}^{\#}(\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{Y}})+\frac{1}{\beta}\mathbf{\dot{\Lambda}}_{2,k},\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\right\rangle$.
-Hence, the update of $\mathbf{\dot{W}}_{k+1}$ is given by
-$\mathbf{\dot{W}}_{k+1}=\mathbf{\dot{W}}_{k}-\frac{\mathcal{W}\mathcal{A}^{\#}(\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{Y}})+\frac{1}{\beta}\mathbf{\dot{\Lambda}}_{2,k})+\mathbf{\dot{W}}_{k}-\mathbf{\dot{Z}}_{k+1}-\frac{1}{\beta}\mathbf{\dot{\Lambda}}_{1,k}}{\eta}-\frac{\lambda}{\beta\eta}\nabla p(\mathbf{\dot{W}}_{k}).$ (54)
-Step 7 and 8 is used to update the Lagrange multipliers. Finally, the recovered image ${\mathbf{\dot{X}}}_{*}$ is given by ${\mathbf{\dot{X}}}_{*}=\mathcal{W}^{\#}(\mathbf{\dot{Z}}_{*})$.
+4.2 Global Convergence for Algorithm 3
+The global convergence is established for a generalization of SLRQA-NF, i.e. Model (11). The augmented Lagrangian function of (11) is
+$\mathcal{L}_{\beta_{1},\beta_{2}}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}}_{1},\mathbf{\dot{\Lambda}}_{2})=f({\mathbf{\dot{X}}})+g(\mathbf{\dot{W}})+\left\langle\mathbf{\dot{\Lambda}}_{1},\mathcal{C}_{1}({\mathbf{\dot{X}}})+\mathcal{B}_{1}(\mathbf{\dot{W}})-\mathbf{\dot{B}}_{1}\right\rangle+\left\langle\mathbf{\dot{\Lambda}}_{2},\mathcal{B}_{2}(\mathbf{\dot{W}})-\mathbf{\dot{B}}_{2}\right\rangle$
+$\qquad\qquad\qquad\qquad\qquad\qquad\quad+\frac{\beta_{1}}{2}\|\mathcal{C}_{1}({\mathbf{\dot{X}}})+\mathcal{B}_{1}(\mathbf{\dot{W}})-\mathbf{\dot{B}}_{1}\|^{2}_{\mathrm{F}}+\frac{\beta_{2}}{2}\|\mathcal{B}_{2}(\mathbf{\dot{W}})-\mathbf{\dot{B}}_{2}\|_{\mathrm{F}}^{2}$
+$\qquad\qquad=f({\mathbf{\dot{X}}})+g(\mathbf{\dot{W}})+\frac{\beta_{1}}{2}\left\|\mathcal{C}_{1}({\mathbf{\dot{X}}})+\mathcal{B}_{1}(\mathbf{\dot{W}})-\mathbf{\dot{B}}_{1}+\frac{\mathbf{\dot{\Lambda}}_{1}}{\beta_{1}}\right\|_{\mathrm{F}}^{2}+\frac{\beta_{2}}{2}\left\|\mathcal{B}_{2}(\mathbf{\dot{W}})-\mathbf{\dot{B}}_{2}+\frac{\mathbf{\dot{\Lambda}}_{2}}{\beta_{2}}\right\|^{2}$
+$\qquad\qquad\qquad\quad-\frac{1}{2\beta_{1}}\left\|\mathbf{\dot{\Lambda}}_{1}\right\|_{\mathrm{F}}^{2}-\frac{1}{2\beta_{2}}\left\|\mathbf{\dot{\Lambda}}_{2}\right\|_{\mathrm{F}}^{2}.$
+The iteration scheme for the proposed PL-ADMM-NF is
+${\mathbf{\dot{X}}}_{k+1}$ $\in\mathop{\mathrm{arg\,min}}_{{\mathbf{\dot{X}}}}f({\mathbf{\dot{X}}})+\frac{\beta_{1}}{2}\left\|\mathcal{C}_{1}({\mathbf{\dot{X}}})+\mathcal{B}_{1}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}_{1}+\frac{\mathbf{\dot{\Lambda}}_{k,1}}{\beta_{1}}\right\|_{\mathrm{F}}^{2}+\frac{L_{k,1}}{2}\|{\mathbf{\dot{X}}}-{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}^{2},$ (59)
+$\mathbf{\dot{W}}_{k+1}$ $\in\mathop{\mathrm{arg\,min}}_{\mathbf{\dot{W}}}\left\langle\nabla g(\mathbf{\dot{W}}_{k}),\mathbf{\dot{W}}\right\rangle+\frac{\beta_{1}}{2}\left\|\mathcal{C}_{1}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}_{1}(\mathbf{\dot{W}})-\mathbf{\dot{B}}_{1}+\frac{\mathbf{\dot{\Lambda}}_{k,1}}{\beta_{1}}\right\|_{\mathrm{F}}^{2}$
+$\qquad\qquad\qquad+\frac{\beta_{2}}{2}\|\mathcal{B}_{2}(\mathbf{\dot{W}})-\mathbf{\dot{B}}_{2}+\frac{\mathbf{\dot{\Lambda}}_{k,2}}{\beta_{2}}\|_{\mathrm{F}}^{2}+\frac{L_{k,2}}{2}\left\|\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\right\|^{2}_{\mathrm{F}},$
+$\mathbf{\dot{\Lambda}}_{k+1,1}$ $=\mathbf{\dot{\Lambda}}_{k,1}+\mu\beta_{1}(\mathcal{C}_{1}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}_{1}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}_{1}),$
+$\mathbf{\dot{\Lambda}}_{k+1,2}$ $=\mathbf{\dot{\Lambda}}_{k,2}+\mu\beta_{2}(\mathcal{B}_{2}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}_{2}).$
+The following assumptions are made for the convergence of PL-ADMM-NF.
+Assumption 4.2.
+  1. B1
+    $f({\mathbf{\dot{X}}})$ and $g(\mathbf{\dot{W}})$ are bounded from below and have $L_{f}$ and $L_{g}$ Lipschitz continuous gradients respectively, i.e., for every ${\mathbf{\dot{X}}}_{1},{\mathbf{\dot{X}}}_{2},\mathbf{\dot{W}}_{1}$, and $\mathbf{\dot{W}}_{2}$, it holds that
+    $\|\nabla g(\mathbf{\dot{W}}_{1})-\nabla g(\mathbf{\dot{W}}_{2})\|_{\mathrm{F}}\leq L_{g}\|\mathbf{\dot{W}}_{1}-\mathbf{\dot{W}}_{2}\|_{\mathrm{F}},$ (60)
+    $\|\nabla f({\mathbf{\dot{X}}}_{1})-\nabla f({\mathbf{\dot{X}}}_{2})\|_{\mathrm{F}}\leq L_{f}\|{\mathbf{\dot{X}}}_{1}-{\mathbf{\dot{X}}}_{2}\|_{\mathrm{F}}.$
+  2. B2
+    The matrix $\mathcal{C}_{1}^{\#}\mathcal{C}_{1}$ is full rank or $f({\mathbf{\dot{X}}})$ is coercive, $\mathcal{B}_{2}^{\#}\mathcal{B}_{2}$ is full rank, ${\rm range}(\mathcal{B}_{1})\subseteq{\rm range}(\mathcal{C}_{1})$, $\mathbf{\dot{B}}_{1}\in{\rm range}(\mathcal{C}_{1})$, and $\mathbf{\dot{B}}_{2}\in{\rm range}(\mathcal{B}_{2})$.
+  3. B3
+    The parameters $\beta_{1},\beta_{2}>0$, $\mu\in(0,2)$, and there exist three constants $a_{1}>0,a_{2}>0$, and $r>1$ such that
+    $(q_{1}^{-}-3r\theta_{1})\mathbf{I}\succeq a_{1}\mathbf{I}\,\quad\mbox{and}\,\quad q_{2}^{-}\mathbf{I}+\beta_{2}\mathcal{B}_{2}^{\#}\mathcal{B}_{2}+\beta_{1}\mathcal{B}_{1}^{\#}\mathcal{B}_{1}-(3r\theta_{2}+L_{g})\mathbf{I}\succeq a_{2}\mathbf{I},$ (61)
+    where $\mathbf{I}$ is the identity matrix with correct size and $\theta_{1}=\theta_{1,1}+\theta_{2,1},\theta_{2}=\theta_{1,2}+\theta_{2,2}$,
+    $\theta_{1,1}=\frac{4(q_{1}+L_{f})^{2}\mu}{\rho(\mu)^{2}\beta_{1}\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}},\quad\theta_{1,2}=\frac{4\beta_{1}\|\mathcal{B}_{1}\|^{2}\|\mathcal{C}_{1}\|^{2}}{\rho(\mu)\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}},$ (62)
+    $\theta_{2,1}=\frac{3\mu\|\mathcal{B}_{1}\|^{2}(4q_{1}^{2}+|1-\mu|^{2}(4(q_{1}+L_{f})^{2}))}{\beta_{2}\rho(\mu)^{4}\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}\lambda_{+}^{\mathcal{B}_{2}^{\#}\mathcal{B}_{2}}},$
+    $\theta_{2,2}=\frac{6(L_{g}+q_{2})^{2}\mu}{\beta_{2}\rho(\mu)^{2}\lambda_{+}^{\mathcal{B}_{2}^{\#}\mathcal{B}_{2}}}+\frac{3\mu\|\mathcal{B}_{1}\|^{2}(|1-\mu|^{2}+1)(4\beta_{1}^{2}\|\mathcal{B}_{1}\|^{2}\|\mathcal{C}_{1}\|^{2})}{\beta_{2}\rho(\mu)^{4}\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}\lambda_{+}^{\mathcal{B}_{2}^{\#}\mathcal{B}_{2}}}.$
+  4. B4
+    The parameters $\beta_{1},\beta_{2}$ satisfies $\frac{\kappa}{2L_{f}}>\theta_{3,0}+\theta_{4,0},\frac{1}{2L_{g}}>\theta_{4,0},$ where $\kappa\in(0,1)$ is a given constant and
+    $\theta_{3,0}=\frac{3}{2\beta_{1}\rho(\mu)\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}},\quad\theta_{4,0}=\max\left\{\frac{3}{2\beta_{2}\rho(\mu)\lambda_{+}^{\mathcal{B}_{2}^{\#}\mathcal{B}_{2}}},\frac{3\|\mathcal{B}_{1}\|^{2}}{2\beta_{2}\rho(\mu)^{2}\lambda_{+}^{\mathcal{B}_{2}^{\#}\mathcal{B}_{2}}\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}}\right\}.$ (63)
 Remark 4.1.
-{ The combination of two constraints in (4) with respect to $\mathbf{\dot{W}}$ and $\mathbf{\dot{Z}}$ can be represented by $(\mathbf{0};\mathbf{I})$ and $(\mathbf{T};\mathbf{I})$, where $\mathbf{T}$ is the matrix form of $\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#})$.
-It follows that Assumption A3 is not satisfied.
-However, this assumption is essential in all known proofs for nonconvex ADMM and its variants [29, 36, 22] for the global convergence.
+For (B1), the differentiability of $f({\mathbf{\dot{X}}})$ and Lipschitz continuity of $\nabla f({\mathbf{\dot{X}}})$ in SLRQA-NF (4) comes from the differentiability of $\phi$ and Lipschitz differentiability of $\nabla\phi$.
+For more discussions on the Lipschitz continuity for spectral function, we refer the readers to [23, Corollary 2.5] and [13, Theorem 3.3].
+It is straightforward to verify that (B2) holds for SLRQA-NF (4).
+As mentioned in Section 1.2, the combination of two constraints in (4) with respect to $\mathbf{\dot{W}}$ and $\mathbf{\dot{Z}}$ can be represented by $(\mathbf{I};\mathbf{0})$ and $(\mathbf{I};\mathbf{T})$, where $\mathbf{T}$ is the matrix form of $\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#})$.
+In this case, A3 in Assumption 4.1 does not hold for SLRQA-NF.
+However, this assumption is required for all known proofs for the global convergence of nonconvex ADMM and its variants [35, 43, 27].
 Hence, existing techniques have difficulties proving the convergence of Algorithm 3.
-It is challenging to prove the global convergence result of variants of nonconvex ADMM without Assumption A3.
-We leave it as future works.}
+For (B3) and (B4), if $q_{1}^{-}$ is large enough, then the first condition in (61) is satisfied.
+Since $\mathcal{B}_{2}^{\#}\mathcal{B}_{2}$ is full rank, if $\beta_{1}$ and $\beta_{2}$ is large enough such that
+$\beta_{1}>\max\left\{\frac{3L_{f}}{\kappa\rho(\mu)\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}},1\right\}+\max\left\{\frac{3L_{f}}{\rho(\mu)\lambda_{+}^{\mathcal{B}_{2}^{\#}\mathcal{B}_{2}}},\frac{3L_{f}\|\mathcal{B}_{1}\|^{2}}{\rho(\mu)^{2}\lambda_{+}^{\mathcal{B}_{2}^{\#}\mathcal{B}_{2}}\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}}\right\},$ (64)
+$\beta_{2}>\max\left\{\frac{3L_{g}}{\rho(\mu)(\lambda_{+}^{\mathcal{B}_{2}^{\#}\mathcal{B}_{2}})^{2}},\frac{3L_{g}\|\mathcal{B}_{1}\|^{2}}{\rho(\mu)^{2}(\lambda_{+}^{\mathcal{B}_{2}^{\#}\mathcal{B}_{2}})^{2}\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}},1\right\}+\frac{6L_{g}\mu\|\mathcal{B}_{1}\|^{2}(4q_{1}^{2}+|1-\mu|(4(q_{1}+L_{f})^{2}))}{\beta_{2}\rho(\mu)^{4}\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}(\lambda_{+}^{\mathcal{B}_{2}^{\#}\mathcal{B}_{2}})^{2}}$
+$\qquad+\frac{12L_{g}(L_{g}+q_{2})^{2}\mu}{\beta_{2}\rho(\mu)^{2}(\lambda_{+}^{\mathcal{B}_{2}^{\#}\mathcal{B}_{2}})^{2}}+\frac{6L_{g}\mu\|\mathcal{B}_{1}\|^{2}(|1-\mu|+1)(4\beta_{1}^{2}\|\mathcal{B}_{1}\|^{2}\|\mathcal{C}_{1}\|^{2})}{\beta_{2}\rho(\mu)^{4}\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}(\lambda_{+}^{\mathcal{B}_{2}^{\#}\mathcal{B}_{2}})^{2}},$
+hold, then (B3) and (B4) are satisfied.
+Lemmas 4.11 and 4.12 present the sufficient decrease property of $\mathcal{L}_{\beta_{1},\beta_{2}}$ for ${\mathbf{\dot{X}}}$ and $\mathbf{\dot{W}}$ update respectively and are used in Lemma 4.13.
+The proofs follow the same spirit in the existing analysis [43, 35].
+Lemma 4.11 (Sufficient descent of $\mathcal{L}_{\beta_{1},\beta_{2}}$ for ${\mathbf{\dot{X}}}$ update).
+The sequence $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k,1},\mathbf{\dot{\Lambda}}_{k,2})\}_{k\geq 0}$ generated by (59) satisfies:
+$\mathcal{L}_{\beta_{1},\beta_{2}}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k,1},\mathbf{\dot{\Lambda}}_{k,2})-\mathcal{L}_{\beta_{1},\beta_{2}}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k,1},\mathbf{\dot{\Lambda}}_{k,2})\geq\frac{L_{k,1}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2}_{\mathrm{F}}.$
+Proof.
+According to the ${\mathbf{\dot{X}}}_{k+1}$ update in (59), the following equality holds:
+$f({\mathbf{\dot{X}}}_{k+1})-f({\mathbf{\dot{X}}}_{k})+\frac{\beta_{1}}{2}\|\mathcal{C}_{1}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}_{1}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}_{1}+\mathbf{\dot{\Lambda}}_{k,1}/\beta_{1}\|_{\mathrm{F}}^{2}+\frac{L_{k,1}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}$ (65)
+$\leq$ $\frac{\beta_{1}}{2}\left\|\mathcal{C}_{1}({\mathbf{\dot{X}}}_{k})+\mathcal{B}_{1}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}_{1}+\mathbf{\dot{\Lambda}}_{k,1}/\beta_{1}\right\|_{\mathrm{F}}^{2}.$
+It follows that
+$\mathcal{L}_{\beta_{1},\beta_{2}}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k,1},\mathbf{\dot{\Lambda}}_{k,2})-\mathcal{L}_{\beta_{1},\beta_{2}}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k,1},\mathbf{\dot{\Lambda}}_{k,2})$
+$=$ $f({\mathbf{\dot{X}}}_{k})-f({\mathbf{\dot{X}}}_{k+1})-\frac{\beta_{1}}{2}\|\mathcal{C}_{1}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}_{1}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}_{1}+\mathbf{\dot{\Lambda}}_{k,1}/\beta_{1}\|_{\mathrm{F}}^{2}$
+$+\frac{\beta_{1}}{2}\|\mathcal{C}_{1}({\mathbf{\dot{X}}}_{k})+\mathcal{B}_{1}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}_{1}+\mathbf{\dot{\Lambda}}_{k,1}/\beta_{1}\|_{\mathrm{F}}^{2}\geq\frac{L_{k,1}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2},$
+which completes the proof. ∎
+Lemma 4.12 (Sufficient descent of $\mathcal{L}_{\beta_{1},\beta_{2}}$ for $\mathbf{\dot{W}}$ update).
+Suppose B1 in Assumption 4.2 holds. The sequence $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k,1},\mathbf{\dot{\Lambda}}_{k,2})\}_{k\geq 0}$ generated by (59) satisfies:
+$\mathcal{L}_{\beta_{1},\beta_{2}}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k,1},\mathbf{\dot{\Lambda}}_{k,2})-\mathcal{L}_{\beta_{1},\beta_{2}}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k,1},\mathbf{\dot{\Lambda}}_{k,2})\geq\|\Delta\mathbf{\dot{W}}_{k+1}\|^{2}_{B_{k}},$
+where $B_{k}:=L_{k,2}\mathbf{I}-L_{g}\mathbf{I}+\frac{\beta_{1}}{2}\mathcal{B}_{1}\mathcal{B}_{1}^{\#}+\frac{\beta_{2}}{2}\mathcal{B}_{2}\mathcal{B}_{2}^{\#}$.
+Proof.
+According to the optimality condition of $\mathbf{\dot{W}}_{k+1},$ the following equality holds:
+$\nabla g(\mathbf{\dot{W}}_{k})+\beta_{1}\mathcal{C}_{1}^{\#}(\mathcal{C}_{1}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}_{1}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}_{1}+\mathbf{\dot{\Lambda}}_{k,1}/\beta_{1})$ (66)
+$\qquad\qquad\qquad\qquad+\beta_{2}\mathcal{B}_{2}^{\#}(\mathcal{B}_{2}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}_{2}+\mathbf{\dot{\Lambda}}_{k,2}/\beta_{2})+L_{k,2}\Delta\mathbf{\dot{W}}_{k+1}=0.$
+It follows that
+$\mathcal{L}_{\beta_{1},\beta_{2}}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k,1},\mathbf{\dot{\Lambda}}_{k,2})-\mathcal{L}_{\beta_{1},\beta_{2}}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k,1},\mathbf{\dot{\Lambda}}_{k,2})$
+$=$ $\frac{\beta_{1}}{2}\|\mathcal{C}_{1}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}_{1}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}_{1}+\frac{\mathbf{\dot{\Lambda}}_{k}}{\beta_{1}}\|_{\mathrm{F}}^{2}-\frac{\beta_{1}}{2}\|\mathcal{C}_{1}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}_{1}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}_{1}+\frac{\mathbf{\dot{\Lambda}}_{k}}{\beta_{1}}\|_{\mathrm{F}}^{2}$
+$+g(\mathbf{\dot{W}}_{k})-g(\mathbf{\dot{W}}_{k+1})+\frac{\beta_{2}}{2}\left\|\mathcal{B}_{2}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}_{2}+\frac{\mathbf{\dot{\Lambda}}_{k,2}}{\beta_{2}}\right\|_{\mathrm{F}}^{2}-\frac{\beta_{2}}{2}\left\|\mathcal{B}_{2}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}_{2}+\frac{\mathbf{\dot{\Lambda}}_{k,2}}{\beta_{2}}\right\|_{\mathrm{F}}^{2}$
+$=$ $g(\mathbf{\dot{W}}_{k})-g(\mathbf{\dot{W}}_{k+1})-\beta_{1}\left\langle\mathcal{C}_{1}^{\#}(\mathcal{C}_{1}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}_{1}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}_{1}+\frac{\mathbf{\dot{\Lambda}}_{k,1}}{\beta_{1}}),\Delta\mathbf{\dot{W}}_{k+1}\right\rangle$
+$-\beta_{2}\left\langle\mathcal{B}_{2}^{\#}(\mathcal{B}_{2}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}_{2}+\frac{\mathbf{\dot{\Lambda}}_{k,2}}{\beta_{2}}),\Delta\mathbf{\dot{W}}_{k+1}\right\rangle+\frac{\beta_{1}}{2}\|\mathcal{B}_{1}(\Delta\mathbf{\dot{W}}_{k+1})\|_{\mathrm{F}}^{2}+\frac{\beta_{2}}{2}\|\mathcal{B}_{2}(\Delta\mathbf{\dot{W}}_{k+1})\|_{\mathrm{F}}^{2}$
+${=}$ $g(\mathbf{\dot{W}}_{k})-g(\mathbf{\dot{W}}_{k+1})+\left\langle\nabla g(\mathbf{\dot{W}}_{k}),\Delta\mathbf{\dot{W}}_{k+1}\right\rangle+L_{k,2}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}$
+$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+\frac{\beta_{1}}{2}\|\mathcal{B}_{1}\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}+\frac{\beta_{2}}{2}\|\mathcal{B}_{2}\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}\,$
+[…]
+which completes the proof. ∎
+Lemma 4.13 is used in Lemma 4.14 for constructing a merit function.
+Lemma 4.13.
+Suppose B1 in Assumption 4.2 holds. The sequence $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k,1},\mathbf{\dot{\Lambda}}_{k,2})\}_{k\geq 0}$ generated by (59) satisfies:
+$\mathcal{L}_{\beta_{1},\beta_{2}}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1,1},\mathbf{\dot{\Lambda}}_{k+1,2})\leq\mathcal{L}_{\beta_{1},\beta_{2}}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k,1},\mathbf{\dot{\Lambda}}_{k,2})$
+$\qquad\qquad\qquad\qquad-\|\Delta\mathbf{\dot{W}}_{k+1}\|^{2}_{B_{k}}-\frac{L_{k,1}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2}_{\mathrm{F}}+\frac{1}{\beta_{1}\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1,1}\|^{2}_{\mathrm{F}}+\frac{1}{\beta_{2}\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1,2}\|^{2}_{\mathrm{F}}.$
+Proof.
+According to the update of $\mathbf{\dot{\Lambda}}_{1}$ and $\mathbf{\dot{\Lambda}}_{2}$ in (59), Lemma 4.11, and Lemma 4.12, we have
+$\mathcal{L}_{\beta_{1},\beta_{2}}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1,1},\mathbf{\dot{\Lambda}}_{k+1,2})$
+$=$ $\mathcal{L}_{\beta_{1},\beta_{2}}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k,1},\mathbf{\dot{\Lambda}}_{k,2})+\frac{1}{\beta_{1}\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1,1}\|^{2}_{\mathrm{F}}+\frac{1}{\beta_{2}\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1,2}\|_{\mathrm{F}}^{2}$
+$\leq$ $\mathcal{L}_{\beta_{1},\beta_{2}}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k,1},\mathbf{\dot{\Lambda}}_{k,2})-\|\Delta\mathbf{\dot{W}}_{k+1}\|_{B_{k}}^{2}-\frac{L_{k,1}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}$
+$+\frac{1}{\beta_{1}\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1,1}\|^{2}_{\mathrm{F}}+\frac{1}{\beta_{2}\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1,2}\|^{2}_{\mathrm{F}}.$
+The proofs in Lemma 4.14 and Theorem 4.5 are the key differences to the existing convergence analyses of ADMM-type algorithms.
+Lemma 4.14 establishes upper bounds of $\|\Delta\mathbf{\dot{\Lambda}}_{k+1,1}\|$ and $\|\Delta\mathbf{\dot{\Lambda}}_{k+1,2}\|$ for constructing a merit function and Theorem 4.5 proves the boundedness of the sequence generated by (59).
+Lemma 4.14 and Theorem 4.5 are proven without using the widely-used “range assumption” but the newly proposed assumption in B2 of Assumption 4.2.
+Lemma 4.14.
+Suppose B1 and B2 in Assumption 4.2 hold, the sequence $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k,1},\mathbf{\dot{\Lambda}}_{k,2})\}_{k\geq 0}$ generated by (59) satisfies:
+$\frac{1}{\beta_{1}\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1,1}\|^{2}_{\mathrm{F}}\leq$ $\theta_{1,1}(\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2}_{\mathrm{F}}+\|\Delta{\mathbf{\dot{X}}}_{k}\|^{2}_{\mathrm{F}})+\theta_{1,2}(\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}+\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}^{2})$ (67)
+$+\theta_{1,3}(\|\mathcal{C}_{1}^{\#}(\Delta\mathbf{\dot{\Lambda}}_{k,1})\|^{2}_{\mathrm{F}}-\|\mathcal{C}_{1}^{\#}(\Delta\mathbf{\dot{\Lambda}}_{k+1,1})\|^{2}_{\mathrm{F}}),$
+$\frac{1}{\beta_{2}\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1,2}\|^{2}_{\mathrm{F}}\leq$ $\theta_{2,1}(\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}+\|\Delta{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}^{2}+\|\Delta{\mathbf{\dot{X}}}_{k-1}\|_{\mathrm{F}}^{2})$
+$+\theta_{2,2}(\|\Delta\mathbf{\dot{W}}_{k}$ $\|_{\mathrm{F}}^{2}+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}+\|\Delta\mathbf{\dot{W}}_{k-1}\|_{\mathrm{F}}^{2})-\theta_{2,3}(\|\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1,1}\|_{\mathrm{F}}^{2}-\|\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k,1}\|_{\mathrm{F}}^{2})$
+$-\theta_{2,3}(\|\mathcal{C}_{1}^{\#}$ $\Delta\mathbf{\dot{\Lambda}}_{k,1}\|_{\mathrm{F}}^{2}-\|\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k-1,1}\|_{\mathrm{F}}^{2})-\theta_{2,4}(\|\mathcal{B}_{2}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1,2}\|_{\mathrm{F}}^{2}-\|\mathcal{B}_{2}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k,2}\|_{\mathrm{F}}^{2}).$
+Furthermore, the following inequality holds:
+$\mathcal{L}_{\beta_{1},\beta_{2}}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1,1},\mathbf{\dot{\Lambda}}_{k+1,2})+\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2}_{\frac{L_{k,1}}{2}-r\theta_{1}\mathbf{I}}+r\theta_{1}\|\Delta{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}^{2}+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{B_{k}-r\theta_{2}\mathbf{I}}^{2}$ (68)
+$+r\theta_{2}\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}^{2}+r\theta_{3}\|\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1,1}\|^{2}_{\mathrm{F}}+r\theta_{4}\|\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k,1}\|^{2}_{\mathrm{F}}+r\theta_{5}\|\mathcal{B}_{2}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1,2}\|_{\mathrm{F}}^{2}$
+$+\frac{r-1}{\beta_{1}\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1,1}\|^{2}_{\mathrm{F}}+\frac{r-1}{\beta_{2}\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1,2}\|^{2}_{\mathrm{F}}$
+$\leq$ $\mathcal{L}_{\beta_{1},\beta_{2}}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k,1},\mathbf{\dot{\Lambda}}_{k,2})+2r\theta_{1}\|\Delta{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}^{2}+r\theta_{1}\|\Delta{\mathbf{\dot{X}}}_{k-1}\|_{\mathrm{F}}^{2}+2r\theta_{2}\|\Delta\mathbf{\dot{W}}_{k}\|^{2}_{\mathrm{F}}$
+$+r\theta_{2}\|\Delta\mathbf{\dot{W}}_{k-1}\|^{2}_{\mathrm{F}}+r\theta_{3}\|\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k,1}\|^{2}_{\mathrm{F}}+r\theta_{4}\|\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k-1,1}\|^{2}_{\mathrm{F}}+r\theta_{5}\|\mathcal{B}_{2}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k,2}\|^{2}_{\mathrm{F}},$
+where $r>1,\theta_{3}=\theta_{1,3}+\theta_{2,3},\theta_{4}=\theta_{2,4},\theta_{5}=\theta_{2,5},\theta_{1},\theta_{2}$ are defined in (62), and
+$\theta_{1,3}=\frac{|1-\mu|}{\beta_{1}\mu\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}\rho(\mu)},\theta_{2,3}=\frac{3|1-\mu|\|\mathcal{B}_{1}\|^{2}}{\beta_{2}\rho(\mu)^{3}\mu\lambda_{+}^{\mathcal{B}_{2}^{\#}\mathcal{B}_{2}}\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}},$ (69)
+$\theta_{2,4}=\frac{3|1-\mu|^{3}\|\mathcal{B}_{1}\|^{2}}{\beta_{2}\rho(\mu)^{3}\mu\lambda_{+}^{\mathcal{B}_{2}^{\#}\mathcal{B}_{2}}\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}},\theta_{2,5}=\frac{|1-\mu|}{\beta_{2}\rho(\mu)\mu\lambda_{+}^{\mathcal{B}_{2}^{\#}\mathcal{B}_{2}}}.$
+Proof.
+Let $k\geq 1$ be fixed and define the matrix
+$\mathbf{\dot{H}}_{k+1,1}:=-\nabla f({\mathbf{\dot{X}}}_{k+1})-L_{k,1}\Delta{\mathbf{\dot{X}}}_{k+1}+\beta_{1}\mathcal{C}_{1}^{\#}\mathcal{B}_{1}\Delta\mathbf{\dot{W}}_{k+1},\quad\mathbf{\dot{H}}_{k+1,2}:=-L_{k,2}\Delta\mathbf{\dot{W}}_{k+1}-\nabla g(\mathbf{\dot{W}}_{k}).$ (70)
+Hence the following inequality holds:
+$\Delta\mathbf{\dot{H}}_{k+1,1}=L_{k-1,1}\Delta{\mathbf{\dot{X}}}_{k}-L_{k,1}\Delta{\mathbf{\dot{X}}}_{k+1}+\nabla f({\mathbf{\dot{X}}}_{k})-\nabla f({\mathbf{\dot{X}}}_{k+1})+\beta_{1}\mathcal{C}_{1}^{\#}\mathcal{B}_{1}(\Delta\mathbf{\dot{W}}_{k+1}-\Delta\mathbf{\dot{W}}_{k}),$
+$\Delta\mathbf{\dot{H}}_{k+1,2}=L_{k-1,2}\Delta\mathbf{\dot{W}}_{k}-L_{k,2}\Delta\mathbf{\dot{W}}_{k+1}+\nabla g(\mathbf{\dot{W}}_{k-1})-\nabla g(\mathbf{\dot{W}}_{k}).$
+It follows from the triangle inequality that:
+$\|\Delta\mathbf{\dot{H}}_{k+1,1}\|_{\mathrm{F}}\leq(L_{k,1}+L_{f})\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}+L_{k-1,1}\|\Delta{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}+\beta_{1}\|\mathcal{B}_{1}\|\|\mathcal{C}_{1}\|(\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}),$
+$\|\Delta\mathbf{\dot{H}}_{k+1,2}\|_{\mathrm{F}}\leq\|\nabla g(\mathbf{\dot{W}}_{k-1})-\nabla g(\mathbf{\dot{W}}_{k})\|_{\mathrm{F}}+L_{k,2}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+L_{k-1,2}\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}.$
+By B1 in Assumption 4.2, $\nabla g(\mathbf{\dot{W}})$ and $\nabla f({\mathbf{\dot{X}}})$ are $L_{g}$ and $L_{f}$ Lipschitz continuous respectively.
+According to $q_{1}=\sup_{k\geq 0}L_{k,1}<\infty,q_{2}=\sup_{k\geq 0}L_{k,2}<\infty$, we have $\|\Delta\mathbf{\dot{H}}_{k+1,2}\|_{\mathrm{F}}\leq(L_{g}+q_{2})\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}+q_{2}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}.$ Hence it follows that
+$\|\Delta\mathbf{\dot{H}}_{k+1,1}\|^{2}_{\mathrm{F}}\leq 4(q_{1}+L_{f})^{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}+4q_{1}^{2}\|\Delta{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}^{2}$ (71)
+$\qquad\qquad\qquad\qquad\qquad\qquad\qquad+4\beta_{1}^{2}\|\mathcal{B}_{1}\|^{2}\|\mathcal{C}_{1}\|^{2}(\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}+\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}^{2}),$
+$\|\Delta\mathbf{\dot{H}}_{k+1,2}\|^{2}_{\mathrm{F}}\leq 2(L_{g}+q_{2})^{2}\|\Delta\mathbf{\dot{W}}_{k}\|^{2}_{\mathrm{F}}+2q_{2}^{2}\|\Delta\mathbf{\dot{W}}_{k+1}\|^{2}_{\mathrm{F}}.$
+Expressing the optimality condition of ${\mathbf{\dot{X}}}$ and $\mathbf{\dot{W}}$ subproblems using $\mathbf{\dot{H}}_{k+1,1}$ and $\mathbf{\dot{H}}_{k+1,2}$ respectively, we have
+$\mathbf{\dot{H}}_{k+1,1}=\beta_{1}\mathcal{C}_{1}^{\#}(\mathcal{C}_{1}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}_{1}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}_{1}+\mathbf{\dot{\Lambda}}_{k,1}/\beta_{1}),$
+$\mathbf{\dot{H}}_{k+1,2}=\beta_{1}\mathcal{B}_{1}^{\#}(\mathcal{C}_{1}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}_{1}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}_{1}+\mathbf{\dot{\Lambda}}_{k,1}/\beta_{1})+\beta_{2}\mathcal{B}_{2}^{\#}(\mathcal{B}_{2}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}_{2}+\mathbf{\dot{\Lambda}}_{k,2}/\beta_{2}).$
+Combining this with the $\mathbf{\dot{\Lambda}}_{1}$ and $\mathbf{\dot{\Lambda}}_{2}$ updates, it follows that
+$\mathcal{C}_{1}^{\#}\mathbf{\dot{\Lambda}}_{k+1,1}$ $=\mu\mathbf{\dot{H}}_{k+1,1}+(1-\mu)\mathcal{C}_{1}^{\#}\mathbf{\dot{\Lambda}}_{k,1},$ (72)
+$\mathcal{B}_{2}^{\#}\mathbf{\dot{\Lambda}}_{k+1,2}$ $=\mu\mathbf{\dot{H}}_{k+1,2}+(1-\mu)\mathcal{B}_{2}^{\#}\mathbf{\dot{\Lambda}}_{k,2}-\mathcal{B}_{1}^{\#}(\Delta\mathbf{\dot{\Lambda}}_{k+1,1})-\mu\mathcal{B}_{1}^{\#}(\mathbf{\dot{\Lambda}}_{k,1}).$
+Since $\mu\in(0,2),$ we have
+$\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1,1}$ $=\rho(\mu)\frac{\mu}{\rho(\mu)}\Delta\mathbf{\dot{H}}_{k+1,1}+|1-\mu|(\text{sign}(1-\mu)\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k,1}),$ (73)
+$\mathcal{B}_{2}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1,2}$ $=\frac{\rho(\mu)}{3}\frac{3\mu}{\rho(\mu)}\Delta\mathbf{\dot{H}}_{k+1,2}+|1-\mu|\text{sign}(1-\mu)\mathcal{B}_{2}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k,2}$
+$-\frac{\rho(\mu)}{3}\frac{3}{\rho(\mu)}\mathcal{B}_{1}^{\#}(\Delta\mathbf{\dot{\Lambda}}_{k+1,1})+\frac{\rho(\mu)}{3}\frac{3|1-\mu|\text{sign}(1-\mu)}{\rho(\mu)}(\mathcal{B}_{1}^{\#}(\Delta\mathbf{\dot{\Lambda}}_{k,1})).$
+By the convexity of $\|\cdot\|^{2}_{\mathrm{F}}$, the update of $\mathbf{\dot{\Lambda}}_{1}$ and $\mathbf{\dot{\Lambda}}_{2}$ in (59) and B2 in Assumption 4.2, it follows that $\Delta\mathbf{\dot{\Lambda}}_{k+1,1}\in\text{{\rm range}}(\mathcal{C}_{1})$, $\Delta\mathbf{\dot{\Lambda}}_{k+1,2}\in\text{{\rm range}}(\mathcal{B}_{2})$.
+The following inequalities hold:
+[…]
+(${\leq}$ $\frac{\mu^{2}}{\rho(\mu)}\|\Delta\mathbf{\dot{H}}_{k+1,1}\|^{2}_{\mathrm{F}}+|1-\mu|\|\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k,1}\|^{2}_{\mathrm{F}}-|1-\mu|\|\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1,1}\|^{2}_{\mathrm{F}},$
+and
+[…]
+$+\frac{3}{\rho(\mu)}\|\mathcal{B}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1,1}\|^{2}+\frac{3|1-\mu|^{2}}{\rho(\mu)}\|\mathcal{B}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k,1}\|_{\mathrm{F}}^{2}-|1-\mu|\|\mathcal{B}_{2}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1,2}\|^{2}_{\mathrm{F}}$
+(${\leq}$ $\frac{3\mu^{2}}{\rho(\mu)}\|\Delta\mathbf{\dot{H}}_{k+1,2}\|^{2}_{\mathrm{F}}+|1-\mu|\|\mathcal{B}_{2}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k,2}\|^{2}_{\mathrm{F}}-|1-\mu|\|\mathcal{B}_{2}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1,2}\|^{2}_{\mathrm{F}}$
+$+\frac{3\mu^{2}}{\rho(\mu)^{3}\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}}\|\mathcal{B}_{1}\|^{2}\|\Delta\mathbf{\dot{H}}_{k+1,1}\|_{\mathrm{F}}^{2}+\frac{3|1-\mu|\|\mathcal{B}_{1}\|^{2}}{\rho(\mu)^{2}\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}}\|\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k,1}\|_{\mathrm{F}}^{2}-\frac{3|1-\mu|\|\mathcal{B}_{1}\|^{2}}{\rho(\mu)^{2}\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}}\|\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1,1}\|_{\mathrm{F}}^{2}$
+$+\frac{3\mu^{2}|1-\mu|^{2}\|\mathcal{B}_{1}\|^{2}}{\rho(\mu)^{3}\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}}\|\Delta\mathbf{\dot{H}}_{k,1}\|^{2}_{\mathrm{F}}+\frac{3|1-\mu|^{3}\|\mathcal{B}_{1}\|^{2}}{\rho(\mu)^{2}\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}}\|\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k-1,1}\|_{\mathrm{F}}^{2}-\frac{3|1-\mu|^{3}\|\mathcal{B}_{1}\|^{2}}{\rho(\mu)^{2}\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}}\|\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k,1}\|^{2}$
+(${\leq}$ $\frac{3\mu^{2}}{\rho(\mu)}\left(2(L_{g}+q_{2})^{2}\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}^{2}+2q_{2}^{2}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}\right)$
+$+\frac{3\mu^{2}\|\mathcal{B}_{1}\|^{2}}{\rho(\mu)^{3}\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}}(4(q_{1}+L_{f})^{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}+4q_{1}^{2}\|\Delta{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}^{2}+4\beta_{1}^{2}\|\mathcal{B}_{1}\|^{2}\|\mathcal{C}_{1}\|^{2}(\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}+\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}^{2}))$
+$+\frac{3\mu^{2}|1-\mu|^{2}\|\mathcal{B}_{1}\|^{2}}{\rho(\mu)^{3}\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}}(4(q_{1}+L_{f})^{2}\|\Delta{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}^{2}+4q_{1}^{2}\|\Delta{\mathbf{\dot{X}}}_{k-1}\|_{\mathrm{F}}^{2})$
+$+4\beta_{1}^{2}\|\mathcal{B}_{1}\|^{2}\|\mathcal{C}_{1}\|^{2}(\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}^{2}+\|\Delta\mathbf{\dot{W}}_{k-1}\|_{\mathrm{F}}^{2})+\frac{3|1-\mu|\|\mathcal{B}_{1}\|^{2}}{\rho(\mu)^{2}\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}}(\|\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k,1}\|_{\mathrm{F}}^{2}-\|\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1,1}\|_{\mathrm{F}}^{2})$
+$+|1-\mu|\left(\|\mathcal{B}_{2}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k,2}\|^{2}_{\mathrm{F}}-\|\mathcal{B}_{2}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1,2}\|^{2}_{\mathrm{F}}\right)+\frac{3|1-\mu|^{3}\|\mathcal{B}_{1}\|^{2}}{\rho(\mu)^{2}\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}}\left(\|\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k-1,1}\|_{\mathrm{F}}^{2}-\|\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k,1}\|_{\mathrm{F}}^{2}\right).$
+Consequently, Inequality (68) follows from (71), B2 in Assumption 4.2, and the definition of $\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}$ and $\lambda_{+}^{\mathcal{B}_{2}^{\#}\mathcal{B}_{2}}$.
+Finally, Inequality (68) follows from multiplying Inequality (67) by $r>1$ and combining it with Lemma 4.13.
+∎
+Let $\mathcal{T}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}}_{1},\mathbf{\dot{\Lambda}}_{2},{\mathbf{\dot{X}}}^{\prime},\mathbf{\dot{W}}^{\prime},\mathbf{\dot{\Lambda}}_{1}^{\prime},\mathbf{\dot{\Lambda}}_{2}^{\prime},{\mathbf{\dot{X}}}^{\prime\prime},\mathbf{\dot{W}}^{\prime\prime},\mathbf{\dot{\Lambda}}_{1}^{\prime\prime}):=\mathcal{L}_{\beta_{1},\beta_{2}}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}}_{1},\mathbf{\dot{\Lambda}}_{2})+2r\theta_{1}\|{\mathbf{\dot{X}}}-{\mathbf{\dot{X}}}^{\prime}\|_{\mathrm{F}}^{2}+r\theta_{1}\|{\mathbf{\dot{X}}}^{\prime}-{\mathbf{\dot{X}}}^{\prime\prime}\|_{\mathrm{F}}^{2}+2r\theta_{2}\|\mathbf{\dot{W}}-\mathbf{\dot{W}}^{\prime}\|^{2}_{\mathrm{F}}+r\theta_{2}\|\mathbf{\dot{W}}^{\prime}-\mathbf{\dot{W}}^{\prime\prime}\|_{\mathrm{F}}^{2}+r\theta_{3}\|\mathcal{C}_{1}^{\#}(\mathbf{\dot{\Lambda}}_{1}-\mathbf{\dot{\Lambda}}_{1}^{\prime})\|^{2}_{\mathrm{F}}+r\theta_{4}\|\mathcal{C}_{1}^{\#}(\mathbf{\dot{\Lambda}}_{1}^{\prime}-\mathbf{\dot{\Lambda}}_{1}^{\prime\prime})\|^{2}_{\mathrm{F}}+r\theta_{5}\|\mathcal{B}_{2}^{\#}(\mathbf{\dot{\Lambda}}_{2}-\mathbf{\dot{\Lambda}}_{2}^{\prime})\|^{2}_{\mathrm{F}}$.
+The merit function $\mathcal{T}_{k}$ is defined by
+$\mathcal{T}_{k}:=\mathcal{T}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k,1},\mathbf{\dot{\Lambda}}_{k,2},{\mathbf{\dot{X}}}_{k-1},\mathbf{\dot{W}}_{k-1},\mathbf{\dot{\Lambda}}_{k-1,1},\mathbf{\dot{\Lambda}}_{k-1,2},{\mathbf{\dot{X}}}_{k-2},\mathbf{\dot{W}}_{k-2},\mathbf{\dot{\Lambda}}_{k-2,1})$ (75)
+$=\mathcal{L}_{\beta_{1},\beta_{2}}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k,1},\mathbf{\dot{\Lambda}}_{k,2})+2r\theta_{1}\|\Delta{\mathbf{\dot{X}}}_{k}\|^{2}_{\mathrm{F}}+r\theta_{1}\|\Delta{\mathbf{\dot{X}}}_{k-1}\|^{2}_{\mathrm{F}}+2r\theta_{2}\|\Delta\mathbf{\dot{W}}_{k}\|^{2}_{\mathrm{F}}+r\theta_{2}\|\Delta\mathbf{\dot{W}}_{k-1}\|^{2}_{\mathrm{F}}$
+$+r\theta_{3}\|\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k,1}\|^{2}_{\mathrm{F}}+r\theta_{4}\|\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k-1,1}\|_{\mathrm{F}}^{2}+r\theta_{5}\|\mathcal{B}_{2}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k,2}\|^{2}_{\mathrm{F}}.$
+It follows from (68), the definition of $\mathcal{T}_{k}$, and Assumption 4.2 that
+$\mathcal{T}_{k+1}+a\left(\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1,1}\|^{2}_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1,2}\|^{2}_{\mathrm{F}}\right)\leq\mathcal{T}_{k+1}+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{B_{k}-3r\theta_{2}\mathbf{I}}^{2}$ (76)
+$+\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2}_{\frac{L_{k,1}}{2}-3r\theta_{1}\mathbf{I}}+\frac{r-1}{\beta_{1}\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1,1}\|^{2}_{\mathrm{F}}+\frac{r-1}{\beta_{2}\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1,2}\|^{2}_{\mathrm{F}}\leq\mathcal{T}_{k}\leq\mathcal{T}_{k_{0}},$
+where $k\geq k_{0},a=\min\{a_{1},a_{2},\frac{r-1}{\beta_{1}\mu},\frac{r-1}{\beta_{2}\mu}\}$ and $a_{1},a_{2}$ are defined in Assumption 4.2, the first inequality follows from B3 and B4 in Assumption 4.2, the second inequality follows from Lemma 4.14, and the third inequality is due to the induction of $\mathcal{T}_{k+1}\leq\mathcal{T}_{k}$ for any $k\geq k_{0}$.
+Theorem 4.5 (Bounded sequence of $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k,1},\mathbf{\dot{\Lambda}}_{k,2})\}_{k\geq 0}$).
+Assume Assumption 4.2 holds.
+The sequence $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k,1},\mathbf{\dot{\Lambda}}_{k,2})\}_{k\geq 0}$ generated by (59) is bounded.
+Consequently, the limiting point set of $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k,1},\mathbf{\dot{\Lambda}}_{k,2})\}_{k\geq 0}$ is nonempty.
+Proof.
+According to (76), there exists $k_{0}\geq 1$ such that $\mathcal{T}_{k+1}\leq\mathcal{T}_{k_{0}}$ for all $k\geq k_{0}.$ Therefore, it holds that
+$f({\mathbf{\dot{X}}}_{k+1})+g(\mathbf{\dot{W}}_{k+1})+\frac{\beta_{1}}{2}\left\|\mathcal{C}_{1}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}_{1}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}_{1}+\frac{\mathbf{\dot{\Lambda}}_{k+1,1}}{\beta_{1}}\right\|_{\mathrm{F}}^{2}$ (77)
+$+\frac{\beta_{2}}{2}\left\|\mathcal{B}_{2}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}_{2}+\frac{\mathbf{\dot{\Lambda}}_{k+1,2}}{\beta_{2}}\right\|_{\mathrm{F}}^{2}-\frac{1}{2\beta_{1}}\left\|\mathbf{\dot{\Lambda}}_{k+1,1}\right\|_{\mathrm{F}}^{2}-\frac{1}{2\beta_{2}}\left\|\mathbf{\dot{\Lambda}}_{k+1,2}\right\|_{\mathrm{F}}^{2}$
+$+(r\theta_{2}+a)(\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2})+(r\theta_{1}+a)\left(\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}+\|\Delta{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}^{2}\right)+a(\|\Delta\mathbf{\dot{\Lambda}}_{k+1,1}\|_{\mathrm{F}}^{2}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1,2}\|_{\mathrm{F}}^{2})$
+$+r\theta_{3}\|\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1,1}\|_{\mathrm{F}}^{2}+r\theta_{4}\|\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k,1}\|_{\mathrm{F}}^{2}+r\theta_{5}\|\mathcal{B}_{2}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1,2}\|_{\mathrm{F}}^{2}\leq\mathcal{T}_{k_{0}}.$
+It follows from (72) that
+$\mathcal{C}_{1}^{\#}\mathbf{\dot{\Lambda}}_{k+1,1}$ $=\mu\mathbf{\dot{H}}_{k+1,1}+(1-\mu)\mathcal{C}_{1}^{\#}\mathbf{\dot{\Lambda}}_{k,1},$ (78)
+$\mathcal{B}_{2}^{\#}\mathbf{\dot{\Lambda}}_{k+1,2}$ $=\mu\mathbf{\dot{H}}_{k+1,2}+(1-\mu)\mathcal{B}_{2}^{\#}\mathbf{\dot{\Lambda}}_{k,2}-(1-\mu)\mathcal{B}_{1}^{\#}(\Delta\mathbf{\dot{\Lambda}}_{k+1,1})-\mu\mathcal{B}_{1}^{\#}(\mathbf{\dot{\Lambda}}_{k+1,1}).$
+Since $\mu\in(0,2),$ Equation (78) can be rewritten as:
+$\mu\mathcal{C}_{1}^{\#}\mathbf{\dot{\Lambda}}_{k+1,1}=$ $\rho(\mu)\frac{\mu\mathbf{\dot{H}}_{k+1}}{\rho(\mu)}+|1-\mu|\left(\text{sign}(1-\mu)\mathcal{C}_{1}^{\#}(\mathbf{\dot{\Lambda}}_{k}-\mathbf{\dot{\Lambda}}_{k+1})\right),$ (79)
+$\mu\mathcal{B}_{2}^{\#}\mathbf{\dot{\Lambda}}_{k+1,2}=$ $\frac{\rho(\mu)}{3}\frac{3\mu}{\rho(\mu)}\mathbf{\dot{H}}_{k+1,2}+|1-\mu|\text{sign}(1-\mu)\mathcal{B}_{2}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1,2}$
+$-\frac{\rho(\mu)}{3}\frac{3(1-\mu)}{\rho(\mu)}\mathcal{B}_{1}^{\#}(\Delta\mathbf{\dot{\Lambda}}_{k+1,1})-\frac{\rho(\mu)}{3}\frac{3\mu}{\rho(\mu)}(\mathcal{B}_{1}^{\#}(\mathbf{\dot{\Lambda}}_{k+1,1})).$
+It follows from the convexity of $\|\cdot\|^{2}$ that
+$\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}\mu^{2}\|\mathbf{\dot{\Lambda}}_{k+1,1}\|_{\mathrm{F}}^{2}\leq$ $\frac{\mu^{2}}{\rho(\mu)}\|\mathbf{\dot{H}}_{k+1,1}\|_{\mathrm{F}}^{2}+|1-\mu|\|\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1,1}\|_{\mathrm{F}}^{2},$
+$\lambda_{+}^{\mathcal{B}_{2}^{\#}\mathcal{B}_{2}}\mu^{2}\|\mathbf{\dot{\Lambda}}_{k+1,2}\|_{\mathrm{F}}^{2}\leq$ $\frac{3\mu^{2}}{\rho(\mu)}\|\mathbf{\dot{H}}_{k+1,2}\|_{\mathrm{F}}^{2}+|1-\mu|\|\mathcal{B}_{2}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1,2}\|_{\mathrm{F}}^{2}$
+$+\frac{3|1-\mu|^{2}}{\rho(\mu)}\|\mathcal{B}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1,1}\|_{\mathrm{F}}^{2}+\frac{3\mu^{2}}{\rho(\mu)}\|\mathcal{B}_{1}^{\#}\mathbf{\dot{\Lambda}}_{k+1,1}\|_{\mathrm{F}}^{2}.$
+According to (70), it holds that
+$\|\mathbf{\dot{H}}_{k+1,1}\|_{\mathrm{F}}^{2}\leq 3q_{1}^{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}+3\|\nabla f({\mathbf{\dot{X}}}_{k+1})\|_{\mathrm{F}}^{2}+3\beta_{1}^{2}\|\mathcal{C}_{1}\|^{2}\|\mathcal{B}_{1}\|^{2}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2},$
+$\|\mathbf{\dot{H}}_{k+1,2}\|_{\mathrm{F}}^{2}\leq 2(q_{2}+L_{g})^{2}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}+2\|\nabla g(\mathbf{\dot{W}}_{k+1})\|_{\mathrm{F}}^{2}.$
+Hence it follows that
+$-\frac{1}{2\beta_{1}}\|\mathbf{\dot{\Lambda}}_{k+1,1}\|_{\mathrm{F}}^{2}\geq-\theta_{3,0}\|\nabla f({\mathbf{\dot{X}}}_{k+1})\|_{\mathrm{F}}^{2}-\theta_{3,1}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}-\theta_{3,2}\|\Delta\mathbf{\dot{W}}_{k+1}\|^{2}-\theta_{3,3}\|\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1,1}\|_{\mathrm{F}}^{2},$ (80)
+$-\frac{1}{2\beta_{2}}\|\mathbf{\dot{\Lambda}}_{k+1,2}\|_{\mathrm{F}}^{2}\geq-\theta_{4,0}\|\nabla g(\mathbf{\dot{W}}_{k+1})\|_{\mathrm{F}}^{2}-\theta_{4,0}\|\nabla f({\mathbf{\dot{X}}}_{k+1})\|_{\mathrm{F}}^{2}-\theta_{4,2}\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}-\theta_{4,2}\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}^{2}$
+$-\theta_{4,5}\|\mathcal{B}_{2}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1,2}\|_{\mathrm{F}}^{2}-\theta_{4,3}\|\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1,1}\|_{\mathrm{F}}^{2}+\theta_{4,4}\|\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k,1}\|_{\mathrm{F}}^{2}-\theta_{4,1}\|\Delta{\mathbf{\dot{X}}}_{k}\|^{2}-\theta_{4,1}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2},$
+where
+$\theta_{3,0}=\frac{3}{2\beta_{1}\rho(\mu)\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}},~{}~{}\theta_{3,1}=\frac{3q_{1}^{2}}{2\beta_{1}\rho(\mu)\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}},~{}~{}\theta_{3,2}=\frac{3\beta_{1}\|\mathcal{C}_{1}\|^{2}\|\mathcal{B}_{1}\|^{2}}{2\rho(\mu)\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}},\theta_{3,3}=\frac{|1-\mu|}{2\beta_{1}\mu^{2}\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}},$
+$\theta_{4,0}=\max\left\{\frac{3}{\beta_{2}\rho(\mu)\lambda_{+}^{\mathcal{B}_{2}^{\#}\mathcal{B}_{2}}},\frac{9\|\mathcal{B}_{1}\|^{2}}{2\beta_{2}\rho(\mu)^{2}\lambda_{+}^{\mathcal{B}_{2}^{\#}\mathcal{B}_{2}}\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}}\right\},$
+$\theta_{4,1}=\frac{9q_{1}^{2}\|\mathcal{B}_{1}\|^{2}}{2\beta_{2}\rho(\mu)^{2}\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}\lambda_{+}^{\mathcal{B}_{2}^{\#}\mathcal{B}_{2}}}+\frac{6|1-\mu|^{2}\|\mathcal{B}_{1}\|^{2}(q_{1}+L_{f})^{2}}{\beta_{2}\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}\lambda_{+}^{\mathcal{B}_{2}^{\#}\mathcal{B}_{2}}\rho(\mu)^{3}},$
+$\theta_{4,2}=\frac{3(L_{g}+q_{2})^{2}}{\beta_{2}\rho(\mu)\lambda_{+}^{\mathcal{B}_{2}^{\#}\mathcal{B}_{2}}}+\frac{6\|\mathcal{B}_{1}\|^{2}(|1-\mu|^{2})(\beta_{1}^{2}\|\mathcal{B}_{1}\|^{2}\|\mathcal{C}_{1}\|^{2})}{\beta_{2}\rho(\mu)^{3}\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}\lambda_{+}^{\mathcal{B}_{2}^{\#}\mathcal{B}_{2}}}+\frac{9\beta_{1}^{2}\|\mathcal{B}_{1}\|^{4}\|\mathcal{C}_{1}\|^{2}}{2\beta_{2}\rho(\mu)^{2}\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}\lambda_{+}^{\mathcal{B}_{2}^{\#}\mathcal{B}_{2}}},$
+$\theta_{4,3}=\frac{3|1-\mu|\|\mathcal{B}_{1}\|^{2}}{2\beta_{2}\rho(\mu)^{2}\mu^{2}\lambda_{+}^{\mathcal{B}_{2}^{\#}\mathcal{B}_{2}}\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}},\theta_{4,4}=\frac{3|1-\mu|^{3}\|\mathcal{B}_{1}\|^{2}}{2\beta_{2}\rho(\mu)^{2}\mu^{2}\lambda_{+}^{\mathcal{B}_{2}^{\#}\mathcal{B}_{2}}\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}},\theta_{4,5}=\frac{|1-\mu|}{2\beta_{2}\mu^{2}\lambda_{+}^{\mathcal{B}_{2}^{\#}\mathcal{B}_{2}}}.$
+Define $\theta_{6}=\theta_{3,0}+\theta_{4,0},\theta_{7}=\theta_{3,1}+\theta_{4,1},\theta_{8}=\theta_{3,2}+\theta_{4,2},\theta_{9}=\theta_{3,3}+\theta_{4,3},\theta_{10}=\theta_{4,4},\theta_{11}=\theta_{4,5}$.
+Using Inequalities (80) and (77), we obtain
+$(1-\kappa)f({\mathbf{\dot{X}}}_{k+1})+\frac{\beta_{1}}{2}\left\|\mathcal{C}_{1}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}_{1}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}_{1}+\frac{\mathbf{\dot{\Lambda}}_{k+1,1}}{\beta_{1}}\right\|_{\mathrm{F}}^{2}+\frac{\beta_{2}}{2}\left\|\mathcal{B}_{2}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}_{2}+\frac{\mathbf{\dot{\Lambda}}_{k+1,2}}{\beta_{2}}\right\|_{\mathrm{F}}^{2}$
+$+a(\|\Delta\mathbf{\dot{\Lambda}}_{k+1,1}\|_{\mathrm{F}}^{2}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1,2}\|_{\mathrm{F}}^{2})+(r\theta_{2}+a-\theta_{8})\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}^{2}+(r\theta_{1}+a-\theta_{7})\left(\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}\right)$
+$+(r\theta_{3}-\theta_{9})\|\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1,1}\|_{\mathrm{F}}^{2}+(r\theta_{4}-\theta_{10})\|\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k,1}\|_{\mathrm{F}}^{2}+(r\theta_{5}-\theta_{11})\|\mathcal{B}_{2}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1,2}\|_{\mathrm{F}}^{2}$
+$\leq\mathcal{T}_{k_{0}}-\inf_{\mathbf{\dot{W}}}\left\{g(\mathbf{\dot{W}})-\theta_{4,0}\|\nabla g(\mathbf{\dot{W}})\|_{\mathrm{F}}^{2}\right\}-\inf_{{\mathbf{\dot{X}}}}\left\{\kappa f({\mathbf{\dot{X}}})-\theta_{6}\|\nabla f({\mathbf{\dot{X}}})\|_{\mathrm{F}}^{2}\right\},$ (81)
+where $\kappa\in(0,1)$.
+According to (60) in B1 of Assumption 4.2, setting ${\mathbf{\dot{X}}}_{1}={\mathbf{\dot{X}}}-\delta_{2}\nabla f({\mathbf{\dot{X}}})$ and ${\mathbf{\dot{X}}}_{2}={\mathbf{\dot{X}}}$, it follows that $\kappa f({\mathbf{\dot{X}}}_{k}-\delta_{1}\nabla f({\mathbf{\dot{X}}}_{k}))\leq\kappa f({\mathbf{\dot{X}}}_{k})-\kappa(\delta_{2}-\frac{L_{f}\delta_{2}^{2}}{2})\|\nabla f({\mathbf{\dot{X}}}_{k})\|_{\mathrm{F}}^{2}.$ Since $g$ and $f$ are bounded from below, there exists $M$ such that
+$-M<\inf\{g(\mathbf{\dot{W}})-(\delta_{1}-\frac{L_{g}\delta_{1}^{2}}{2})\|\nabla g(\mathbf{\dot{W}})\|_{\mathrm{F}}^{2}:\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}\},$ (82)
+$-M<\inf\{\kappa f({\mathbf{\dot{X}}})-\kappa(\delta_{2}-\frac{L_{f}\delta_{2}^{2}}{2})\|\nabla f({\mathbf{\dot{X}}})\|_{\mathrm{F}}^{2}:{\mathbf{\dot{X}}}\in\mathbb{H}^{m\times n}\}.$
+We choose $\delta_{1}=\frac{1}{L_{g}},\delta_{2}=\frac{1}{L_{f}}$.
+According to B4 in Assumption 4.2, we have $\theta_{4,0}<\frac{1}{2L_{g}}=(\delta_{1}-\frac{L_{g}\delta_{1}^{2}}{2}),\theta_{6}<\frac{\kappa}{2L_{f}}=\kappa(\delta_{2}-\frac{L_{f}\delta_{2}^{2}}{2}).$ Since $r>1$ and $\mu\in(0,2),$ according to the denifition of $\theta_{i},i=1,\cdots,9$, it holds that $r\theta_{1}+a-\theta_{7}>0$,$r\theta_{2}+a-\theta_{8}>0,r\theta_{3}-\theta_{9}>0,r\theta_{4}-\theta_{10}>0,r\theta_{5}-\theta_{11}>0$.
+It follows from (81) that
+$(1-\kappa)f({\mathbf{\dot{X}}}_{k+1})+\frac{\beta_{2}}{2}\|\mathcal{B}_{2}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}_{2}+\frac{\mathbf{\dot{\Lambda}}_{k+1,2}}{\beta_{2}}\|_{\mathrm{F}}^{2}$ (83)
+$+\frac{\beta_{1}}{2}\left\|\mathcal{C}_{1}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}_{1}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}_{1}+\frac{\mathbf{\dot{\Lambda}}_{k+1,1}}{\beta_{1}}\right\|_{\mathrm{F}}^{2}+a(\|\Delta\mathbf{\dot{\Lambda}}_{k+1,1}\|_{\mathrm{F}}^{2}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1,2}\|_{\mathrm{F}}^{2})<\mathcal{T}_{k_{0}}+2M.$
+According to (83), $\Delta\mathbf{\dot{\Lambda}}_{k+1,1},\Delta\mathbf{\dot{\Lambda}}_{k+1,2}$ are bounded.
+By the $\mathbf{\dot{\Lambda}}_{2}$ update in (59), we have $\mathcal{B}_{2}\mathbf{\dot{W}}_{k+1}=\frac{1}{\beta_{2}\mu}\Delta\mathbf{\dot{\Lambda}}_{k+1,2}+\mathbf{\dot{B}}_{2}$.
+If $\mathcal{B}_{2}^{\#}\mathcal{B}_{2}$ is full rank, $\{\mathbf{\dot{W}}_{k}\}$ is bounded.
+By the $\mathbf{\dot{\Lambda}}_{1}$ update in (59), we have $\mathcal{C}_{1}{\mathbf{\dot{X}}}_{k+1}=\frac{1}{\beta_{1}\mu}\Delta\mathbf{\dot{\Lambda}}_{k+1,1}-\mathcal{B}_{1}\mathbf{\dot{W}}_{k+1}+\mathbf{\dot{B}}_{1}.$ Since $\{\mathcal{B}_{1}\mathbf{\dot{W}}_{k}\}_{k\geq k_{0}}$ is bounded, if $\mathcal{C}_{1}^{\#}\mathcal{C}_{1}$ is full rank or $f({\mathbf{\dot{X}}}_{k+1})$ is coercive, it follows that $\{{\mathbf{\dot{X}}}_{k}\}_{k\geq k_{0}}$ is bounded.
+From the fact that $\frac{\beta_{1}}{2}\left\|\mathcal{C}_{1}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}_{1}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}_{1}+\frac{\mathbf{\dot{\Lambda}}_{k+1,1}}{\beta_{1}}\right\|_{\mathrm{F}}^{2}$ and $\frac{\beta_{2}}{2}\left\|\mathcal{B}_{2}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}_{2}+\frac{\mathbf{\dot{\Lambda}}_{k+1,2}}{\beta_{2}}\right\|_{\mathrm{F}}^{2}$ are bounded, it follows that $\{\mathbf{\dot{\Lambda}}_{k_{1}}\}_{k\geq k_{0}}$ and $\{\mathbf{\dot{\Lambda}}_{k_{2}}\}_{k\geq k_{0}}$ are bounded.
+As a consequence, $\{{\mathbf{\dot{X}}}_{k}\}_{k\geq 1},\{\mathbf{\dot{W}}_{k}\}_{k\geq 1},\{\mathbf{\dot{\Lambda}}_{k,1}\}_{k\geq 1},\{\mathbf{\dot{\Lambda}}_{k,2}\}_{k\geq 1}$ are bounded.
+∎
+Lemma 4.15 is used in Lemma 4.16 to show the limiting behavior of sequence $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k,1},\mathbf{\dot{\Lambda}}_{k,2})\}_{k\geq 0}$.
+Lemma 4.15 (Subgradient bound).
+Suppose that B1 in Assumption holds. Let $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k,1},\mathbf{\dot{\Lambda}}_{k,2})\}_{k\geq 0}$ be a sequence generated by (59). Then
+$\widetilde{\mathbf{\dot{D}}}_{k+1}:=(\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k+1}},\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k+1}},\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k+1,1}},\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k+1,2}})\in\partial\mathcal{L}_{\beta_{1},\beta_{2}}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1,1},\mathbf{\dot{\Lambda}}_{k+1,2}),$
+where
+$\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k+1,1}}$ $:=\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}-L_{k,1}\Delta{\mathbf{\dot{X}}}_{k+1}-\beta_{1}\mathcal{C}_{1}^{\#}\mathcal{B}_{1}\Delta\mathbf{\dot{W}}_{k+1},\quad\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k+1}}:=\frac{1}{\beta_{1}\mu}\Delta\mathbf{\dot{\Lambda}}_{k+1,1},\,$
+$\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k+1}}$ $:=\nabla g(\mathbf{\dot{W}}_{k+1})-\nabla g(\mathbf{\dot{W}}_{k})+\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1}-L_{k,2}\Delta\mathbf{\dot{W}}_{k+1},\quad\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k+1,2}}:=\frac{1}{\beta_{2}\mu}\Delta\mathbf{\dot{\Lambda}}_{k+1,2}.$
+Furthermore, it holds that
+$|\|\widetilde{\mathbf{\dot{D}}}_{k+1}\||\leq\pi\left(\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1,1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1,2}\|_{\mathrm{F}}\right),$
+where $\pi=\max\left\{q_{1},\beta_{1}\|\mathcal{C}_{1}\|\|\mathcal{B}_{1}\|+q_{2}+L_{g},\|\mathcal{C}_{1}\|+\|\mathcal{B}_{1}\|+\frac{1}{\beta_{1}\mu},\|\mathcal{B}_{2}\|+\frac{1}{\beta_{2}\mu}\right\}.$
+Proof.
+According to the optimality condition of the update of ${\mathbf{\dot{X}}}_{k+1},$ it follows that
+$-\beta_{1}\mathcal{C}_{1}^{\#}(\mathcal{C}_{1}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}_{1}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}_{1}+\mathbf{\dot{\Lambda}}_{k,1}/\beta_{1})-L_{k,1}\Delta{\mathbf{\dot{X}}}_{k+1}=\nabla f({\mathbf{\dot{X}}}_{k+1}),$
+which yields
+$\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k+1}}=-\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1,1}-L_{k,1}\Delta{\mathbf{\dot{X}}}_{k+1}+\beta_{1}\mathcal{C}_{1}^{\#}\mathcal{B}_{1}\Delta\mathbf{\dot{W}}_{k+1}\in\partial_{{\mathbf{\dot{X}}}}\mathcal{L}_{\beta_{1},\beta_{2}}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1,1},\mathbf{\dot{\Lambda}}_{k+1,2}).$ (84)
+According to the optimality condition of the update of $\mathbf{\dot{W}}_{k+1},$ the following equality holds:
+$\nabla g(\mathbf{\dot{W}}_{k})+\beta_{1}\mathcal{C}_{1}^{\#}(\mathcal{C}_{1}{\mathbf{\dot{X}}}_{k+1}-\mathcal{B}_{1}\mathbf{\dot{W}}_{k+1}-\mathbf{\dot{B}}_{1}+\frac{\mathbf{\dot{\Lambda}}_{k,1}}{\beta_{1}})+\beta_{2}\mathcal{B}_{2}^{\#}(\mathcal{B}_{2}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}_{2}+\frac{\mathbf{\dot{\Lambda}}_{k,2}}{\beta_{2}})+L_{k,2}\Delta\mathbf{\dot{W}}_{k+1}=0.$ (85)
+It follows from (85) and the update of $\mathbf{\dot{\Lambda}}_{k+1,1},\mathbf{\dot{\Lambda}}_{k+1,2}$ that
+$\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k+1}}$ $=\nabla g(\mathbf{\dot{W}}_{k+1})-\nabla g(\mathbf{\dot{W}}_{k})+\mathcal{C}_{1}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1,1}+\mathcal{B}_{2}^{\#}\Delta\mathbf{\dot{\Lambda}}_{k+1,2}-L_{k,2}\Delta\mathbf{\dot{W}}_{k+1},$ (86)
+$\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k+1,1}}$ $=\frac{1}{\mu\beta_{1}}\Delta\mathbf{\dot{\Lambda}}_{k+1,1}\in\partial_{\mathbf{\dot{\Lambda}}_{1}}\mathcal{L}_{\beta_{1},\beta_{2}}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1,1},\mathbf{\dot{\Lambda}}_{k+1,2}),$
+$\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k+1,2}}$ $=\frac{1}{\mu\beta_{2}}\Delta\mathbf{\dot{\Lambda}}_{k+1,2}\in\partial_{\mathbf{\dot{\Lambda}}_{2}}\mathcal{L}_{\beta_{1},\beta_{2}}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1,1},\mathbf{\dot{\Lambda}}_{k+1,2}).$
+Hence, it follows that $\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k+1}}\in\partial_{\mathbf{\dot{W}}}\mathcal{L}_{\beta_{1},\beta_{2}}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1,1},\mathbf{\dot{\Lambda}}_{k+1,2}).$ Combining (84), (86), we obtain
+$\|\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k+1}}\|$ $\leq\|\mathcal{C}_{1}\|\|\Delta\mathbf{\dot{\Lambda}}_{k+1,1}\|_{\mathrm{F}}+q_{1}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}+\beta_{1}\|\mathcal{C}_{1}\|\|\mathcal{B}_{1}\|\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}},$
+$\leq\|\mathcal{B}_{1}\|\|\Delta\mathbf{\dot{\Lambda}}_{k+1,1}\|_{\mathrm{F}}+\|\mathcal{B}_{2}\|\|\Delta\mathbf{\dot{\Lambda}}_{k+1,2}\|_{\mathrm{F}}+(q_{2}+L_{g})\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}},$
+$\|\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k+1,1}}\|$ $=\frac{1}{\beta_{1}\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1,1}\|_{\mathrm{F}},\|\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k+1,2}}\|=\frac{1}{\beta_{2}\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1,2}\|_{\mathrm{F}}.$
+Therefore, the following inequality holds:
+$|\|\widetilde{\mathbf{\dot{D}}}_{k+1}\||\leq\|\mathbf{\dot{D}}_{\mathbf{\dot{W}}_{k+1}}\|_{\mathrm{F}}+\|\mathbf{\dot{D}}_{{\mathbf{\dot{X}}}_{k+1}}\|_{\mathrm{F}}+\|\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k+1,1}}\|_{\mathrm{F}}+\|\mathbf{\dot{D}}_{\mathbf{\dot{\Lambda}}_{k+1,2}}\|_{\mathrm{F}}$ (87)
+$\leq\pi(\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1,1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1,2}\|_{\mathrm{F}}),$
+which completes the proof. ∎
+Lemma 4.16.
+Suppose that Assumption 4.2 holds.
+Any limit point $({\mathbf{\dot{X}}}_{*},\mathbf{\dot{W}}_{*},\mathbf{\dot{\Lambda}}_{*,1},\mathbf{\dot{\Lambda}}_{*,2})$ of the sequence $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k,1},\mathbf{\dot{\Lambda}}_{k,2})\}_{k\geq 0}$ generated by (59) is a stationary point of (11), i.e.
+[…]
+Proof.
+Let $\{({\mathbf{\dot{X}}}_{k_{j}},\mathbf{\dot{W}}_{k_{j}},\mathbf{\dot{\Lambda}}_{k_{j},1},\mathbf{\dot{\Lambda}}_{k_{j},2})\}_{j\geq 0}$ be a subsequence of $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k,1},\mathbf{\dot{\Lambda}}_{k,2})\}_{k\geq 0}$ such that $({\mathbf{\dot{X}}}_{*},\mathbf{\dot{W}}_{*},\mathbf{\dot{\Lambda}}_{*,1},\mathbf{\dot{\Lambda}}_{*,2})=\lim_{j\rightarrow\infty}({\mathbf{\dot{X}}}_{k_{j}},\mathbf{\dot{W}}_{k_{j}},\mathbf{\dot{\Lambda}}_{k_{j},1},\mathbf{\dot{\Lambda}}_{k_{j},2})$.
+By the continuity of $\mathcal{L}_{\beta_{1},\beta_{2}}$, $\mathcal{L}_{\beta_{1},\beta_{2}}({\mathbf{\dot{X}}}_{k_{j}},\mathbf{\dot{W}}_{k_{j}},\mathbf{\dot{\Lambda}}_{k_{j},1},\mathbf{\dot{\Lambda}}_{k_{j},2})\rightarrow\mathcal{L}_{\beta_{1},\beta_{2}}({\mathbf{\dot{X}}}_{*},\mathbf{\dot{W}}_{*},\mathbf{\dot{\Lambda}}_{*,1},\mathbf{\dot{\Lambda}}_{*,2})$ as $j\rightarrow\infty$.
+Let $\mathbf{\dot{D}}_{k_{j}}\in\partial\mathcal{L}_{\beta_{1},\beta_{2}}({\mathbf{\dot{X}}}_{k_{j}},\mathbf{\dot{W}}_{k_{j}},\mathbf{\dot{\Lambda}}_{k_{j},1},\mathbf{\dot{\Lambda}}_{k_{j},2})$ and according to Lemma 4.15, it follows that $\|\mathbf{\dot{D}}_{k_{j}}\|_{\mathrm{F}}\leq\rho\left(\|\Delta\mathbf{\dot{W}}_{k_{j}}\|_{\mathrm{F}}+\|\Delta{\mathbf{\dot{X}}}_{k_{j}}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k_{j},1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k_{j},2}\|_{\mathrm{F}}\right).$ According to Theorem 4.5, it follows that $\mathbf{\dot{D}}_{k_{j}}\rightarrow 0.$ By the closeness criterion of the limiting subdifferential, $({\mathbf{\dot{X}}}_{*},\mathbf{\dot{W}}_{*},\mathbf{\dot{\Lambda}}_{*,1},\mathbf{\dot{\Lambda}}_{*,2})\in\text{crit}\,\mathcal{L}_{\beta_{1},\beta_{2}}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}}_{1},\mathbf{\dot{\Lambda}}_{2})$.
+The proof is completed.
+∎
+Since we have established the key properties, i.e.
+sufficient decrease property (Lemma 4.14), the existence of limiting point (Theorem 4.5 and Lemma 4.16) and subgradient bound (Lemma 4.15), the global convergence of the algorithm is similar to the proof method in the literatures [2, 3, 43].
+We present the convergence result in Theorem 4.6, but omit its proof.
+Theorem 4.6.
+Let $\tilde{\varGamma}$ denote the set of the limit points of the sequence $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k,1},\mathbf{\dot{\Lambda}}_{k,2})\}_{k\geq 0}$.
+Suppose that Assumption 4.2 holds and $\mathcal{T}$ defined in (75) satisfies the KL property on $({\mathbf{\dot{X}}}_{+},\mathbf{\dot{W}}_{+},\mathbf{\dot{\Lambda}}_{+,1},\mathbf{\dot{\Lambda}}_{+,2},{\mathbf{\dot{X}}}_{+},\mathbf{\dot{W}}_{+},\mathbf{\dot{\Lambda}}_{+,1},\mathbf{\dot{\Lambda}}_{+,2},{\mathbf{\dot{X}}}_{+},\mathbf{\dot{W}}_{+},\mathbf{\dot{\Lambda}}_{+,1})$ for any $({\mathbf{\dot{X}}}_{+},\mathbf{\dot{W}}_{+},\mathbf{\dot{\Lambda}}_{+,1},\mathbf{\dot{\Lambda}}_{+,2})\in\tilde{\varGamma}$, then $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k,1},\mathbf{\dot{\Lambda}}_{k,2})\}_{k\geq 0}$ satisfies the finite length property: $\sum_{k=0}^{\infty}\|\Delta{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k,1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k,2}\|_{\mathrm{F}}<\infty,$ and consequently converges to a stationary point of (11).
+According to Theorem 4.6, we have the following convergence corollary of Algorithm 3 for SLRQA-NF.
+Corollary 4.2.
+Suppose the penalty parameter $\beta_{1}$ and $\beta_{2}$ are large enough, and $L_{1,k},L_{2,k}$ are set as given constants for all $k$ such that B3 and B4 in Assumption 4.2 hold, then the sequence $\{({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k,1},\mathbf{\dot{\Lambda}}_{k,2})\}_{k\geq 0}$ generated by Algorithm 3 converges to a stationary point of SLRQA-NF in (4).
+Proof.
+By the definition of SLRQA (2), B1 and B2 in Assumption 4.2 are satisfied.
+Furthermore, it follows from the proof of Corollary 4.1 that $\mathcal{T}$ defined in (75) satisfy the KL property.
+Hence, the final result follows from Theorem 4.6.
+∎
 5 Numerical experiment
@@
 In Section 5.1, we give the parameter setting for SLRQA and SLRQA-NF.
-In Section 6.2, we compare SLRQA and SLRQA-NF models based on quaternion and RGB representation to demonstrate the rationality of quaternion representation.
-Furthermore, in Section 6.3 and 6.4, comparisons of SLRQA and SLRQA-NF with other state-of-the-art methods are also presented to respectively show the superiority of SLRQA and SLRQA-NF.
+In Section 5.2, we compare SLRQA and SLRQA-NF based on quaternion and RGB representation to demonstrate the rationality of quaternion representation.
+Furthermore, in Section 5.3 and 5.4, comparisons of SLRQA and SLRQA-NF with other state-of-the-art methods are also presented to respectively show the superiority of SLRQA and SLRQA-NF.
 All the experiments are performed using MATLAB R2022b running on a desktop with an Intel Core R7-5800H CPU (3.90 GHz) and 16 GB of RAM.
@@
 The models for color image denoising and image inpainting are respectively given by
-$\min_{{\mathbf{\dot{X}}},\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}}$ $\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)+\lambda p(\mathbf{\dot{W}})+\frac{1}{2\tau^{2}}\|{\mathbf{\dot{X}}}-\mathbf{\dot{Y}}\|_{\mathrm{F}}^{2},\quad\mathrm{s.t.}\quad\mathcal{W}({\mathbf{\dot{X}}})=\mathbf{\dot{W}},$ (55)
+$\min_{{\mathbf{\dot{X}}},\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}}$ $\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma)+\lambda p(\mathbf{\dot{W}})+\frac{1}{2\tau^{2}}\|{\mathbf{\dot{X}}}-\mathbf{\dot{Y}}\|_{\mathrm{F}}^{2},\quad\mathrm{s.t.}\quad\mathcal{W}({\mathbf{\dot{X}}})=\mathbf{\dot{W}},$ (88)
 and
-$\min_{\mathbf{\dot{Z}},\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}}$ $\sum_{i}\phi(\sigma_{i}(\mathbf{\dot{Z}}),\gamma)+\lambda p(\mathbf{\dot{W}}),\quad\mathrm{s.t.}\quad\mathcal{P}_{\Omega}(\mathcal{W}^{\#}(\mathbf{\dot{W}}))=\mathcal{P}_{\Omega}(\mathbf{\dot{Y}}),\quad\mathbf{\dot{Z}}=\mathbf{\dot{W}},$ (56)
-where $\mathcal{W}$ is the QDCT [12], $p(\mathbf{\dot{W}})$ is chosen as the Huber function, i.e., $(p(\mathbf{\dot{W}}))_{ij}=\frac{1}{2\delta}|\mathbf{\dot{W}}_{ij}|^{2}$ if $|\mathbf{\dot{W}}_{ij}|<\delta$; and $(p(\mathbf{\dot{W}}))_{ij}=|\mathbf{\dot{W}}_{ij}|-\frac{1}{2}\delta$ otherwise, $\delta>0$ is a prescribed constant, and the function $\phi$ are chosen as the Schatten-$\gamma$, Laplace, or Weighted Schatten-$\gamma$ functions.
-Note that the color image denoising problem (55) is the SLRQA in (2) with $\mathcal{A}$ being identity, $\Omega$ being the entire indices of pixels, and $p$ being the Huber function.
-The color image inpainting problem (56) is the SLRQA-NF in (4) with $\mathcal{A}$ being identity and $p$ being the Huber function.
+$\min_{\mathbf{\dot{Z}},\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}}$ $\sum_{i}(\sigma_{i}(\sqrt{{\mathbf{\dot{Z}}}^{*}\mathbf{\dot{Z}}+\varepsilon^{2}\mathbf{I}}),\gamma)+\lambda p(\mathbf{\dot{W}}),\quad\mathrm{s.t.}\quad\mathcal{P}_{\Omega}(\mathcal{W}^{\#}(\mathbf{\dot{W}}))=\mathcal{P}_{\Omega}(\mathbf{\dot{Y}}),\quad\mathbf{\dot{Z}}=\mathbf{\dot{W}},$ (89)
+where $\mathcal{W}$ is the QDCT, see details in [14], $p(\mathbf{\dot{W}})$ is chosen as the Huber function, and the function $\phi$ is chosen as the Schatten-$\gamma$, Laplace, or Weighted Schatten-$\gamma$ functions.
+Note that the color image denoising problem (88) is SLRQA in (2) with $\mathcal{A}$ being identity, $\Omega$ being the entire indices of pixels, and $p$ being the Huber function.
+The color image inpainting problem (89) is SLRQA-NF in (4) with $\mathcal{A}$ being identity and $p$ being the Huber function.
 Problem (88) with the Schatten-$\gamma$, Laplace, and Weighted Schatten-$\gamma$ functions for $\phi$ is respectively denoted by SLRQA-1, SLRQA-2, and SLRQA-3.
@@
 Specifically, $\delta=1$ if $\epsilon_{k}\geq 10^{-2}$, $\delta=10^{-2}$ if $10^{-3}\leq\epsilon_{k}<10^{-2}$, and $\delta=10^{-4}$ otherwise, where $\epsilon_{k}=\|{\mathbf{\dot{X}}}_{k+1}-{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}+\|\mathbf{\dot{W}}_{k+1}-\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}+\|\mathcal{W}({\mathbf{\dot{X}}}_{k+1})-\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}$ in Algorithm 1 and $\epsilon_{k}=\|\mathbf{\dot{Z}}_{k+1}-\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+\|\mathcal{P}_{\Omega}(\mathcal{W}^{\#}(\mathbf{\dot{W}}_{k+1}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}$ in Algorithm 3.
+Furthermore, for PL-ADMM-NF, the objective function of the subproblem of $\mathbf{\dot{W}}$ is differentiable, we solve the subproblem using the gradient descent method.
 The parameters in Algorithm 1, 2 and 3 are set as $\mu=1.1$, $\beta=10$, $L_{1,k}=L_{2,k}=1,\forall k$, $\eta_{C}=10^{-10}$, and $\eta=10^{-4}$.
 Note that $\eta_{C}=10^{-10}$ is for the high accuracy of the ${\mathbf{\dot{X}}}$ or $\mathbf{\dot{Z}}$-subproblem.
+The threshold $\varepsilon$ of $\phi$ is set to $10^{-2}$ in our experiments.
 The qtfm [1] package is used for the quaternion computations except the QSVD.
@@
 Peak Signal-to-Noise Ration (PSNR) and Structure Similarity (SSIM) [36] are chosen to measure the quality of the images recovered by all methods.
-Specifically, the PSNR of two color image $\mathbf{X}$ and $\mathbf{Y}\in\mathbb{R}^{m\times n\times 3}$ is calculated as: ${\rm PSNR}(\mathbf{X},\mathbf{Y})=10\log_{10}\frac{\textit{peakval}}{\rm{MSE}},$ where $\text{MSE}=\frac{1}{3mn}\sum_{i=1}^{m}\sum_{j=1}^{n}\sum_{k=1}^{3}|\mathbf{X}_{i,j,k}-\mathbf{Y}_{i,j,k}|^{2},$ and peakval is taken from the range of the image data type.
-Furthermore, the SSIM of $\mathbf{X}$ and $\mathbf{Y}$ is given by
-$\text{SSIM}(\mathbf{X}_{ij},\mathbf{Y}_{ij})=\frac{1}{MN}\sum_{i=1}^{M}\sum_{j=1}^{N}\text{SSIM}(\mathbf{X}_{ij},\mathbf{Y}_{ij})$ (57)
-$\text{SSIM}(\mathbf{X}_{ij},\mathbf{Y}_{ij})=l(\mathbf{X}_{ij},\mathbf{Y}_{ij})\,c(\mathbf{X}_{ij},\mathbf{Y}_{ij})\,s(\mathbf{X}_{ij},\mathbf{Y}_{ij}),$ (58)
-where $\mathbf{X}_{ij}$ and $\mathbf{Y}_{ij}$ denote the local image patches, $M$ and $N$ denote the number of local patches along horizontal and vertical direction respectively, the three functions $l(\mathbf{X}_{ij},\mathbf{Y}_{ij})$, $c(\mathbf{X}_{ij},\mathbf{Y}_{ij})$, and $s(\mathbf{X}_{ij},\mathbf{Y}_{ij})$ respectively denote luminance item, contrast item, and structure item, and they are defined as follows:
-$l(\mathbf{X}_{ij},\mathbf{Y}_{ij})=\frac{2\mu_{\mathbf{X}_{ij}}\mu_{\mathbf{Y}_{ij}}+C_{1}}{\mu_{\mathbf{X}_{ij}}^{2}+\mu_{\mathbf{Y}_{ij}}^{2}+C_{1}},\quad c(\mathbf{X}_{ij},\mathbf{Y}_{ij})=\frac{2\sigma_{\mathbf{X}_{ij}}\sigma_{\mathbf{Y}_{ij}}+C_{2}}{\sigma_{\mathbf{X}_{ij}}^{2}+\sigma_{\mathbf{Y}_{ij}}^{2}+C_{2}},\quad s(\mathbf{X}_{ij},\mathbf{Y}_{ij})=\frac{\sigma_{\mathbf{X}_{ij}\mathbf{Y}_{ij}}+C_{3}}{\beta_{\mathbf{X}_{ij}}\sigma_{\mathbf{Y}_{ij}}+C_{3}},$ (59)
-with $\mu_{\mathbf{X}_{ij}},\mu_{\mathbf{Y}_{ij}},\sigma_{\mathbf{X}_{ij}},\sigma_{\mathbf{Y}_{ij}},\sigma_{{\mathbf{X}_{ij}}{\mathbf{Y}_{ij}}},C_{1},C_{2}$ and $C_{3}$ being the mean, standard deviation and cross-covariance of image patches $\mathbf{X}_{ij}$ and $\mathbf{Y}_{ij}$, $10^{-4},9\times 10^{-4}$ and $4.5\times 10^{-4}$ respectively.
 MATLAB commands psnr and ssim are used to compute the PSNR and SSIM for color images.
@@
 The figure and table show that SLRQA with quaternion representation achieves higher PSNR and SSIM compared with RGB representation in all cases.
-This observation coincides with the intuition that quaternion representation utilizes the information between channels to achieve better performance.
+This observation coincides with the intuition that quaternion representation uses the information between channels to achieve better performance.
 Figure 1: The 10 color images (512 $\times$ 512 $\times$ 3) for numerical experiments.
@@
 5.3 Color image denoising
-NSS is a significant technique in color image denoising problems.
+Non-local self-similarity (NSS) is a significant technique in color image denoising problems.
 The comparison without NSS directly reflects the superiority of the compared models, and the comparison with NSS better reflects the robustness of the models in practice.
-To fully compare the SLRQA with other state-of-the-art color image denoising methods, we tested their performance without and with the usage of non-local self-similarity (NSS) respectively in Section 6.3.1 and Section 6.3.2.
+To fully compare the SLRQA with other state-of-the-art color image denoising methods, we tested their performance without and with the usage of NSS respectively in Section 5.3.1 and Section 5.3.2.
 5.3.1 Color image denoising without NSS
-The compared methods include SRRC[39], WNNM[14], LRQA[10], and DWT[5]333 The implementations that we use are available from:
+The compared methods include pQSTV [37], QWSNM [48], MCNNFNM [33], SRRC [49], QWNNM [20], LRQA [12], and DWT [7]333 The implementations that we use are available at
+pQSTV: https://github.com/Huang-chao-yan/pQSTV.git
+QWSNM: https://github.com/qiuxuanzhizi/QWSNM.git
+MCNNFNM: https://github.com/wangzhi-swu/MCNNFNM.git
 SRRC: https://github.com/zt9877/SRRC.git
-WNNM: http://www4.comp.polyu.edu.hk/ cslzhang/papers.htm
 LRQA: https://www.fst.um.edu.mo/personal/wp-content/uploads/2021/05/LRQA.zip
 .
+QWSNM, MCNNFNM, SRRC, QWNNM, and LRQA are low-rank minimization models.
+QWSNM extends the WSNM into the quaternion domain for color restoration.
+MCNNFNM is a multi-channel nuclear norm minus Frobenius norm minimization model.
 SRRC exploits the weighted Schatten $p$-norm as the regularizer for the rank residual to obtain a new rank minimization model.
-WNNM uses the weighted nuclear norm as an alternative function to the rank function.
+QWNNM uses the weighted nuclear norm as an alternative function to the rank function.
 LRQA is a quaternion-based model that makes use of the low-rankness property by a nonconvex surrogate function.
-Compared with SLRQA, none of SRRC, WNNM, and LRQA exploits the sparsity feature of the image.
+Compared with SLRQA, none of SRRC, QWNNM, and LRQA exploit the sparsity feature of the image.
+pQSTV is a dictionary method for color image denoising that combines total variation regularization and pure quaternion representation.
 DWT [7] is a vanilla analysis-based approach (5) where low-rankness prior is not used and is implemented on our own.
-It is noted that SRRC, WNNM, and DWT are RGB-based methods, and LRQA and SLRQA are quaternion-based models.
+It is noted that MCNNFNM, SRRC, and DWT are monochromatic methods, while QWNNM, pQSTV, QWSNM, LRQA, and SLRQA are quaternion-based models.
 LRQA-1 and LRQA-2 denote the LRQA models using Schatten-$\gamma$ and Laplace function, respectively.
+The comparisons with LRQA and QWNNM can be regarded as ablation experiments to present the significance of sparsity and low-rankness prior respectively.
 For the compared methods, we use the default parameters stated in the corresponding papers.
@@
 It can be seen that SLRQA has better results than the other methods in terms of PSNR and SSIM in most cases.
-Furthermore, The PSNR value of SRRC, WNNM, and DWT are much lower than that of SLRQA when $\tau$ is large.
-Note that since NSS is not utilized in these methods, the results in Table 4 underscore the effectiveness and robustness of the SLRQA model and further validate the advanced nature and adaptability of its intrinsic architecture.
+Furthermore, the PSNR values of SRRC, QWNNM, and DWT are much lower than that of SLRQA when $\tau$ is large.
+Note that since NSS is not used in these methods, the results in Table 4 underscore the effectiveness and robustness of the SLRQA and further validate the advanced nature and adaptability of its intrinsic architecture.
 We also give the visual comparison between all competing methods and SLRQA as shown in Figure 4.
@@
 In particular, in terms of hair details, the images denoised by the DWT and LRQA-1 methods are still noisy.
-The images recovered by LRQA and SRRC methods contain blurry parts and hence the detail features of the images are lost.
-In contrast, the images recovered by SLRQA models have better results both numerically and visually.
+The images recovered by LRQA and SRRC methods contain blurry parts, and hence the detailed features of the images are lost.
+In contrast, the images recovered by SLRQAs have better results both numerically and visually.
 To further analyze the details of PL-ADMM, we plot the empirical convergence of PL-ADMM for Problem (88) on ten tested images with $\tau=30$.
@@
 We can observe that $\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}\rightarrow 0$.
-Hence, it follows from (42) that $|\|\widetilde{\mathbf{\dot{D}}}_{k+1}\||\rightarrow 0$.
-According to the update of $\mathbf{\dot{\Lambda}}_{k+1}$ and Lemma 4.7, it follows that $({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})$ converges to a stationary point of $\mathcal{L}_{\beta}$ and hence a KKT point of the original problem in (55).
+Hence, it follows that $({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})$ converges to a stationary point of $\mathcal{L}_{\beta}$ and hence a KKT point of the original problem in (88) [43, Theorem 2].
 Table 4: PSNR/SSIM results on color image denoising problem of different methods without NSS.
-Image Algorithm SRRC[39] WNNM[14] DWT[5] LRQA-1[10] LRQA-2[10] SLRQA-1 SLRQA-2 SLRQA-3
-$\tau=10$ Image1 29.23/0.70 28.06/0.73 27.67/0.61 28.48/0.61 29.21/0.65 31.55/0.76 31.45/0.78 32.86/0.81
-Image2 28.56/0.73 28.01/0.70 27.98/0.71 28.44/0.72 28.42/0.72 30.36/0.80 30.19/0.80 31.20/0.82
-Image3 28.03/0.60 27.81/0.61 28.21/0.61 28.49/0.64 28.56/0.65 31.33/0.77 31.79/0.78 32.03/0.84
-Image4 28.12/0.60 27.90/0.65 28.29/0.66 28.53/0.64 28.55/0.68 30.09/0.74 31.45/0.75 30.36/0.75
-Image5 28.82/0.66 28.72/0.64 28.67/0.64 28.85/0.65 28.71/0.67 32.13/0.85 31.18/0.80 32.50/0.87
-Image6 28.32/0.60 27.89/0.56 28.02/0.61 28.72/0.65 28.66/0.65 30.75/0.76 30.92/0.80 30.77/0.76
-Image7 28.45/0.62 27.32/0.58 28.02/0.59 28.65/0.60 29.04/0.94 30.95/0.73 30.82/0.75 31.10/0.80
-Image8 28.01/0.60 28.11/0.59 28.46/0.58 28.56/0.63 28.24/ 0.63 31.23/0.76 30.75/0.74 31.92/0.76
-Image9 28.02/0.74 27.98/0.73 27.62/0.71 28.82/0.76 29.10/0.76 30.14/0.70 29.66/0.78 30.80/0.79
-Image10 30.52/0.74 29.72/0.72 29.62/0.70 30.90/0.76 31.10/0.79 31.14/0.78 31.26/0.80 31.32/0.81
-$\tau=30$ Image1 20.32/0.30 20.41/0.25 20.36/0.28 20.49/0.27 20.80/0.30 25.47/0.66 24.30/0.61 26.40/0.67
-Image2 20.01/0.32 20.73/0.32 20.09/0.33 20.44/0.37 20.40/0.37 23.90/0.61 22.77/0.64 24.02/0.64
-Image3 00.12/0.21 20.24/0.22 20.55/0.25 20.50/0.27 21.10/0.31 26.15/0.65 25.21/0.64 26.22/0.68
-Image4 20.22/0.29 20.10/0.26 20.01/0.29 20.68/0.29 20.20/0.31 23.81/0.59 22.74/0.57 24.21/0.64
-Image5 21.02/0.32 20.58/0.33 20.65/0.37 21.23/0.30 21.30/0.33 26.60/0.68 25.32/0.58 26.34/0.68
-Image6 20.87/0.33 21.01/0.37 20.37/0.37 20.76/0.36 20.61/0.35 23.95/0.66 21.72/0.64 22.50/0.63
-Image7 20.33/0.29 20.71/0.29 20.94/0.22 20.93/0.26 20.21/0.21 25.12/0.63 24.39/0.68 26.42/0.68
-Image8 20.56/0.22 20.42/0.24 21.08/0.28 20.77/0.29 20.90/0.29 25.14/0.64 24.06/0.62 26.12/0.67
-Image9 24.82/0.58 24.22/0.58 23.62/0.55 24.90/0.58 25.10/0.61 25.30/0.64 25.45/0.68 25.80/0.68
-Image10 24.32/0.64 23.82/0.62 23.62/0.61 24.34/0.66 24.90/0.68 24.97/0.66 25.02/0.70 25.68/0.70
-$\tau=50$ Image1 17.92/0.17 17.01/0.14 18.01/0.18 17.66/0.17 17.30/0.18 23.33/0.50 23.50/0.51 23.90/0.51
-Image2 16.02/0.21 17.03/0.20 17.94/0.25 17.53/0.25 17.10/0.26 22.04/0.48 22.77/0.52 22.06/0.45
-Image3 17.75/0.15 17.40/0.21 17.24/0.19 17.56/0.17 17.90/0.18 24.21/0.51 24.91/0.58 24.99/0.60
-Image4 17.08/0.14 16.60/0.16 16.87/0.17 17.82/0.19 17.20/0.14 22.23/0.46 22.34/0.47 23.51/0.49
-Image5 18.04/0.17 18.08/0.17 18.31/0.18 18.39/0.19 18.30/0.17 21.84/0.52 21.62/0.67 22.72/0.65
-Image6 17.53/0.25 18.01/0.26 17.11/0.22 17.75/0.26 18.61/0.29 21.24/0.50 21.02/0.51 22.00/0.53
-Image7 17.60/0.17 18.21/0.18 17.11/0.13 18.08/0.17 18.71/0.17 23.57/0.48 23.09/0.45 23.92/0.53
-Image8 17.02/0.14 17.42/0.18 17.62/0.19 17.90/0.18 18.10/0.29 23.14/0.48 23.26/0.52 24.32/0.56
-Image9 22.02/0.51 21..92/0.49 21.62/0.46 22.34/0.52 22.67/0.55 22.78/0.56 23.02/0.62 23.04/0.63
-Image10 21.82/0.50 21.42/0.48 21.32/0.45 22.32/0.51 22.10/0.51 22.74/0.54 22.60/0.52 23.10/0.58
+Image Algorithm pQSTV[37] QWSNM[48] MCNNFNM[33] SRRC[49] QWNNM[15] DWT[7] LRQA-1[12] LRQA-2[12] SLRQA-1 SLRQA-2 SLRQA-3
+$\tau=10$ Image1 30.32/0.71 29.00/0.79 30.09/0.79 29.23/0.70 28.06/0.73 27.67/0.61 28.48/0.61 29.21/0.65 31.55/0.76 31.45/0.78 32.86/0.81
+Image2 29.21/0.80 30.10/0.79 28.10/0.75 28.56/0.73 28.01/0.70 27.98/0.71 28.44/0.72 28.42/0.72 30.36/0.80 30.19/0.80 31.20/0.82
+Image3 30.20/0.78 29.02/0.75 30.12/0.77 28.03/0.60 27.81/0.61 28.21/0.61 28.49/0.64 28.56/0.65 31.33/0.77 31.79/0.78 32.03/0.84
+Image4 28.56/0.70 29.20/0.72 28.90/0.69 28.12/0.60 27.90/0.65 28.29/0.66 28.53/0.64 28.55/0.68 30.09/0.74 31.45/0.75 30.36/0.75
+Image5 30.92/0.80 30.08/0.78 30.81/0.80 28.82/0.66 28.72/0.64 28.67/0.64 28.85/0.65 28.71/0.67 32.13/0.85 31.18/0.80 32.50/0.87
+Image6 28.09/0.70 30.01/0.76 29.40/0.71 28.32/0.60 27.89/0.56 28.02/0.61 28.72/0.65 28.66/0.65 30.75/0.76 30.92/0.80 30.77/0.76
+Image7 30.08/0.75 29.60/0.74 30.23/0.75 28.45/0.62 27.32/0.58 28.02/0.59 28.65/0.60 29.04/0.94 30.95/0.73 30.82/0.75 31.10/0.80
+Image8 29.20/0.70 30.91/0.71 30.70/0.75 28.01/0.60 28.11/0.59 28.46/0.58 28.56/0.63 28.24/ 0.63 31.23/0.76 30.75/0.74 31.92/0.76
+Image9 30.00/0.75 29.80/0.71 29.32/0.72 28.02/0.74 27.98/0.73 27.62/0.71 28.82/0.76 29.10/0.76 30.14/0.70 29.66/0.78 30.80/0.79
+Image10 30.86/0.75 30.90/0.75 31.02/0.76 30.52/0.74 29.72/0.72 29.62/0.70 30.90/0.76 31.10/0.79 31.14/0.78 31.26/0.80 31.32/0.81
+$\tau=30$ Image1 23.98/0.56 24.80/0.55 25.00/0.57 20.32/0.30 20.41/0.25 20.36/0.28 20.49/0.27 20.80/0.30 25.47/0.66 24.30/0.61 26.40/0.67
+Image2 22.80/0.55 23.01/0.57 21.32/0.55 20.01/0.32 20.73/0.32 20.09/0.33 20.44/0.37 20.40/0.37 23.90/0.61 22.77/0.64 24.02/0.64
+Image3 22.90/0.54 23.43/0.58 22.03/0.52 20.12/0.21 20.24/0.22 20.55/0.25 20.50/0.27 21.10/0.31 26.15/0.65 25.21/0.64 26.22/0.68
+Image4 22.39/0.43 23.09/0.55 23.10/0.55 20.22/0.29 20.10/0.26 20.01/0.29 20.68/0.29 20.20/0.31 23.81/0.59 22.74/0.57 24.21/0.64
+Image5 23.03/0.59 24.43/0.60 24.60/0.60 21.02/0.32 20.58/0.33 20.65/0.37 21.23/0.30 21.30/0.33 26.60/0.68 25.32/0.58 26.34/0.68
+Image6 21.30/0.59 21.01/0.55 22.30/0.60 20.87/0.33 21.01/0.37 20.37/0.37 20.76/0.36 20.61/0.35 23.95/0.66 21.72/0.64 22.50/0.63
+Image7 23.09/0.56 24.33/0.60 24.90/0.61 20.33/0.29 20.71/0.29 20.94/0.22 20.93/0.26 20.21/0.21 25.12/0.63 24.39/0.68 26.42/0.68
+Image8 23.33/0.57 24.01/0.58 24.40/0.61 20.56/0.22 20.42/0.24 21.08/0.28 20.77/0.29 20.90/0.29 25.14/0.64 24.06/0.62 26.12/0.67
+Image9 24.90/0.60 24.23/0.59 24.89/0.61 24.82/0.58 24.22/0.58 23.62/0.55 24.90/0.58 25.10/0.61 25.30/0.64 25.45/0.68 25.80/0.68
+Image10 23.90/0.60 23.84/0.55 24.01/0.62 22.32/0.60 23.82/0.62 23.62/0.61 24.34/0.66 24.90/0.68 24.97/0.66 25.02/0.70 25.68/0.70
+$\tau=50$ Image1 18.02/0.32 19.03/0.45 20.30/0.48 17.92/0.17 17.01/0.14 18.01/0.18 17.66/0.17 17.30/0.18 23.33/0.50 23.50/0.51 23.90/0.51
+Image2 18.30/0.40 19.02/0.45 18.46/0.42 16.02/0.21 17.03/0.20 17.94/0.25 17.53/0.25 17.10/0.26 22.04/0.48 22.77/0.52 22.06/0.45
+Image3 20.01/0.49 20.30/0.52 21.33/0.50 17.75/0.15 17.40/0.21 17.24/0.19 17.56/0.17 17.90/0.18 24.21/0.51 24.91/0.58 24.99/0.60
+Image4 20.30/0.39 21.02/0.42 20.80/0.40 17.08/0.14 16.60/0.16 16.87/0.17 17.82/0.19 17.20/0.14 22.23/0.46 22.34/0.47 23.51/0.49
+Image5 20.20/0.45 21.03/0.50 20.39/0.51 18.04/0.17 18.08/0.17 18.31/0.18 18.39/0.19 18.30/0.17 21.84/0.52 21.62/0.67 22.72/0.65
+Image6 19.32/0.48 19.80/0.50 20.20/0.48 17.53/0.25 18.01/0.26 17.11/0.22 17.75/0.26 18.61/0.29 21.24/0.50 21.02/0.51 22.00/0.53
+Image7 20.90/0.49 21.02/0.50 21.30/0.50 17.60/0.17 18.21/0.18 17.11/0.13 18.08/0.17 18.71/0.17 23.57/0.48 23.09/0.45 23.92/0.53
+Image8 21.93/0.50 22.03/0.49 22.30/0.51 17.02/0.14 17.42/0.18 17.62/0.19 17.90/0.18 18.10/0.29 23.14/0.48 23.26/0.52 24.32/0.56
+Image9 22.30/0.50 21.79/0.51 22.80/0.52 22.02/0.51 21..92/0.49 21.62/0.46 22.34/0.52 22.67/0.55 22.78/0.56 23.02/0.62 23.04/0.63
+Image10 22.01/0.50 21.90/0.48 22.34/0.55 21.82/0.50 21.42/0.48 21.32/0.45 22.32/0.51 22.10/0.51 22.75/0.54 22.59/0.52 23.11/0.58
 (a) Image5
 (b) Noisy Image
-(c) SRRC
-(d) WNNM
-(e) DWT
-(f) LRQA-1
-(g) LRQA-2
-(h) SLRQA-1
-(i) SLRQA-2
-(j) SLRQA-3
-Figure 4: Color image denoising results of different methods on “Image5” with $\tau=10$.
+(c) pQSTV
+(d) QWSNM
+(e) MCNNFNM
+(f) SRRC
+(g) QWNNM
+(h) DWT
+(i) LRQA-1
+(j) LRQA-2
+(k) SLRQA-1
+(l) SLRQA-2
+(m) SLRQA-3
+Figure 4: Color image denoising results of different methods on ”Image5” with $\tau=10$.
 Figure 5: The error convergence of SLRQA-1 for image denoising problems.
 5.3.2 Color image denoising with NSS
-For color image denoising problems with NSS, in addition to the methods compared in Section 6.3.1, we also compare SLRQA with CBM3D [11] [^4] The implementation is available from http://www.cs.tut.fi/~foi/GCF-BM3D..
+For color image denoising problems with NSS, in addition to the methods compared in Section 5.3.1, we also compare SLRQA with CBM3D [44] [^4] The implementation is available from https://webpages.tuni.fi/foi/GCF-BM3D/ released 30 June 2024..
 Instead of constructing a denoising model, CBM3D employs block matching and sparse 3D collaborative filtering techniques to denoise the image, a process that cannot be achieved without using NSS.
@@
 The visual comparisons of Image5 with $\tau=30$ are shown in Figure 6 as a typical example.
-It is shown in Figure 6 that the images recovered by CBM3D and WNNM methods are still noisy.
+It is shown in Figure 6 that the images recovered by CBM3D and QWNNM methods are still noisy.
 In addition, the images recovered by DWT, LRQA-1, and LRQA-2 are overly smooth, resulting in the loss of the features of the original images.
@@
 From the above numerical experiments, we can conclude that the SLRQA has better denoising performance compared to other methods with or without using NSS in the sense of PSNR and SSIM.
-Table 5: PSNR/SSIM results on color image denoising problem of different methods with NSS.
-Image Algorithm CBM3D[11] SRRC[39] WNNM[14] DWT[5] LRQA-1[10] LRQA-2[10] SLRQA-1 SLRQA-2 SLRQA-3
-$\tau=10$ Image1 35.71/0.92 35.67/0.92 34.85/0.91 34.62/0.91 35.19/0.92 35.34/0.92 35.69/0.92 35.57/0.92 35.86/0.93
-Image2 32.22/0.86 32.02/0.85 31.44/0.86 30.89/0.85 31.94/0.87 31.88/0.87 32.20/0.87 32.19/0.87 32.40/0.87
-Image3 34.67/0.88 34.72/0.88 34.37/0.88 34.22/0.88 34.63/0.88 34.72/0.88 34.75/0.88 34.79/0.88 35.03/0.88
-Image4 34.36/0.91 34.12/0.91 33.95/0.91 33.89/0.91 34.01/0.91 34.05/0.91 34.43/0.91 34.45/0.91 34.36/0.91
-Image5 37.08/0.95 36.82/0.94 35.94/0.95 35.87/0.94 36.48/0.95 36.71/0.95 37.10/0.95 37.18/0.95 37.50/0.95
-Image6 37.02/0.97 36.32/0.96 36.04/0.96 36.02/0.96 36.65/0.97 36.66/0.97 37.39/0.97 37.22/0.97 37.27/0.97
-Image7 34.25/0.91 34.01/0.89 33.80/0.90 33.92/0.90 34.25/0.90 34.24/0.90 34.74/0.91 34.82/0.91 35.00/0.91
-Image8 35.94/0.94 34.82/0.93 34.71/0.93 34.66/0.93 35.14/0.94 35.24/ 0.94 35.93/0.94 35.75/0.94 35.92/0.94
-Image9 33.25/0.81 32.81/0.79 32.60/0.79 32.32/0.75 33.75/0.80 34.14/0.88 34.24/0.90 34.92/0.89 35.01/0.90
-Image10 32.94/0.78 33.52/0.80 32.33/0.83 32.85/0.82 32.84/0.89 32.24/ 0.80 33.96/0.84 35.45/0.82 34.08/0.84
-$\tau=30$ Image1 30.82/0.85 30.76/0.84 30.41/0.85 30.36/0.85 30.66/0.86 31.00/0.86 31.30/0.87 31.30/0.87 31.40/0.88
-Image2 27.76/0.74 27.01/0.73 26.73/0.73 26.99/0.74 27.30/0.73 27.40/0.73 27.67/0.74 27.77/0.74 28.00/0.74
-Image3 31.02/0.81 30.82/0.80 30.74/0.80 30.55/0.80 30.93/0.81 31.10/0.81 31.15/0.81 31.21/0.81 31.22/0.81
-Image4 29.45/0.80 29.22/0.79 29.10/0.81 29.00/0.80 29.37/0.81 29.20/0.81 29.57/0.81 29.74/0.81 29.81/0.81
-Image5 32.21/0.85 31.02/0.85 31.08/0.85 30.65/0.84 31.34/0.85 31.30/0.85 31.50/0.86 31.32/0.85 31.34/0.85
-Image6 31.55/0.93 30.87/0.92 31.01/0.92 30.37/0.93 31.56/0.94 31.61/0.94 31.68/0.94 31.72/0.94 31.50/0.94
-Image7 29.23/0.80 28.81/0.79 28.71/0.79 28.94/0.79 29.09/0.80 29.21/0.80 29.32/0.80 29.39/0.80 29.42/0.80
-Image8 29.92/0.82 29.56/0.82 29.42/0.81 28.68/0.82 29.82/0.82 29.90/0.82 29.80/0.82 30.06/0.82 30.12/0.83
-Image9 29.25/0.80 29.01/0.79 28.80/0.74 28.12/0.69 29.05/0.80 29.24/0.80 29.74/0.81 29.82/0.81 30.02/0.82
-Image10 30.94/0.94 30.82/0.93 30.71/0.93 30.66/0.93 31.14/0.94 31.34/ 0.94 31.93/0.94 31.75/0.94 31.92/0.94
-$\tau=50$ Image1 28.32/0.82 27.92/0.81 28.01/0.81 28.15/0.82 28.31/0.82 28.30/0.82 28.60/0.82 28.50/0.82 28.60/0.82
-Image2 25.23/0.67 25.02/0.66 25.03/0.66 24.94/0.65 25.22/0.67 25.10/0.67 25.47/0.67 25.77/0.67 25.60/0.67
-Image3 28.82/0.28 28.75/0.77 28.74/0.76 28.24/0.75 28.85/0.77 28.90/0.77 28.95/0.78 28.91/0.78 28.99/0.78
-Image4 27.12/0.74 27.08/0.74 26.60/0.74 26.87/0.74 27.06/0.74 27.20/0.74 27.27/0.75 27.34/0.75 27.51/0.75
-Image5 28.56/0.76 28.34/0.76 28.08/0.76 28.01/0.76 28.41/0.77 28.30/0.77 28.50/0.77 28.62/0.78 28.72/0.78
-Image6 28.77/0.90 28.53/0.90 28.01/0.89 28.11/0.90 28.52/0.92 28.61/0.92 28.98/0.93 29.02/0.93 29.00/0.93
-Image7 26.53/0.75 26.43/0.75 26.21/0.74 26.11/0.74 26.74/0.75 26.71/0.74 26.82/0.75 26.89/0.75 26.92/0.75
-Image8 28.12/0.72 27.90/0.71 27.42/0.71 27.62/0.72 28.02/0.72 28.10/0.73 28.10/0.73 28.26/0.73 28.32/0.73
-Image9 26.65/0.66 26.71/0.66 26.80/0.68 26.02/0.63 27.05/0.69 27.24/0.70 27.34/0.70 27.62/0.71 28.70/0.74
-Image10 27.34/0.70 27.02/0.68 26.81/0.64 26.67/0.62 27.34/0.65 27.64/ 0.70 27.73/0.69 27.75/0.70 28.02/0.72
+Table 5: PSNR/SSIM results on color image denoising problem of different methods with NSS. (To highlight the gap between methods, we keep three decimal digits for SSIM.)
+Image Algorithm pQSTV [37] QWSNM [48] MCNNFNM [33] CBM3D [44] SRRC [49] QWNNM [20] DWT [7] LRQA-1 [12] LRQA-2 [12] SLRQA-1 SLRQA-2 SLRQA-3
+$\tau=10$ Image1 34.78/0.917 35.48/0.921 35.10/0.913 35.71/0.920 35.67/0.919 34.16/0.895 34.62/0.913 35.19/0.919 35.34/0.922 35.68/0.924 35.58/0.921 35.86/0.926
+Image2 30.84/0.862 31.98/0.844 31.25/0.839 32.22/0.855 32.02/0.849 31.37/0.862 30.89/0.850 31.94/0.868 31.88/0.866 32.13/0.865 32.19/0.869 32.40/0.872
+Image3 36.85/0.934 38.50/0.975 37.89/0.960 37.67/0.948 37.72/0.943 36.37/0.940 36.22/0.934 37.51/0.955 38.62/0.972 38.71/0.973 38.79/0.978 38.83/0.984
+Image4 39.79/0.969 40.10/0.975 39.51/0.956 39.36/0.952 39.12/0.948 38.24/0.936 38.89/0.946 39.67/0.955 40.87/0.972 40.99/0.974 41.05/0.976 40.36/0.976
+Image5 36.49/0.920 37.05/0.951 36.29/0.943 37.08/0.951 36.82/0.944 35.00/0.921 35.87/0.942 36.48/0.947 36.71/0.945 37.01/0.949 37.18/0.951 37.50/0.954
+Image6 35.56/0.924 37.12/0.962 37.08/0.963 37.02/0.969 36.32/0.962 35.42/0.942 36.02/0.959 36.65/0.966 36.66/0.966 37.37/0.969 37.22/0.968 37.27/0.969
+Image7 34.11/0.901 35.08/0.916 34.66/0.907 34.25/0.905 34.01/0.894 33.69/0.882 33.92/0.899 34.25/0.904 34.24/0.904 34.68/0.904 34.82/0.905 35.00/0.906
+Image8 34.27/0.923 35.77/0.938 35.46/0.934 35.94/0.942 34.82/0.932 34.25/0.910 34.66/0.928 35.14/0.935 35.24/ 0.935 35.84/0.941 35.75/0.940 35.92/0.942
+Image9 34.15/0.829 34.83/0.870 33.37/0.826 33.25/0.805 32.81/0.794 32.60/0.788 32.32/0.749 33.75/0.804 34.14/0.884 34.24/0.895 34.92/0.892 35.01/0.902
+Image10 33.05/0.812 33.32/0.826 32.80/0.811 32.94/0.780 33.52/0.802 32.84/0.833 32.85/0.818 32.84/0.893 32.24/ 0.804 33.50/0.836 35.45/0.820 34.08/0.842
+$\tau=30$ Image1 30.35/0.801 30.21/0.767 29.90/0.817 30.82/0.854 30.76/0.844 30.41/0.854 30.36/0.848 30.66/0.859 31.00/0.862 31.30/0.871 31.30/0.871 31.40/0.877
+Image2 27.28/0.722 27.71/0.743 26.67/0.710 27.76/0.740 27.01/0.729 26.73/0.729 26.99/0.737 27.30/0.730 27.40/0.730 27.67/0.740 27.77/0.741 28.00/0.743
+Image3 30.05/0.818 30.12/0.802 30.67/0.801 31.02/0.810 30.82/0.801 30.74/0.802 30.55/0.801 30.93/0.811 31.10/0.813 31.15/0.813 31.21/0.814 31.22/0.814
+Image4 29.26/0.764 29.52/0.770 28.85/0.767 29.45/0.802 29.22/0.791 29.10/0.806 29.00/0.798 29.37/0.805 29.203/0.81 29.57/0.806 29.74/0.807 29.81/0.809
+Image5 30.64/0.846 30.02/0.832 30.35/0.816 32.21/0.847 31.02/0.846 31.08/0.846 30.65/0.837 31.34/0.853 31.30/0.853 31.50/0.859 31.32/0.853 31.34/0.854
+Image6 31.35/0.902 30.19/0.823 30.90/0.898 31.55/0.932 30.87/0.923 31.01/0.922 30.37/0.927 31.56/0.941 31.61/0.941 31.68/0.940 31.72/0.941 31.50/0.937
+Image7 28.72/0.762 29.02/0.773 28.99/0.778 29.23/0.801 28.81/0.792 28.71/0.790 28.94/0.794 29.09/0.799 29.21/0.801 29.32/0.801 29.39/0.801 29.42/0.802
+Image8 30.04/0.808 30.29/0.834 29.36/0.802 29.92/0.820 29.56/0.815 29.42/0.814 28.68/0.817 29.82/0.819 29.90/0.822 29.80/0.818 30.06/0.823 30.12/0.827
+Image9 29.00/0.770 29.12/0.782 28.87/0.712 29.25/0.795 29.01/0.794 28.80/0.738 28.12/0.690 29.05/0.799 29.24/0.804 29.74/0.806 29.82/0.805 30.02/0.816
+Image10 29.95/0.925 28.71/0.845 29.19/0.713 30.94/0.942 30.82/0.932 30.71/0.929 30.66/0.928 31.14/0.935 31.34/ 0.935 31.93/0.942 31.75/0.940 31.92/0.942
+$\tau=50$ Image1 28.07/0.816 28.21/0.767 27.33/0.731 28.32/0.815 27.92/0.810 28.01/0.814 28.15/0.815 28.31/0.818 28.30/0.818 28.60/0.821 28.50/0.819 28.60/0.819
+Image2 25.57/0.631 25.36/0.625 24.66/0.630 25.23/0.670 25.02/0.661 25.03/0.660 24.94/0.653 25.22/0.666 25.10/0.666 25.47/0.669 25.77/0.674 25.60/0.670
+Image3 28.55/0.728 28.95/0.782 28.57/0.779 28.82/0.778 28.75/0.770 28.74/0.761 28.24/0.754 28.85/0.769 28.90/0.773 28.95/0.781 28.91/0.781 28.99/0.784
+Image4 27.01/0.736 27.15/0.712 26.67/0.707 27.12/0.739 27.08/0.740 26.60/0.736 26.87/0.737 27.06/0.742 27.20/0.741 27.27/0.746 27.34/0.746 27.51/0.749
+Image5 27.06/0.604 27.82/0.760 27.47/0.697 28.56/0.762 28.34/0.760 28.08/0.756 28.01/0.755 28.41/0.769 28.30/0.767 28.50/0.769 28.62/0.776 28.72/0.775
+Image6 28.90/0.907 28.84/0.912 27.81/0.823 28.77/0.900 28.53/0.895 28.01/0.892 28.11/0.899 28.52/0.916 28.61/0.918 28.98/0.930 29.02/0.931 29.00/0.931
+Image7 26.04/0.681 26.19/0.692 26.55/0.683 26.53/0.749 26.43/0.745 26.21/0.738 26.11/0.737 26.74/0.747 26.71/0.744 26.82/0.750 26.89/0.751 26.92/0.753
+Image8 27.91/0.712 27.74/0.721 26.73/0.702 28.12/0.723 27.90/0.714 27.42/0.714 27.62/0.719 28.02/0.723 28.10/0.729 28.10/0.730 28.26/0.732 28.32/0.732
+Image9 27.63/0.664 27.90/0.687 26.19/0.650 26.65/0.655 26.71/0.664 26.80/0.682 26.02/0.629 27.05/0.694 27.24/0.700 27.34/0.701 27.62/0.705 28.70/0.736
+Image10 27.07/0.665 27.71/0.702 27.06/0.646 27.34/0.702 27.02/0.682 26.81/0.639 26.67/0.618 27.34/0.652 27.64/ 0.701 27.73/0.687 27.75/0.702 28.02/0.721
 (a) Image5
 (b) Noisy Image
-(c) CBM3D
-(d) SRRC
-(e) WNNM
-(f) DWT
+(c) pQSTV
+(d) QWSNM
+(e) MCNNFNM
+(f) CBM3D
+(g) SRRC
+(h) QWNNM
+(i) DWT
+(j) LRQA-1
+(k) LRQA-2
+(l) SLRQA-1
+(m) SLRQA-2
+(n) SLRQA-3
+Figure 6: Color image denoising results of tested methods on “Image5” with $\tau=30$.
+5.4 Color Image Inpainting
+For color image inpainting problems, the compared methods include Bilinear Factorization (BF) methods such as NCALR [25], and Rank Minimization (RM) methods such as SLRI [26], LRQA [12], and QWNNM [20].
+555 The implementations that we use are available from:
+LRQA: https://www.fst.um.edu.mo/personal/wp-content/uploads/2021/05/LRQA.zip
+QWNNM: https://github.com/Huang-chao-yan/QWNNM
+NCARL: https://github.com/hyzhang98/NCARL.
+We note that the BF methods require a rank estimate as a prior.
+For fairness, the SLRI is extended to quaternion representation, which is referred to as QSLRI and performs better than the SLRI method.
+To fully demonstrate the performance of NCARL, we choose the best PSNR and SSIM from the initial rank range [50, 80, 100, 120, 150].5080100120150[50,\,80,\,100,\,120,\,150].[ 50 , 80 , 100 , 120 , 150 ] .
+The comparisons with LRQA and QSLRI can be regarded as ablation experiments to present the significance of sparsity and low-rankness prior respectively.
+The results of all tested algorithms are shown in Table 6.
+It is shown therein that SLRQA-NF-3 always enjoys the best performance compared with other methods in terms of PSNR and SSIM.
+The visual comparisons between SLRQA-NF and all competing inpainting methods on the Image2 with $\chi=0.7$ is shown in Figure 7 as a typical example.
+Although the NCARL method completes the image, the details of the recovered image are lost, and the recovered images are blurry.
+The images recovered by LRQA-1 and LRQA-2 methods have a relatively good effect, but the images are still slightly noisy or blurry.
+In addition, the images recovered by QWNNM are still little noisy.
+In contrast, SLRQA-NF methods all demonstrate considerable inpainting results and have the highest PSNR and SSIM values.
+(a) Image2
+(b) Corrupted
+(c) NCARL
+(d) QWSNM
+(e) QWNNM
+(f) QSLRI
 (g) LRQA-1
 (h) LRQA-2
-(i) SLRQA-1
-(j) SLRQA-2
-(k) SLRQA-3
-Figure 6: Color image denoising result of tested methods on “Image5” with $\tau=30$.
-6.4 Color Image Inpainting
-For color image inpainting problems, the compared methods include Bilinear Factorization (BF) methods such as LMafit [31], MCNMF [33], and Rank Minimization (RM) methods such as LADM [21], LRQA [10], and WNNM [13]555 The implementations that we use are available from:
-LMafit: https://github.com/optsuite/LMaFit
-MCNMF: https://xu-yangyang.github.io/codes/MCNF.zip
-LRQA: https://www.fst.um.edu.mo/personal/wp-content/uploads/2021/05/LRQA.zip
-WNNM: https://www4.comp.polyu.edu.hk/cslzhang/code/WNNM_MC_code.zip
-.
-We note that the BF methods require a rank estimate as a prior.
-For fairness, the LADM is extended to quaternion representation, which is referred to as QLADM and performs better than the LADM method.
-To fully demonstrate the performance of LMaFit and MCNMF, we choose the best PSNR and SSIM from the initial rank range [50, 80, 100, 120, 150].5080100120150[50,\,80,\,100,\,120,\,150].[ 50 , 80 , 100 , 120 , 150 ] .
-The results of all tested algorithms are shown in Table 6.
-It is shown therein that SLRQA-NF-3 always enjoys the best performance compared with other methods in terms of PSNR and SSIM.
-The visual comparisons between SLRQA-NF and all competing inpainting methods on the Image2 with $\chi=0.7$ are shown in Figure 7 as a typical example.
-Although the LMaFit and MCNMF methods complete the image, the details of the recovered image are lost, and the recovered images are blurry.
-The images recovered by LRQA-1 and LRQA-2 methods have a relatively good effect, but the images are still slightly noisy or blurry.
-In addition, the images recovered by WNNM are noisy.
-In contrast, SLRQA-NF methods all demonstrate considerable inpainting results and have the highest PSNR and SSIM values.
-(a) Image2
-(b) Corrupted image
-(c) WNNM
-(d)LMaFit
-(e) MCNMF
-(f) QLADM
-(g) LRQA-1
-(h) LRQA-2
-(i) SLRQA-NF-1
-(j) SLRQA-NF-2
-(k) SLRQA-NF-3
+(i) SLRQA-NF1
+(j) SLRQA-NF2
+(k) SLRQA-NF3
 Figure 7: Color image inpainting result of tested methods on “Image2” with $\chi=0.7$.
+To give a more detailed analysis, the errors of PL-ADMM-NF are also demonstrated for tested images.
+Figure 8 shows the empirical convergence results of SLRQA-NF-1, which indicate that even though the errors fluctuate at the early stage of iteration, they all converge in later iterations.
+We can observe that $\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1,1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1,2}\|_{\mathrm{F}}\rightarrow 0$.
+Hence, it follows from (87) that $|\|\widetilde{\mathbf{\dot{D}}}_{k+1}\||\rightarrow 0$.
+According to the update of $\mathbf{\dot{\Lambda}}_{k+1}$ and Lemma 4.16, it follows that $({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}}_{1},\mathbf{\dot{\Lambda}}_{2})$ converges to a stationary point of $\mathcal{L}_{\beta_{1},\beta_{2}}$ and hence a KKT point of the original problem in (89).
 Table 6: PSNR/SSIM results on image inpainting problem of different methods.
-Image Algorithm LMaFit[31] WNNM [13] MCNMF [33] QLADM[21] LRQA-1 [10] LRQA-2 [10] SLRQA-NF-1 SLRQA-NF-2 SLRQA-NF-3
-$\chi=0.5$ Image1 24.69/0.74 25.25/0.60 29.35/0.83 31.16/0.95 30.87/0.89 30.25/0.87 32.32/0.94 31.26/0.94 32.47/0.94
-Image2 22.14/0.80 23.55/0.81 24.22/0.86 30.89/0.85 26.63/0.91 26.20/0.90 28.04/0.94 27.54/0.94 28.16/0.94
-Image3 23.43/0.80 25.23/0.75 27.05/0.88 29.65/0.97 28.37/0.90 27.94/0.89 30.40/0.96 29.65/0.97 30.60/0.97
-Image4 29.09/0.91 31.16/0.85 34.73/0.96 36.39/0.99 35.99/0.97 35.51/0.97 37.69/0.99 36.27/0.98 37.90/0.99
-Image5 26.79/0.85 30.09/0.87 30.02/0.92 33.73/0.98 33.27/0.96 32.72/0.95 34.82/0.98 33.75/0.98 34.98/0.98
-Image6 19.64/0.61 22.01/0.61 22.04/0.87 26.56/0.92 25.27/0.79 25.72/0.80 28.25/0.93 26.62/0.91 27.97/0.92
-Image7 25.00/0.81 24.96/0.75 25.14/0.82 28.84/0.94 27.59/0.87 27.86/0.88 27.82/0.93 28.92/0.94 29.33/0.94
-Image8 24.18/0.87 25.17/0.86 27.21/0.92 30.03/0.97 29.62/0.95 29.06/0.94 31.06/0.97 30.16/0.97 31.22/0.97
-Image9 24.93/0.94 26.02/0.94 27.04/0.95 30.99/0.98 30.12/0.98 29.56/0.97 31.58/0.98 31.09/0.98 31.71/0.98
-Image10 24.36/0.93 26.18/0.94 23.73/0.92 30.69/0.98 29.86/0.98 29.20/0.97 31.28/0.98 30.84/0.98 31.43/0.98
-$\chi=0.7$ Image1 22.32/0.66 23.05/0.52 24.40/0.67 27.29/0.89 25.97/0.74 25.47/0.72 28.40/0.88 27.30/0.88 28.62/0.89
-Image2 20.19/0.73 20.13/0.67 21.52/0.77 24.19/0.89 22.51/0.79 22.44/0.79 25.05/0.90 24.30/0.89 25.21/0.90
-Image3 21.20/0.72 21.50/0.63 22.78/0.75 26.09/0.93 23.69/0.77 23.67/0.77 26.87/0.93 26.00/0.92 27.10/0.94
-Image4 25.87/0.85 28.34/0.80 28.50/0.89 31.89/0.97 31.07/0.93 30.44/0.92 33.37/0.97 31.80/0.97 33.63/0.98
-Image5 24.25/0.78 25.75/0.77 26.19/0.83 29.58/0.95 28.53/0.88 27.85/0.86 30.95/0.96 29.68/0.95 31.14/0.96
-Image6 17.49/0.52 18.13/0.44 19.66/0.67 22.78/0.82 20.28/0.57 19.86/0.55 23.78/0.81 22.70/0.80 24.00/0.83
-Image7 22.59/0.74 22.40/0.67 21.17/0.82 26.24/0.90 24.62/0.78 24.58/0.78 26.63/0.88 26.19/0.89 26.57/0.88
-Image8 21.79/0.81 22.34/0.79 23.17/0.84 26.34/0.94 24.87/0.87 24.54/0.87 27.41/0.94 26.42/0.94 27.60/0.95
-Image9 22.66/0.91 23.33/0.90 24.77/0.93 27.83/0.97 26.13/0.94 25.69/0.94 28.56/0.97 27.81/0.97 28.74/0.97
-Image10 21.55/0.89 23.22/0.90 24.18/0.92 27.54/0.96 25.95/0.95 25.34/0.94 28.46/0.97 27.58/0.96 28.64/0.97
-$\chi=0.8$ Image1 21.13/0.59 20.77/0.43 19.95/0.44 25.05/0.85 22.87/0.61 22.76/0.60 26.08/0.84 24.54/0.79 26.30/0.85
-Image2 19.32/0.68 17.72/0.54 17.87/0.58 22.34/0.84 19.44/0.65 20.27/0.69 23.36/0.86 22.25/0.84 23.53/0.86
-Image3 20.26/0.65 18.71/0.53 18.93/0.59 23.98/0.90 20.52/0.65 21.35/0.69 24.91/0.90 23.61/0.88 25.13/0.91
-Image4 24.62/0.79 25.48/0.72 23.53/0.75 29.09/0.95 27.75/0.87 27.20/0.85 30.74/0.96 28.73/0.94 31.04/0.96
-Image5 23.37/0.74 22.52/0.65 21.25/0.66 27.10/0.93 25.53/0.80 25.05/0.78 28.70/0.93 27.09/0.92 28.97/0.94
-Image6 16.86/0.46 15.41/0.30 15.94/0.34 20.82/0.74 17.93/0.45 14.65/0.31 21.61/0.74 20.10/0.70 21.82/0.76
-Image7 21.53/0.69 20.67/0.60 19.35/0.57 24.79/0.87 22.81/0.71 22.64/0.71 25.59/0.85 24.52/0.86 25.43/0.86
-Image8 20.53/0.77 19.88/0.71 19.22/0.71 24.26/0.91 21.69/0.79 21.93/0.80 25.30/0.92 23.90/0.90 25.50/0.92
-Image9 21.93/0.90 20.78/0.84 21.23/0.85 25.89/0.95 23.43/0.91 23.24/0.91 26.74/0.96 25.56/0.95 26.98/0.96
-Image10 20.47/0.87 20.48/0.84 20.34/0.85 25.52/0.95 23.10/0.90 22.85/0.90 26.67/0.96 25.32/0.95 26.90/0.96
+Image Algorithm NCALR [25] QWSNM [48] QWNNM  [20] QSLRI [26] LRQA-1  [12] LRQA-2 [12] SLRQA-NF-1 SLRQA-NF-2 SLRQA-NF-3
+$\chi=0.5$ Image1 26.48/0.85 28.16/0.86 25.25/0.60 31.16/0.95 30.87/0.89 30.25/0.87 32.32/0.94 31.26/0.94 32.48/0.94
+Image2 25.55/0.90 24.56/0.87 23.55/0.81 30.89/0.85 26.63/0.91 26.20/0.90 28.04/0.94 27.54/0.94 28.17/0.94
+Image3 26.75/0.91 27.87/0.90 25.23/0.75 29.65/0.97 28.37/0.90 27.94/0.89 30.40/0.96 29.65/0.97 30.61/0.97
+Image4 28.47/0.93 28.56/0.92 31.16/0.85 36.39/0.99 35.99/0.97 35.51/0.97 37.69/0.99 36.27/0.98 37.91/0.99
+Image5 33.46/0.94 34.23/0.91 30.09/0.87 33.73/0.98 33.27/0.96 32.72/0.95 34.82/0.98 33.75/0.98 34.98/0.98
+Image6 30.26/0.90 31.03/0.87 22.01/0.61 26.56/0.92 25.27/0.79 25.72/0.80 28.25/0.93 26.62/0.91 27.97/0.92
+Image7 27.48/0.91 28.65/0.92 24.96/0.75 28.84/0.94 27.59/0.87 27.86/0.88 27.82/0.93 28.92/0.94 29.33/0.94
+Image8 28.12/0.93 29.42/0.94 25.17/0.86 30.03/0.97 29.62/0.95 29.06/0.94 31.06/0.97 30.16/0.97 31.22/0.97
+Image9 29.83/0.94 28.89/0.95 26.02/0.94 30.99/0.98 30.12/0.98 29.56/0.97 31.58/0.98 31.09/0.98 31.71/0.98
+Image10 29.47/0.95 28.43/0.93 26.18/0.94 30.69/0.98 29.86/0.98 29.20/0.97 31.28/0.98 30.84/0.98 31.42/0.98
+$\chi=0.7$ Image1 27.39/0.84 27.41/0.85 23.05/0.52 27.29/0.89 25.97/0.74 25.47/0.72 28.40/0.88 27.30/0.88 28.62/0.89
+Image2 25.66/0.87 24.11/0.85 20.13/0.67 24.19/0.89 22.51/0.79 22.44/0.79 25.05/0.90 24.30/0.89 25.22/0.90
+Image3 26.44/0.93 24.12/0.92 21.50/0.63 26.09/0.93 23.69/0.77 23.67/0.77 26.87/0.93 26.00/0.92 27.10/0.94
+Image4 31.32/0.96 30.47/0.95 28.34/0.80 31.89/0.97 31.07/0.93 30.44/0.92 33.37/0.97 31.80/0.97 33.63/0.98
+Image5 26.55/0.91 27.46/0.92 25.75/0.77 29.58/0.95 28.53/0.88 27.85/0.86 30.95/0.96 29.68/0.95 31.14/0.96
+Image6 23.77/0.74 20.57/0.75 18.13/0.44 22.78/0.82 20.28/0.57 19.86/0.55 23.78/0.81 22.70/0.80 24.00/0.83
+Image7 24.01/0.83 25.43/0.85 22.40/0.67 26.24/0.90 24.62/0.78 24.58/0.78 26.63/0.88 26.19/0.89 26.57/0.88
+Image8 26.08/0.93 25.45/0.92 22.34/0.79 26.34/0.94 24.87/0.87 24.54/0.87 27.41/0.94 26.42/0.94 27.60/0.95
+Image9 22.36/0.87 24.15/0.93 23.33/0.90 27.83/0.97 26.13/0.94 25.69/0.94 28.56/0.97 27.81/0.97 28.74/0.97
+Image10 23.14/0.91 23.76/0.92 23.22/0.90 27.54/0.96 25.95/0.95 25.34/0.94 28.46/0.97 27.58/0.96 28.64/0.97
+$\chi=0.8$ Image1 19.43/0.79 23.16/0.81 20.77/0.43 25.05/0.85 22.87/0.61 22.76/0.60 26.08/0.84 24.54/0.79 26.30/0.85
+Image2 17.65/0.80 22.47/0.85 17.72/0.54 22.34/0.84 19.44/0.65 20.27/0.69 23.36/0.86 22.25/0.84 23.53/0.86
+Image3 22.65/0.84 23.45/0.86 18.71/0.53 23.98/0.90 20.52/0.65 21.35/0.69 24.91/0.90 23.61/0.88 25.13/0.91
+Image4 28.41/0.93 27.48/0.92 25.48/0.72 29.09/0.95 27.75/0.87 27.20/0.85 30.74/0.96 28.73/0.94 31.04/0.96
+Image5 26.08/0.88 26.45/0.89 22.52/0.65 27.10/0.93 25.53/0.80 25.05/0.78 28.70/0.93 27.09/0.92 28.97/0.94
+Image6 18.04/0.65 19.76/0.74 15.41/0.30 20.82/0.74 17.93/0.45 14.65/0.31 21.61/0.74 20.10/0.70 21.82/0.76
+Image7 22.18/0.80 23.16/0.79 20.67/0.60 24.79/0.87 22.81/0.71 22.64/0.71 25.60/0.85 24.53/0.86 25.42/0.86
+Image8 21.38/0.84 24.74/0.90 19.88/0.71 24.26/0.91 21.69/0.79 21.93/0.80 25.31/0.92 23.91/0.90 25.50/0.92
+Image9 24.07/0.89 25.19/0.91 20.78/0.84 25.89/0.95 23.43/0.91 23.24/0.91 26.74/0.96 25.56/0.95 26.97/0.96
+Image10 22.92/0.90 23.79/0.92 20.48/0.84 25.52/0.95 23.10/0.90 22.85/0.90 27.67/0.96 25.33/0.95 26.91/0.96
+Figure 8: The error convergence of SLRQA-NF for image inpainting problems.
 6 Conclusion
-In this paper, we propose a novel SLRQA model for the color image process problems.
-Different from most existing models which only consider one or two properties from low-rankness, sparsity, and quaternion representation, SLRQA utilizes these properties all.
-Furthermore, SLRQA does not need an initial rank esamite.
-A PL-ADMM algorithm is proposed to solve the SLRQA model.
-Furthermore, we prove that any limit point of PL-ADMM for SLRQA is a stationary point under mild assumptions.
-In addition, the global convergence analysis of the sequence generated by PL-ADMM under the KL-inequality assumption is presented.
-When the observation is noise-free, a SLRQA-NF model of the limiting case of SLRQA model is proposed.
-Subsequently, a nonconvex PL-ADMM-NF algorithm is also proposed to solve the SLRQA-NF model.
+In this paper, we propose a novel SLRQA for the color image process problems.
+Different from most existing models which only consider one or two properties from low-rankness, sparsity, and quaternion representation, SLRQA uses these properties all.
+Furthermore, SLRQA does not need an initial rank estimate.
+A PL-ADMM algorithm is proposed to solve the SLRQA and its global convergence is guaranteed.
+When the observation is noise-free, an SLRQA-NF of the limiting case of SLRQA is proposed.
+Subsequently, a PL-ADMM-NF algorithm is also proposed to solve the SLRQA-NF.
+Under a newly proposed assumption, the global convergence of PL-ADMM-NF is established.
+To the best of our knowledge, this is the first ADMM-type algorithm without the “range assumption” and still guarantees global convergence.
 Extensive experiments for color image denoising and inpainting problems have demonstrated the robustness and effectiveness of SLRQA and SLRQA-NF.
+Data availability statement
+The data that support the findings of this study are openly available at the following URL/DOI: https://github.com/dengzhanwang/SLRQA/tree/main.